A Complete Bibliography of Publications in Electronic Transactions on Numerical Analysis (ETNA)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.20

Title word cross-reference

(0, +∞) [DO18]. (0, 2) [Lén06]. (1, 1) [GS06].

(l) [SF93]. (n) [BR08]. 1 [Bar06, NP20, XS20]. 2 [BLM06, CP03, HKKR18, Mar07, Mar11]. 3 [BLM06, BB01, BRC^{+}16, BHRT03, TV07]. 3500 [JC06]. 9 [Kou97]. [−1, 1] [BCGAM08]. 1ψ_{1} [Sch06]. _{δ}^{γ}Φ [Lei19]. _{q}(1, 1) [Arv06]. _{q}(2) [Arv06]. A [AR07]. A, V [AR07]. α [JN07, VC14]. C^{0} [BW12]. C^{n} [PP06]. e^{z} [VCL08]. ℓ^{p} [CRZT20]. ℓ_{1} [DdSV20]. exp(−τA)b [SSS10]. F [GN16, Zha04]. f(A)b [MR17]. H(curl) [Cal15]. H(div) [Hip97]. H − LU [LO07]. H_{2}([0, T], μ, R^{d}) [Sch03]. H_{∞} [CKP15]. k [MSV95, PQRT06]. L [CHR02, KR20]. L^{2} [Bar06]. l_{1} [Ger17]. Λ [COSV93]. M [BU07]. R^{4} [JO14]. MR^{3} [WL11, WL12]. N [ÁNCQ06]. N(II) [BMSR06]. N ≤ 4 [Meu12]. P [ZB10, Car11, Car12, Pry14]. ψ [AR07]. Q [MM99, Arv06, AK06, EMJ16, IL06, Ost10, Psa04, Roz06]. q > 1 [Ost10]. Q_{1} [MC05]. qd [WL11]. QR [AMVW15, FXG19]. QZ [VW13]. R [Lub06]. R^{3} [DGH14]. R_{d} [She04]. s [IE17]. Θ [BB00]. U_{q}(so(5)) [Roz06]. V [Zha04]. x^{T} A^{−1}y [FMR15a]. Z [BHRT03].

-Bernstein [Ost10]. -biharmonic [Pry14]. -classical [MM99]. -continued [IL06]. -control [CKP15]. -curve [CHR02, KR20]. -cycle [Zha04]. -D [BRC^{+}16, XS20]. -eigenpairs [CRZT20]. -fractal [VC14]. -Hahn [Arv06]. -interpolation [EMJ16].

-matrices [BU07]. -methods [BB00]. -numerical [Psa04]. -orthogonal [Roz06]. -pinch [BHRT03]. -point [Kou97]. -polynomials [AK06]. -regular [ZB10]. -regularization [Ger17]. -smoothing [GN16]. -step [IE17, MSV95]. -th [ÁNCQ06]. -type [DdSV20, Lén06, WL11, Lei19].

/9 [dJvdPDV17].

4th [Mar07, Mar11].

5- [dJvdPDV17]. 5-/9-point [dJvdPDV17].

75th [GOR^{+}08].

9-point [dJvdPDV17].

a-posteriori [Gat04]. Abel [MG14, Mok15, Mok16]. Abel-type [MG14]. abstracts [Ano09a, Ano10b, Ano13a, Ano14c, Ano15b, Ano15d, Ano16, Ano17b, Ano17c, Ano18c, Ano19a, Ano20c]. accelerated [BGP04, Che99, DdSV20, SSVS12]. accelerating [WNK18]. Acceleration [AS09, AS05a, BK13, Lóp16, Val19, WO97]. Accumulation [Ore10, NU05]. accuracy [Ful03, Ves96]. accurate [Bra00, EBJ21, MACI06, TC08, WLR02]. ACMS [HKKR18]. acoustic [HL20]. action [AS05a]. adapted [Val19]. Adaptive [BFK^{+}10, DDS^{+}18, JOT08, KKR18, MSV95, Wac13, BMS08a, BMM97, BRS08, CW16, FFK^{+}15, Ful97, JG14, KCW18, KRR16, MR18, NP96, PD17, KNS19]. Addendum [BBF^{+}00a]. Additive [BGyS21, MR18, Bre03, DHI03, FSS97, Han06, Mar07, MP08, Par21]. Addressing [BC05]. ADI [BMS08b, BKS15, Kür19]. adjoint [BFM12, GS15, GHS14, KK08, Neu16]. admissible [DTM17]. ADMM [FTRH20].

ADMM-Softmax [FTRH20]. Advancing [PLM03]. against [BP07]. aggregation [BDH15, Ema10, MMP09, NP19b, Not10]. aggregation-based [NP19b, Not10]. aggregation/disaggregation [MMP09]. aggressive [BFH^{+}15, TV18]. ahead [GH94]. air [LHL^{+}20]. airfoil [RFHS08]. Alan [Pot05]. algebra [AAS21, DENP09, FMR15a, Gau02, Sus07]. Algebraic [BBKL15, EMS09, NP19a, Pop08, AO10, ABM93, Aré09, Bra00, EMS13, HS08b, HJ09, JN07, JC06, KLRW19, KM96, KM07, LC14, LT18, Man03, MN08, Meu01, Not10, PMLFT09, PGH11, Rom07, Sch93]. algebras [Arv06, JO14, JO17, Opf15]. algorithm [ABM93, AMVW15, BEG15, BR14, Bar05, BK13, BP03, BW12, Bun97, BMX07, CL03, Die97, Dis97, Faß96, Faß07, For02, Gan04, HMRS04, HJ09, JC17, Kal20, Kou97, Kre05, LV08, LDB13a, LDB13b, LC14, Man08, Osb08, PPS18, Ste97b, Tif11, Tu05, Tu07, TW16, VW13, WNK18, WLHH18, WL11, WL12, YNYF15, ZC19]. algorithmic [Koe99]. Algorithms [LZ08a, Aur15, BR08, CW16, Cha11, CS94, DDR95, ENH10, Els05, FI20, GKM16, IN11, JVV12, KCW18, MR18, MH01, Pav99, PCP06, RKvdDA14, Ves96, ZX17]. aligned [Dis97]. all-at-once [GW19]. Almost [Ade17]. along [Dos08, Sch05]. Alternating [HNP10, WNK18, ZM13]. Alternative [Che06]. American [Oos03]. American-style [Oos03]. AMG [Ema10, GN16, Haa00]. AMGr [GN16]. Ampère [DG06, Nei14]. amplitude [GSW08]. anaerobic [SCCJC09]. Analysis [AR07, BSS09, BDH15, Bre03, CG18, ELT21, FCA21, FK08, HLP10, LMN10, NV19, Sch93, Xu06, AM11, Ahm19, AAS21, ACS00, AK10, BBD^{+}08, BJR^{+}09, BDR09, BF12, CMO18, CCDR08, CC00, Che97, DNV05, DN11, DSC08, FJMS06, For02, FH13, GS07, Gau02, GGM12, GO10, HHP13, HLNT18, HR14, JS18, Kin11, MB06, MC05, MH01, MX96, MX97, MG14, NP19a, Nei14, PLM03, SAA09, SG94, WD08, XS20, YNYF15, Yan98]. analytic [DDPS20, JJ10, MY02, MR93, MPS18, Ost10]. analytical [ÁNCQ06]. Andoyer [Bre94]. Anger [LP17]. angle [GMT16]. anisotropic [ADM07, BBKL15, CL03, DGP99, MC05, Mav97, NC06]. annihilation [Lor99]. announcement [PQRT06]. ANSYS [PLM03]. anti [Ahm19, Cab07]. anti-palindromic [Ahm19]. anti-periodic [Cab07]. Any [DTM17, Sch16]. AOR [FMP96]. Apostol [KS18]. Apostol-type [KS18]. Appell [LP08]. Application [OR11, ZK17, Bor09, CMM95, GW10, HLC10, KM08, KLS17, LP17, LKF08, Oos03, Sim10, LP08]. Applications [Bar11, BGSC20b, BGSC20a, BJR^{+}09, BB01, BCDI15, BRC^{+}16, BCDV06, CMP01, EMJ16, FXG19, FLMT05, FMR15a, KZ07, NNR99, OS98, RÁRP06, She10, YHL^{+}20, ZM13]. applied [BDG04, DS11, EMS13, FS08, JO05, LR16, MST08, PGH11, TV07]. approach [AM06, Arv06, Bor09, DG06, FJMS06, FMSV08, FTRH20, GW10, HHP19, KKR20, KS18, MM99, MS08, PMLFT09, TMS97, Tif13]. approximants [Bar06, Zho06]. Approximate [Kür19, SV10a, BL08, CFM^{+}10, ET09, LN10, MPV13, MPV15, WA08]. Approximating [Ash17, vBHS14, BR08]. Approximation [GSW08, GRS00, VCK08, Yam06a, AR09, AMRT06, ABH17, BEG15, BY98, BK06, BLM06, BSS09, BCDO18, BEJ^{+}18, BKK^{+}18, Bor09, Del02, ER01, ERS17, Faß96, FFK^{+}15, FP01, FMR15b, GCPP99, HLX14, LT14, Lub06, Man03, MJ11, Mas02, MMN18, Mok16, Nal19, NP20, Nav05, Ost10, Rap06, ST08, Sch03, SG98, VB95, VC15]. Approximations [BFdP^{+}18, FLMT05, Fro09, Kha06, RV08, WA08]. arbitrary [LO13, LWH20]. arcs [BDD06]. area [Leo06, Leo09, VA09]. argument [GRS00]. arguments [Pea15]. arising [CCC07, DE06, EMS13, NB08a, NRS17, Woh00]. arithmetic [DDR95, Lau08]. Arnoldi [EMS09, EHJM17, GN13, HS97, HJ09, Mee98, Meu15]. arrays [Dam02]. aspects [FM02, SAA09]. assessment [BBD18]. assignment [CC00]. associated [Arv06, EV93, HS12, MS06, SG99, Sta02]. asymptotes [BEG15]. Asymptotic [AD04, BKS18, DB08, FLMT05, Kou97, Dam02, DT09]. Asymptotics [HC05, MFMGO05, Sta02, Man06]. Asynchronous [FSS97, BMPS95, SME02]. atmospheric [Auz18, RS21a, RS21b]. atom [RR07]. atomic [TT05]. atoms [ÁNCQ06]. Augmented [Neu19, AVV14, DG06, Gat07, RS21a, RS21b, Zít08]. automatic [Sch08b]. autonomous [HHO14]. average [MR18]. averaged [DRST16]. avoiding [CKD15]. aware [Sil03]. axisymmetric [KMC11].

B [KK08, ZK17]. B-spline [KK08].

B-splines [ZK17]. backward

[AD09, GJV13]. bacteria [SCCJC09]. Bakhvalov [NV19]. Bakhvalov-type [NV19]. balancing [BM05, Wat06a]. balayage [Göt06]. ball [BLM06, Xu06]. Banach [CMO18]. bang [Wac13]. bang-bang [Wac13]. barrier [NV19]. barrier-function [NV19]. barycenter [OR11]. Barzilai [HNP10]. based [Bar11, BHN13, BB08, BBK18, BDR09, BDH15, BFK^{+}10, BD20, CP00, DK13, EL05, EHJM17, FCA21, Gat04, GSV12, GN16, Gut12, HKKR18, HRT08, HK14, HLP10, HLP15, HLC10, JC17, KL21, Kin11, KLRW19, KL18, Lam10, LMR15, MJ11, NP19b, NP96, Not10, Tif11, YG08]. bases [Fer05, MS06, PCP06]. basic [FM02]. basis [GSV12, HOR08, KCW18, KKR18, KKR20, LW06, WE13, WA08]. BDDC [BPS17, CW16, HKLW20a, HKLW20b, KCW18, KKR18, KLRW19, KKR20, PD17, Tu05, Tu07, TW16]. Behavior [LM00, Böt04, DB08, KS06]. behaviour [FP01, GJV13]. Beijing [LHL^{+}20]. BEM [HLP15, Kol12]. BEM-based [HLP15]. Benchmarking [Ema10]. bending [SG94]. Beowulf [FO03]. Bernoulli [HMRS04]. Bernoulli-like [HMRS04]. Bernstein [BCGAM08, DRST16, Gau10, Ita06, Lub06, Nal19, Ost10, WA08]. Besov [Tik05]. Bessel [Man06, Rap06, Ste02]. best [BLM06, Bar06, Kha06, Mas02, Ran07, Van07]. beta [BH20, LWH20]. Beta-type [LWH20]. between [Gau02, LMS09, Mee98, SK06, ZY13]. beyond [VW13]. Bezoutians [ER14]. BFGS [ABHK09, SHH^{+}12]. biased [BCDO18]. BiCGSTAB [EJS03, SF93]. bidiagonal [WL12]. bidiagonalization [CHR02, HRT08, ZX17]. Bidomain [GGM12]. bifurcation [SY09]. biharmonic [BGS10, CPS18, KRZ16, Pry14, Zha04]. bilinear [FMR15a, Kim07b]. binary [KM08]. Binet [GM18]. Binet-type [GM18]. Bingham [AVV14]. binomial [JN10]. bio [LKF08]. bio-heat-transfer-equation [LKF08]. biofilm [DE06]. Biorthogonal [VVV19]. Biot [HK18]. biperiodic [HR14]. birthday [GOR^{+}08, RS14]. bisection [DDR95, EH15]. bisection-like [DDR95]. Bivariate [BCDV06, MM09a]. Block [BGSC20b, BGSC20a, Bar15, BU07, CFM^{+}10, FLS17, Val19, WL11, AHS17, ABA20, AK10, CS99, DN11, Dub01, ELT21, EJS03, EMS09, EHJM17, Fat98, GGM12, HK17, HJ09, HN08b, KN03, KLS17, Lam08b, MPV15, MZLG14, NZ14, SS09, Soo17, TZ17, TLF07, LZ08b]. Block-proximal [Val19]. block-triangular [GGM12]. block-tridiagonal [CS99]. blocks [Gre04, GS06]. blow [NB08b]. blow-up [NB08b]. blur [Bar05]. bodies [Kro06]. Boltzmann [CL03]. Boosting [PMLFT09]. Borges [JST18]. Borsuk [AFHM04]. Borwein [HNP10]. Bouligand [CN19]. bound [COSV93]. boundaries [VK95]. Boundary [ET09, KL12, Ash17, BHN13, BK16, ČF18, Che99, Che09, DB08, ER01, FFK^{+}15, GN14, HJ08b, HL20, HLP10, KK08, Lam05, LMN10, Mar18, MST08, Nas15, Sch08a, SZ20]. boundary-value [Che09]. bounded [AR07]. Bounds [Sch02, Vos99, AD04, Bea98, BR08, CHR02, CGT05, CKP15, FSS18, Leo09, LR16, MPS18, PP08, TV18]. Bregman [DdSV20]. broadcasting [YHL^{+}20]. Bruijn [COSV93]. bubbly [TV07]. bucket [AO10]. bucket-sorted [AO10]. bulges [Kre05].

Cache [DHK^{+}00, Sil03]. Calculation [Kan08, BL08, HLC10]. calculations [SK06]. capabilities [PLM03]. Carathéodory [ST08]. Cascadic [LH12]. case [BGSC20b, BGSC20a, BL08, HS08a, HS08b, LDB13a, LDB13b, LT05, MX96, MX97, Ost10, TLF07, Wat06a]. cases [TT05]. Cassini [VK99]. Cauchy [BMP09, CR96, Drm15, JR02, JR04]. Cauchy-like [BMP09]. Cayley [AMVW17]. CCFD [FCA21]. cell [MM09b, VA09]. Center [Bra97]. centrosymmetric [ER14]. certain [ABBS07, Arv06, BMP09, Nov05]. CFD [Ema10]. CG [BK02, LT05]. CGNE [Neu19]. chains [BBD^{+}08, BU07, BKK^{+}18, CG18, FMSV08, HMRS04, Jac10, Tif11, Tif13]. Chan [Zei99]. change [KCW18]. changing [Fra21]. channel [Kan19a, Kan19b]. Characteristics [Tik05]. characterization [GS94]. characterizing [SS08]. Chebyshev [BK13, FP01, GN16, HH93, WR09]. Chebyshev-like [HH93]. Chinese [YHL^{+}20]. choice [CW16, Kin11, Kin13, Wu03]. Cholesky [BBF^{+}00a, BBF^{+}00b, GM04, RT05]. Cholesky-like [BBF^{+}00a, BBF^{+}00b]. CholeskyQR2 [YNYF15]. Christoffel [DK10]. Ciarlet [YG96]. CIP [Sch08a]. circle [Bar06, BDD06, BCDI15, CBGV05, Gau06, Gla08, MS14, SL12, Sin95]. circles [Roş09]. circuits [KJ09]. Circulant [Ng97]. circular [CMR12]. citation [BDR09, BF12]. citation-based [BDR09]. class [ABBS07, ENH10, JG14, LT18, MFMGO05, Tik05]. classical [Gau02, HK18, LZ07, MM99, RS12]. clusters [FO03]. Coarse [KKR20, HKKR18, HKLW20a, HKLW20b, KRR16, KLRW19, MR18]. Coarsening [AO10, Bol08, Bra00, BFH^{+}15, Dis97, Mav97, Spr02, TV18]. coefficient [Lam05, SCCJC09]. coefficients [LDB13a, LDB13b, LH12]. Coercive [SZ20]. coin [BCDO18]. Collocation [GSV12, JR04, JKM14, AS14, ACS00, ER01, JR02, KK08, KL12, LR16, PGH11, YG08, ZK17]. column [PCP06]. Combinatorial [PCP06, Tif13]. combined [BKK^{+}18, PPN13, SSVS12]. Communication [BM05, CKD15]. communication-avoiding [CKD15]. commutator [ET09, Pea15]. commuting [She10]. Comonotone [VC15]. compact [GGCT09, PSC97, SSVS12]. compact-WENO [GGCT09]. companion [BDG04]. Comparing [Spr02]. Comparison [LMS09, Sta03, Ful97, JVV12, KRR16, LMM12, Mee98, RV08, VW97]. complement [KLS16, Zha00]. complementarity [Oos03]. complete [Meu12]. completion [LM20, RKvdDA14]. complex [AM11, BMX99, Ber04, KP13, KM17, PPS18, SF93, SHZF08, TT05, VCK08]. complexity [BMS08a]. component [HK14]. composite [SMS^{+}19]. compressible [Ema10]. compression [MB06]. Computation [ACR99, Car12, HRV13, KMC11, Nar18, Bra97, KLS17, MS14, MVV09, MR93, MACI06, Nov05, Sch08b]. computational [SAA09]. computations [BC05, DS11, EH15, FO03, HS08b, KS00, ST02]. computer [NND06]. Computing [AMVW17, JO07, LY14, MPV13, MPV15, SSS10, BFM20, IN11, JVV12, KQT13, KM08, LWH20, MV15, MH01, NNR99, OS15, PCP06, TC08, VVM09]. concerning [EP07]. Condition [AK10, MS06, BP07, DB08, Sim10, Zha01]. Conditional [AS14, BBK18, Dis97]. conditioned [Kin13, Peñ04]. conditions [BHN13, BKO17, ČF18, Che09, ET09, Gat07, HL20, MST08, Neu16, Sch08a]. conductors [BHRT03]. confidence [NO02]. configurations [vBHS14]. confluent [DT09, LR04]. Conformal [CMR12, LT94, SL16]. conjecture [KV95, Zei99]. conjugate [CKD15, Dub01, GJV13, KN03, NP96, ST02]. connected [CMR12, DK10, HS12, SL16]. connections [KKR20]. considered [Lau08]. consistent [Aré09, BKS18]. consolidation [HK18]. constant [COSV93, HR06, LDB13a]. constitutive [Mál08]. constrained [Bar05, Bar11, QvGvW^{+}15, Ran07, VC14]. constraining [LMSY14]. constraint [CLRS04, JOT08, RZ12]. constraints [CW16, KCW18, LMM12, Ram08, RR10]. Construction [JST18, XBC19]. constructions [HR06]. constructive [BMFJ^{+}06, MJ11]. contemporary [JVV12]. contents [Ano09a, Ano10b, Ano13a, Ano14c, Ano15b, Ano15d, Ano16, Ano17b, Ano17c, Ano18c, Ano19a, Ano20c]. context [LO13]. continued [IL06, Khr06, LB96]. Continuous [BB00, BMSR06, Fra21, CKP15, HJ09, KQT13, Lor99, Man06]. continuous-time [CKP15, HJ09]. contour [ST08]. control [CKP15, DSSS13, Kol12, PW13, Pea15, Wac13]. controlled [BKO17, DJ06a]. convection [BZ03, ELST18, FK08, Kou97, MST08, PSC97, PW13, RV08, WLR02]. convection-diffusion [ELST18, FK08, Kou97, MST08, PSC97, PW13, RV08, WLR02]. Convergence [BEN94, Che09, ELST18, FS08, Ger17, HK17, HHP13, HP20, HR14, JS18, Kin11, KL18, MC05, MMP09, MG14, RR10, Yan98, Zha04, Ade17, BK02, Bog05, ENH10, EBJ21, Fat98, FFK^{+}15, FP01, Fro09, Kin13, LT05, Mas06, MN08, NR14, OS98, Par21, PSS99, Sch16, SG94, Ste02, TV18]. convergent [BPP15, BMS08a, Bra06, HMRS04]. converse [CC97]. convex [BEG15, BSDMGFTR09, BSS09, For02, Kro06, RZ12]. convolution [NND06, Ng97, Yam06a]. convolution-like [Ng97]. coordinate [Gut12, Yam06b]. coordinates [SJ08, SSVS12]. coquaternionic [JO14]. coquaternions [JO17]. correction [GK14]. correctness [DDR95]. corrector [PPN13]. countable [AK06]. counterexample [SS08]. Counting [KP13]. coupled [GSV12, HJ16, HR14, HC06, NRS17]. coupling [GS07]. course [Gan08]. CPU [JO08]. CPU-GPU [JO08]. Creating [AH12]. Creation [Lor99]. criteria [ABBS07, AL08, HJV18]. critical [Kan08, Kan15, Kan19a, Kan19b]. Cross [BB08, GS16]. Cross-Gramian [BB08]. Cross-points [GS16]. Crout [LS05]. Crouzeix [LMR15]. CT [Bar11]. Cubature [OSLS20, Van19]. cubic [MYZ18, VC15]. current [AR07, Kol12]. curvature [CHR02]. curve [CHR02, KR20, Sch16]. curved [AR09]. curvilinear [Yam06b]. cut [SG99]. cycle [Zha04]. cyclic [HK17]. cylinder [SG99].

D [BLM06, BB01, BRC^{+}16, BHRT03, CP03, HKKR18, Mar07, Mar11, TV07, XS20]. DAE [BKO17]. DAE-index [BKO17]. damped [ZY13]. Darboux [BD04]. Darcy [AM06, GS07, HJ16, LMS09]. Data [RKvdDA14, BB08, DdSV20, DdAR17, FFK^{+}15, GK08, GK14, Sil03]. data-sparse [BB08]. dataflow [EL05]. Davidson [HN08a, SvdVM98, SS98]. Davison [Sad15]. deblurring [BHN13, BD20, HJ08a]. Decay [BR08, FSS18]. decaying [PMLFT09]. decomposition [BBD18, BP03, BGyS21, CW16, CS94, Dub99, FXG19, FKSV07, Gan04, GS16, Haa00, HJV18, Hen94, KRZ16, LO07, LNG15, MM04, Pav99]. Decompositional [BBD^{+}08]. decompositions [MR93, Peñ04, RY15, Ste95]. decrease [Dos08]. dedicated [GOR^{+}08, RS14]. Dedication [Pot05]. deferred [GK14]. deficient [SGG15]. defined [Cot06]. definite [CMP01, FSS18, GMT16, LT05, LZ07, Pan19, ZB10]. definition [Kno08]. deflated [MZLG14, WE13]. deflation [CKD15, Gut12, KKR20, MX15, TV07]. degenerate [DE06]. degree [Bar06, Gla08, Ran07]. delay [CDN09, FMMS08, FL07, GN14, KS06]. denoising [BBK18, FCA21, FM02, LMSY14]. dense [Che99]. dependency [WD08]. dependent [BD17, HOR08, Lam08a]. deposition [MM09b]. Derivation [Lam08a]. derivative [AG19, Car11, Car12]. derivatives [BK06, LWH20, PQRT06]. derived [Gre04]. descending [TDLC09]. descent [AEEG08, LMM12, NZ14]. described [Mál08]. detection [FL07]. detective [Lau08]. determinants [Sch02]. development [GRS00, Pea15]. developments [FLMT05]. devices [SMS^{+}19]. DGFEM [FK08]. diagonal [VVM09]. Diameter [Leo09, Leo06]. difference [ÁNACS07, Bra06, Cab07, CP00, CP03, FMR15b, Mar18, MS07, RR07, RV08, SK06]. differences [BRC^{+}16]. different [DJJ11, LMS09, Zha01]. differential [Aré09, BSS16, Die97, EBJ21, FL07, GJV13, JN07, KS06, Khr06, KM96, KM07, LT18, MYZ18, MFMGO05, NRS17, Ore10, Sch93, Sch05, SCCJC09]. differential-algebraic [Aré09, JN07, KM96, KM07, LT18, Sch93]. differentiation [Aur15]. diffusion [BZ03, BHRT03, BBKL15, DE06, ELST18, EG06, FK08, FMR15b, Kou97, LDB13a, LDB13b, LR16, LK21, MST08, PSC97, PW13, RV08, WLR02]. dimensional [ÁNCQ06, Bre03, HK14, NP20, RV08]. dimensions [Cal15, DG06, Hip97, Kim07a, KRR16, KL12]. Direct [CDN09, BFM20]. directed [CEHT10]. direction [ZM13]. Directional [Mav97]. Dirichlet [DG06, For02, GKM16, Gat07]. disaggregation [MMP09]. Discontinuous [EDG^{+}15, Pry14, HLNT18, JS08, LDB13a, LDB13b, SMS^{+}19, TW16]. discrepancy [VCL08]. Discrete [Bre04, Che99, DGH14, FKK10, LO13, AL09, BL06, Bor09, CR15, Faß96, Lor99, Pry14, STO08, VB95]. Discretization [Kin13, BHRT03, DE06, GS07, GS16, GLMT21, HL20, LMS97, Lis08, LMR15, Mar11, PSC97, Tu07, Wac13, WLR02]. discretizations [EDG^{+}15, KM07, LMS09, Mar07, TW16, Woh00]. discretized [ELST18]. discretizing [Che94]. disordered [BFK^{+}10]. dispersion [Kle08]. Displacement [CNP94, LM20]. distance [BBKL15, MS14, NPR08, Smi10]. distributed [AH03, FMR15b, Kol12]. Distribution [Pri06, Dam02, LMSY14, SK06, SK08]. distributions [Nar18, TT05]. divergence [NC06, Ste02]. division [AAS21]. dLV [IN11]. Domain [CS94, Pav99, AR07, AH12, BBD18, CW16, Dob12, GS16, Haa00, HJV18, HLX14, LO07, MRT03, Mit97, Xu18, Yam06b]. domain-decomposition [LO07]. domain-robust [Dob12]. domains [AR09, CMR12, DK10, EV93, KP13, Kol12, LMN10, SL16]. Double [GMT16, CDN09, LP08]. downdating [Bar15]. DP [Bre03, HKLW20a, HKLW20b, Kim07a, KCW18, KKR18, KKR20, Mar11]. DR [MZLG14]. Dual [AMRT06, BBK18, CMO18, CCC07, LV14, Par21, SM08]. dual-mixed [CMO18]. Duality [AK06, LO13]. duct [HLC10]. due [Bre94]. dynamic [ABBS07, BC05, BRS08, DH07, EP07, KZ07]. dynamical [LR03, Man08, VCL08]. dynamics [OR11].

Ed [She06]. eddy [AR07, Kol12, NW03]. Edge [CKH14, LMR15, KL21, NP19a, NP19b]. Edge-based [LMR15]. Edge-Preserving [CKH14]. effects [APL04, Kan15, Kan19a, Kan19b]. efficiency [Gat04]. Efficient [BHN13, Ber04, Che94, FG17, GK14, KN03, QvGvW^{+}15, SvdVM98, TV07, Van19, BKS15, CKD15, Cha11, HRT08, NN16, PV09, SGB07]. Eigenfunction [MYZ18]. eigenfunctions [PP06]. eigenpairs [CRZT20]. eigenproblem [FS08, Pan19]. eigenproblems [FMMS08, JG14, OS98, SvdVM98]. eigenreduction [Ves96]. eigensolver [TC08]. eigensolvers [KL21, Kny98]. eigenstructure [BFS08, CC00, Ita06]. eigentriplets [TC08]. eigenvalue [Ahm19, AR09, AH10, BMX99, BFS08, CRS94, CKV04, DDR95, EH15, HS97, HP08, KLS16, Kal20, KN03, KV13, KL18, MW02, NZ14, Sor98, SV10b, SS98, TLL15, TZ17, Var01, Vos99, WSS98, ZC19]. eigenvalues [AD04, Ash17, AMVW17, KP13, Lei19, MX15, MACI06, NNR99, VVM09]. eigenvectors [HS12]. elastic [GLMT21, KMC11]. elasticity [BBD18, CMM95, Gat07, Lis08, OR11]. electric [SMS^{+}19]. electro [PV09]. electro-physiology [PV09]. electrocardiology [GGM12]. electrochemical [MM09b]. electromagnetic [HR14]. electronic [DS11, KJ09]. electrostatic [GAM05]. Electrostatics [Van07]. Element [HLP15, AMRT06, AM06, AL08, AD04, BPP15, BMS08a, BB01, BRS08, CW96, Che99, DGP99, DHI03, FFK^{+}15, GS16, Gat04, Gat07, HK14, HL20, HLP10, HLX14, Kim07b, KL12, KS00, LMN10, LMR15, MC05, Mar07, Mar11, Nei14, NC06, SJ08, SG94, SZ20, Tu07, Woh00, YG96, Zha04]. elements [Bun97, Che97, GH99, HK10, NP19a, NP19b, Sch08a]. elimination [HLC10, KL21, NU05]. ellipses [DDPS20]. Elliptic [VA09, AH04, AH10, BRC^{+}16, Bog05, CP00, CS94, DG06, GJV13, Hoh06, LH12, LMR15, Mar07, NZ14, Nov05, TW16]. Emden [ABBS07]. empirical [LHL^{+}20]. endpoint [Ill03]. Energy [GJV13, Pri06, Ran07, SSVS12]. Enhanced [ERS17]. Enhancement [Lam08b]. ensemble [SHH^{+}12]. entire [Ped07]. entrance [SK06, SK08]. entropies [BFdP^{+}18]. entropy [Man08, Raj08]. equal [Leo06, Leo09, LRL08]. Equal-order [LRL08]. equally [Roş09]. equation [ÁNACS07, ACS00, AC06, BB00, BP21, BMP09, Bol08, CL03, CQ01, DG06, DB08, DE06, EHJM17, EN08, FFK^{+}15, FMR15b, FH13, HJV18, HJ09, KL12, LKF08, MACI06, NND06, Nei14, Pry14, RR07, RV08, SSVS12, ZK17]. equations [AHS17, ABJE15, ABBS07, ABM93, AS14, ADW15, Aré09, AM06, AH10, BZ03, BPP15, BBD18, BMS08b, BB14, BKS15, Boc97, BL97, BSS16, Cab07, CMM95, CR96, CCDR08, CDN09, Che94, Cot06, DH07, Die97, Dob12, DHI03, EMS13, EDG^{+}15, ET09, EP07, EBJ21, FI20, FL07, GS07, Gan04, GSV12, GJV13, HMRS04, HHP13, HJ16, HL20, JN07, JR02, JR04, JKM14, KS06, KZ07, Kle08, Koe07, Kou97, KM96, KM07, LMS97, LDB13a, LDB13b, LMW20, LC14, LT18, LH12, MYZ18, MSV95, MFMGO05, MS08, MG14, Mok15, Mok16, NB08b, Nas15, Ng97, NRS17, Ore10, PV09, Per00, PGH11, Sch93, Sch05, SY09, SF93, Sta03, Sty08, WD08, YG08]. equidistributed [GN14]. equilibrium [BDD06, BCDO18]. equispaced [GK14]. equivalence [SM08]. Error [BY98, BB11, DJ06a, DDPS20, HK14, MH01, MPS18, SAA09, YG08, BMM97, CS14, DGP99, Ful03, Gat04, GJV13, KR20, LR16, MACI06, NV19, NC06, Ore10, PGH11, ST02, XS20, YNYF15]. errors [SME02]. estimate [YG08]. estimated [FKSV07]. Estimates [FMR15a, BB01, BB11, BMM97, DDPS20, Ful03, HK14, Kha06, MR17, XS20, ZN17]. Estimating [CS14, BCDO18]. estimation [AH03, BY98, FvdMS14, MACI06, NC06, PGH11, PT13, Psa04, ST02]. Estimations [BFM12]. estimator [KR20]. estimators [DGP99, Gat04, Man03]. ETNA [Ano20a]. Euler [BFdP^{+}18, Khr06]. European [PPN13]. Evaluating [BDR09, Car11, ST08]. Evaluation [SG99, FG17, KS18, KM17, LW06, SJ08]. even [AM11]. evolution [AS14, Gan04, HOR08, HHP13, HLNT18, MS07]. evolutionary [GW19]. exact [Kha06]. examples [Zho06]. existence [AFHM04]. exotic [PPN13]. expansion [HR14, JN10, LZ07, MYZ18, SvdVM98]. expansions [LWH20, LW06, Mas06]. Experiences [Boc97]. experimental [CT03]. experiments [BJM99, Meu01]. explained [Gut10]. Explicit [BCDI15, APL04, BEG15, HOR08, JS18, MPV13, MPV15, PPN13]. exploiting [FJK^{+}20]. exponential [BKS18, Dam02, FvdMS14, HHO14, KM17, Mas06, PT13, ST08, Sta02]. exponential-type [BKS18]. exponentially [Mar18]. exponentials [Böt04, KLS17, LR04, PMLFT09]. exponents [MH01]. extended [BEJ^{+}18, HJ09, MPV13]. extension [CR15, JO14, MPS18, VW13]. Extensions [JN07, BCDV06, CS14, MPV15, Sch06]. exterior [HRV13, KL12]. extraction [HP08]. extrapolated [CRZT20]. extrapolation [BFM12, DS11, EMS13, FMOS21, MR17]. Extremal [RS12, Göt06, HC05]. extreme [EV93].

Faber [EV93, HS97]. factor [DSC08]. factored [DH07]. Factorization [ÁNACS07, BY99, FXG19, GM04, HNP10, Koe07, LS05, RT05, SAA09, SG06]. factorization/solver [VVM08]. factorizations [BBF^{+}00a, BBF^{+}00b, OB05, SGB07, WL11]. Factors [FO03, Kro06]. Fast [Aur15, AMVW15, BCR98, BMP09, JO05, KM17, LW06, Nas15, Ng94, PW13, PK08, TMS97, AC06, BHRT03, BP03, CCDR08, EH15, FGV09, LV08, SG06]. FDM [Ade17]. Fejér [SL12, ST08, RS12]. Fekete [BL08, SV10a]. FEM [Ade17, FKK10, HR14, Kol12]. FEM-FDM [Ade17]. FETI [Bre03, HKLW20a, HKLW20b, Kim07a, KCW18, KKR18, KKR20, Mar11]. FETI-DP [Bre03, HKLW20a, HKLW20b, Kim07a, KCW18, KKR18, KKR20, Mar11]. FFT [NP96]. FFT-based [NP96]. FGMRES [AD09]. fidelities [BB11]. field [BdPNdP14, HK18, KP13, Opf10, RR07, Tos00]. Fields [Hoc11]. fifty [Not16]. Filter [DSC08]. filtering [ABHK09, BP03, Bre16, NP96, SHH^{+}12]. Finding [PV05, BGP04, Jac10]. Fine [Sim06, dJvdPDV17]. Fine-grain [dJvdPDV17]. Finite [HL20, HLX14, AMRT06, AM06, AL08, AD04, BPP15, BMS08a, BB01, BRC^{+}16, BRS08, Bun97, CP00, CP03, CW96, Che97, DGP99, FMR15b, GS16, Gat04, Gat07, HK14, HLP10, HK10, Kim07b, KS00, LO13, LMN10, LMR15, LKF08, Mar07, Nei14, NC06, Sch08a, SJ08, Sch16, SZ20, ST02, Tu07, Woh00, WLR02, YG96, HLP15, SK06]. finite-element [AM06]. First [CMM95, BMM97, BDR09, CL03, LB96, Pav99].

First-order [CMM95, BMM97, CL03, LB96, Pav99]. fitted [KS06, KK08, Mar18]. five [RV08]. five-point [RV08]. fixed [CG18, JKM14, Van07]. fixed-sized [CG18]. flexible [MZLG14]. Flip [FXG19]. Flip-flop [FXG19]. floating [DDR95, SME02]. flop [FXG19]. flow [AL09, AMRT06, BKO17, Bir12, Dis97, DDS^{+}18, HLX14, HLC10, Kan15, Kan19a, Kan19b, PSS99, RFHS08, SK06, SK08, TV07, Tu05, Tu07]. flows [CR09, Ema10, GGCT09, Kan08, LRL08, Mál08, STO08, Ste97a]. fluid [AVV14, AMRT06, CR09, OR11, SV10b, TDLC09, Vos03]. fluid-solid [SV10b, Vos03]. fluids [Mál08]. flux [HJ16]. FOM [Sch16, Soo17]. footnote [CS99]. Forchheimer [LMS09]. form [CMM95, Drm15, FMR15a, FGV09, HJ16, NC06, Wat06b]. forms [AK10, Zha01]. formula [BFdP^{+}18, Bre94, Ill03, MPS18]. formulae [DDPS20, EMJ16, Not16, OSLS20]. formulas [BCDI15, CS14, RGVP99, Roş09]. formulated [AR09]. formulation [AR07, BRC^{+}16, CCC07, DHI03, HK18, KCW18, Tu05]. formulations [HJV18]. FOSLS [BMM97]. four [RV08]. four- [RV08]. Fourier [CQ01, HR14, LZ07, Man06, Ste02]. Fourier-mode [HR14]. Fourth [Kle08, BFR19, Han06, MYZ18, SSVS12]. fourth-order [BFR19, Han06, MYZ18, SSVS12]. Fowler [ABBS07]. Fractal [Nav05, Nal19, VC14, VC15]. Fractional [ZZW20, BSS16, Die97, EBJ21]. fractions [IL06, Khr06, LB96]. frame [BD20, GLMT21]. framework [Bre03, PD17]. Fréchet [Car11, Car12]. Fredholm [NNR99, YG08]. free [Bol08, BK16, Che09, GM06, Kin11, Kin13, LPS20, LY14, LT18, SV10b]. frequency [DNV05, DdAR17, Gan04]. Freud [GAM05]. Freud-type [GAM05]. frog [LMW20]. frugal [HKLW20a, HKLW20b]. full [Bor09, Mit97, Nal19, WLHH18]. full-Newton [Bor09]. full-space [WLHH18]. fully [HC06]. function [BMFJ^{+}06, BEJ^{+}18, Dos08, ERS17, LZ08a, LW06, LP08, NV19, Sta02, VK95]. functional [LMW20, Mas02, RZ12]. functionals [ABH17, CGM10, LMM12]. functions [AEEG08, ÁNCQ06, BY98, BR08, Bre04, CFÁN09, DT09, DDPS20, DJ06b, FM02, Fro09, FLS17, GSV12, GRS00, HHP19, Kha06, Koe99, KS18, LWH20, LP17, Man06, MJ11, MPS18, Ost10, Ped07, Pri96, Rap06, RGVP99, RS12, ST08, Sch08b, Sch05, SG99, SG98, Sin95, Tik05, ZZW20, Zho06, Sha12]. Future [Ano20a].

Galerkin [BPS17, EG06, EDG^{+}15, GW10, HHP13, HLNT18, JS08, Mok16, Pry14, SMS^{+}19, TW16, Yam06b]. gas [DDS^{+}18]. Gauss [ACR99, Ash17, Bra97, CS14, DRST16, Gau06, LWH20, Not16, Osb08, PR20, PPS18, Sil03]. Gauss-type [Gau06]. Gaussian [CDN09, DV12, DDPS20, OSLS20, Sin95, VV18]. Gaussian-type [DDPS20]. GCDs [WA08]. GCV [CNO08]. Gegenbauer [Kei08]. Gene [Gau02, GOR^{+}08, Man08]. General [Bra00, Sch03, BB11, Dam02, Erh95, KMFO05, LMN10, Pop08, Soo16, Zho06]. generalised [FS08]. generalization [AEEG08, MJ11, PV09]. Generalizations [Hoc05, Bre94, MPS18]. Generalized [RÁRP06, AL09, BGSC20b, BGSC20a, BFM20, BD17, DRST16, Kal20, KKR18, KKR20, KL18, LC14, MFMGO05, Nas15, OS15, SvdVM98, Sor98, Sty08]. GeneRank [BK13, SEH14]. generated [FGV09]. Generating [BSDMGFTR09, BKS15]. generators [Els05]. generic [JJ10]. geometric [JVV12, Lóp16]. George [Pot05]. Gerschgorin [KV95, VK95]. Gersgorin [FKSV07, VK99, Var01, VCK08]. Gersgorin-type [VK99, Var01]. Gershgorin [CKV04]. Gershgorin-type [CKV04]. GEW [ZK17]. ghost [Van07]. Gilbert [BPP15]. Givens [JO05]. GK [WL12]. global [BBS19, BEJ^{+}18, Gra16, KM96, LDB13a, LDB13b, Ore10]. global-in-time [LDB13a, LDB13b]. globalized [CR09]. GMRES [DGH14, DTM17, EMS09, Erh95, IE17, LZ08b, MZLG14, Meu12, Neu19, Sim10, Soo16, Soo17, TLF07, WE13, Zít08]. GMRES-Type [DGH14, Neu19, Soo16]. Golub [Gau02, GOR^{+}08, Man08, ZX17, ZC19]. Google [DN11]. Google-like [DN11]. Gordon [BKS18, GSV12]. governing [SV10b]. GPU [JO08, dJvdPDV17]. GPU-type [dJvdPDV17]. graded [Ste95]. Gradient [LMM12, LR03, BBK18, CKD15, Dos08, GJV13, KN03, Lam10, NP96, Pan19, ST02]. gradients [Dub01]. grain [dJvdPDV17]. Gram [Bar15]. Gramian [BB08]. graph [BFM20, Els05, NN16]. graphics [NND06]. graphs [BFdP^{+}18, BF12, HS12]. gravity [MM09b]. grid [Bol08, DHK^{+}00]. grids [ADM07, BSDMGFTR09, Bun97, GN14, LP09, VA09]. group [CL03, GK08, Ore10, Roz06]. grouping [Auz18]. growing [CG18]. growth [KM08, SCCJC09]. Grüss [MJ11].

H [Gau02]. Hadamard [KM08]. Hahn [Arv06]. half [BK06, CBGV05]. half-line [BK06, CBGV05]. Halperin [Lóp16]. Hamiltonian [ADW15, ABA20, BMX99, BFS08, Faß07, MW02, Wat06b]. Hammerstein [Mok16]. hand [BBS19, BK02, EJS03, JST05, LZ08b, MZLG14]. Hankel [Drm15, ER14, Ng94]. harmful [Wat06a]. Harmonic [Bea98, HP08, DTM17, EDG^{+}15, Hoc05, ZC19]. having [Ped07]. heat [HJV18, KJ09, LKF08, NB08b]. Hellan [KRZ16]. Helmholtz [EN08, KL12]. help [Ger17]. Hermite [Ash17, BCDI15, MM09a, MJM16, She04, Sta02, Sta06, VV18, ZZW20]. Hermite-type [BCDI15]. Hermitian [ABH17, CMP01, EH15, Fro09, FSS18, HS97, KL21, KL18, PP08, ZB10]. Herrmann [KRZ16]. Hessenberg [AHS17, JO05, TT05, VW13]. Hessenberg-upper [VW13]. heterogeneous [BHRT03, HKLW20a, HKLW20b, KKR18, KKR20]. HHT [JN07]. HHT- [JN07]. Hierarchical [Bol08, DFV09]. hierarchy [DDS^{+}18]. High [CP00, Mar18, Mok15, Aré09, GK14, Kim07b, Lam08a, NP20, PSC97, Ste97a]. high-dimensional [NP20]. High-order [CP00, Mar18, Mok15, GK14, Kim07b, Lam08a]. higher [Bun97, Ng97, Sch08a]. higher-order [Ng97]. highly [Bra00, BFK^{+}10, FLS20, GS06]. Hilbert [GHS14, MF06, Sch03, Yam06a]. Homogeneous [HN08a, ČF18, Gat07, Yat06]. homotopy [ZY13]. Householder [SAA09]. hull [TLF07]. hulls [Gre04]. hurricane [Ful97]. Hybrid [GGCT09, AMRT06, Per00, Tu07, CNO08]. hybridizable [TW16]. hydrogen [RR07]. hydrogenlike [ÁNCQ06]. Hyper [FFK^{+}15]. Hyper-singular [FFK^{+}15]. hyperbolic [HNS04]. hypergeometric [ÁNACS07, CFÁN09, Cot06, DT09, DJ06b, GRS00]. hypergeometric-type [ÁNACS07, CFÁN09]. Hypergraph [KL21]. hypersingular [DO18]. hypocycloidal [EV93].

ideal [She10, TLF07]. identification [HHP19, KJ09]. Identifying [DdAR17, FMSV08, Tif11]. IDR [Gut10, Zem17]. II [FFK^{+}15, FLS20, KV95, Lui11, MX97, Sha12]. III [CG18, VK95]. ill [CN19, Kin11, Kin13, NR14, Neu19]. ill-conditioned [Kin13]. ill-posed [CN19, Kin11, NR14, Neu19]. ILU [BJM99, LS05]. Image [NO02, BHN13, BBK18, BD20, CCC07, FCA21, HJ08a, JC17, LMSY14, LPS20, ZM13]. images [Bar05, NO02]. imaging [Bar11, DENP09]. imbalance [NRS17]. imbedding [MRT03]. Impact [MM09b]. implementation [CT03]. Implementing [JO08]. Implicit [APL04, PPN13, AS09, FMR15b, LN10, Mál08, Sch93]. Implicit-explicit [PPN13]. Implicitly [BR14, CR09, Bre16, CRS94]. Improved [AS12, BGP04, CKP15, TV18, PPN13, WD08]. Including [DGH14, KS18, WR09]. inclusion [CKV04, Hoc11, Lei19, Var01]. Incomplete [BH20]. incompressible [BBD18, LRL08, Mál08, OR11, Per00, RFHS08, STO08]. inconsistent [HH93]. Incremental [KLS17]. indefinite [SG06]. independent [AO10, Kin13]. index [Aré09, BKO17]. induced [Nar18]. inducing [MH17]. inequalities [Bre04, LO13, MJ11, Mas02, Rap06]. inequality [Gau10, Ger17, KZ07, Yat06]. inertia [KV13]. Inexact [DSSS13, WSS98, BBMP06, FS08, HLC10, Yan98]. inf [LRL08]. inf-sup [LRL08]. infants [LKF08]. infinite [BEN94, MX15]. influence [Els05, Xu18]. Information [DGH14]. initial [Aré09, BK16, GN14, HJ08b, Lam05, Sch16, SZ20]. initial-boundary [BK16, HJ08b, Lam05]. initial-boundary-value [GN14]. initialization [BGP04]. inner [BEN94, BMSR06, Mee98, VB95]. inner-outer [BEN94]. input [CC00, MX96, MX97]. integrable [MJ11]. integral [CCDR08, CDN09, Che94, DJ06b, FFK^{+}15, FLS20, HLP10, JR02, JR04, JKM14, MG14, Mok15, Mok16, Nas15, Ng97, PGH11, Sus07, YG08]. integral-algebraic [PGH11]. integrals [Car12, DO18, FLMT05, LP08, Nov05, ST08]. integrands [Atk04, Ill03]. Integrating [Sha12]. integration [DO18, Fra21, GK14, KL18]. integration-based [KL18]. integrator [HHO14]. integrators [BKS18, ST08]. Interaction [RFHS08, JS18]. Interconnecting [HLP15]. interest [HHP19]. interface [Ade17]. intergrid [Che97]. interior [BOS08, BGS10, BW12, BPS17, JO08]. interlacing [DJJ11]. internal [Ema10]. Internality [DRST16]. interplay [Gau02]. interpolation [BB14, BCDI15, BCDV06, Bre94, Dam02, Del02, EMJ16, GM06, GCPP99, Gla08, GPT11, GK08, KS18, Lén06, Lén10, LRL08, MM09a, Opf15, PMLFT09, She04, SV10a, Van07, VC14, ZZW20, vBHS14]. interpolatory [HNS04, RS12]. interpretation [GAM05, GJV13]. interval [BCGAM08, JR02, JR04]. intervals [BDD06]. introduction [FLMT05]. invariance [AO10]. invariant [KQT13, SS08]. invariants [KM96]. Inverse [BLM06, DGH14, BdPNdP14, CGM10, CFM^{+}10, CN19, FS08, Göt06, LN10, LY14, PMLFT09]. inverse-free [LY14]. inverses [BFM20, FSS18, PCP06]. Inversion [ER14, DdAR17, LR04, MPV13, MPV15, RKvdDA14]. invert [Mee98]. involved [FO03]. involving [DT09, SF93]. irreducible [NPR08]. irregular [BSDMGFTR09, Cal15]. Isogeometric [CPS18, XS20, HLNT18]. Isotropic [NC06]. issues [MMP09]. iterated [BD17]. iteration [BMS08b, BDG04, Bre16, BEN94, CN19, Dub99, Fat98, FS08, Kür19, Lam08b, MS06, MSS15, NZ14, VVM09, WR09, Yan98]. iterations [CNP94, HJ08a, NP96, Pan19, Sch16, SME02, ZN17]. Iterative [CCC07, LNG15, NND06, Bar11, Bog05, BW12, Che94, DN11, DSC08, JG14, KRR16, MSV95, MMP09, NNR99, Ng94, PW13, Sty08, ZB10].

J [ABA20]. J-Lanczos [ABA20]. Jacobi [SS98, DJJ11, Gau10, HK17, HN08a, KMFO05, SvdVM98]. Jacobian [NU05]. Johnson [KRZ16]. Joint [LO07, DdSV20]. Jordan [Gre04, TLF07]. jump [LH12].

Kaczmarz [Pop08]. Kahan [ZX17, ZC19]. Kalman [ABHK09, SHH^{+}12]. Kantorovich [AFHM04]. kernel [Ita06, Nas15, LT94]. kind [Mok16, NNR99]. kinetic [HL20]. Klein [BKS18, GSV12]. Korteweg [Kle08]. Kronecker [BBD^{+}08]. Kronrod [ACR99, MPS18, Not16]. Krylov [KNS19, BBS19, BSS09, BFS08, Cha11, DSSS13, EN08, FP01, FLS17, FJK^{+}20, GS15, GNR15, GHS14, Gut12, Lam05, Lam08b, LY14, MPV13, MPV15, MS06, Meu17, VVV19, WLHH18, WO97, Wei94, ZN17]. Kutta [AS14, Bir12, GS94, LT18].

L [Peñ04]. lacunary [Lén10]. Lagrange [MRT03, WLHH18]. Lagrangian [DG06]. Laguerre [CC97, Mas02, MMN18, SSS10]. Lamé [MFMGO05]. Laminar [Kan15, Kan19a, Kan19b, Kan08]. Laminar-turbulent [Kan15, Kan19a, Kan19b, Kan08]. Lanczos [ABA20, BY98, BEJ^{+}18, Bre16, CRS94, CHR02, CNO08, Fro09, HRT08, Lam08b, Mee98, PPS18, TZ17, TC08, ZX17, ZC19]. Lanczos-hybrid [CNO08]. Landau [BPP15]. Langenhop [KZ07]. Laplace [LP08]. Laplacian [BFM20, EN08, NN16]. Large [ABHK09, LV14, NW03, ABJE15, AS09, AS12, ABH17, BMS08b, BFS08, BKS15, Böt04, CRS94, Cha11, HS97, HJ09, LY14, LR03, OS98, SG98, Sor98, WSS98]. Large-eddy [NW03]. Large-scale [ABHK09, LV14, ABJE15, BMS08b, BFS08, HJ09, LR03]. larger [Kre05]. Larsen [PV09]. laser [JS18]. latitudinal [Roş09]. lattice [ÁNACS07, Che09]. lattices [NP20]. Laurent [CGM10, CBGV05, Pri96]. Laurent-type [Pri96]. Lavrentiev [Neu16]. law [HS12, KV13, Mál08]. layer [KS06, KMC11, MM09b]. layers [Auz18]. laying [Sil03]. LBB [WD08]. LDU [Peñ04]. Leap [LMW20]. Leap-frog [LMW20]. Least [HJ16, Aco06, BMM97, BY99, Boc97, Bor09, CMM95, CNP94, CL03, Faß96, GPT11, KS00, LV08, LV14, MS08, Ng94, Osb08, Pav99, VB95, WNK18, ZM13]. least-square [Boc97]. least-squares [CL03, Faß96, GPT11, KS00, WNK18, ZM13]. Left [LZ07]. Left-definite [LZ07]. Legendre [BK16, LWH20, PGH11, SG99]. Lehmann [Bea98]. Lehmer [COSV93]. Leja [BCR98]. lemma [CGT05]. lemniscatic [SL16]. level [BBD18, Cal15, Han06, HLC10, Kin11, Kin13, KLRW19, LMS97, NP19a, TV18]. level-free [Kin11, Kin13]. Levenberg [CN19]. Levinson [GH94, VMV07]. Levinson- [GH94]. Levinson-like [VMV07]. Lie [Ore10]. Lifshitz [BPP15]. like [BBF^{+}00a, BBF^{+}00b, BGP04, BMP09, DN11, DDR95, HH93, HMRS04, LP09, Mál08, MVV09, Ng97, VMV07]. limited [ABHK09, SHH^{+}12]. line [BK06, CBGV05, VA09]. linear [Aco06, Ade17, AS12, AS05a, BPP15, Ber04, CMM95, CGM10, CC00, CFM^{+}10, Che94, DENP09, ER01, EJS03, Ema10, FMR15a, FJK^{+}20, Gat07, Gau02, GSW08, GS06, HH93, HHP13, JST05, JO08, KQT13, Kin11, KM96, LB96, MSV95, MJ11, MZLG14, Meu17, MS08, Neu19, OR11, Oos03, Ore10, Par21, Pop08, SGG15, SF93, SHZF08, Soo16, SME02, Sta03, TLL15, TZ17, VMV07, ZB10, ZM13, ZC19]. linearities [NU05]. linearly [Sch93]. link [BF12]. liquid [BKO17]. LMS [NP96]. Local [BMM97, KM96, LRL08, Man03, APL04, BEG15, EV93, Fat98, MST08]. Locality [RT05]. Localized [Fer05, BGyS21].

Locally [HS12, BGSC20b, BGSC20a, KN03]. location [Meu15]. logistic [FTRH20]. long [Ste97b]. Look [GH94]. Look-ahead [GH94]. loop [BC05]. Lothar [RS14]. Low [BSS16, Sty08, BKK^{+}18, GO10, HP20, Kle08, Kür19, LM20, SGB07, TDLC09, WA08]. low-order [HP20]. Low-rank [BSS16, Sty08, BKK^{+}18, GO10, Kür19, LM20]. lower [AD04, Cab07, COSV93]. LSQR [BR14]. LU [BY99, SGB07]. LU-factorizations [SGB07]. lumpings [Jac10]. Lyapunov [BKS15, FH13, MH01, Sch05, Sty08]. Lyapunov-type [Sch05]. Lyndon [KS18].

machines [JOT08]. Maclaurin [BFdP^{+}18]. magnetic [BHRT03, DdAR17, RR07]. magnetostatic [AG19]. Majorization [PP08]. Man [Sad15]. manifold [ER01]. many [Man06, RKvdDA14]. Mapping [CEHT10, CMR12, LT94, Lui11]. mappings [AH12, FP01]. Maps [Gra16, DK10, Rom07, SL16]. Markov [BBD^{+}08, BU07, BKK^{+}18, CR15, FMSV08, HMRS04, Jac10, Tif11, Tif13]. Marquardt [CN19]. mass [BKO17, WR09]. massive [BFH^{+}15, TV18]. matching [GS07]. Mathematical [Mál08]. Matlab [Sha12]. Matrices [AR14, AAS21, ABA20, ABH17, Aur15, AMVW17, BSS09, BBF^{+}00a, BBF^{+}00b, BU07, BR08, BDG04, BMP09, BDH15, Böt04, BEN94, CMP01, DN11, Drm15, ELT21, EH15, Erh95, Faß07, FGV09, Fro09, FLS17, FSS18, HS12, HS02, HN08b, JO05, Kei08, KV95, KM08, LM20, LS05, LY14, LT05, LT14, LR04, MPV15, MSS15, MVV09, Opf10, PP08, She10, SG98, Sim10, SF93, Smi10, Ste95, Ste97b, VVM09, VVM08, VK95, XBC19]. matricial [GRS00]. Matrix [LM20, LB96, LR04, RY15, AHS17, AEEG08, ABJE15, AM11, ABH17, AK10, BJR^{+}09, BEJ^{+}18, BM05, BK16, CP03, Car11, Car12, CGT05, Dub99, EKPR19, EST06, ERS17, FI20, Fro09, FH13, GW10, GMT16, HNP10, HMRS04, Hoc11, JVV12, KLS17, LZ08a, MS14, MY02, MACI06, NPR08, Pan19, PV05, Psa04, QvGvW^{+}15, ST08, SS08, Sin95, TMS97, TT05, TC08, VCK08, Vos99, WR09, Wat06b, WA08]. matrix-free [BK16]. matrix-vector [BM05]. Max [LT14]. Max-min [LT14]. maximum [FKK10, OS15]. maximum/minimum [OS15]. Maxwell [BPP15, EDG^{+}15]. MC [FCA21]. MC-based [FCA21]. MCMC [BHN13]. MCMC-based [BHN13]. mdLVs [IN11]. mean [JVV12]. meander [LP09]. meander-like [LP09]. means [BDR09]. measure [BDD06, SL12]. measurements [RKvdDA14]. measures [BCGAM08, EST06, HC05, Man06]. mechanics [JN07]. media [AMRT06, Tu05, Tu07]. Meixner [dMPPR99]. Mellin [JKM14]. membrane [PV09]. memory [ABHK09, SHH^{+}12]. meromorphic [BLM06]. mesh [ELST18, KS06, NV19]. meshes [BC09, CS94, DGP99, LO13, MC05, OR11]. meta [Jac10, Tif11]. meta-stable [Tif11]. metastable [FMSV08]. method [AEEG08, ABJE15, AS05a, ABA20, AH03, Ash17, ACS00, AH10, ABHK09, BZ03, Bar11, BMS08a, BFS08, BBK18, BEJ^{+}18, BBMP06, BL97, BOS08, BGS10, BW12, BPS17, BGyS21, BK16, BD20, Cab07, CRS94, Cal15, CFÁN09, CKD15, CW96, CR09, CQ01, Che09, CNO08, CG18, DdSV20, DGP99, DSC08, Dub01, ELST18, ELT21, EKPR19, EH15, EL05, EG06, EHJM17, EN08, FCA21, Gat07, GJV13, HOR08, HS08b, HJ16, HK14, HS97, Hip97, HLP10, Jac10, JO07, JS18, JG14, JN07, JO08, JR02, JS08, KS06, KK08, Kim07b, KR20, KL12, KN03, KRZ16, LMS97, LMW20, LMN10, LY14, LP08, Lóp16, LP17, LKF08, LR03, MYZ18, Mar07, Mar11, Mee98, MZLG14, Meu17, Mit97, MACI06, MG14, Mok15, MP08, NND06, NN16]. method [NNR99, NC06, Not10, OS98, Par21, PV09, Per00, PR20, PGH11, Sad15, SMS^{+}19, SSVS12, SvdVM98, STO08, ST02, TZ17, Woh00, Wu03, Yam06b, YG08, ZK17, ZY13, ZM13, Zít08]. methodology [CL03]. Methods [DGH14, AMRT06, AM06, AD04, BEG15, BBS19, BB00, BSS09, BB14, BKS15, BFM20, BRS08, Boc97, Bog05, BDH15, BKK^{+}18, BFH^{+}15, BMPS95, CCDR08, CDN09, CMP01, CCC07, Che97, CRZT20, DN11, DSSS13, EDG^{+}15, EMS09, EBJ21, FFK^{+}15, FP01, FSS97, FLS17, FJK^{+}20, Ful97, FMP96, Gan08, GS16, Gat04, GS15, GN13, GNR15, GW10, GSV12, GS94, GN16, GLMT21, GHS14, Gut12, HNP10, Han06, HK17, HKKR18, Hoh06, HLC10, JST05, JR04, Khr06, Kim07a, KKR18, KLRW19, KKR20, KS00, KL18, Lam05, Lam08a, Lam08b, LNG15, LT18, LM00, LMR15, MSV95, MMP09, MR93, MN08, NW03, NP19a, Nei14, Neu19, Ng94, Ore10, Pav99, PSS99, PPN13, Pry14, RT05, SG06, Sch05, SdS06, SGG15, Soo16, Sor98, SS98, SZ20, Sty08, TV07, TV18, Tos00]. methods [Val19, VVV19, VW97, Wei94, Xu18, ZB10, Zha04, HLP15, KNS19]. MHD [LMS97]. Micchelli [MPS18]. microscopy [EKPR19]. microwave [DENP09]. mid [Sch93]. mid-point [Sch93]. Mimetic [BC09, BRC^{+}16]. min [LT14]. min-max [LT14]. Minimal [Gla08, KV95, VK95, ABJE15, PR20, VCK08]. minimal-norm [PR20]. minimax [BCDO18]. Minimization [HS08a, Wei94, Kin11, Par21, RZ12, SHZF08]. minimization-based [Kin11]. minimizing [Kür19]. minimum [Del02, Kan08, OS15, Vos99]. minors [KM08]. MinRes [Kol12]. Miranda [AFHM04]. mixed [AMRT06, AM06, AL08, AD09, Bra06, CMO18, Gat04, Gat07, HS08a, Han06, HJV18, HK10, JS08, LMN10, MST08, NC06, Tu05, YG96]. mixed-hybrid [AMRT06]. mode [HK14, HR14]. Model [Auz18, BB08, BF12, Cha11, DDS^{+}18, DHI03, EL05, HK18, LMS09, MJ11, PK08]. Modeling [GLMT21, DE06, Raj08]. modelling [NB08a]. models [BDR09, JC17, SCCJC09]. modes [KMC11]. modification [SL12]. modifications [GO10, KL21, SGB07]. Modified [SG94, DT09, Lam08a, Mok15, Rap06]. moduli [HRV13]. Moment [EST06]. moments [BFM12, Lam08a]. Monge [DG06, Nei14]. monic [Psa04]. monomial [FvdMS14]. monomial-exponential [FvdMS14]. Monotone [Fro09, VC15, Bog05, MP08]. monotonicity [CR15, VK95]. Morley [Mar07, Zha04]. mortar [CW96, GS07, Kim07a, Mar07, Mar11, Woh00]. mortaring [HJ08b]. motion [VCL08]. moving [BEG15, HLX14, RFHS08]. MR [LT05]. MSC [TMS97]. MSC/NASTRAN [TMS97]. Mühlbach [Bre94]. Multi [CQ01, GN13, CC00, CL03, DdAR17, FL07, MX97]. multi-frequency [DdAR17]. multi-group [CL03]. multi-input [CC00, MX97]. Multi-parameter [GN13]. Multi-symplectic [CQ01]. multi-term [FL07]. Multidimensional [HNS04, BGSC20a]. multifrontal [EL05]. Multigrid [AVV14, BKK^{+}18, Dis97, FMOS21, Hip97, HS02, LT94, AO10, AH03, BZ03, BDH15, BL97, BFK^{+}10, BD20, Bun97, CS94, CL03, Che97, Den97, DHK^{+}00, Ful03, GW10, JC06, KLRW19, KN03, Kou97, LH12, LM00, Man03, Mit97, NW03, NN16, NP19a, NP19b, Not10, Oos03, PSC97, Per00, PSS99, Pop08, SSVS12, Ste97a, WO97, Woh00, Zha04]. multiharmonic [Kol12]. Multilevel [KNS19, KS00, LK21, MRT03, BBKL15, BFH^{+}15, DSC08, EN08, Ful97, JC17, LW06, MN08, Meu01, QvGvW^{+}15, Spr02, TMS97, VW97]. multinomial [FTRH20]. multiparameter [BF12, HP08]. multipatch [HLNT18]. Multiple [VV18, BBS19, EJS03, JST05, LZ08b, MYZ18, MZLG14]. multiplication [BM05]. multiplicative [ELST18, ELT21]. multipliers [MRT03]. multiply [CMR12, DK10, SL16]. multiresolution [MB06]. Multiscale [HKKR18, Bra97, FM02, MR18]. multishift [ABM93]. multisplitting [BMPS95, FMP96]. multivariate [BP07, EKPR19, Kro06, PT13, Zho06, vBHS14]. Müntz [Mok16, Nal19].

Nash [BCDO18]. Nash-equilibrium [BCDO18]. NASTRAN [TMS97]. nature [BC05]. Navier [Boc97, ET09, Mav97, Per00]. near [Van07]. nearly [BFK^{+}10, Smi10, Tif13, VA09]. negative [Boc97, Ped07]. network [AM06]. networks [BKO17, CEHT10, DDS^{+}18]. Neumann [Lóp16, BHN13, BFdP^{+}18, Che09, GKM16, KL12, Nas15, Tos00]. Newbery [Faß96]. Newman [COSV93]. Newton [BMS08b, WLHH18, BBMP06, Bor09, Bre94, HLC10, JO07, Kal20, NP96, Osb08, PR20, WE13, WSS98, ZY13]. Newton-ADI [BMS08b]. Nikiforov [CFÁN09]. Nikol’skiı [Tik05]. Nitsche [HJ08b]. nodal [SJ08]. nodes [Roş09]. Noise [HJ08a, Kin11, Kin13]. Non [FSS18, FMP96, GS07, dMPPR99, ACR99, AVV14, BC09, BKS18, CMO18, CN19, Dis97, Gat07, Haa00, HS97, HHO14, HLP10, Jac10, JJ10, MN08, Mok15, RZ12, Sim10, Sta03, XBC19, ZZW20, ZB10]. non-aligned [Dis97]. non-autonomous [HHO14]. non-convex [RZ12]. non-generic [JJ10]. non-homogeneous [Gat07]. non-linear [Sta03]. Non-matching [GS07]. non-overlapping [Haa00]. non-positive [ACR99]. non-regularized [AVV14]. non-relativistic [BKS18]. non-reversible [Jac10]. non-selfadjoint [XBC19]. non-smooth [CN19, Mok15, ZZW20]. non-stagnation [Sim10]. Non-standard [dMPPR99, CMO18, HLP10].

Non-stationary [FMP96]. non-symmetric [MN08]. Non-Toeplitz [FSS18]. non-uniform [BC09]. Noncommutative [Sch06]. nonconforming [AD04, Che97]. nondivision [Opf15]. nonharmonic [For02]. nonlinear [Ahm19, AS09, AG19, Bor09, CQ01, Che09, DENP09, DB08, FKK10, FK08, HP20, HLC10, JG14, Kle08, KS00, KV13, MP08, WO97]. nonlinearity [MYZ18]. nonlocal [APL04]. nonnegative [HNP10]. nonnegatively [Bar05, Bar11]. nonnegativity [CLRS04]. nonoverlapping [PV05]. nonsmooth [Sch08b]. nonstandard [BB01]. nonsymmetric [BP21, CFM^{+}10, FS08, LZ08b, MSV95, NB08a]. Nonuniform [AR14]. norm [ABJE15, Boc97, CGT05, Del02, HS08a, PR20, SG98]. normal [BSS09, LT14, SK06, SK08, Smi10]. normality [NPR08, TT05]. normally [FKSV07]. norms [EH15, Kro06]. note [Aré09, BMX99, Faß96, Gat04, GW19, GHS14, Neu16, SS09, Ves96, Yat06, ZY13]. null [MVV09]. number [AK10, BP07, Els05, JO17, Kan08, Kan15, Kan19a, Kan19b, Ste97a, TDLC09]. numbers [KS18, MS06]. Numerical [ACS00, BBD18, BJM99, BFR19, CMO18, DENP09, KS18, MR93, Meu01, MACI06, NB08b, SY09, SK06, SK08, STO08, TDLC09, WD08, Zha01, ÁNCQ06, ABM93, AH04, BMX99, CCDR08, Che09, DG06, DB08, Die97, FM02, FJK^{+}20, Gau02, GN14, GJV13, Gre04, GLMT21, LMS09, LMM12, Psa04, Rap06, Sch03, Sch05, TLF07]. numerically [Aré09]. numerics [FL07].

O [BR08]. Oblique [JST05, Lam10]. observer [EHJM17]. obstacle [BRS08]. obtain [AD09]. Obtaining [CGT05]. occasion [GOR^{+}08, RS14]. occluded [AC06]. ocean [DHI03]. odd [AM11]. ODEs [AS09, AS12]. off [SG99]. old [OSLS20]. once [GW19]. one [AL09, EH15, Ill03, JO14, MS08, RV08, She06, Yam06b]. one-dimensional [RV08]. one-sided [JO14]. onto [AS05a, CMR12, SL16]. Open [BMFJ^{+}06, HS08b]. Operational [Mok16, MG14]. operator [BD17, HS08a, JG14, KK08, Koe07, Lam10, NZ14]. operator-dependent [BD17]. operators [Aco06, ABH17, BFM12, Che97, FG17, HOR08, HLP10, Lor99, NNR99, NC06, RY09, Sus07]. Optimal [ADM07, Lis08, Ade17, Auz18, BMS08a, DSSS13, FFK^{+}15, JST18, KN03, Kol12, Meu17, Neu16, SAA09, Wac13, vBHS14]. Optimality [NU05]. Optimality-preserving [NU05]. optimization [BSS09, DHK^{+}00, JC17, LR03, QvGvW^{+}15, WLHH18]. optimization-based [JC17]. Optimized [YHL^{+}20, EDG^{+}15, LDB13a, LDB13b, Xu18]. Optimizing [Bir12]. optimum [Dub99]. option [KLS17]. options [Oos03, PPN13]. order [ABBS07, BMM97, BFR19, BMSR06, Bun97, CMM95, CP00, Cha11, CL03, Die97, EH15, FMR15b, GK14, Han06, HHP13, HP20, Kim07b, Kle08, Lam08a, LB96, MYZ18, Mar07, Mar11, Mar18, Mok15, Ng97, NC06, PSC97, Pav99, PK08, Sch08a, Sch08b, SSVS12, TW16, Zho06, LRL08]. orderings [BJM99]. ordinary [MYZ18, Ore10, Sch05]. Orthogonal [Aco06, CBGV05, Gau99, IL06, Koe07, LP09, AK06, BGyS21, Che06, Cot06, FLMT05, GCPP99, GAM05, Koe99, Lor99, MM99, Nar18, RÁRP06, Roz06, SL12, Sch02, Sim06]. Orthogonality [KMFO05, dMPPR99]. Orthonormal [VB95, MY02, Mas06]. orthonormalization [Zem17]. Oscillation [ABBS07, DH07]. oscillators [ÁNCQ06]. oscillatory [FLS20, Sha12]. Ostrowski [MJ11]. other [JO14, Sch05]. outer [BEN94, LNG15]. ovals [VK99]. over-penalized [BOS08, BGS10]. overlapping [Cal15, Haa00, HKKR18, MP08]. oversmoothing [HP20]. overview [BFH^{+}15]. oxymoron [Kny98].

Padé [BY98, GH94, Sta02, Sta06, Zho06]. page [Ano09b, Ano10c, Ano13b, Ano14d, Ano15c, Ano15e, Ano17d, Ano17e, Ano18d, Ano19b, Ano20d]. pair [COSV93]. pairs [GMT16, Pan19]. palindromic [Ahm19]. pantograph [BB00]. parabolic [Ade17, AS14, Bog05, DB08, DSSS13, FKK10, GKM16, GN14, HJ08b, HHO14, HLNT18, MP08, SG99]. Parallel [BMPS95, HC06, BZ03, BJM99, BM05, CMP01, CG18, DDR95, Erh95, FMP96, GW19, Haa00, HRT08, Mit97, SK06, SME02, VVM08, dJvdPDV17]. Parallelism [WE13]. Parameter [FvdMS14, HK18, KS06, KK08, PT13, AH03, BCDO18, BMS08b, BFR19, BD04, GN13, HOR08, HHP19, Kin11, Kin13, Kno08, LPS20, Pea15, Wu03]. parameter-dependent [HOR08]. parameter-free [LPS20]. Parameter-robust [HK18, Pea15]. Parameter-uniform [KS06, KK08]. parameterizations [YHL^{+}20]. parameterized [SY09]. parameters [BKS15, ENH10, KMFO05, Kür19]. parametric [Ber04, MJ11]. parametrized [Faß07]. Parareal [DSSS13]. Parareal-in-time [DSSS13]. Part [CG18, FFK^{+}15, LDB13a, LDB13b]. partial [GJV13, VCL08, Zem17]. particular [JKM14]. partition [Leo06, Mit97, VK95]. partitioned [KV95, TLL15, VK95]. partitioning [Els05, MH17]. partitions [Leo09]. Past [Ano20a]. path [Dos08, JJ10]. PDE [Lam08a, QvGvW^{+}15, RKvdDA14]. PDE-constrained [QvGvW^{+}15]. PDEs [GW19, Hoh06]. penalized [BOS08, BGS10]. penalty [BOS08, BGS10, BW12, BPS17, HP20]. pencil [EKPR19, VW13]. pencils [BMX07, Hoc11, MX15]. Performance [BFM20, FO03, Xu18]. perilous [Lau08]. period [DV12]. Periodic [Rom07, Cab07, FL07, Hen94]. permeability [DdAR17]. perspective [Pry14]. perspectives [HS08b]. Perturbation [AM11, Ahm19, AAS21, CC00, SME02, TLL15, CKP15]. perturbations [CGM10]. perturbed [Bog05, Bra06, BFR19, ELST18, EG06, GN14, KS06, KK08, LR16, LK21, Mar18, XS20]. PET [Bar11]. phase [AL09, KL21, LP17]. photosynthetic [SCCJC09]. physical [BFK^{+}10]. physiology [PV09]. Pick [Ped07]. pictures [Sim06]. Piecewise [ER01, Kim07b, Bre04, HR06]. piecewise-constant [HR06]. pinch [BHRT03]. pipe [Kan08, Kan15, SK08]. pivoting [LS05, SG06]. placement [MX96, MX97]. plane [LM00]. plasma [JS18]. plate [SG94]. plates [SK06]. Play [CKH14]. Plug [CKH14]. Plug-and-Play [CKH14]. plus [ER14, Ng94, VVM09]. PML [Lis08]. POD [HHP13]. Poincaré [Bre04, LO13]. point [BBS19, BL06, DDR95, GS06, GO10, HC06, JO08, KK08, Kou97, LO07, LN10, Mar18, NB08a, RV08, Sch93, SdS06, Sim10, SME02, Woh00, dJvdPDV17, vBHS14]. points [AK06, BCR98, BCDV06, BL08, EV93, GS16, JJ10, JO14, MS14, Rom07, SV10a]. Poisson [BB01, Bol08, Yam06b]. polar [Dub99, SSVS12]. pole [MX96, MX97, Van07]. poles [Van07]. Pollaczeck [MMN18]. Pollaczeck-Laguerre [MMN18]. pollution [LHL^{+}20]. polygonal [LMN10]. polygons [For02]. Polynomial [Mas02, MMN18, MW02, Opf15, Ran07, SEH14, BGP04, Bre04, BFH^{+}15, CGM10, FMMS08, Fer05, FP01, Gre04, HMRS04, JR02, MYZ18, MFMGO05, MV15, Nar18, SV10a, TV18, TLF07, VB95, vBHS14]. Polynomials [Opf10, AM11, Arv06, AK06, BH20, BP07, Che06, Cot06, CBGV05, DJJ11, EV93, Faß96, FLMT05, GCPP99, GAM05, Gau99, Gau10, GM18, GN16, GRS00, HC05, IL06, JO14, JO17, Kei08, Koe99, Koe07, Kro06, KS18, KMFO05, Lor99, MS14, MM99, MY02, MF06, Ost10, Psa04, RÁRP06, Roz06, SL12, Sch02, SSS10, Sim06, Sta02, Sta06, VV18, VC15, WA08, Yat06, Zha01, dMPPR99]. porous [AMRT06, Tu05, Tu07]. posed [CN19, Kin11, NR14, Neu19]. positive [ACR99, BFM12, CMP01, FSS18, LT05, ZB10]. Positivity [IN11]. possible [DTM17, Sch16]. posteriori [DGP99, Gat04, HJV18, LR16, NC06, PGH11]. potential [AR07, AMRT06]. power [CRZT20, HS12, KS18, Mál08]. power-law [HS12]. power-law-like [Mál08]. powers [Böt04, BFM12]. practice [MMP09]. precision [AD09, Lau08, ST02]. Preconditioned [BBS19, GS15, Kny98, Pan19, CR09, EN08, LY14, NZ14]. preconditioner [BPS17, CW96, CNP94, DHI03, GGM12, Kim07a, LN10, LH12, Mar11]. Preconditioners [BY99, DK13, GS06, AK10, BBD18, BJM99, BU07, BBMP06, Bre03, BGyS21, CP00, CFM^{+}10, Che99, CT03, CPS18, Dob12, ET09, GO10, GHS14, Han06, HC06, HLC10, KLRW19, KN03, LO07, LK21, MRT03, Meu01, Ng97, Pea15, QvGvW^{+}15, SM08]. Preconditioning [CP03, HN08b, KLRW19, NB08a, SEH14, AVV14, Ber04, CMP01, GW19, HS02, Kim07b, Soo16, WR09, WSS98, Zha00]. prediction [Ful97]. predictor [AS12, PPN13]. predictor-corrector [PPN13]. Preface [Ano10a, Ano15a, Ano14a, Ano14b, Ano17a, Ano18a, Ano18b, Ano20b, GOR^{+}08]. premature [LKF08]. prescribed [XBC19]. Present [Ano20a]. Preserving [CKH14, BFS08, LC14, MX15, NU05]. pressure [CMM95, SK06, SK08]. pricing [KLS17, PPN13]. primal [BBK18, CW16, KCW18, SM08]. primal-dual [BBK18]. primes [Pri06]. principle [LPS20]. principles [FKK10, Sel02]. Prior [DGH14]. Probability [BCDO18, BP07]. Probing [SdS06]. Problem [SEH14, AR07, Ade17, AG19, AL09, AVV14, AMRT06, AH10, BLM06, BdPNdP14, BMS08b, Bra06, BGS10, CMO18, CGM10, Che09, CPS18, CKP15, CN19, DG06, EG06, HP08, KLRW19, KL12, KM08, Mar07, Mar11, MX96, MX97, PMLFT09, Pri06, QvGvW^{+}15, RR07, RS12, SV10b, TZ17, Vos03, WLR02, Yam06b, Zha04, ZC19]. Problems [DGH14, Ahm19, AR09, AL08, AH03, Ash17, ADM07, AH04, BBS19, BMFJ^{+}06, BB01, BMX99, BFS08, Bir12, BY99, BRC^{+}16, BRS08, BL06, Bog05, Bor09, BBKL15, BFR19, BK16, CRS94, CP00, CS94, DSSS13, ELST18, FCA21, FKK10, FK08, FMR15a, GKM16, GS15, Göt06, GN14, GHS14, HKLW20a, HKLW20b, HJ08b, HS97, HHO14, HLNT18, HC06, KK08, KLS16, Kal20, KJ09, Kin11, Kin13, KKR18, KKR20, KN03, Kol12, KV13, KL18, KRZ16, Lam05, LV08, LO07, LN10, LMN10, LT14, LR16, Lis08, LMR15, Lub06, LR03, MR18, MM04, Mar18, MST08, Mav97, MW02, MS07, MP08, NR14, Neu19, NZ14, NB08a, Ng94, NRS17, Oos03, PW13, Pea15, PSS99, RKvdDA14, SdS06, Sor98, SZ20, TV07, TLL15, TMS97, Tos00, TW16, VK99]. problems [Wac13, WLHH18, WR09, Woh00, WSS98, XS20, ZM13, dJvdPDV17]. process [AHS17, BY98, Meu15]. processes [Raj08, Sch03]. processors [JC06, dJvdPDV17]. producing [Raj08]. product [ABH17, DO18, Drm15, LNG15, Ste97b, VB95]. products [BDR09, BEN94, BMSR06, GW10, GHS14, Mee98, Ste95, Ste97b]. programming [AM06]. programs [JO08]. projected [Dos08, HNP10, Sty08]. projected-gradient [Dos08]. projecting [AS05a]. projection [DJ06a, GNR15, JST05, KS00, Lam10, Lóp16, LRL08, MST08, Osb08, STO08]. projectors [She10]. Proof [Zei99]. propagation [HJ08a, PK08]. properties [EMS09, FJK^{+}20, KM07, Mál08, MY02, Ram08, Rap06, SMS^{+}19, Wei94]. property [Pop08]. proximal [FG17, Val19]. Pseudo [Par21, Zho06]. Pseudo-linear [Par21]. pseudo-multivariate [Zho06]. Pseudospectral [Lui11, CQ01, EBJ21]. pumps [BKO17]. pyramids [GH99].

Q [Meu17]. Q-OR [Meu17]. QR [BGP04, BDG04, Kre05, MH01, OB05, Ste95, VVM08]. QR-factorization [VVM08]. QR-factorization/solver [VVM08]. QR-like [BGP04]. Quadratic [ČF18, BFS08, Dos08, Fat98, FJK^{+}20, MS14, MV15, Sta02, Sta06]. quadratically [HMRS04]. Quadrature [Atk04, AS05b, GM06, RGVP99, ACR99, BCGAM08, CS14, DV12, DNV05, DDPS20, ER01, Gau99, Gau06, Ill03, JST18, MJ11, MJM16, MPS18, Ng97, Not16, PPS18, Roş09, Sin95, VV18, CDN09]. Quadrature-free [GM06]. quadratures [Che06, CBGV05]. quadrilateral [HK10]. quadrilaterals [HRV13]. quality [BSDMGFTR09]. Quantum [Arv06, Man08, Roz06]. Quasi [BBMP06, FFK^{+}15, GM06, HK17, NR14, SHZF08, WLHH18]. quasi-cyclic [HK17]. quasi-interpolation [GM06]. quasi-Lagrange [WLHH18]. Quasi-Newton [BBMP06]. Quasi-optimal [FFK^{+}15]. quasi-residual [SHZF08]. quasi-solutions [NR14]. quasilinear [MS07]. quasiseparable [EH15, VVM08]. quaternion [AAS21, CS99, JO05]. quaternionic [JO07]. quaternions [Opf10]. question [She06]. questions [HS08b]. quotient [Fat98, Hen94, OS98]. QZ [MS06, Sor98].

radial [ÁNCQ06, GSV12, LW06]. radiation [Sta03]. radii [KQT13]. radiosity [ACS00, AC06, NND06]. Ramanujan [IL06, Sch06]. Random [SCCJC09, BP07, Els05]. Randomized [SGG15, EKPR19]. range [FJK^{+}20, Psa04]. rank [BKK^{+}18, BSS16, DS11, GO10, GM04, KL21, Kür19, LM20, MS08, NP20, SGG15, SGB07, Sty08, WA08]. rank- [NP20]. rank-1 [KL21]. rank-deficient [SGG15]. rank-one [MS08]. rate [HHP19]. rates [EBJ21, FFK^{+}15, Ger17, HP20, Kin13, NR14, Neu16, RR10]. Rational [BB14, Bar06, Bor09, FI20, Gla08, GPT11, GK14, Ill03, MPV15, Pri96, RGVP99, SL12, VVV19, Van07, VC14, Vos03]. rationally [FGV09]. Raviart [LMR15, YG96]. Ray [Pri96, BL97]. Rayleigh [Fat98, Hoc05, HP08, OS98]. RCWA [HS08b]. reaction [DE06, EG06, LR16, LK21]. reaction-diffusion [DE06, EG06, LK21]. real [BDD06, CBGV05, KM17, Mas02, MMN18, NPR08, TC08, XBC19]. reconstructor [RS21a, RS21b]. Recovery [AR14, DdSV20, Lam10]. rectangles [Hoh06]. rectangular [MC05, Pop08, SGB07]. recurrence [ÁNCQ06, CFÁN09, Koe07, LB96]. recurrences [GH94, Wei94]. Recursive [Nov05, Soo16]. Recycling [GS15]. Reduced [DS11, HOR08]. Reducibility [GS94]. reduction [Auz18, BB08, BFK^{+}10, Cha11, FMOS21, FGV09, JOT08, PK08, Ran07, Wat06b]. reduction-based [BFK^{+}10]. reductions [JO05]. reference [RT05]. refined [Hoc05, OR11, TC08]. refinement [BMM97, DDS^{+}18]. reflectors [SS09]. regimes [BKS18]. region [Bar05, FLS20, LHL^{+}20]. regions [BSDMGFTR09, Hoc11, LP09, Leo06]. regression [FTRH20]. regular [ZB10]. Regularization [BL06, CKH14, Ram08, BD17, CLRS04, CNO08, DJ06a, GNR15, Ger17, HHP19, HP20, Kin13, Neu16, RR10, RY09, RY15, VW97, Wac13, Wu03]. regularized [AVV14, LV08, LV14, PR20, WNK18, ZM13]. regularizers [DdSV20]. regularizing [DSC08, HJ08a]. Reichel [RS14]. related [FMMS08, Gau06, Göt06, JO17, LB96, Ped07, RR07, Roz06, TT05]. relation [GN16, KM96, ZY13]. relations [ÁNCQ06, CFÁN09, Mál08]. relationship [Soo17, VCL08]. relativistic [BKS18, GGCT09]. relaxation [ENH10, Gan04, GKM16, LM00, Pop08]. remark [Bar06]. Remarks [PP06, YG96, Zít08]. Remez [Yat06]. Remez-type [Yat06]. Renyi [BFdP^{+}18]. representation [DFV09, DJ06b]. represented [ABH17]. research [Bra97, Not16]. Residual [LPS20, ABJE15, CR09, Gat04, Kür19, LMSY14, SHZF08]. residual-based [Gat04]. residual-minimizing [Kür19]. resolvent [SG98]. response [TLL15, TZ17, ZC19]. restarted [Bre16, CRS94, HRT08, Sch16, ZN17, Zít08]. Restarting [SS98, BR14, WE13, ZC19]. restoration [Bar05, CCC07, LPS20, NO02, ZM13]. restricted [OB05]. restriction [PP06]. result [KRZ16]. resultant [WA08]. results [BCDV06, BBKL15, DDPS20, EP07, GK08, Gre04, HP20, MN08, PQRT06]. retards [LR03]. Retooling [Dub01]. revealing [FXG19, GM04]. reversible [Jac10, Tif13]. revisited [Arv06, LT14, LT18, Sad15, Sel02]. Revisiting [BdPNdP14, MV15, NP19b]. Reynolds [Kan08, Kan15, Kan19a, Kan19b, Ste97a, TDLC09]. Riccati [ABM93, BMS08b, BP21, BMP09, EMS13, HJ09, LC14]. Richardson [FMOS21, MP08]. Riemann [MF06]. right [BBS19, BK02, EJS03, JST05, LZ08b, MZLG14]. right-hand [BBS19, BK02, EJS03, JST05, LZ08b, MZLG14]. Ritz [Bea98, DTM17, Hoc05, HP08, Meu15, PP08, ZN17]. Rivlin [MPS18]. Robbins [Zei99]. Robust [GPT11, LR16, BZ03, BGP04, BHRT03, Dob12, GN14, HRT08, HK18, Jac10, Kol12, Pea15]. robustness [WE13]. role [Sch08a]. root [BGP04, Car11, Car12]. root-finding [BGP04]. roots [Gra16, JO07, MV15]. rotated [MC05]. rotating [STO08]. rotation [GK08]. rotations [OB05]. Roundoff [YNYF15]. row [AS05a]. RQ [Yan98]. RQ-iteration [Yan98]. RRB [dJvdPDV17]. RRB-solver [dJvdPDV17]. rule [DO18, LWH20, Sch93]. rules [ACR99, Bar11, BCGAM08, DRST16, Gau06, JST18, Kin11, Kin13, LMM12, MJM16, Ng97, OSLS20, Van19]. Runge [AS14, Bir12, GS94, LT18]. Rush [PV09].

saddle [BBS19, BL06, GS06, GO10, HC06, LO07, LN10, NB08a, SdS06, Sim10, Woh00]. saddle-point [BL06, SdS06]. Saff [She06]. sampling [Ash17, BP07]. satisfying [ČF18]. Scalable [JC06, MR18]. scalar [GHS14]. scale [ABJE15, ABHK09, BMS08b, BFS08, Cha11, HJ09, LV14, LR03, Sor98]. scaled [FG17]. scales [EP07]. scattered [GK08]. scattering [CL03, GSW08, HR14]. scheduling [BC05]. scheme [BPP15, Bra06, GGCT09, GN14, SSVS12]. schemes [AS09, AS12, ADW15, BC09, BEN94, CP00, HOR08, Mar18, MS07, WO97]. Schmidt [Bar15]. Schrödinger [BKS18, CQ01, GSV12, Kle08]. Schur [GH94, KLS16, Kal20, Wat06b, Zha00]. Schur-type [GH94]. Schwarz [Bre03, BGyS21, Cal15, CPS18, CG18, DK10, DHI03, ELST18, ELT21, EDG^{+}15, FSS97, Gan08, Han06, HKKR18, HC06, LDB13a, LDB13b, LMR15, LK21, Mar07, MR18, MP08, Par21, Xu18]. scientific [BC05, BDR09, Bra97]. second [ABBS07, HHP13, LP08, Mok16, NNR99, NC06, Sch08b, TW16]. second-order [Sch08b, TW16]. sections [JN10]. sector [LZ08a]. segmentation [JC17]. Seidel [Sil03]. selection [BMS08b]. Self [BKS15, BFM12, EST06, GS15, GHS14, KK08]. self-adjoint [BFM12, GS15, GHS14, KK08]. Self-generating [BKS15]. self-similar [EST06]. selfadjoint [XBC19]. Semi [ENH10, LC14, WR09]. Semi-convergence [ENH10]. semi-iteration [WR09]. semi-stabilizing [LC14]. semiaxis [Mas02, MMN18]. Semicoarsening [Den97]. semiconductor [SMS^{+}19]. semiiteration [HH93]. semilinear [AH04, Bog05, DB08, Gan04, HL20, NB08b]. seminorms [BD17]. semiseparable [Kei08, QvGvW^{+}15, VVM09]. sense [JST18, MS08]. sensitivities [SJ08]. sensivity [BD04]. Separable [Osb08, Bor09, SY09]. separation [Gra16]. sequences [BGSC20b, BGSC20a, Ber04, CP03, DJJ11, Pri96]. sequentially [QvGvW^{+}15]. series [DT09, For02, GRS00, KS18, LWH20, Ste02]. set [CKV04, DTM17, JST18, Lei19, OB05, VCK08]. sets [AO10, AK06, FKSV07, KV95, VK95]. shallow [DHI03]. shape [MH17]. Sharp [ZN17, Gau10]. sharpness [Var01, Yat06, VK95]. shift [BKS15, Kür19, Mee98]. shift-invert [Mee98]. Shifted [CRZT20, BDG04, EN08, SHZF08, Soo16]. Shishkin [ELST18]. shocked [HLC10]. short [ČF18, Wei94]. sided [JO14]. sides [BBS19, BK02, EJS03, JST05, LZ08b, MZLG14]. signal [For02]. Signorini [BB01]. similar [EST06]. Simple [RY09]. Simpler [LZ08b]. simplex [Xu06]. simplification [LP08, LP17]. simplified [KR20]. simulation [DDS^{+}18, GLMT21, NW03]. simulations [BHRT03, Ema10, JS18, SMS^{+}19]. Simultaneous [HHP19, VV18]. sinc [Hoh06, NND06, Yam06a, YG08, EG06, Yam06b]. sinc-convolution [NND06]. single [MX96]. single-input [MX96]. Singular [FKSV07, Atk04, BFK^{+}10, Che94, FFK^{+}15, FXG19, GS06, Hen94, Hoh06, IN11, JO14, JR02, JR04, JKM14, LY14, MS14, Man06, MM04, MR93, OS15]. singularities [JKM14]. singularity [Ill03]. singularly [Bog05, Bra06, BFR19, ELST18, EG06, GN14, KS06, KK08, LR16, LK21, Mar18, XS20]. singularly-perturbed [EG06]. SIRT [ENH10]. sixtieth [RS14]. size [LMM12]. sized [CG18]. sizes [Sch05]. skew [AM11, BMX99, BBF^{+}00a, BBF^{+}00b, BMX07]. skew-Hamiltonian [BMX99]. skew-symmetric [BBF^{+}00a, BBF^{+}00b, BMX07]. skew-symmetric/symmetric [BMX07]. slices [BLM06]. Slit [DK10, CMR12]. slope [Sch08b]. small [FL07, Leo06]. smooth [CN19, LW06, Mok15, ZZW20]. smoothed [BDH15]. smoother [Sil03, Ste97a]. smoothers [Bir12, LM00]. smoothing [BFH^{+}15, GN16, HNS04, Mav97, Pop08, RY09, TV18]. Sobolev [Bre04, BMSR06, GCPP99, PQRT06, RÁRP06]. Softmax [FTRH20]. Software [Koe99]. solid [SV10b, Vos03]. Solution [Hoh06, NRS17, ABM93, AL09, AH04, AC06, BMX99, BP21, BMP09, Bol08, CR96, Che94, DG06, Die97, EHJM17, FKK10, FMR15b, GJV13, Kal20, KN03, KL18, Nas15, Neu19, Sta03, TMS97]. solutions [Aco06, BEG15, Cab07, DB08, FL07, HJ09, LC14, MYZ18, MFMGO05, Mok15, NB08b, NR14]. solvability [BKO17, MM09a]. solver [HRT08, Kol12, PLM03, VMV07, VVM08, dJvdPDV17]. solvers [BHRT03, BSS16, Gut12, PW13, Sta03]. Solving [BFS08, VMV07, ZM13, AM06, BBS19, CCC07, CL03, CR09, EG06, EDG^{+}15, GGCT09, HMRS04, KL12, LV08, Meu17, MS08, NND06, Ng94, Pea15, PV09, SHZF08]. Some [BB01, Gre04, She10, Zít08, AH04, BSS09, DDR95, DJ06b, FKK10, Gat04, MS14, NB08b, PR20, Rom07, Sch08b, OSLS20]. sonic [Dis97]. SOR [HS08a]. sorted [AO10]. source [BLM06, MM04, Neu16]. Space [BRS08, AS14, Bol08, GHS14, Gut12, HJV18, HKKR18, HKLW20a, HKLW20b, HLNT18, MVV09, PCP06, Sch03, SZ20, WLHH18]. space-time [AS14, HJV18, HLNT18, SZ20]. spaced [Roş09]. spaces [CMO18, HKKR18, KRR16, KKR20, MS06, MR18, Mas02, PQRT06, RÁRP06]. spacing [VA09]. Sparse [AR14, BB08, BR08, BM05, Bun97, CFM^{+}10, Che99, DdSV20, Erh95, GW10, LS05, PV05, PCP06, RT05, SG06, TMS97]. Sparsity [MH17, LMM12, Ram08, RR10, RZ12]. Sparsity-inducing [MH17]. spatially [JG14, LDB13a, LDB13b, Val19]. Specht [SG94]. Special [BJR^{+}09, RS14, BK02, ELT21, HK14, Koe99, KS18, GOR^{+}08]. SPECT [Bar11]. spectra [XBC19]. Spectral [AR09, GGM12, Gut12, KLS16, AH10, Aur15, BK16, DHI03, HS08a, HJ16, Jac10, Kal20, KV95, Lam05, Lam08a, Lam08b, Mee98, PPN13, Vos03]. spectrum [FXG19, SF93]. spectrum-revealing [FXG19]. sphere [AS05b, BP03, Fer05, GM06, Leo06, Leo09, TDLC09, Yat06]. Spherical [Roş09, FM02, SSVS12]. spheroidal [MACI06]. spin [KM07]. spin-stabilized [KM07]. spline [ČF18, KK08, Kha06, LR16]. splines [VC14, VC15, ZK17]. split [DdSV20]. splitting [CGT05, JS18, ZB10]. square [Boc97, RY09, VCK08]. squares [Aco06, BMM97, BY99, Bor09, CMM95, CNP94, CL03, Faß96, GPT11, HJ16, KS00, LV08, LV14, MS08, Ng94, Osb08, Pav99, VB95, WNK18, ZM13]. SR [Faß07, SAA09].

Stability [BD04, CCDR08, GK08, KM07, Sch05, AS14, AD09, BFM20, FH13, HK18, Kou97, MV15, Ste97b]. stabilizability [KQT13]. Stabilization [MST08, BL06, LRL08, Sch08a]. stabilized [KM07]. stabilizing [LC14, RV08]. Stable [MB06, APL04, AMVW15, Jac10, LRL08, PSC97, Tif11]. stage [BMPS95, CMP01]. Stagnation [Soo17, Meu12, Sim10]. staircase [BMX07]. standard [Bre03, CMO18, HLP10, KL18, SvdVM98, dMPPR99]. standing [BL97]. state [DN11, Per00]. states [FMSV08, Jac10, Tif11]. stationary [FMP96, LMS97, LP17, RV08]. statistically [Bar11]. Steady [DN11, CR09, Per00]. Steady-state [DN11, Per00]. steepest [AEEG08, NZ14]. stencil [dJvdPDV17]. step [AL09, IE17, LMM12, MSV95, Sch05]. stepped [FJMS06]. stepping [Kle08]. Stieltjes [DJJ11]. still [Lub06]. stochastic [ADW15, BB00, BC05, DN11, LMW20, RKvdDA14, Sch03, Sch05]. Stokes [AL09, BBD18, Boc97, CMM95, Dob12, ET09, GS07, HJ16, Mav97, Pea15, Per00, WD08]. Stopping [AL08, Bar11, HJV18]. story [Lau08]. strain [Ran07]. strangeness [LT18]. strangeness-free [LT18]. strategies [Bre16, CP03, DDS^{+}18, Spr02]. strategy [PSC97, SHZF08]. stratified [TDLC09]. streaming [GW10]. Strong [GM04, DK13, RR07]. strongly [TDLC09]. Structural [CFÁN09]. Structure [MX15, BFS08, DS11, ELT21, LC14, MW02, QvGvW^{+}15, Sim06]. structure-preserving [BFS08, LC14]. Structured [FMMS08, WA08, BBD^{+}08, BSDMGFTR09, BC09, BKK^{+}18, BMX07, DHK^{+}00, LT18, MM04, MX15, NPR08, Smi10]. structures [FJMS06, GLMT21, HR14, SV10b]. study [ÁNCQ06, AC06, CT03, HS08b, LHL^{+}20, SK06, SK08, STO08, TDLC09]. style [Oos03, PPN13]. subdomains [Cal15, CG18]. Subgaussian [AR14]. subgraphs [DK13]. subimages [NO02]. subintervals [DV12]. subject [SGB07]. subspace [BBS19, BSS09, BFS08, DdSV20, DSSS13, FP01, FLS17, GS15, GHS14, Lam05, Lam08b, LY14, MSS15, Meu17, SS08, VVV19, WO97, Wei94, ZN17]. subspace-accelerated [DdSV20]. subspaces [AS05a, MPV13, MPV15, SvdVM98]. Substitution [FI20]. substructures [PV05]. substructuring [BW12, KRR16, SM08]. suitable [dJvdPDV17]. sum [OS15, PMLFT09]. summation [Sch06]. sums [FvdMS14, KM17, PT13, VCL08]. sup [LRL08]. Superlinear [BK02]. superresolution [EKPR19]. SUPG [Kno08]. support [BDD06, ČF18, JOT08]. supported [HS12]. surface [MB06, YHL^{+}20]. surfaces [Atk04]. survey [DDPS20, JVV12, MMN18, Not16, RÁRP06]. SVD [Cha11, Drm15, FMSV08, HRT08, JJ10, Ste97b, Tif11, WL12]. SVD-based [Tif11]. SVD/Krylov [Cha11]. swallowtail [FLS20]. Sylvester [ABJE15, BB14, BKS15, EHJM17, KV13, WA08]. Sylvester-observer [EHJM17]. symbolic [KL21]. symmetric [AM11, AMVW17, BBF^{+}00a, BBF^{+}00b, BB14, Ber04, BOS08, BGS10, BPS17, BMX07, CRS94, FGV09, KLS16, Kal20, KN03, LS05, LNG15, LT05, LK21, MPV15, MN08, SG06, SHZF08, SS98, Ves96, Vos99, WSS98]. symmetric/symmetric [BMX07]. symmetrized [BMSR06]. symplectic [ADW15, CQ01, GS94, MSS15, SS09, SAA09]. synchronous [BMPS95]. synthesis [HK14]. system [BMM97, CMM95, CL03, GGM12, JO08, Pav99]. Systems [Cot06, AS09, AS12, BKS18, BB08, Ber04, BFK^{+}10, CR96, CC00, CFM^{+}10, Cha11, Che94, Che99, CS99, Den97, EJS03, Ema10, FMMS08, Fra21, FJK^{+}20, GSW08, GS06, GO10, HH93, JST05, KQT13, LB96, LZ08b, LK21, MSV95, MZLG14, MN08, Meu17, MS08, NN16, NB08a, Pop08, SG06, SGG15, SHZF08, Soo16, VMV07, ZB10]. Szego [DNV05, DRST16, LT94, MF06].

T [BP21]. T-Riccati [BP21]. Table

[Ano09a, Ano10b, Ano13a, Ano14c, Ano15b, Ano15d, Ano16, Ano17b, Ano17c, Ano18c, Ano19a, Ano20c, GH94]. tale [Sim06]. Tau [MG14, Mok15]. Taylor [Sta06]. Tearing [HLP15]. technique [CKD15, KM08, NV19]. techniques [KLS16, SS98, WSS98]. technology [PLM03]. Tensor [BRC^{+}16, BKK^{+}18, CRZT20, Lei19, LNG15, OS15]. tensor-structured [BKK^{+}18]. tensors [WNK18]. term [FL07, HP20]. th [ÁNCQ06, Car11, Car12]. their [CS14, DRST16, EMJ16, FL07, GM18, GN16, GRS00, KM96, Man06, She10, Var01]. theorem [CR15, Gau06, LZ07, Lui11, PQRT06, CC97]. theorems [AFHM04, Gau06, GMT16, Nal19, Sch03, Var01]. theoretical [Gre04, Mee98]. Theory [FL07, LMM12, BMFJ^{+}06, BGSC20b, BGSC20a, BSS09, EMS13, FM02, FS08, Gut12, MMP09, VCL08]. thermal [KJ09]. Thick [ZC19]. thin [MM09b]. thin-layer [MM09b]. Three [TT05, Hip97, HK18, Kim07a, KLRW19, KL12, Nei14, XBC19]. three-field [HK18]. three-level [KLRW19]. Tianjin [LHL^{+}20]. tidbits [She10]. Tikhonov [BD17, CLRS04, GN13, GNR15, HP20, LMM12, Ram08, RZ12, RY15, Wu03]. Time [DE06, HLNT18, AS14, AL09, CKP15, EDG^{+}15, EP07, FMOS21, FMMS08, FMR15b, Fra21, Gan08, HJV18, HJ09, KQT13, Kle08, Lam08a, LDB13a, LDB13b, SZ20, DSSS13]. time-delay [FMMS08]. time-dependent [Lam08a]. Time-discretization [DE06]. time-harmonic [EDG^{+}15]. time-invariant [KQT13]. Time-multipatch [HLNT18]. time-step [AL09]. time-stepping [Kle08].

Title [Ano09b, Ano10c, Ano13b, Ano14d, Ano15c, Ano15e, Ano17d, Ano17e, Ano18d, Ano19b, Ano20d]. Toeplitz [BGSC20b, BGSC20a, BDH15, Böt04, CP00, CNP94, ER14, FGV09, FSS18, HS02, HN08b, LT05, MVV09, Ng94, Vos99]. Toeplitz-like [MVV09]. Toeplitz-plus-Hankel [Ng94, ER14]. tomography [Auz18, RS21a, RS21b]. Topological [AG19, BKO17]. torsional [KMC11]. total [BBK18, LV08, LV14, Par21, VW97]. trace [BFM12]. track [Ful97]. training [JOT08]. trajectory [WLHH18]. transboundary [LHL^{+}20]. Transfer [SG98, BY98, Che97, LKF08]. transform [KJ09, Yam06a]. transformation [AMVW17, BD04, JO05, KKR18, KKR20, Mee98, YG08]. transformation-of-basis [KKR20]. transformations [SAA09, WL11]. Transformed [NP20]. Transforming [DFV09]. transforms [Che99]. Transient [Böt04]. transition [Kan08, Kan15, Kan19a, Kan19b]. transmission [LMN10]. transmission/mixed [LMN10]. transport [EMS13, Sta03]. treatment [BFR19]. tree [EL05]. tree-based [EL05]. trends [MY02]. triangles [MB06]. triangular [BU07, GGM12, KLS17, VW13]. tribute [Gau02]. tridiagonal [AMVW17, CS99, FGV09, FSS18, LT05, NPR08, XBC19]. Trigonometric [DV12, BP07, Faß96, Nav05]. Truncated [Sor98, DSSS13, MS06, Yan98]. truncation [Ful03, NV19, Xu18]. truncations [DRST16]. trust [Bar05]. tuples [Sch08b]. Turán [CS14]. turbulent [Auz18, Kan08, Kan15, Kan19a, Kan19b]. Two [Cha11, Han06, HLC10, Soo16, AL09, BBD18, BFR19, Bre03, BMPS95, Cal15, CMP01, CGT05, DG06, HK14, HHP19, KK08, KRR16, LMS97, Mar18, NP19a, Sim06, TV18]. two-dimensional [Bre03, HK14]. Two-level [Han06, HLC10, BBD18, Cal15, LMS97, NP19a, TV18]. two-parameter [BFR19, HHP19]. two-point [KK08, Mar18]. two-stage [BMPS95, CMP01]. Type [DGH14, ÁNACS07, BKS18, BCDI15, Bra06, CFÁN09, Cot06, CKV04, DdSV20, DDPS20, Fra21, FJK^{+}20, GAM05, Gau06, GM18, GH94, Ill03, JKM14, KS18, Lei19, Lén06, LWH20, MJ11, MST08, MG14, Neu19, NV19, Nov05, Pri96, RS12, Sch02, Sch05, Soo16, VK99, Var01, WL11, Yam06b, Yat06, dJvdPDV17]. types [Bog05].

U [Peñ04]. unbounded [Kol12].

Uncertainty [Sel02]. unconditionally [APL04]. uncoupled [Tif13]. unidimensional [BGSC20b]. unified [Nei14, PD17]. Uniform [Bog05, Del02, ÁNACS07, BC09, KS06, KK08, VA09]. Uniformly [Bra06, WLR02]. unilateral [JO17]. unimodular [Gla08]. union [Hoh06]. Unique [MM09a]. uniqueness [Bar06]. unit [BDD06, CBGV05, Gla08, Leo06, Leo09, MS14, SL12, Sin95, Xu06]. unitary [AMVW15, DFV09]. unitary-weight [DFV09]. univariate [BL08]. unknown [Bar05]. unsteady [Bir12]. unstructured [CS94, DHK^{+}00]. unsymmetric [EL05, SF93, SV10b, TC08]. unweighted [HS12]. update [SGB07]. updating [MS08]. upper [Cab07, VW13]. use [KJ09, Kre05, Yam06a]. Using [AD09, Ash17, ACS00, ABHK09, BEG15, BFdP^{+}18, BFS08, CL03, CS14, DDS^{+}18, For02, FH13, HNS04, Hoh06, KN03, Lam08a, Mit97, NP96, OB05, PMLFT09, PSC97, SHH^{+}12, TMS97, Zha04]. Uvarov [CFÁN09].

Vaidya [CT03]. validation [BDR09]. value [Ash17, BK16, Che09, DTM17, FXG19, FKSV07, GN14, HJ08b, Hen94, KK08, Lam05, LMN10, MM04, Mar18, MR93, SZ20, ZN17]. valued [JO05, Sin95]. values [Aré09, BdPNdP14, Hoc11, IN11, LY14, Meu15, OS15, PP08]. Vandermonde [Drm15, LR04, Opf10, Sch02]. vanilla [PPN13]. Variable [Lau08, AMRT06, Lam05, LDB13a, LDB13b, Osb08, Sch05]. variable-coefficient [Lam05]. Variable-precision [Lau08]. variables [Lor99]. Variants [Zem17, PR20]. variation [BBK18, Par21, VW97]. variation-based [BBK18]. Variational [LMSY14, SHH^{+}12, BL06, Ger17, JC17, MH17, Pry14]. variationally [AR09]. variations [LZ07]. various [AK10]. Varying [IE17, HC05]. Vector [EMS13, MR17, BM05, JOT08, Tos00]. vectors [VB95]. velocity [CMM95, HJ16]. velocity-flux [HJ16]. velocity-vorticity-pressure [CMM95]. Verified [FH13]. version [EJS03, Erh95]. versions [LS05]. versus [Neu19]. vertex [KLRW19]. vertex-based [KLRW19]. very [ABH17, MM99]. via [AMVW17, BY98, BC05, BSS09, CHR02, FP01, LT94, MR17, SAA09, ST08, Yam06a]. vibration [FJMS06, Vos03]. vibrations [SV10b]. view [MF06]. virtual [YHL^{+}20]. viscoelastic [HLX14, NB08a]. volatility [HHP19]. Volterra [CCDR08, CDN09, MG14, PGH11, YG08]. volume [BJR^{+}09, GOR^{+}08, LO13, LMR15, LKF08, RS14, WLR02]. vorticity [CMM95]. Vries [Kle08]. vs [LRL08].

wall [Kan15, Kan19a, Kan19b]. water [DHI03]. Wave [BL97, ÁNCQ06, HL20, LMW20, MACI06, PK08]. Wave-ray [BL97]. waveform [Gan04, GKM16]. wavelet [BP03, Che99, ER01, RS21a, RS21b]. wavelets [ČF18, HNS04, HR06]. Weak [ADW15]. weakly [BOS08, BGS10]. Weber [LP17]. Weierstrass [PQRT06]. weight [DFV09, OSLS20, RS12]. Weighted [BK06, MJM16, ZX17, CNO08, FSS97, Lén06, Lén10, Lub06, PQRT06, Pri06, RÁRP06, SHZF08, SV10a, ZC19]. weighted-GCV [CNO08]. weights [ACR99, Dam02, DRST16, Mas06, Mas02, MMN18]. well [Peñ04]. WENO [GGCT09]. where [Wat06a]. while [LMSY14]. whiteness [LPS20]. Whitney [GH99]. window [MF06]. without [BD04, Ger17, MPV13, MPV15]. words [KS18]. work [Koe99]. works [ST02]. worst [LT05, TLF07]. worst-case [LT05, TLF07].

Xu [BCDV06].

years [Ano20a, Not16]. Yuen [Zei99].

Zeros [EV93, JO14, JN10, CR15, COSV93, DJJ11, GAM05, GM18, JO17, Ped07, Sim06, VCL08]. zonal [GRS00].

References

Ahmad:2021:PAM

[AAS21] Sk. Safique Ahmad, Istkhar Ali, and Ivan Slapničar. Perturbation analysis of matrices over a quaternion division algebra. Electronic Transactions on Numerical Analysis (ETNA), 54:128–149, 2021. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                                       http://etna.mcs.kent.
                                                       edu/vol.54.2021/pp128-
                                                       149.dir/pp128-149.pdf;
                                                        http://etna.mcs.kent.
                                                       edu/volumes/2021-2030/
                                                       vol54/abstract.php?vol=
                                                       54&pages=128-149.

Archid:2020:BJL

[ABA20] Atika Archid, Abdeslem Hafid Bentbib, and Said Agoujil. A block J-Lanczos method for Hamiltonian matrices. Electronic Transactions on Numerical Analysis (ETNA), 52:26–42, 2020. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.52.2020/pp26-
42.dir/pp26-42.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol52/abstract.php?vol=
52&pages=26-42.

Akin-Bohner:2007:OCC

[ABBS07] Elvan Akin-Bohner, Martin Bohner, and Samir H. Saker. Oscillation criteria for a certain class of second order Emden–Fowler dynamic equations. Electronic Transactions on Numerical Analysis (ETNA), 27:1–12, 2007. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
27.2007/pp1-12.dir/pp1-
12.pdf.

August:2017:AFV

[ABH17] Moritz August, Mari Carmen Bañuls, and Thomas Huckle. On the approximation of functionals of very large Hermitian matrices represented as matrix product operators. Electronic Transactions on Numerical Analysis (ETNA), 46:215–232, 2017. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.46.2017/pp215-
232.dir/pp215-232.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol46/abstract.php?vol=
46&pages=215-232.

Auvinen:2009:LSK

[ABHK09] H. Auvinen, J. M. Bardsley, H. Haario, and T. Kauranne. Largescale Kalman filtering using the limited memory BFGS method. Electronic Transactions on Numerical Analysis (ETNA), 35: 217–233, 2009. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
35.2009/pp217-233.dir/
pp217-233.pdf.

Agoujil:2014:MRN

[ABJE15] Said Agoujil, Abdeslem H. Bentbib, Khalide Jbilou, and Mostafa El Sadek. A minimal residual norm method for large-scale Sylvester matrix equations. Electronic Transactions on Numerical Analysis (ETNA), 43:45–59, 2014– 2015. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.43.2014-2015/
pp45-59.dir/pp45-59.pdf;
 http://etna.mcs.kent.
edu/vol.43.2014/pp45-
59.dir/pp45-59.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol43/abstract.php?vol=
43&pages=45-59.

Ammar:1993:MAN

[ABM93] Gregory Ammar, Peter Benner, and Volker Mehrmann. A multishift algorithm for the numerical solution of algebraic Riccati equations. Electronic Transactions on Numerical Analysis (ETNA), 1:33–48, 1993. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.1.1993/pp33-48.
dir/pp33-48.pdf.

Atkinson:2006:SFS

[AC06] Kendall Atkinson and David Chien. A study of the fast solution of the occluded radiosity equation. Electronic Transactions on Numerical Analysis (ETNA), 23: 219–250, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
23.2006/pp219-250.dir/
pp219-250.pdf.

Acosta:2006:OLS

[Aco06] Eva Acosta. Orthogonal least squares solutions for linear operators. Electronic Transactions on Numerical Analysis (ETNA), 24:1–6, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        24.2006/pp1-6.dir/pp1-
                                        6.pdf.

Ammar:1999:CGK

[ACR99] G. S. Ammar, D. Calvetti, and L. Reichel. Computation of Gauss–Kronrod quadrature rules with nonpositive weights. Electronic Transactions on Numerical Analysis (ETNA), 9:26–38, 1999. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.9.1999/pp26-38.
                                        dir/pp26-38.pdf.

Atkinson:2000:NAR

[ACS00] Kendall Atkinson, David Da-Kwun Chien, and Jaehoon Seol. Numerical analysis of the radiosity equation using the collocation method. Electronic Transactions on Numerical Analysis (ETNA), 11: 94–120, 2000. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.

11.2000/pp94-120.dir/
pp94-120.pdf.

Armentano:2004:ALB

[AD04] María G. Armentano and Ricardo G. Durán. Asymptotic lower bounds for eigenvalues by nonconforming finite element methods. Electronic Transactions on Numerical Analysis (ETNA), 17:93–101, 2004. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
17.2004/pp93-101.dir/
pp93-101.pdf.

Arioli:2008:UFO

[AD09] M. Arioli and I. S. Duff.

Using FGMRES to obtain backward stability in mixed precision. Electronic Transactions on Numerical Analysis (ETNA), 33: 31–44, 2008–2009. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
33.2008-2009/pp31-44.
dir/pp31-44.pdf.

Adewole:2017:AOC

[Ade17] Matthew O. Adewole. Almost optimal convergence of FEM-FDM for a linear parabolic interface problem. Electronic Transactions on Numerical Analysis (ETNA), 46:337–358, 2017. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.46.2017/pp337-
                                        358.dir/pp337-358.pdf;
                                         http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol46/abstract.php?vol=
                                        46&pages=337-358.

Asvadurov:2007:OGA

[ADM07] S. Asvadurov, V. Druskin, and S. Moskow. Optimal grids for anisotropic problems. Electronic Transactions on Numerical Analysis (ETNA), 26:55–81, 2007. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.26.2007/pp55-
                                        81.dir/pp55-81.pdf.

Anton:2014:WSS

[ADW15] Cristina Anton, Jian Deng, and Yau Shu Wong. Weak symplectic schemes for stochastic Hamiltonian equations. Electronic Transactions on Numerical Analysis (ETNA), 43:1–20, 2014– 2015. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.43.2014/pp1-20.
                                        dir/pp1-20.pdf;   http:
                                        //etna.mcs.kent.edu/volumes/
                                        2011-2020/vol43/abstract.
                                        php?vol=43&pages=1-20.

Afanasjew:2007:GSD

[AEEG08] M. Afanasjew, M. Eiermann, O. G. Ernst, and S. Güttel. A generalization of the steepest descent method for matrix functions. Electronic Transactions on Numerical Analysis (ETNA), 28: 206–222, 2007–2008. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
28.2007-2008/pp206-222.
dir/pp206-222.pdf. Spe-

cial volume for Gene Golub.

Alefeld:2004:ETK

[AFHM04] Götz Alefeld, Andreas Frommer, Gerhard Heindl, and Jan Mayer. On the existence theorems of Kantorovich, Miranda and Borsuk. Electronic Transactions on Numerical Analysis (ETNA), 17:102–111, 2004. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
17.2004/pp102-111.dir/
pp102-111.pdf.

Amstutz:2019:TDN

[AG19] Samuel Amstutz and Peter Gangl. Topological derivative for the nonlinear magnetostatic problem. Electronic Transactions on Numerical Analysis (ETNA), 51:169–218, 2019. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.51.2019/pp169-
                                        218.dir/pp169-218.pdf;
                                         http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol51/abstract.php?vol=
                                        51&pages=169-218.

Ascher:2003:MMD

[AH03] U. M. Ascher and E. Haber. A multigrid method for distributed parameter estimation problems. Electronic Transactions on Numerical Analysis (ETNA), 15:1–17, 2003. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        15.2003/pp1-17.dir/pp1-
                                        17.pdf.

Atkinson:2004:NSS

[AH04] Kendall Atkinson and Weimin Han. On the numerical solution of some semilinear elliptic problems. Electronic Transactions on Numerical Analysis (ETNA), 17:206–217, 2004. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        17.2004/pp206-217.dir/
                                        pp206-217.pdf.

Atkinson:2010:SME

[AH10] Kendall Atkinson and Olaf Hansen. A spectral method for the eigenvalue problem for elliptic equations. Electronic Transactions on Numerical Analysis (ETNA), 37:386–412, 2010. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic).

Atkinson:2012:CDM

[AH12] Kendall Atkinson and Olaf Hansen. Creating domain mappings. Electronic Transactions on Numerical Analysis (ETNA), 39: 202–230, 2012. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
39.2012/pp202-230.dir/.

Ahmad:2019:PAP

[Ahm19] Sk. Safique Ahmad. Perturbation analysis for palindromic and anti-palindromic nonlinear eigenvalue problems. Electronic Transactions on Numerical Analysis (ETNA), 51:151–168, 2019. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.51.2019/pp151-
168.dir/pp151-168.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol51/abstract.php?vol=
51&pages=151-168.

Addam:2017:BHP

[AHS17] M. Addam, M. Heyouni, and H. Sadok. The block Hessenberg process for matrix equations. Electronic Transactions on Numerical Analysis (ETNA), 46:460–473, 2017. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.46.2017/pp460-
473.dir/pp460-473.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol46/abstract.php?vol=
46&pages=460-473.

Atakishiyev:2006:DPO

[AK06] N. M. Atakishiyev and U. Klimyk. Duality of q-polynomials, orthogonal on countable sets of points. Electronic Transactions on Numerical Analysis (ETNA), 24:108–180, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
24.2006/pp108-180.dir/
pp108-180.pdf.

Axelsson:2009:CNA

[AK10] Owe Axelsson and János Karátson. Condition number analysis for various forms of block matrix preconditioners. Electronic Transactions on Numerical Analysis (ETNA), 36: 168–194, 2009–2010. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
36.2009-2010/pp168-194.
dir/pp168-194.pdf;   http:
//etna.mcs.kent.edu/volumes/
2001-2010/vol36/abstract.
php?vol=36&pages=168-
194.

Arioli:2007:SCM

[AL08] M. Arioli and D. Loghin.

Stopping criteria for mixed finite element problems. Electronic Transactions on Numerical Analysis (ETNA), 29:178–192, 2007–2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
29.2007-2008/pp178-192.
dir/pp178-192.pdf.

Angulo:2008:DSG

[AL09] Wilfredo Angulo and Hilda López. On the discrete solution of the generalized Stokes problem in one time-step for two phase flow. Electronic Transactions on Numerical Analysis (ETNA), 34: 170–186, 2008–2009. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
34.2008-2009/pp170-186.
dir/pp170-186.pdf.

Arioli:2006:NPA

[AM06] M. Arioli and G. Manzini. A network programming approach in solving Darcy’s equations by mixed finiteelement methods. Electronic Transactions on Numerical Analysis (ETNA), 22:41–70, 2006. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.22.2006/pp41-
70.dir/pp41-70.pdf.

Ahmad:2011:PAC

[AM11] Sk. Safique Ahmad and Volker Mehrmann. Perturbation analysis for complex symmetric, skew symmetric, even and odd matrix polynomials. Electronic Transactions on Numerical Analysis (ETNA), 38: 275–302, 2011. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
38.2011/pp275-302.dir/
pp275-302.pdf.

Arioli:2006:DVM

[AMRT06] M. Arioli, J. Maryška, M. Rozložník, and M. Tůma. Dual variable methods for mixed-hybrid finite element approximation of the potential fluid flow problem in porous media. Electronic Transactions on Numerical Analysis (ETNA), 22:17–40, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.22.2006/pp17-
40.dir/pp17-40.pdf.

Aurentz:2015:FSU

[AMVW15] Jared L. Aurentz, Thomas Mach, Raf Vandebril, and David S. Watkins. Fast and stable unitary QR algorithm. Electronic Transactions on Numerical Analysis (ETNA), 44:327–341, 2015. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.44.2015/pp327-
341.dir/pp327-341.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol44/abstract.php?vol=
44&pages=327-341.

Aurentz:2017:CES

[AMVW17] Jared L. Aurentz, Thomas Mach, Raf Vandebril, and David S. Watkins. Computing the eigenvalues of symmetric tridiagonal matrices via a Cayley transformation. Electronic Transactions on Numerical Analysis (ETNA), 46:447–459, 2017. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.46.2017/pp447-
459.dir/pp447-459.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol46/abstract.php?vol=
46&pages=447-459.

Alvarez-Nodarse:2007:FHT

[ÁNACS07] R. Álvarez-Nodarse, N. M. Atakishiyev, and R. S. Costas-Santos. Factorization of the hypergeometrictype difference equation on the uniform lattice. Electronic Transactions on Numerical Analysis (ETNA), 27:34–50, 2007. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.27.2007/pp34-
                                        50.dir/pp34-50.pdf.

Alvarez-Nodarse:2006:RRR

[ÁNCQ06] R. Álvarez-Nodarse, J. L. Cardoso, and N. R. Quintero. On recurrence relations for radial wave functions for the N-th dimensional oscillators and hydrogenlike atoms: analytical and numerical study. Electronic Transactions on Numerical Analysis (ETNA), 24:7–23, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        24.2006/pp7-23.dir/pp7-
                                        23.pdf.

Anonymous:2009:TCAa

[Ano09a] Anonymous. Table of contents and abstracts. Electronic Transactions on Numerical Analysis (ETNA), 35:??, 2009. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
35.2009/contents.pdf.

Anonymous:2009:TPa

[Ano09b] Anonymous. Title page.

Electronic Transactions on Numerical Analysis (ETNA), 35:??, 2009. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
35.2009/title_page.pdf.

Anonymous:2009:P

[Ano10a] Anonymous. Preface. Electronic Transactions on Numerical Analysis (ETNA), 36:??, 2009–2010. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
36.2009-2010/preface.
pdf.

Anonymous:2009:TCAb

[Ano10b] Anonymous. Table of contents and abstracts. Electronic Transactions on Numerical Analysis (ETNA), 36:??, 2009–2010. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
36.2009-2010/contents.
pdf.

Anonymous:2009:TPb

[Ano10c] Anonymous. Title page. Electronic Transactions on Numerical Analysis (ETNA), 36:??, 2009–2010. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.36.2009-2010/
                                        title_page.pdf.

Anonymous:2013:TCA

[Ano13a] Anonymous. Table of contents and abstracts. Electronic Transactions on Numerical Analysis (ETNA), 40:??, 2013. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic).

Anonymous:2013:TP

[Ano13b] Anonymous. Title page. Electronic Transactions on Numerical Analysis (ETNA), 40:??, 2013. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic).

Anonymous:2014:Pa

[Ano14a] Anonymous. Preface. Electronic Transactions on Numerical Analysis (ETNA), 42:??, 2014. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        42.2014/preface.pdf.

Anonymous:2014:Pb

[Ano14b] Anonymous. Preface. Electronic Transactions on Numerical Analysis (ETNA), 42:??, 2014. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
42.2014/preface.pdf.

Anonymous:2014:TCAa

[Ano14c] Anonymous. Table of contents and abstracts. Electronic Transactions on Numerical Analysis (ETNA), 41:??, 2014. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
41.2014/contents.pdf.

Anonymous:2014:TPa

[Ano14d] Anonymous. Title page.

Electronic Transactions on Numerical Analysis (ETNA), 41:??, 2014. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
41.2014/title_page.pdf.

Anonymous:2014:P

[Ano15a] Anonymous. Preface. Electronic Transactions on Numerical Analysis (ETNA), 43:??, 2014–2015. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
43.2014-2015/preface.
pdf.

Anonymous:2014:TCAb

[Ano15b] Anonymous. Table of contents and abstracts. Electronic Transactions on Numerical Analysis (ETNA), 43:??, 2014–2015. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        43.2014-2015/contents.
                                        pdf.

Anonymous:2014:TPb

[Ano15c] Anonymous. Title page. Electronic Transactions on Numerical Analysis (ETNA), 43:??, 2014–2015. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.43.2014-2015/
                                        title_page.pdf.

Anonymous:2015:TCA

[Ano15d] Anonymous. Table of contents and abstracts. Electronic Transactions on Numerical Analysis (ETNA), 44:??, 2015. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        44.2015/contents.pdf.

Anonymous:2015:TP

[Ano15e] Anonymous. Title page. Electronic Transactions on Numerical Analysis (ETNA), 44:??, 2015. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
44.2015/title_page.pdf.

Anonymous:2016:TCA

[Ano16] Anonymous. Table of contents and abstracts. Electronic Transactions on Numerical Analysis (ETNA), 45:??, 2016. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
45.2016/contents.pdf.

Anonymous:2017:P

[Ano17a] Anonymous. Preface. Electronic Transactions on Numerical Analysis (ETNA), 47:??, 2017. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
47.2017/preface.pdf.

Anonymous:2017:TCAa

[Ano17b] Anonymous. Table of contents and abstracts. Electronic Transactions on Numerical Analysis (ETNA), 46:??, 2017. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
46.2017/contents.pdf.

Anonymous:2017:TCAb

[Ano17c] Anonymous. Table of contents and abstracts. Electronic Transactions on Numerical Analysis (ETNA), 47:??, 2017. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        47.2017/contents.pdf.

Anonymous:2017:TPa

[Ano17d] Anonymous. Title page. Electronic Transactions on Numerical Analysis (ETNA), 46:??, 2017. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        46.2017/title_page.pdf.

Anonymous:2017:TPb

[Ano17e] Anonymous. Title page. Electronic Transactions on Numerical Analysis (ETNA), 47:??, 2017. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        47.2017/title_page.pdf.

Anonymous:2018:Pa

[Ano18a] Anonymous. Preface. Electronic Transactions on Numerical Analysis (ETNA), 49:??, 2018. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
49.2018/preface.pdf.

Anonymous:2018:Pb

[Ano18b] Anonymous. Preface. Electronic Transactions on Numerical Analysis (ETNA), 50:??, 2018. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
50.2018/preface.pdf.

Anonymous:2018:TCA

[Ano18c] Anonymous. Table of contents and abstracts. Electronic Transactions on Numerical Analysis (ETNA), 50:??, 2018. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
50.2018/contents.pdf.

Anonymous:2018:TP

[Ano18d] Anonymous. Title page.

Electronic Transactions on Numerical Analysis (ETNA), 50:??, 2018. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
50.2018/title_page.pdf.

Anonymous:2019:TCA

[Ano19a] Anonymous. Table of contents and abstracts. Electronic Transactions on Numerical Analysis (ETNA), 51:??, 2019. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        51.2019/contents.pdf.

Anonymous:2019:TP

[Ano19b] Anonymous. Title page. Electronic Transactions on Numerical Analysis (ETNA), 51:??, 2019. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        51.2019/title_page.pdf.

Anonymous:2020:YEP

[Ano20a] Anonymous. 25 years of ETNA: Past, present, and future. Electronic Transactions on Numerical Analysis (ETNA), 53:??, 2020. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        53.2020/ETNAhistory.pdf.

Anonymous:2020:P

[Ano20b] Anonymous. Preface. Electronic Transactions on Numerical Analysis (ETNA), 53:??, 2020. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        53.2020/preface.pdf.

Anonymous:2020:TCA

[Ano20c] Anonymous. Table of contents and abstracts. Electronic Transactions on Numerical Analysis (ETNA), 53:??, 2020. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
53.2020/contents.pdf.

Anonymous:2020:TP

[Ano20d] Anonymous. Title page.

Electronic Transactions on Numerical Analysis (ETNA), 53:??, 2020. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
53.2020/title_page.pdf.

Alber:2010:CIB

[AO10] David M. Alber and Luke N.

Olson. Coarsening invariance and bucket-sorted independent sets for algebraic multigrid. Electronic Transactions on Numerical Analysis (ETNA), 37:367–385, 2010. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic).

Anitescu:2004:ILE

[APL04] Mihai Anitescu, Faranak Pahlevani, and William J. Layton. Implicit for local effects and explicit for nonlocal effects is unconditionally stable. Electronic Transactions on Numerical Analysis (ETNA), 18:174–187, 2004. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

                                        /etna.mcs.kent.edu/vol.
                                        18.2004/pp174-187.dir/
                                        pp174-187.pdf.

Acevedo:2007:APF

[AR07] Ramiro Acevedo and Rodolfo Rodríguez. Analysis of the A, V – A – ψ potential formulation for the eddy current problem in a bounded domain. Electronic Transactions on Numerical Analysis (ETNA), 26:270–284, 2007. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        26.2007/pp270-284.dir/
                                        pp270-284.pdf.

Alonso:2009:SAV

[AR09] Ana Alonso and Anahí Dello Russo. Spectral approximation of variationally formulated eigenvalue problems on curved domains. Electronic Transactions on Numerical Analysis (ETNA), 35:69–87, 2009. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.35.2009/pp69-
                                        87.dir/pp69-87.pdf.

Ayaz:2014:NSR

[AR14] Ulaş Ayaz and Holger Rauhut. Nonuniform sparse recovery with subgaussian matrices. Electronic Transactions on Numerical Analysis (ETNA), 41:167–178, 2014. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.41.2014/pp167-
178.dir/pp167-178.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol41/abstract.php?vol=
41&pages=167-178.

Arevalo:2008:NNC

[Aré09] Carmen Arévalo. A note on numerically consistent initial values for high index differential-algebraic equations. Electronic Transactions on Numerical Analysis (ETNA), 34: 14–19, 2008–2009. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
34.2008-2009/pp14-19.
dir/pp14-19.pdf.

Arvesu:2006:QAS

[Arv06]  Jorge Arvesú.  Quan-

tum algebras su_{q}(2) and su_{q}(1, 1) associated with certain q-Hahn polynomials: a revisited approach. Electronic Transactions on Numerical Analysis (ETNA), 24:24–44, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.

                                        edu/vol.24.2006/pp24-
                                        44.dir/pp24-44.pdf.

Appleby:2005:LAR

[AS05a] Glenn Appleby and Dennis C. Smolarski. A linear acceleration row action method for projecting onto subspaces. Electronic Transactions on Numerical Analysis (ETNA), 20:253–275, 2005. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        20.2005/pp253-275.dir/
                                        pp253-275.pdf.

Atkinson:2005:QS

[AS05b] Kendall Atkinson and Alvise Sommariva. Quadrature over the sphere. Electronic Transactions on Numerical Analysis (ETNA), 20: 104–118, 2005. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        20.2005/pp104-118.dir/
                                        pp104-118.pdf.

AlSayedAli:2009:AIS

[AS09] Mouhamad Al Sayed Ali and Miloud Sadkane. Acceleration of implicit schemes for large systems of nonlinear ODEs. Electronic Transactions on Numerical Analysis (ETNA), 35: 104–117, 2009. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
35.2009/pp104-117.dir/
pp104-117.pdf.

AlSayedAli:2012:IPS

[AS12] Mouhamad Al Sayed Ali and Miloud Sadkane. Improved predictor schemes for large systems of linear ODEs. Electronic Transactions on Numerical Analysis (ETNA), 39:253–270, 2012. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
39.2012/pp253-270.dir/.

Andreev:2014:CST

[AS14] Roman Andreev and Julia Schweitzer. Conditional space-time stability of collocation Runge–Kutta for parabolic evolution equations. Electronic Transactions on Numerical Analysis (ETNA), 41:62–80, 2014. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.41.2014/pp62-
80.dir/pp62-80.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol41/abstract.php?vol=
41&pages=62-80.

Asharabi:2017:AEB

[Ash17] Rashad M. Asharabi. Approximating eigenvalues of boundary value problems by using the Hermite–Gauss sampling method. Electronic Transactions on Numerical Analysis (ETNA), 46:359–374, 2017. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.46.2017/pp359-
374.dir/pp359-374.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol46/abstract.php?vol=
46&pages=359-374.

Atkinson:2004:QSI

[Atk04] Kendall Atkinson. Quadrature of singular integrands over surfaces. Electronic Transactions on Numerical Analysis (ETNA), 17: 133–150, 2004. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
17.2004/pp133-150.dir/
pp133-150.pdf.

Aurentz:2015:FAS

[Aur15] Jared L. Aurentz. Fast algorithms for spectral differentiation matrices. Electronic Transactions on Numerical Analysis (ETNA), 44:281–288, 2015. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.44.2015/pp281-

288.dir/pp281-288.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol44/abstract.php?vol=
44&pages=281-288.

Auzinger:2018:MRA

[Auz18] Günter Auzinger. Model reduction in atmospheric tomography by optimal grouping of turbulent layers. Electronic Transactions on Numerical Analysis (ETNA), 48:286–309, 2018. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.48.2018/pp286-
309.dir/pp286-309.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol48/abstract.php?vol=
48&pages=286-309.

Aposporidis:2014:MPN

[AVV14] Alexis Aposporidis, Panayot S.

Vassilevski, and Alessandro Veneziani. Multigrid preconditioning of the non-regularized augmented Bingham fluid problem. Electronic Transactions on Numerical Analysis (ETNA), 41:42–61, 2014. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu//vol.41.2014/pp42-
61.dir/pp42-61.pdf.

Bardsley:2005:NCT

[Bar05] Johnathan M. Bardsley. A nonnegatively constrained trust region algorithm for the restoration of images with an unknown blur. Electronic Transactions on Numerical Analysis (ETNA), 20:139–153, 2005. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
20.2005/pp139-153.dir/
pp139-153.pdf.

Baratchart:2006:RUB

[Bar06] L. Baratchart. A remark on uniqueness of best rational approximants of degree 1 in L^{2} of the circle. Electronic Transactions on Numerical Analysis (ETNA), 25:54–66, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.25.2006/pp54-
66.dir/pp54-66.pdf.

Bardsley:2011:ANC

[Bar11] Johnathan M. Bardsley. Applications of a nonnegatively constrained iterative method with statistically based stopping rules to CT, PET, and SPECT imaging. Electronic Transactions on Numerical Analysis (ETNA), 38:34–43, 2011. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic).

Barlow:2014:BGS

[Bar15] Jesse L. Barlow. Block Gram–Schmidt downdating. Electronic Transactions on Numerical Analysis (ETNA), 43:163–187, 2014– 2015. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.43.2014/pp163-
187.dir/pp163-187.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol43/abstract.php?vol=
43&pages=163-187.

Baker:2000:CMS

[BB00] Christopher T. H. Baker and Evelyn Buckwar. Continuous Θ-methods for the stochastic pantograph equation. Electronic Transactions on Numerical Analysis (ETNA), 11:131–151, 2000. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
11.2000/pp131-151.dir/
pp131-151.pdf.

BenBelgacem:2001:SNF

[BB01] Faker Ben Belgacem and Susanne C. Brenner. Some nonstandard finite element estimates with applications to 3D Poisson and Signorini problems. Electronic Transactions on Numerical Analysis (ETNA), 12: 134–148, 2001. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
12.2001/pp134-148.dir/
pp134-148.pdf.

Baur:2008:CGB

[BB08] Ulrike Baur and Peter Benner. Cross-Gramian based model reduction for datasparse systems. Electronic Transactions on Numerical Analysis (ETNA), 31: 256–270, 2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
31.2008/pp256-270.dir/
pp256-270.pdf.

Benning:2011:EEG

[BB11] Martin Benning and Martin Burger. Error estimates for general fidelities. Electronic Transactions on Numerical Analysis (ETNA), 38:44–68, 2011. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic).

Benner:2014:RIM

[BB14] Peter Benner and Tobias Breiten. Rational interpolation methods for symmetric Sylvester equations. Electronic Transactions on Numerical Analysis (ETNA), 42:147–164, 2014. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.42.2014/pp147-
164.dir/pp147-164.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol42/abstract.php?vol=
42&pages=147-164.

Bao:2008:DAK

[BBD^{+}08] Yujuan Bao, İlker N.

Bozkurt, Tugrul Dayar, Xiaobai Sun, and Kishor S. Trivedi. Decompositional analysis of Kronecker structured Markov chains. Electronic Transactions on Numerical Analysis (ETNA), 31:271–294, 2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
31.2008/pp271-294.dir/
pp271-294.pdf.

Barrenechea:2018:NAT

[BBD18] Gabriel R. Barrenechea, Michał Bosy, and Victorita Dolean. Numerical assessment of two-level domain decomposition preconditioners for incompressible Stokes and elasticity equations. Electronic Transactions on Numerical Analysis (ETNA), 49:41–63, 2018. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.

                                        edu/vol.49.2018/pp41-
                                        63.dir/pp41-63.pdf;
                                        http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol49/abstract.php?vol=
                                        49&pages=41-63.

Benner:2000:ACL

[BBF^{+}00a] Peter Benner, Ralph Byers, Heike Fassbender, Volker Mehrmann, and David Watkins. Addendum to Cholesky-like factorizations of skew-symmetric matrices. Electronic Transactions on Numerical Analysis (ETNA), 11:93A, 2000. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.11.2000/pp93A.
                                        dir/pp93A.pdf.

Benner:2000:CLF

[BBF^{+}00b] Peter Benner, Ralph Byers, Heike Fassbender, Volker Mehrmann, and David Watkins. Choleskylike factorizations of skewsymmetric matrices. Electronic Transactions on Numerical Analysis (ETNA), 11:85–93, 2000. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.11.2000/pp85-
                                        93.dir/pp85-93.pdf.

Bentbib:2018:CGM

[BBK18] Abdeslem Hafid Bentbib, Abderrahman Bouhamidi, and Karim Kreit. A conditional gradient method for primal-dual total variationbased image denoising. Electronic Transactions on Numerical Analysis (ETNA), 48:310–328, 2018. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.48.2018/pp310-
328.dir/pp310-328.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol48/abstract.php?vol=
48&pages=310-328.

Brandt:2015:ADA

[BBKL15] Achi Brandt, James Brannick, Karsten Kahl, and Irene Livshits. Algebraic distance for anisotropic diffusion problems: multilevel results. Electronic Transactions on Numerical Analysis (ETNA), 44:472–496, 2015. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.44.2015/pp472-
496.dir/pp472-496.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol44/abstract.php?vol=
44&pages=472-496.

Bergamaschi:2006:QNP

[BBMP06] L. Bergamaschi, R. Bru, A. Martínez, and M. Putti. Quasi-Newton preconditioners for the inexact Newton method. Electronic Transactions on Numerical Analysis (ETNA), 23:76–87, 2006. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.23.2006/pp76-
87.dir/pp76-87.pdf.

Badahmane:2019:PGK

[BBS19] A. Badahmane, A. H. Bentbib, and H. Sadok. Preconditioned global Krylov subspace methods for solving saddle point problems with multiple right-hand sides. Electronic Transactions on Numerical Analysis (ETNA), 51:495–511, 2019. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.51.2019/pp495-
511.dir/pp495-511.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol51/abstract.php?vol=
51&pages=495-511.

Banicescu:2005:ASN

[BC05] Ioana Banicescu and Ricolindo L. Cariño. Addressing the stochastic nature of scientific computations via dynamic loop scheduling. Electronic Transactions on Numerical Analysis (ETNA), 21:66–80, 2005. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.21.2005/pp66-
80.dir/pp66-80.pdf.

Batista:2008:MSN

[BC09] E. D. Batista and J. E.

Castillo. Mimetic schemes on non-uniform structured meshes. Electronic Transactions on Numerical Analysis (ETNA), 34:152–162, 2008– 2009. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.34.2008-2009/
pp152-162.dir/pp152-162.
pdf.

Berriochoa:2015:EFH

[BCDI15] Elías Berriochoa, Alicia Cachafeiro, Jaime Díaz, and Jesús Illán. Explicit formulas for Hermite-type interpolation on the circle and applications. Electronic Transactions on Numerical Analysis (ETNA), 44:140–152, 2015. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.44.2015/pp140-
152.dir/pp140-152.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol44/abstract.php?vol=
44&pages=140-152.

Benko:2018:PMA

[BCDO18] David Benko, Dan Coroian, Peter Dragnev, and Ramón Orive. Probability, minimax approximation, and Nash-equilibrium. Estimating the parameter of a biased coin. Electronic Transactions on Numerical Analysis (ETNA), 50:109–128, 2018. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.50.2018/pp109-
                                        128.dir/pp109-128.pdf;
                                         http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol50/abstract.php?vol=
                                        50&pages=109-128.

Bos:2006:BIX

[BCDV06] Len Bos, Marco Caliari, Stefano De Marchi, and Marco Vianello. Bivariate interpolation at Xu points: results, extensions and applications. Electronic Transactions on Numerical Analysis (ETNA), 25:1–16, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        25.2006/pp1-16.dir/pp1-
                                        16.pdf.

Berriochoa:2008:NQR

[BCGAM08] Elías Berriochoa, Alicia Cachafeiro, José M. García-Amor, and Francisco Marcellán. New quadrature rules for Bernstein measures on the interval [−1, 1]. Electronic Transactions on Numerical Analysis (ETNA), 30:278–290, 2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
30.2008/pp278-290.dir/
pp278-290.pdf.

Baglama:1998:FLP

[BCR98] J. Baglama, D. Calvetti, and L. Reichel. Fast Leja points. Electronic Transactions on Numerical Analysis (ETNA), 7:124–140, 1998. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
7.1998/pp124-140.dir/
pp124-140.pdf.  Large

scale eigenvalue problems (Argonne, IL, 1997).

Bueno:2004:SSD

[BD04] M. Isabel Bueno and Froilán M. Dopico. Stability and sensivity of Darboux transformation without parameter. Electronic Transactions on Numerical Analysis (ETNA), 18: 101–136, 2004. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
18.2004/pp101-136.dir/
pp101-136.pdf.

Bianchi:2017:GIT

[BD17] Davide Bianchi and Marco Donatelli. On generalized iterated Tikhonov regularization with operatordependent seminorms. Electronic Transactions on Numerical Analysis (ETNA), 47:73–99, 2017. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.47.2017/pp73-
                                        99.dir/pp73-99.pdf;
                                        http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol47/abstract.php?vol=
                                        47&pages=73-99.

Buccini:2020:MFB

[BD20] Alessandro Buccini and Marco Donatelli. A multigrid frame based method for image deblurring. Electronic Transactions on Numerical Analysis (ETNA), 53:283–312, 2020. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.53.2020/pp283-
                                        312.dir/pp283-312.pdf;
                                         http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol53/abstract.php?vol=
                                        53&pages=283-312.

Benko:2006:SEM

[BDD06] D. Benko, S. B. Damelin, and P. D. Dragnev. On the support of the equilibrium measure for arcs of the unit circle and for real intervals. Electronic Transactions on Numerical Analysis (ETNA), 25:27–40, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.25.2006/pp27-
40.dir/pp27-40.pdf.

Bini:2004:SQI

[BDG04] Dario A. Bini, Francesco Daddi, and Luca Gemignani. On the shifted QR iteration applied to companion matrices. Electronic Transactions on Numerical Analysis (ETNA), 18:137–152, 2004. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
18.2004/pp137-152.dir/
pp137-152.pdf.

Bolten:2015:ASA

[BDH15] Matthias Bolten, Marco Donatelli, and Thomas Huckle. Analysis of smoothed aggregation multigrid methods based on Toeplitz matrices. Electronic Transactions on Numerical Analysis (ETNA), 44:25–52, 2015. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.44.2015/pp25-

                                        52.dir/pp25-52.pdf;
                                        http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol44/abstract.php?vol=
                                        44&pages=25-52.

Bebiano:2014:RIF

[BdPNdP14] Natália Bebiano, João da Providência, Ana Nata, and João P. da Providência. Revisiting the inverse field of values problem. Electronic Transactions on Numerical Analysis (ETNA), 42:1–12, 2014. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        42.2014/pp1-12.dir/pp1-
                                        12.pdf;   http://etna.
                                        mcs.kent.edu/volumes/
                                        2011-2020/vol42/abstract.
                                        php?vol=42&pages=1-12.

Bini:2008:ESP

[BDR09] Dario A. Bini, Gianna M. Del Corso, and Francesco Romani. Evaluating scientific products by means of citation-based models: a first analysis and validation. Electronic Transactions on Numerical Analysis (ETNA), 33:1–16, 2008– 2009. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.33.2008-2009/
                                        pp1-16.dir/pp1-16.pdf.

Beattie:1998:HRL

[Bea98] Christopher Beattie. Harmonic Ritz and Lehmann bounds. Electronic Transactions on Numerical Analysis (ETNA), 7:18–39, 1998. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.7.1998/pp18-39.
dir/pp18-39.pdf. Large

scale eigenvalue problems (Argonne, IL, 1997).

Bachar:2014:MAA

[BEG15] Mostafa Bachar, Thierry Estebenet, and Allal Guessab. A moving asymptotes algorithm using new local convex approximation methods with explicit solutions. Electronic Transactions on Numerical Analysis (ETNA), 43:21–44, 2014– 2015. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.43.2014/pp21-
44.dir/pp21-44.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol43/abstract.php?vol=
43&pages=21-44.

Bentbib:2018:EGL

[BEJ^{+}18] A. H. Bentbib, M. El Ghomari, C. Jagels, K. Jbilou, and L. Reichel. The extended global Lanczos method for matrix function approximation. Electronic Transactions on Numerical Analysis (ETNA), 50:144–163, 2018. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.50.2018/pp144-
                                        163.dir/pp144-163.pdf;
                                         http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol50/abstract.php?vol=
                                        50&pages=144-163.

Bru:1994:CIP

[BEN94] Rafael Bru, L. Elsner, and M. Neumann. Convergence of infinite products of matrices and inner-outer iteration schemes. Electronic Transactions on Numerical Analysis (ETNA), 2: 183–193, 1994. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        2.1994/pp183-193.dir/
                                        pp183-193.pdf.

Bertaccini:2004:EPS

[Ber04] Daniele Bertaccini. Efficient preconditioning for sequences of parametric complex symmetric linear systems. Electronic Transactions on Numerical Analysis (ETNA), 18:49–64, 2004. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.

edu/vol.18.2004/pp49-
64.dir/pp49-64.pdf.

Bozzo:2012:MML

[BF12] Enrico Bozzo and Dario Fasino. A multiparameter model for link analysis of citation graphs. Electronic Transactions on Numerical Analysis (ETNA), 39:464–475, 2012. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.39.2012/pp464-
475.dir/pp464-475.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol39/abstract.php?vol=
39&pages=464-475.

Bebiano:2018:ANR

[BFdP^{+}18] Natália Bebiano, Susana Furtado, João da Providência, Wei-Ru Xu, and João P. da Providência. Approximations for von Neumann and Renyi entropies of graphs using the Euler– Maclaurin formula. Electronic Transactions on Numerical Analysis (ETNA), 48:227–242, 2018. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.48.2018/pp227-
242.dir/pp227-242.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/

                                        vol48/abstract.php?vol=
                                        48&pages=227-242.

Brousek:2015:OMM

[BFH^{+}15] Jan Brousek, Pavla Franková, Milan Hanuš, Hana Kopincová, Roman Kužel, Radek Tezaur, Petr Vaněk, and Zbyněk Vastl. An overview of multilevel methods with aggressive coarsening and massive polynomial smoothing. Electronic Transactions on Numerical Analysis (ETNA), 44:401–442, 2015. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.44.2015/pp401-
                                        442.dir/pp401-442.pdf;
                                         http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol44/abstract.php?vol=
                                        44&pages=401-442.

Brannick:2010:ARB

[BFK^{+}10] J. Brannick, A. Frommer, K. Kahl, S. MacLachlan, and L. Zikatanov. Adaptive reduction-based multigrid for nearly singular and highly disordered physical systems. Electronic Transactions on Numerical Analysis (ETNA), 37:276–295, 2010. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic).

Brezinski:2012:ETP

[BFM12] Claude Brezinski, Paraskevi Fika, and Marilena Mitrouli.

Estimations of the trace of powers of positive selfadjoint operators by extrapolation of the moments. Electronic Transactions on Numerical Analysis (ETNA), 39:144–155, 2012. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
39.2012/pp144-155.dir/.

Benzi:2020:PSD

[BFM20] Michele Benzi, Paraskevi Fika, and Marilena Mitrouli. Performance and stability of direct methods for computing generalized inverses of the graph Laplacian. Electronic Transactions on Numerical Analysis (ETNA), 53:439–458, 2020. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.53.2020/pp439-
458.dir/pp439-458.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol53/abstract.php?vol=
53&pages=439-458.

Brdar:2019:NTS

[BFR19] Mirjana Brdar, Sebastian Franz, and Hans-Görg Roos. Numerical treatment of singularly perturbed fourth-order two-parameter problems. Electronic Transactions on Numerical Analysis (ETNA), 51:50–62, 2019. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.51.2019/pp50-
62.dir/pp50-62.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol51/abstract.php?vol=
51&pages=50-62.

Benner:2007:SLS

[BFS08] Peter Benner, Heike Fassbender, and Martin Stoll. Solving large-scale quadratic eigenvalue problems with Hamiltonian eigenstructure using a structurepreserving Krylov subspace method. Electronic Transactions on Numerical Analysis (ETNA), 29: 212–229, 2007–2008. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
29.2007-2008/pp212-229.
dir/pp212-229.pdf.

Bini:2004:IIA

[BGP04] Dario A. Bini, Luca Gemignani, and Victor Y. Pan. Improved initialization of the accelerated and robust QR-like polynomial rootfinding. Electronic Transactions on Numerical Analysis (ETNA), 17:195–205, 2004. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.

17.2004/pp195-205.dir/
pp195-205.pdf.

Brenner:2010:WPS

[BGS10] Susanne C. Brenner, Thirupathi Gudi, and Li-Yeng Sung. A weakly overpenalized symmetric interior penalty method for the biharmonic problem. Electronic Transactions on Numerical Analysis (ETNA), 37:214–238, 2010. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
37.2010/pp214-238.dir/
pp214-238.pdf.

Barbarino:2020:BGLb

[BGSC20a] Giovanni Barbarino, Carlo Garoni, and Stefano Serra-Capizzano. Block generalized locally Toeplitz sequences: theory and applications in the multidimensional case. Electronic Transactions on Numerical Analysis (ETNA), 53:113–216, 2020. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.53.2020/pp113-
216.dir/pp113-216.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol53/abstract.php?vol=
53&pages=113-216.

Barbarino:2020:BGLa

[BGSC20b] Giovanni Barbarino, Carlo Garoni, and Stefano Serra-Capizzano. Block generalized locally Toeplitz sequences: theory and applications in the unidimensional case. Electronic Transactions on Numerical Analysis (ETNA), 53:28–112, 2020. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.53.2020/pp28-
112.dir/pp28-112.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol53/abstract.php?vol=
53&pages=28-112.

Brenner:2021:ASP

[BGyS21] Susanne C. Brenner, José C. Garay, and Li yeng Sung. Additive Schwarz preconditioners for a localized orthogonal decomposition method. Electronic Transactions on Numerical Analysis (ETNA), 54:234–255, 2021. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.54.2021/pp234-
255.dir/pp234-255.pdf;
 http://etna.mcs.kent.
edu/volumes/2021-2030/
vol54/abstract.php?vol=
54&pages=234-255.

Bello-Hernandez:2020:IBP

[BH20] Manuel Bello-Hernández.

Incomplete beta polynomials. Electronic Transactions on Numerical Analysis (ETNA), 52:195–202, 2020. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.52.2020/pp195-
202.dir/pp195-202.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol52/abstract.php?vol=
52&pages=195-202.

Bardsley:2013:EMB

[BHN13] Johnathan M. Bardsley, Marylesa Howard, and James G. Nagy. Efficient MCMC-based image deblurring with Neumann boundary conditions. Electronic Transactions on Numerical Analysis (ETNA), 40:476–488, 2013. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu//vol.
40.2013/pp476-488.dir/
pp476-488.pdf.

Bochev:2003:TRP

[BHRT03] Pavel B. Bochev, Jonathan J.

Hu, Allen C. Robinson, and Raymond S. Tuminaro. Towards robust 3D Z-pinch simulations: discretization and fast solvers for magnetic diffusion in heterogeneous conductors. Electronic Transactions on Numerical Analysis (ETNA), 15:186–210, 2003. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        15.2003/pp186-210.dir/
                                        pp186-210.pdf.

Birken:2012:ORK

[Bir12] Philipp Birken. Optimizing Runge–Kutta smoothers for unsteady flow problems. Electronic Transactions on Numerical Analysis (ETNA), 39:298–312, 2012. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

                                        /etna.mcs.kent.edu/vol.
                                        39.2012/pp298-312.dir/.

Benzi:1999:NEP

[BJM99] Michele Benzi, Wayne Joubert, and Gabriel Mateescu. Numerical experiments with parallel orderings for ILU preconditioners. Electronic Transactions on Numerical Analysis (ETNA), 8:88–114, 1999. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

                                        /etna.mcs.kent.edu/vol.
                                        8.1999/pp88-114.dir/pp88-
                                        114.pdf.

Beckermann:2008:SVM

[BJR^{+}09] Bernhard Beckermann, Khalide Jbilou, Lothar Reichel, Yousef Saad, Miloude Sadkane, and Ahmed Salam. Special volume on matrix analysis and applications. Electronic Transactions on Numerical Analysis (ETNA), 33:vii, 2008/2009. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic).

Beckermann:2002:SCC

[BK02] Bernhard Beckermann and Arno B. J. Kuijlaars. Superlinear CG convergence for special right-hand sides. Electronic Transactions on Numerical Analysis (ETNA), 14:1–19, 2002. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
14.2002/pp1-19.dir/pp1-
19.pdf.

Balazs:2006:WAD

[BK06] Katherine Balázs and Theodore Kilgore. Weighted approximation of derivatives on the half-line. Electronic Transactions on Numerical Analysis (ETNA), 25: 121–128, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
25.2006/pp121-128.dir/
pp121-128.pdf.

Benzi:2013:CAG

[BK13] Michele Benzi and Verena Kuhlemann. Chebyshev acceleration of the GeneRank algorithm. Electronic Transactions on Numerical Analysis (ETNA), 40: 311–320, 2013. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu//vol.
40.2013/pp311-320.dir/
pp311-320.pdf.

Brumm:2016:MFL

[BK16] Bernd Brumm and Emil Kieri. A matrix-free Legendre spectral method for initial-boundary value problems. Electronic Transactions on Numerical Analysis (ETNA), 45:283–304, 2016. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.45.2016/pp283-
304.dir/pp283-304.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol45/abstract.php?vol=
45&pages=283-304.

Bolten:2018:MMC

[BKK^{+}18] Matthias Bolten, Karsten Kahl, Daniel Kressner, Francisco Macedo, and Sonja Sokolović. Multigrid methods combined with low-rank approximation for tensor-structured Markov chains. Electronic Transactions on Numerical Analysis (ETNA), 48:348–361, 2018. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.48.2018/pp348-
361.dir/pp348-361.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol48/abstract.php?vol=
48&pages=348-361.

Baum:2017:TSD

[BKO17] Ann-Kristin Baum, Michael Kolmbauer, and Günter Offner. Topological solvability and DAE-index conditions for mass flow controlled pumps in liquid flow networks. Electronic Transactions on Numerical Analysis (ETNA), 46:394–423, 2017. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.46.2017/pp394-
423.dir/pp394-423.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol46/abstract.php?vol=
46&pages=394-423.

Benner:2014:SGE

[BKS15] Peter Benner, Patrick Kürschner, and Jens Saak. Self-generating and efficient shift parameters in ADI methods for large Lyapunov and Sylvester equations. Electronic Transactions on Numerical Analysis (ETNA), 43:142–162, 2014– 2015. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.43.2014/pp142-
162.dir/pp142-162.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol43/abstract.php?vol=
43&pages=142-162.

Baumstark:2018:ACE

[BKS18] Simon Baumstark, Georgia Kokkala, and Katharina Schratz. Asymptotic consistent exponential-type integrators for Klein–Gordon– Schrödinger systems from relativistic to non-relativistic regimes. Electronic Transactions on Numerical Analysis (ETNA), 48:63–80, 2018. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.48.2018/pp63-
80.dir/pp63-80.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol48/abstract.php?vol=
48&pages=63-80.

Brandt:1997:WRM

[BL97] A. Brandt and I. Livshits. Wave-ray multigrid method for standing wave equations. Electronic Transactions on Numerical Analysis (ETNA), 6:162–181, 1997. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
6.1997/pp162-181.dir/

pp162-181.pdf. Special

issue on multilevel methods (Copper Mountain, CO, 1997).

Bochev:2006:RSD

[BL06] P. B. Bochev and R. B.

Lehoucq. Regularization and stabilization of discrete saddle-point variational problems. Electronic Transactions on Numerical Analysis (ETNA), 22: 97–113, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
22.2006/pp97-113.dir/
pp97-113.pdf.

Bos:2008:CAF

[BL08] L. P. Bos and N. Levenberg. On the calculation of approximate Fekete points: the univariate case. Electronic Transactions on Numerical Analysis (ETNA), 30:377–397, 2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
30.2008/pp377-397.dir/
pp377-397.pdf.

Baratchart:2006:ISP

[BLM06] L. Baratchart, J. Leblond, and J.-P. Marmorat. Inverse source problem in a 3D ball from best meromorphic approximation on 2D slices. Electronic Transactions on Numerical Analysis (ETNA), 25:41–53, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.25.2006/pp41-
53.dir/pp41-53.pdf.

Bisseling:2005:CBP

[BM05] Rob H. Bisseling and Wouter Meesen. Communication balancing in parallel sparse matrix-vector multiplication. Electronic Transactions on Numerical Analysis (ETNA), 21:47–65, 2005. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.21.2005/pp47-
65.dir/pp47-65.pdf.

Baratchart:2006:OPC

[BMFJ^{+}06] L. Baratchart, A. Martínez-Finkelshtein, D. Jimenez, D. S. Lubinsky, H. N. Mhaskar, I. Pritsker, M. Putinar, N. Stylianopoulos, V. Totik, P. Varju, and Y. Xu. Open problems in constructive function theory. Electronic Transactions on Numerical Analysis (ETNA), 25:511–525, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
25.2006/pp511-525.dir/
pp511-525.pdf.

Berndt:1997:LEE

[BMM97] Markus Berndt, Thomas A.

Manteuffel, and Stephen F. McCormick. Local error estimates and adaptive refinement for firstorder system least squares (FOSLS). Electronic Transactions on Numerical Analysis (ETNA), 6:35–43, 1997. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.6.1997/pp35-43.
dir/pp35-43.pdf. Special

issue on multilevel methods (Copper Mountain, CO, 1997).

Bini:2008:FSC

[BMP09] Dario A. Bini, Beatrice Meini, and Federico Poloni. Fast solution of a certain Riccati equation through Cauchy-like matrices. Electronic Transactions on Numerical Analysis (ETNA), 33:84–104, 2008–2009. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
33.2008-2009/pp84-104.
dir/pp84-104.pdf.

Bru:1995:PSA

[BMPS95] Rafael Bru, Violeta Migallón, José Penadés, and Daniel B. Szyld. Parallel, synchronous and asynchronous two-stage multisplitting methods. Electronic Transactions on Numerical Analysis (ETNA), 3:24–38, 1995. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.3.1995/pp24-38.
dir/pp24-38.pdf.

Becker:2008:CAF

[BMS08a] Roland Becker, Shipeng Mao, and Zhong-Ci Shi. A convergent adaptive finite element method with optimal complexity. Electronic Transactions on Numerical Analysis (ETNA), 30:291–304, 2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
30.2008/pp291-304.dir/
pp291-304.pdf.

Benner:2007:PSP

[BMS08b] Peter Benner, Hermann Mena, and Jens Saak. On the parameter selection problem in the newton-ADI iteration for large-scale Riccati equations. Electronic Transactions on Numerical Analysis (ETNA), 29: 136–149, 2007–2008. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
29.2007-2008/pp136-149.
dir/pp136-149.pdf.

Bueno:2006:CSS

[BMSR06] M. Isabel Bueno, Francisco Marcellán, and Jorge Sánchez-Ruiz. Continuous symmetrized Sobolev inner products of order N(II). Electronic Transactions on Numerical Analysis (ETNA), 24:55–65, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.24.2006/pp55-
65.dir/pp55-65.pdf.

Benner:1999:NNS

[BMX99] Peter Benner, Volker Mehrmann, and Hongguo Xu. A note on the numerical solution of complex Hamiltonian and skew-Hamiltonian eigenvalue problems. Electronic Transactions on Numerical Analysis (ETNA), 8:115–126, 1999. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
8.1999/pp115-126.dir/
pp115-126.pdf.

Byers:2007:SSA

[BMX07] Ralph Byers, Volker Mehrmann, and Hongguo Xu. A structured staircase algorithm for skew-symmetric/ symmetric pencils. Electronic Transactions on Numerical Analysis (ETNA), 26:1–33, 2007. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        26.2007/pp1-33.dir/pp1-
                                        33.pdf.

Bochev:1997:ENN

[Boc97]  P. Bochev.  Experi-

ences with negative norm least-square methods for the Navier–Stokes equations. Electronic Transactions on Numerical Analysis (ETNA), 6:44–62, 1997. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.6.1997/pp44-62.
                                        dir/pp44-62.pdf. Special

issue on multilevel methods (Copper Mountain, CO, 1997).

Boglaev:2005:UCM

[Bog05] Igor Boglaev. Uniform convergence of monotone iterative methods for semilinear singularly perturbed problems of elliptic and parabolic types. Electronic Transactions on Numerical Analysis (ETNA), 20: 86–103, 2005. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        20.2005/pp86-103.dir/
                                        pp86-103.pdf.

Bolten:2007:HGC

[Bol08] Matthias Bolten. Hierarchical grid coarsening for the solution of the Poisson equation in free space. Electronic Transactions on Numerical Analysis (ETNA), 29:70–80, 2007–2008. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
29.2007-2008/pp70-80.
dir/pp70-80.pdf.

Borges:2009:FNA

[Bor09] Carlos F. Borges. A full-Newton approach to separable nonlinear least squares problems and its application to discrete least squares rational approximation. Electronic Transactions on Numerical Analysis (ETNA), 35:57–68, 2009. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.35.2009/pp57-
68.dir/pp57-68.pdf.

Brenner:2008:WPS

[BOS08] Susanne C. Brenner, Luke Owens, and Li-Yeng Sung. A weakly over-penalized symmetric interior penalty method. Electronic Transactions on Numerical Analysis (ETNA), 30:107–127, 2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.

                                        edu/vol.30.2008/pp107-
                                        127.dir/pp107-127.pdf.

Bottcher:2004:TBP

[Böt04] Albrecht Böttcher. Transient behavior of powers and exponentials of large Toeplitz matrices. Electronic Transactions on Numerical Analysis (ETNA), 18:1–41, 2004. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        18.2004/pp1-41.dir/pp1-
                                        41.pdf.

Bohme:2003:FAF

[BP03] Martin Böhme and Daniel Potts. A fast algorithm for filtering and wavelet decomposition on the sphere. Electronic Transactions on Numerical Analysis (ETNA), 16:70–93, 2003. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.16.2003/pp70-
                                        93.dir/pp70-93.pdf.

Bottcher:2007:PAC

[BP07] Albrecht Böttcher and Daniel Potts. Probability against condition number and sampling of multivariate trigonometric random polynomials. Electronic Transactions on Numerical Analysis (ETNA), 26:178–189, 2007. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
26.2007/pp178-189.dir/
pp178-189.pdf.

Benner:2021:SNR

[BP21] Peter Benner and Davide Palitta. On the solution of the nonsymmetric T-Riccati equation. Electronic Transactions on Numerical Analysis (ETNA), 54:68–88, 2021. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.54.2021/pp68-
88.dir/pp68-88.pdf;
http://etna.mcs.kent.
edu/volumes/2021-2030/
vol54/abstract.php?vol=
54&pages=68-88.

Banas:2015:CLF

[BPP15] Lubomir Baňas, Marcus Page, and Dirk Praetorius. A convergent linear finite element scheme for the Maxwell–Landau–Lifshitz– Gilbert equations. Electronic Transactions on Numerical Analysis (ETNA), 44:250–270, 2015. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.44.2015/pp250-
270.dir/pp250-270.pdf;
 http://etna.mcs.kent.

                                        edu/volumes/2011-2020/
                                        vol44/abstract.php?vol=
                                        44&pages=250-270.

Brenner:2017:BPS

[BPS17] Susanne C. Brenner, Eun-Hee Park, and Li-Yeng Sung. A BDDC preconditioner for a symmetric interior penalty Galerkin method. Electronic Transactions on Numerical Analysis (ETNA), 46:190–214, 2017. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.46.2017/pp190-
                                        214.dir/pp190-214.pdf;
                                         http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol46/abstract.php?vol=
                                        46&pages=190-214.

Benzi:2007:DBA

[BR08] Michele Benzi and Nader Razouk. Decay bounds and O(n) algorithms for approximating functions of sparse matrices. Electronic Transactions on Numerical Analysis (ETNA), 28: 16–39, 2007–2008. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        28.2007-2008/pp16-39.
                                        dir/pp16-39.pdf. Special

volume for Gene Golub. Baglama:2014:IRL

[BR14] James Baglama and Daniel J. Richmond. Implicitly restarting the LSQR algorithm. Electronic Transactions on Numerical Analysis (ETNA), 42:85–105, 2014. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.42.2014/pp85-
105.dir/pp85-105.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol42/abstract.php?vol=
42&pages=85-105.

Brandt:1997:GCR

[Bra97] Achi Brandt. The Gauss Center research in multiscale scientific computation. Electronic Transactions on Numerical Analysis (ETNA), 6:1–34, 1997. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
6.1997/pp1-34.dir/pp1-

34.pdf. Special issue on multilevel methods (Copper Mountain, CO, 1997).

Brandt:2000:GHA

[Bra00]  Achi Brandt.  General

highly accurate algebraic coarsening. Electronic Transactions on Numerical Analysis (ETNA), 10: 1–20, 2000. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
10.2000/pp1-20.dir/pp1-

20.pdf. Multilevel methods (Copper Mountain, CO, 1999).

Brayanov:2006:UCD

[Bra06] Iliya A. Brayanov. Uniformly convergent difference scheme for singularly perturbed problem of mixed type. Electronic Transactions on Numerical Analysis (ETNA), 23:288–303, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

                                        /etna.mcs.kent.edu/vol.
                                        23.2006/pp288-303.dir/
                                        pp288-303.pdf.

Blanco:2016:TFD

[BRC^{+}16] J. Blanco, O. Rojas, C. Chacón, J. M. Guevara-Jordan, and J. Castillo. Tensor formulation of 3-D mimetic finite differences and applications to elliptic problems. Electronic Transactions on Numerical Analysis (ETNA), 45:457–475, 2016. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.45.2016/pp457-
                                        475.dir/pp457-475.pdf;
                                         http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol45/abstract.php?vol=
                                        45&pages=457-475.

Brezinski:1994:GNI

[Bre94] C. Brezinski. The generalizations of Newton’s interpolation formula due to Mühlbach and Andoyer. Electronic Transactions on Numerical Analysis (ETNA), 2:130–137, 1994. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
2.1994/pp130-137.dir/
pp130-137.pdf.

Brenner:2003:ATD

[Bre03] Susanne C. Brenner. Analysis of two-dimensional FETI-DP preconditioners by the standard additive Schwarz framework. Electronic Transactions on Numerical Analysis (ETNA), 16:165–185, 2003. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
16.2003/pp165-185.dir/
pp165-185.pdf.

Brenner:2004:DSP

[Bre04] Susanne C. Brenner. Discrete Sobolev and Poincaré inequalities for piecewise polynomial functions. Electronic Transactions on Numerical Analysis (ETNA), 18:42–48, 2004. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.18.2004/pp42-
48.dir/pp42-48.pdf.

Breuer:2016:NFS

[Bre16] Alex Breuer. New filtering strategies for implicitly restarted Lanczos iteration. Electronic Transactions on Numerical Analysis (ETNA), 45:16–32, 2016. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.45.2016/pp16-
                                        32.dir/pp16-32.pdf;
                                        http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol45/abstract.php?vol=
                                        45&pages=16-32.

Blum:2008:SAF

[BRS08] Heribert Blum, Andreas Rademacher, and Andreas Schröder. Space adaptive finite element methods for dynamic obstacle problems. Electronic Transactions on Numerical Analysis (ETNA), 32:162–172, 2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

                                        /etna.mcs.kent.edu/vol.
                                        32.2008/pp162-172.dir/
                                        pp162-172.pdf.

Barrera-Sanchez:2008:GQS

[BSDMGFTR09] P. Barrera-Sánchez, F. J. Domínguez-Mota, G. González-Flores, and J. G. Tinoco-Ruiz. Generating quality structured convex grids on irregular regions. Electronic Transactions on Numerical Analysis (ETNA), 34:76–89, 2008–2009. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
34.2008-2009/pp76-89.
dir/pp76-89.pdf.

Bellalij:2008:ASK

[BSS09] M. Bellalij, Y. Saad, and H. Sadok. Analysis of some Krylov subspace methods for normal matrices via approximation theory and convex optimization. Electronic Transactions on Numerical Analysis (ETNA), 33:17–30, 2008–2009. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
33.2008-2009/pp17-30.
dir/pp17-30.pdf.

Breiten:2016:LRS

[BSS16] Tobias Breiten, Valeria Simoncini, and Martin Stoll. Low-rank solvers for fractional differential equations. Electronic Transactions on Numerical Analysis (ETNA), 45:107–132, 2016. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.45.2016/pp107-
132.dir/pp107-132.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol45/abstract.php?vol=
45&pages=107-132.

Benzi:2007:BTP

[BU07] Michele Benzi and Bora Uçar. Block triangular preconditioners for M-matrices and Markov chains. Electronic Transactions on Numerical Analysis (ETNA), 26:209–227, 2007. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        26.2007/pp209-227.dir/
                                        pp209-227.pdf.

Bungartz:1997:MAH

[Bun97] Hans-Joachim Bungartz. A multigrid algorithm for higher order finite elements on sparse grids. Electronic Transactions on Numerical Analysis (ETNA), 6:63–77, 1997. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.6.1997/pp63-77.
                                        dir/pp63-77.pdf. Special

issue on multilevel methods (Copper Mountain, CO, 1997).

Brenner:2012:ISA

[BW12] Susanne C. Brenner and Kening Wang. An iterative substructuring algorithm for a C^{0} interior penalty method. Electronic Transactions on Numerical Analysis (ETNA), 39: 313–332, 2012. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
39.2012/pp313-332.dir/.

Bai:1998:EEP

[BY98] Zhaojun Bai and Qiang Ye. Error estimation of the Padé approximation of transfer functions via the Lanczos process. Electronic Transactions on Numerical Analysis (ETNA), 7:1–17, 1998. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
7.1998/pp1-17.dir/pp1-

17.pdf. Large scale eigenvalue problems (Argonne, IL, 1997).

Bjorck:1999:PLS

[BY99] A. Björck and J. Y. Yuan.

Preconditioners for least squares problems by LU factorization. Electronic Transactions on Numerical Analysis (ETNA), 8:26–35, 1999. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.8.1999/pp26-35.
dir/pp26-35.pdf.

Bader:2003:RPM

[BZ03] Michael Bader and Christoph Zenger. A robust and parallel multigrid method for convection diffusion equations. Electronic Transactions on Numerical Analysis (ETNA), 15:122–131, 2003. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
15.2003/pp122-131.dir/
pp122-131.pdf.

Cabada:2007:MLU

[Cab07]  Alberto Cabada.  The

method of lower and upper solutions for periodic and anti-periodic difference equations. Electronic Transactions on Numerical Analysis (ETNA), 27:13–25, 2007. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.27.2007/pp13-
25.dir/pp13-25.pdf.

Calvo:2015:TLO

[Cal15] Juan G. Calvo. A twolevel overlapping Schwarz method for H(curl) in two dimensions with irregular subdomains. Electronic Transactions on Numerical Analysis (ETNA), 44:497–521, 2015. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.44.2015/pp497-
521.dir/pp497-521.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/

vol44/abstract.php?vol=
44&pages=497-521.

Cardoso:2011:EFD

[Car11] João R. Cardoso. Evaluating the Fréchet derivative of the matrix p th root. Electronic Transactions on Numerical Analysis (ETNA), 38:202–217, 2011. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic).

Cardoso:2012:CMT

[Car12] João R. Cardoso. Computation of the matrix p th root and its Fréchet derivative by integrals. Electronic Transactions on Numerical Analysis (ETNA), 39:414–436, 2012. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.39.2012/pp414-
436.dir/pp414-436.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol39/abstract.php?vol=
39&pages=414-436.

Cruz-Barroso:2005:OLP

[CBGV05] Ruymán Cruz-Barroso and Pablo González-Vera. Orthogonal Laurent polynomials and quadratures on the unit circle and the real halfline. Electronic Transactions on Numerical Analysis (ETNA), 19:113–134, 2005. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
19.2005/pp113-134.dir/
pp113-134.pdf.

Craven:1997:CLT

[CC97] Thomas Craven and George Csordas. On a converse of Laguerre’s Theorem. Electronic Transactions on Numerical Analysis (ETNA), 5:7–17, 1997. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
5.1997/pp7-17.dir/pp7-
17.pdf.

Cawood:2000:PAE

[CC00] M. E. Cawood and C. L. Cox. Perturbation analysis for eigenstructure assignment of linear multi-input systems. Electronic Transactions on Numerical Analysis (ETNA), 11:25–42, 2000. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.11.2000/pp25-
42.dir/pp25-42.pdf.

Chan:2007:IMS

[CCC07] Tony F. Chan, Ke Chen, and Jamylle L. Carter. Iterative methods for solving the dual formulation arising from image restoration. Electronic Transactions on Numerical Analysis (ETNA), 26:299–311, 2007.

CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
26.2007/pp299-311.dir/
pp299-311.pdf.

Capobianco:2008:SAF

[CCDR08] G. Capobianco, D. Conte, I. Del Prete, and E. Russo. Stability analysis of fast numerical methods for Volterra integral equations. Electronic Transactions on Numerical Analysis (ETNA), 30:305–322, 2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
30.2008/pp305-322.dir/
pp305-322.pdf.

Cardone:2009:GDQ

[CDN09] Angelamaria Cardone, Ida Del Prete, and Claudia Nitsch. Gaussian Direct Quadrature methods for double delay Volterra integral equations. Electronic Transactions on Numerical Analysis (ETNA), 35: 201–216, 2009. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
35.2009/pp201-216.dir/
pp201-216.pdf.

Crofts:2010:MDN

[CEHT10] Jonathan J. Crofts, Ernesto Estrada, Desmond J. Higham, and Alan Taylor. Mapping directed networks. Electronic Transactions on Numerical Analysis (ETNA), 37:337–350, 2010. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic).

Cerna:2018:QSW

[ČF18] Dana Černá and Václav Finěk. Quadratic spline wavelets with short support satisfying homogeneous boundary conditions. Electronic Transactions on Numerical Analysis (ETNA), 48:15–39, 2018. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.48.2018/pp15-
39.dir/pp15-39.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol48/abstract.php?vol=
48&pages=15-39.

Cardoso:2009:SRR

[CFÁN09] J. L. Cardoso, C. M. Fernandes, and R. Álvarez-Nodarse. Structural and recurrence relations for hypergeometric-type functions by Nikiforov–Uvarov method. Electronic Transactions on Numerical Analysis (ETNA), 35:17–39, 2009. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.

edu/vol.35.2009/pp17-
39.dir/pp17-39.pdf.

Cerdan:2010:BAI

[CFM^{+}10] J. Cerdán, T. Faraj, N. Malla, J. Marín, and J. Mas. Block approximate inverse preconditioners for sparse nonsymmetric linear systems. Electronic Transactions on Numerical Analysis (ETNA), 37:23–40, 2010. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.37.2010/pp23-
40.dir/pp23-40.pdf.

Ciaramella:2018:APS

[CG18] Gabriele Ciaramella and Martin J. Gander. Analysis of the parallel Schwarz method for growing chains of fixed-sized subdomains: Part III. Electronic Transactions on Numerical Analysis (ETNA), 49:210–243, 2018. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.49.2018/pp210-
243.dir/pp210-243.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol49/abstract.php?vol=
49&pages=210-243.

Castillo:2009:LPP

[CGM10] Kenier Castillo, Luis Garza, and Francisco Marcellán. Laurent polynomial perturbations of linear functionals. An inverse problem. Electronic Transactions on Numerical Analysis (ETNA), 36:83–98, 2009– 2010. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.36.2009-2010/
pp83-98.dir/pp83-98.pdf.

Chen:2005:OBT

[CGT05] Doron Chen, John R. Gilbert, and Sivan Toledo. Obtaining bounds on the two norm of a matrix from the splitting lemma. Electronic Transactions on Numerical Analysis (ETNA), 21:28–46, 2005. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.21.2005/pp28-
46.dir/pp28-46.pdf.

Chahlaoui:2011:TES

[Cha11] Younès Chahlaoui. Two efficient SVD/Krylov algorithms for model order reduction of large scale systems. Electronic Transactions on Numerical Analysis (ETNA), 38:113–145, 2011. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic).

Chen:1994:EIS

[Che94]  Ke Chen.  Efficient it-

erative solution of linear systems from discretizing singular integral equations. Electronic Transactions on Numerical Analysis (ETNA), 2:76–91, 1994. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.2.1994/pp76-91.
dir/pp76-91.pdf.

Chen:1997:AIT

[Che97] Zhangxin Chen. The analysis of intergrid transfer operators and multigrid methods for nonconforming finite elements. Electronic Transactions on Numerical Analysis (ETNA), 6:78–96, 1997. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.6.1997/pp78-96.
dir/pp78-96.pdf. Special

issue on multilevel methods (Copper Mountain, CO, 1997).

Chen:1999:DWT

[Che99] Ke Chen. Discrete wavelet transforms accelerated sparse preconditioners for dense boundary element systems. Electronic Transactions on Numerical Analysis (ETNA), 8:138–153, 1999. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.

                                        8.1999/pp138-153.dir/
                                        pp138-153.pdf.

Chelyshkov:2006:AOP

[Che06] Vladimir S. Chelyshkov. Alternative orthogonal polynomials and quadratures. Electronic Transactions on Numerical Analysis (ETNA), 25:17–26, 2006. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.25.2006/pp17-
                                        26.dir/pp17-26.pdf.

Chernov:2009:CLN

[Che09] I. A. Chernov. Convergence of a lattice numerical method for a boundaryvalue problem with free boundary and nonlinear Neumann boundary conditions. Electronic Transactions on Numerical Analysis (ETNA), 35:40–56, 2009. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.35.2009/pp40-
                                        56.dir/pp40-56.pdf.

Calvetti:2002:CCB

[CHR02] D. Calvetti, P. C. Hansen, and L. Reichel. Lcurve curvature bounds via Lanczos bidiagonalization. Electronic Transactions on Numerical Analysis (ETNA), 14:20–35, 2002. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.14.2002/pp20-
35.dir/pp20-35.pdf.

Carson:2014:EDT

[CKD15] Erin Carson, Nicholas Knight, and James Demmel. An efficient deflation technique for the communication-avoiding conjugate gradient method. Electronic Transactions on Numerical Analysis (ETNA), 43:125–141, 2014–2015. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.43.2014/pp125-
141.dir/pp125-141.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol43/abstract.php?vol=
43&pages=125-141.

Chen:2014:PPE

[CKH14] Donghui Chen, Misha E.

Kilmer, and Per Christian Hansen. “Plug-and-Play” edge-preserving regularization. Electronic Transactions on Numerical Analysis (ETNA), 41:465–477, 2014. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.41.2014/pp465-
477.dir/pp465-477.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/

                                        vol41/abstract.php?vol=
                                        41&pages=465-477.

Christov:2014:IPB

[CKP15] Nicolai D. Christov, Mihail M. Konstantinov, and Petko Hr. Petkov. Improved perturbation bounds for the continuous-time H_{∞}- control problem. Electronic Transactions on Numerical Analysis (ETNA), 43: 90–99, 2014–2015. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.43.2014/pp90-
                                        99.dir/pp90-99.pdf;
                                        http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol43/abstract.php?vol=
                                        43&pages=90-99.

Cvetkovic:2004:NGT

[CKV04] Ljiljana Cvetkovic, Vladimir Kostic, and Richard S. Varga. A new Geršhgorintype eigenvalue inclusion set. Electronic Transactions on Numerical Analysis (ETNA), 18:73–80, 2004. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.18.2004/pp73-
                                        80.dir/pp73-80.pdf.

Chang:2003:MAS

[CL03] B. Chang and B. Lee. A multigrid algorithm for solving the multigroup, anisotropic scattering Boltzmann equation using first-order system leastsquares methodology. Electronic Transactions on Numerical Analysis (ETNA), 15:132–151, 2003. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
15.2003/pp132-151.dir/
pp132-151.pdf.

Calvetti:2004:TRN

[CLRS04] Daniela Calvetti, Bryan Lewis, Lothar Reichel, and Fiorella Sgallari. Tikhonov regularization with nonnegativity constraint. Electronic Transactions on Numerical Analysis (ETNA), 18:153–173, 2004. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
18.2004/pp153-173.dir/
pp153-173.pdf.

Cai:1995:FOS

[CMM95] Zhiqiang Cai, Thomas A.

Manteuffel, and Stephen F. McCormick. First-order system least squares for velocity-vorticity-pressure form of the Stokes equations, with application to linear elasticity. Electronic Transactions on Numerical Analysis (ETNA), 3: 150–159, 1995. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        3.1995/pp150-159.dir/
                                        pp150-159.pdf.

Camano:2018:NAD

[CMO18] Jessika Camaño, Cristian Muñoz, and Ricardo Oyarzúa. Numerical analysis of a dual-mixed problem in non-standard Banach spaces. Electronic Transactions on Numerical Analysis (ETNA), 48:114–130, 2018. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.48.2018/pp114-
                                        130.dir/pp114-130.pdf;
                                         http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol48/abstract.php?vol=
                                        48&pages=114-130.

Castel:2001:PTS

[CMP01] M. Jesús Castel, Violeta Migallón, and José Penadés. On parallel two-stage methods for Hermitian positive definite matrices with applications to preconditioning. Electronic Transactions on Numerical Analysis (ETNA), 12:88–112, 2001. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

                                        /etna.mcs.kent.edu/vol.
                                        12.2001/pp88-112.dir/
                                        pp88-112.pdf.

Czapla:2012:CMC

[CMR12] Roman Czapla, Vladimir Mityushev, and Natalia Rylko. Conformal mapping of circular multiply connected domains onto slit domains. Electronic Transactions on Numerical Analysis (ETNA), 39:286–297, 2012. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
39.2012/pp286-297.dir/.

Clason:2019:BLM

[CN19] Christian Clason and Vu Huu Nhu. Bouligand–Levenberg– Marquardt iteration for a non-smooth ill-posed inverse problem. Electronic Transactions on Numerical Analysis (ETNA), 51:274–314, 2019. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.51.2019/pp274-
314.dir/pp274-314.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol51/abstract.php?vol=
51&pages=274-314.

Chung:2007:WGM

[CNO08] Julianne Chung, James G.

Nagy, and Dianne P. O’Leary. A weighted-GCV method for Lanczos-hybrid regularization. Electronic Transactions on Numerical Analysis (ETNA), 28:

149–167, 2007–2008. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
28.2007-2008/pp149-167.
dir/pp149-167.pdf. Spe-

cial volume for Gene Golub. Chan:1994:DPT

[CNP94] Raymond H. Chan, James G. Nagy, and Robert J. Plemmons. Displacement preconditioner for Toeplitz least squares iterations. Electronic Transactions on Numerical Analysis (ETNA), 2:44–56, 1994. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.2.1994/pp44-56.
dir/pp44-56.pdf.

Csordas:1993:NLP

[COSV93] G. Csordas, A. M. Odlyzko, W. Smith, and R. S. Varga. A new Lehmer pair of zeros and a new lower bound for the de Bruijn–Newman constant Λ. Electronic Transactions on Numerical Analysis (ETNA), 1:104–111, 1993. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
1.1993/pp104-111.dir/
pp104-111.pdf.

Cotfas:2006:SOP

[Cot06]  Nicolae Cotfas.  Sys-

tems of orthogonal polynomials defined by hypergeometric type equations. Electronic Transactions on Numerical Analysis (ETNA), 24:45–54, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.24.2006/pp45-
54.dir/pp45-54.pdf.

Capizzano:2000:HOF

[CP00] Stefano Serra Capizzano and Cristina Tablino Possio. High-order finite difference schemes and Toeplitz based preconditioners for elliptic problems. Electronic Transactions on Numerical Analysis (ETNA), 11:55–84, 2000. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.11.2000/pp55-
84.dir/pp55-84.pdf.

Capizzano:2003:PSF

[CP03] Stefano Serra Capizzano and Cristina Tablino Possio. Preconditioning strategies for 2D finite difference matrix sequences. Electronic Transactions on Numerical Analysis (ETNA), 16:1–29, 2003. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
16.2003/pp1-29.dir/pp1-
29.pdf.

Cho:2018:ISP

[CPS18] D. Cho, L. F. Pavarino, and S. Scacchi. Isogeometric Schwarz preconditioners for the biharmonic problem. Electronic Transactions on Numerical Analysis (ETNA), 49:81–102, 2018. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.49.2018/pp81-
                                        102.dir/pp81-102.pdf;
                                        http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol49/abstract.php?vol=
                                        49&pages=81-102.

Chen:2001:MSF

[CQ01] Jing-Bo Chen and Meng-Zhao Qin. Multi-symplectic Fourier pseudospectral method for the nonlinear Schrödinger equation. Electronic Transactions on Numerical Analysis (ETNA), 12:193–204, 2001. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.12.2001/pp193-
                                        204.dir/pp193-204.pdf.

Calvetti:1996:SCS

[CR96] D. Calvetti and L. Reichel. On the solution of Cauchy systems of equations. Electronic Transactions on Numerical Analysis (ETNA), 4:125–137, 1996. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
4.1996/pp125-137.dir/
pp125-137.pdf.

Chehab:2008:IPG

[CR09] Jean-Paul Chehab and Marcos Raydan. Implicitly preconditioned and globalized residual method for solving steady fluid flows. Electronic Transactions on Numerical Analysis (ETNA), 34:136–151, 2008–2009. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
34.2008-2009/pp136-151.
dir/pp136-151.pdf.

Castillo:2015:DEM

[CR15] Kenier Castillo and Fernando R. Rafaeli. On the discrete extension of Markov’s theorem on monotonicity of zeros. Electronic Transactions on Numerical Analysis (ETNA), 44:271–280, 2015. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.44.2015/pp271-
280.dir/pp271-280.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol44/abstract.php?vol=
44&pages=271-280.

Calvetti:1994:IRL

[CRS94] D. Calvetti, L. Reichel, and D. C. Sorensen. An implicitly restarted Lanczos method for large symmetric eigenvalue problems. Electronic Transactions on Numerical Analysis (ETNA), 2:1–21, 1994. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
2.1994/pp1-21.dir/pp1-
21.pdf.

Cipolla:2020:SEP

[CRZT20] Stefano Cipolla, Michela Redivo-Zaglia, and Francesco Tudisco. Shifted and extrapolated power methods for tensor ℓ^{p}-eigenpairs. Electronic Transactions on Numerical Analysis (ETNA), 53:1–27, 2020. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
53.2020/pp1-27.dir/pp1-
27.pdf;   http://etna.
mcs.kent.edu/volumes/
2011-2020/vol53/abstract.
php?vol=53&pages=1-27.

Chan:1994:DDM

[CS94] Tony F. Chan and Barry F. Smith. Domain decomposition and multigrid algorithms for elliptic problems on unstructured meshes. Electronic Transactions on Numerical Analysis (ETNA), 2:171–182, 1994. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
2.1994/pp171-182.dir/
pp171-182.pdf.

Costa:1999:FQB

[CS99] Cecília Costa and Rogério Serôdio. A footnote on quaternion block-tridiagonal systems. Electronic Transactions on Numerical Analysis (ETNA), 9:53–55, 1999. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.9.1999/pp53-55.
dir/pp53-55.pdf.

Cvetkovic:2014:EEG

[CS14] Aleksandar S. Cvetković and Miodrag M. Spalević. Estimating the error of Gauss–Turán quadrature formulas using their extensions. Electronic Transactions on Numerical Analysis (ETNA), 41:1–12, 2014. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
41.2014/pp1-12.dir/pp1-
12.pdf;   http://etna.
mcs.kent.edu/volumes/
2011-2020/vol41/abstract.
php?vol=41&pages=1-12.

Chen:2003:VPI

[CT03] Doron Chen and Sivan Toledo. Vaidya’s preconditioners: implementation and experimental study. Electronic Transactions on Numerical Analysis (ETNA), 16:30–49, 2003. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.16.2003/pp30-
49.dir/pp30-49.pdf.

Casarin:1996:PMF

[CW96] Mario A. Casarin and Olof B. Widlund. A preconditioner for the mortar finite element method. Electronic Transactions on Numerical Analysis (ETNA), 4:75–88, 1996. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.4.1996/pp75-88.
dir/pp75-88.pdf.

Calvo:2016:ACP

[CW16] Juan G. Calvo and Olof B. Widlund. An adaptive choice of primal constraints for BDDC domain decomposition algorithms. Electronic Transactions on Numerical Analysis (ETNA), 45:524–544, 2016. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.45.2016/pp524-
544.dir/pp524-544.pdf;
 http://etna.mcs.kent.

edu/volumes/2011-2020/
vol45/abstract.php?vol=
45&pages=524-544.

Damelin:2002:ADG

[Dam02] S. B. Damelin. The asymptotic distribution of general interpolation arrays for exponential weights. Electronic Transactions on Numerical Analysis (ETNA), 13:12–21, 2002. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.13.2002/pp12-
21.dir/pp12-21.pdf.

Diabate:2008:ABN

[DB08] Nabongo Diabate and Théodore K.

Boni. Asymptotic behavior for numerical solutions of a semilinear parabolic equation with a nonlinear boundary condition. Electronic Transactions on Numerical Analysis (ETNA), 30:237–246, 2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
30.2008/pp237-246.dir/
pp237-246.pdf.

Deidda:2017:IMP

[DdAR17] Gian Piero Deidda, Patricia Díaz de Alba, and Giuseppe Rodriguez. Identifying the magnetic permeability in multi-frequency EM data inversion. Electronic Transactions on Numerical Analysis (ETNA), 47:1–17, 2017. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
47.2017/pp1-17.dir/pp1-
17.pdf;   http://etna.
mcs.kent.edu/volumes/
2011-2020/vol47/abstract.
php?vol=47&pages=1-17.

Djukic:2020:EEG

[DDPS20] D. Lj. Djukić, R. M. Mutavdžić Djukić, A. V. Pejčev, and M. M. Spalević. Error estimates of Gaussiantype quadrature formulae for analytic functions on ellipses — a survey of recent results. Electronic Transactions on Numerical Analysis (ETNA), 53:352–382, 2020. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.53.2020/pp352-
382.dir/pp352-382.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol53/abstract.php?vol=
53&pages=352-382.

Demmel:1995:CSB

[DDR95] James W. Demmel, Inderjit Dhillon, and Huan Ren. On the correctness of some bisection-like parallel eigenvalue algorithms in floating point arithmetic. Electronic Transactions on Numerical Analysis (ETNA), 3:116–149, 1995. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
3.1995/pp116-149.dir/
pp116-149.pdf.

Domschke:2018:ARS

[DDS^{+}18] Pia Domschke, Aseem Dua, Jeroen J. Stolwijk, Jens Lang, and Volker Mehrmann. Adaptive refinement strategies for the simulation of gas flow in networks using a model hierarchy. Electronic Transactions on Numerical Analysis (ETNA), 48:97–113, 2018. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.48.2018/pp97-
113.dir/pp97-113.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol48/abstract.php?vol=
48&pages=97-113.

DeSimone:2020:SAS

[DdSV20] Valentina De Simone, Daniela di Serafino, and Marco Viola. A subspace-accelerated split Bregman method for sparse data recovery with joint ℓ_{1}-type regularizers. Electronic Transactions on Numerical Analysis (ETNA), 53:406–425, 2020. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.53.2020/pp406-
                                        425.dir/pp406-425.pdf;
                                         http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol53/abstract.php?vol=
                                        53&pages=406-425.

Duvnjak:2006:TDD

[DE06] Antonija Duvnjak and Hermann J. Eberl. Timediscretization of a degenerate reaction-diffusion equation arising in biofilm modeling. Electronic Transactions on Numerical Analysis (ETNA), 23:15–37, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.23.2006/pp15-
                                        37.dir/pp15-37.pdf.

Delvos:2002:UAM

[Del02] Franz-Jürgen Delvos. Uniform approximation by minimum norm interpolation. Electronic Transactions on Numerical Analysis (ETNA), 14:36–44, 2002. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.14.2002/pp36-
                                        44.dir/pp36-44.pdf.

Dendy:1997:SMS

[Den97] J. E. Dendy, Jr. Semicoarsening multigrid for systems. Electronic Transactions on Numerical Analysis (ETNA), 6:97–105, 1997. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
6.1997/pp97-105.dir/pp97-

105.pdf. Special issue on multilevel methods (Copper Mountain, CO, 1997).

DiBenedetto:2008:NLA

[DENP09] Fabio Di Benedetto, Claudio Estatico, James G. Nagy, and Matteo Pastorino. Numerical linear algebra for nonlinear microwave imaging. Electronic Transactions on Numerical Analysis (ETNA), 33: 105–125, 2008–2009. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
33.2008-2009/pp105-125.
dir/pp105-125.pdf.

Delvaux:2008:THU

[DFV09] Steven Delvaux, Katrijn Frederix, and Marc Van Barel. Transforming a hierarchical into a unitaryweight representation. Electronic Transactions on Numerical Analysis (ETNA), 33:163–188, 2008–2009. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
33.2008-2009/pp163-188.
dir/pp163-188.pdf.

Dean:2006:ALA

[DG06] E. J. Dean and R. Glowinski. An augmented Lagrangian approach to the numerical solution of the Dirichlet problem for the elliptic Monge–Ampère equation in two dimensions. Electronic Transactions on Numerical Analysis (ETNA), 22:71–96, 2006. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.22.2006/pp71-
                                        96.dir/pp71-96.pdf.

Dong:2014:IPI

[DGH14] Yiqiu Dong, Henrik Garde, and Per Christian Hansen. R^{3}GMRES: Including prior information in GMREStype methods for discrete inverse problems. Electronic Transactions on Numerical Analysis (ETNA), 42:136–146, 2014. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.42.2014/pp136-
                                        146.dir/pp136-146.pdf;
                                         http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol42/abstract.php?vol=
                                        42&pages=136-146.

Dobrowolski:1999:PEE

[DGP99] Manfred Dobrowolski, Steffen Gräf, and Christoph Pflaum. On a posteriori error estimators in the finite element method on anisotropic meshes. Electronic Transactions on Numerical Analysis (ETNA), 8:36–45, 1999. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.8.1999/pp36-45.
dir/pp36-45.pdf.

DeVries:2007:OFD

[DH07] J. DeVries and A. Hulme.

Oscillation of factored dynamic equations. Electronic Transactions on Numerical Analysis (ETNA), 27:26–33, 2007. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.27.2007/pp26-
33.dir/pp26-33.pdf.

Douglas:2003:ASP

[DHI03] Craig C. Douglas, Gundolf Haase, and Mohamed Iskandarani. An additive Schwarz preconditioner for the spectral element ocean model formulation of the shallow water equations. Electronic Transactions on Numerical Analysis (ETNA), 15:18–28, 2003. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.

                                        edu/vol.15.2003/pp18-
                                        28.dir/pp18-28.pdf.

Douglas:2000:COS

[DHK^{+}00] Craig C. Douglas, Jonathan Hu, Markus Kowarschik, Ulrich Rüde, and Christian Weiss. Cache optimization for structured and unstructured grid multigrid. Electronic Transactions on Numerical Analysis (ETNA), 10:21–40, 2000. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.10.2000/pp21-
                                        40.dir/pp21-40.pdf. Mul-

tilevel methods (Copper Mountain, CO, 1999).

Diethelm:1997:ANS

[Die97]  Kai Diethelm.  An al-

gorithm for the numerical solution of differential equations of fractional order. Electronic Transactions on Numerical Analysis (ETNA), 5:1–6, 1997. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        5.1997/pp1-6.dir/pp1-6.
                                        pdf.

Diskin:1997:MAC

[Dis97] Boris Diskin. Multigrid algorithm with conditional coarsening for the nonaligned sonic flow. Electronic Transactions on Numerical Analysis (ETNA), 6:106–119, 1997. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
6.1997/pp106-119.dir/
pp106-119.pdf. Special

issue on multilevel methods (Copper Mountain, CO, 1997).

Dahmen:2006:ECR

[DJ06a] Wolfgang Dahmen and Markus Jürgens. Error controlled regularization by projection. Electronic Transactions on Numerical Analysis (ETNA), 25:67–100, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
25.2006/pp67-100.dir/
pp67-100.pdf.

Driver:2006:IRS

[DJ06b] K. A. Driver and S. J. Johnston. An integral representation of some hypergeometric functions. Electronic Transactions on Numerical Analysis (ETNA), 25:115–120, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
25.2006/pp115-120.dir/
pp115-120.pdf.

Driver:2011:SIZ

[DJJ11] K. Driver, A. Jooste, and K. Jordaan. Stieltjes interlacing of zeros of Jacobi polynomials from different sequences. Electronic Transactions on Numerical Analysis (ETNA), 38: 317–326, 2011. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        38.2011/pp317-326.dir/
                                        pp317-326.pdf.

deJong:2017:FGP

[dJvdPDV17] Martijn de Jong, Auke van der Ploeg, Auke Ditzel, and Kees Vuik. Fine-grain parallel RRB-solver for 5- /9-point stencil problems suitable for GPU-type processors. Electronic Transactions on Numerical Analysis (ETNA), 46:375–393, 2017. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.46.2017/pp375-
                                        393.dir/pp375-393.pdf;
                                         http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol46/abstract.php?vol=
                                        46&pages=375-393.

DeLillo:2009:SMS

[DK10] Thomas K. DeLillo and Everett H. Kropf. Slit maps and Schwarz–Christoffel maps for multiply connected domains. Electronic Transactions on Numerical Analysis (ETNA), 36: 195–223, 2009–2010. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
36.2009-2010/pp195-223.
dir/pp195-223.pdf;   http:
//etna.mcs.kent.edu/volumes/
2001-2010/vol36/abstract.
php?vol=36&pages=195-
223.

Duff:2013:PBS

[DK13] Iain S. Duff and Kamer Kaya. Preconditioners based on strong subgraphs. Electronic Transactions on Numerical Analysis (ETNA), 40:225–248, 2013. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu//vol.
40.2013/pp225-248.dir/
pp225-248.pdf.

deMorales:1999:NSO

[dMPPR99] María Álvarez de Morales, Teresa E. Pérez, Miguel A. Piñar, and André Ronveaux. Non-standard orthogonality for Meixner polynomials. Electronic Transactions on Numerical Analysis (ETNA), 9: 1–25, 1999. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
9.1999/pp1-25.dir/pp1-
25.pdf.

Dayar:2011:SSA

[DN11] Tugrul Dayar and Gökçe N. Noyan. Steady-state analysis of Google-like stochastic matrices with block iterative methods. Electronic Transactions on Numerical Analysis (ETNA), 38:69– 97, 2011. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic).

Daruis:2005:SQF

[DNV05] Leyla Daruis, Olav Njåstad, and Walter Van Assche. Szegő quadrature and frequency analysis. Electronic Transactions on Numerical Analysis (ETNA), 19:48–57, 2005. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.19.2005/pp48-
                                        57.dir/pp48-57.pdf.

DeBonis:2018:PIR

[DO18] Maria Carmela De Bonis and Donatella Occorsio. A product integration rule for hypersingular integrals on (0, +∞). Electronic Transactions on Numerical Analysis (ETNA), 50:129–143, 2018. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.50.2018/pp129-
                                        143.dir/pp129-143.pdf;
                                         http://etna.mcs.kent.
                                        edu/volumes/2011-2020/

vol50/abstract.php?vol=
50&pages=129-143.

Dobrowolski:2012:DRP

[Dob12] Manfred Dobrowolski. On domain-robust preconditioners for the Stokes equations. Electronic Transactions on Numerical Analysis (ETNA), 39:60–74, 2012. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
39.2012/pp60-74.dir.

Dostal:2008:DQF

[Dos08] Zdeněk Dostál. On the decrease of a quadratic function along the projectedgradient path. Electronic Transactions on Numerical Analysis (ETNA), 31:25–29, 2008. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.31.2008/pp25-
29.dir/pp25-29.pdf.

Drmac:2015:SHM

[Drm15]  Zlatko Drmač.  SVD

of Hankel matrices in Vandermonde–Cauchy product form. Electronic Transactions on Numerical Analysis (ETNA), 44:593–623, 2015. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.44.2015/pp593-

                                        623.dir/pp593-623.pdf;
                                         http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol44/abstract.php?vol=
                                        44&pages=593-623.

Djukic:2016:IGA

[DRST16] D. Lj. Djukić, L. Reichel, M. M. Spalević, and J. D. Tomanović. Internality of generalized averaged Gauss rules and their truncations for Bernstein–Szegő weights. Electronic Transactions on Numerical Analysis (ETNA), 45:405–419, 2016. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.45.2016/pp405-
                                        419.dir/pp405-419.pdf;
                                         http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol45/abstract.php?vol=
                                        45&pages=405-419.

Duminil:2011:RRE

[DS11] Sébastien Duminil and Hassane Sadok. Reduced rank extrapolation applied to electronic structure computations. Electronic Transactions on Numerical Analysis (ETNA), 38:347–362, 2011. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

                                        /etna.mcs.kent.edu/vol.
                                        38.2011/pp347-362.dir/
                                        pp347-362.pdf.

Donatelli:2007:FFA

[DSC08] Marco Donatelli and Stefano Serra-Capizzano. Filter factor analysis of an iterative multilevel regularizing method. Electronic Transactions on Numerical Analysis (ETNA), 29: 163–177, 2007–2008. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
29.2007-2008/pp163-177.
dir/pp163-177.pdf.

Du:2013:ITP

[DSSS13] Xiuhong Du, Marcus Sarkis, Christian E. Schaerer, and Daniel B. Szyld. Inexact and truncated Parareal-in-time Krylov subspace methods for parabolic optimal control problems. Electronic Transactions on Numerical Analysis (ETNA), 40:36–57, 2013. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.40.2013/pp36-
57.dir/pp36-57.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol40/abstract.php?vol=
40&pages=36-57.

Deano:2009:MAS

[DT09] Alfredo Deaño and Nico M.

Temme. On modified asymptotic series involving confluent hypergeometric functions. Electronic Transactions on Numerical Analysis (ETNA), 35: 88–103, 2009. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
35.2009/pp88-103.dir/
pp88-103.pdf.

Du:2017:AHR

[DTM17] Kui Du, Jurjen Duintjer Tebbens, and Gérard Meurant. Any admissible harmonic Ritz value set is possible for GM-RES. Electronic Transactions on Numerical Analysis (ETNA), 47:37–56, 2017. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.47.2017/pp37-
56.dir/pp37-56.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol47/abstract.php?vol=
47&pages=37-56.

Dubrulle:1999:OIM

[Dub99] A. A. Dubrulle. An optimum iteration for the matrix polar decomposition. Electronic Transactions on Numerical Analysis (ETNA), 8:21–25, 1999. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.

edu/vol.8.1999/pp21-25.
dir/pp21-25.pdf.

Dubrulle:2001:RMB

[Dub01] A. A. Dubrulle. Retooling the method of block conjugate gradients. Electronic Transactions on Numerical Analysis (ETNA), 12:216–233, 2001. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
12.2001/pp216-233.dir/
pp216-233.pdf.

DaFies:2012:TGQ

[DV12] Gaspare Da Fies and Marco Vianello. Trigonometric Gaussian quadrature on subintervals of the period. Electronic Transactions on Numerical Analysis (ETNA), 39:102–112, 2012. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
39.2012/pp102-112.dir/.

Erfani:2021:NFP

[EBJ21] Shervan Erfani, Esmail Babolian, and Shahnam Javadi. New fractional pseudospectral methods with accurate convergence rates for fractional differential equations. Electronic Transactions on Numerical Analysis (ETNA), 54:150–175, 2021. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.54.2021/pp150-
175.dir/pp150-175.pdf;
 http://etna.mcs.kent.
edu/volumes/2021-2030/
vol54/abstract.php?vol=
54&pages=150-175.

ElBouajaji:2015:DGD

[EDG^{+}15] Mohamed El Bouajaji, Victorita Dolean, Martin J. Gander, Stéphane Lanteri, and Ronan Perrussel. Discontinuous Galerkin discretizations of optimized Schwarz methods for solving the timeharmonic Maxwell’s equations. Electronic Transactions on Numerical Analysis (ETNA), 44:572–592, 2015. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.44.2015/pp572-
592.dir/pp572-592.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol44/abstract.php?vol=
44&pages=572-592.

El-Gamel:2006:SGM

[EG06] Mohamed El-Gamel. The Sinc–Galerkin method for solving singularly-perturbed reaction-diffusion problem. Electronic Transactions on Numerical Analysis (ETNA), 23:129–140, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
23.2006/pp129-140.dir/
pp129-140.pdf.

Eidelman:2015:FBE

[EH15] Yuli Eidelman and Iulian Haimovici. The fast bisection eigenvalue method for Hermitian order one quasiseparable matrices and computations of norms. Electronic Transactions on Numerical Analysis (ETNA), 44:342–366, 2015. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.44.2015/pp342-
366.dir/pp342-366.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol44/abstract.php?vol=
44&pages=342-366.

Elbouyahyaoui:2017:BAB

[EHJM17] L. Elbouyahyaoui, M. Heyouni, K. Jbilou, and A. Messaoudi. A block Arnoldi based method for the solution of the Sylvesterobserver equation. Electronic Transactions on Numerical Analysis (ETNA), 47:18–36, 2017. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.47.2017/pp18-
36.dir/pp18-36.pdf;

                                        http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol47/abstract.php?vol=
                                        47&pages=18-36.

ElGuennouni:2003:BVB

[EJS03] A. El Guennouni, K. Jbilou, and H. Sadok. A block version of BiCGSTAB for linear systems with multiple right-hand sides. Electronic Transactions on Numerical Analysis (ETNA), 16:129–142, 2003. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        16.2003/pp129-142.dir/
                                        pp129-142.pdf.

Ehler:2019:RMM

[EKPR19] Martin Ehler, Stefan Kunis, Thomas Peter, and Christian Richter. A randomized multivariate matrix pencil method for superresolution microscopy. Electronic Transactions on Numerical Analysis (ETNA), 51:63–74, 2019. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.51.2019/pp63-
                                        74.dir/pp63-74.pdf;
                                        http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol51/abstract.php?vol=
                                        51&pages=63-74.

Eisenstat:2005:TBD

[EL05] Stanley C. Eisenstat and Joseph W. H. Liu. A tree-based dataflow model for the unsymmetric multifrontal method. Electronic Transactions on Numerical Analysis (ETNA), 21:1–19, 2005. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
21.2005/pp1-19.dir/pp1-
19.pdf.

Elsner:2005:IRN

[Els05] Ulrich Elsner. The influence of random number generators on graph partitioning algorithms. Electronic Transactions on Numerical Analysis (ETNA), 21: 125–133, 2005. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
21.2005/pp125-133.dir/
pp125-133.pdf.

Echeverria:2018:CMS

[ELST18] Carlos Echeverría, Jörg Liesen, Daniel B. Szyld, and Petr Tichý. Convergence of the multiplicative Schwarz method for singularly perturbed convectiondiffusion problems discretized on a Shishkin mesh. Electronic Transactions on Numerical Analysis (ETNA), 48:40–62, 2018.

CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.48.2018/pp40-
62.dir/pp40-62.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol48/abstract.php?vol=
48&pages=40-62.

Echeverria:2021:AMS

[ELT21] Carlos Echeverría, Jörg Liesen, and Petr Tichý. Analysis of the multiplicative Schwarz method for matrices with a special block structure. Electronic Transactions on Numerical Analysis (ETNA), 54:31–50, 2021. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.54.2021/pp31-
50.dir/pp31-50.pdf;
http://etna.mcs.kent.
edu/volumes/2021-2030/
vol54/abstract.php?vol=
54&pages=31-50.

Emans:2010:BAA

[Ema10] Maximilian Emans. Benchmarking aggregation AMG for linear systems in CFD simulations of compressible internal flows. Electronic Transactions on Numerical Analysis (ETNA), 37: 351–366, 2010. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic).

Eslahchi:2016:IFT

[EMJ16] M. R. Eslahchi and Mohammad Masjed-Jamei. On qinterpolation formulae and their applications. Electronic Transactions on Numerical Analysis (ETNA), 45:58–74, 2016. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.45.2016/pp58-
74.dir/pp58-74.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol45/abstract.php?vol=
45&pages=58-74.

Elbouyahyaoui:2008:APB

[EMS09] L. Elbouyahyaoui, A. Messaoudi, and H. Sadok. Algebraic properties of the block GMRES and block Arnoldi methods. Electronic Transactions on Numerical Analysis (ETNA), 33:207–220, 2008–2009. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
33.2008-2009/pp207-220.
dir/pp207-220.pdf.

El-Moallem:2013:VEA

[EMS13] Rola El-Moallem and Hassane Sadok. Vector extrapolation applied to algebraic Riccati equations arising in transport theory. Electronic Transactions on Numerical Analysis (ETNA), 40: 489–506, 2013. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu//vol.
                                        40.2013/pp489-506.dir/
                                        pp489-506.pdf.

Erlangga:2008:MKM

[EN08] Yogi A. Erlangga and Reinhard Nabben. On a multilevel Krylov method for the Helmholtz equation preconditioned by shifted Laplacian. Electronic Transactions on Numerical Analysis (ETNA), 31:403–424, 2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

                                        /etna.mcs.kent.edu/vol.
                                        31.2008/pp403-424.dir/
                                        pp403-424.pdf.

Elfving:2010:SCR

[ENH10] Tommy Elfving, Touraj Nikazad, and Per Christian Hansen. Semi-convergence and relaxation parameters for a class of SIRT algorithms. Electronic Transactions on Numerical Analysis (ETNA), 37:321–336, 2010. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic).

Erbe:2007:RRC

[EP07] Lynn Erbe and Allan Peterson. Recent results concerning dynamic equations on time scales. Electronic Transactions on Numerical Analysis (ETNA), 27:51–70, 2007. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.27.2007/pp51-
70.dir/pp51-70.pdf.

Ehrich:2001:PLW

[ER01] S. Ehrich and A. Rathsfeld.

Piecewise linear wavelet collocation, approximation of the boundary manifold, and quadrature. Electronic Transactions on Numerical Analysis (ETNA), 12: 149–192, 2001. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
12.2001/pp149-192.dir/
pp149-192.pdf.

Ehrhardt:2014:ICT

[ER14] Torsten Ehrhardt and Karla Rost. Inversion of centrosymmetric Toeplitz-plus-Hankel Bezoutians. Electronic Transactions on Numerical Analysis (ETNA), 42:106–135, 2014. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.42.2014/pp106-
135.dir/pp106-135.pdf;
 http://etna.mcs.kent.

                                        edu/volumes/2011-2020/
                                        vol42/abstract.php?vol=
                                        42&pages=106-135.

Erhel:1995:PGV

[Erh95] Jocelyne Erhel. A parallel GMRES version for general sparse matrices. Electronic Transactions on Numerical Analysis (ETNA), 3:160–176, 1995. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        3.1995/pp160-176.dir/
                                        pp160-176.pdf.

Eshghi:2017:EMF

[ERS17] Nasim Eshghi, Lothar Reichel, and Miodrag M. Spalević. Enhanced matrix function approximation. Electronic Transactions on Numerical Analysis (ETNA), 47:197–205, 2017. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.47.2017/pp197-
                                        205.dir/pp197-205.pdf;
                                         http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol47/abstract.php?vol=
                                        47&pages=197-205.

Escribano:2006:MMS

[EST06] C. Escribano, M. A. Sastre, and E. Torrano. Moment matrix of self-similar measures. Electronic Transactions on Numerical Analysis (ETNA), 24:79–87, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.24.2006/pp79-
87.dir/pp79-87.pdf.

Elman:2009:BCA

[ET09] Howard C. Elman and Ray S. Tuminaro. Boundary conditions in approximate commutator preconditioners for the Navier– Stokes equations. Electronic Transactions on Numerical Analysis (ETNA), 35:257–280, 2009. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.35.2009/pp257-
280.dir/pp257-280.pdf;
 http://etna.mcs.kent.
edu/volumes/2001-2010/
vol35/abstract.php?vol=
35&pages=257-280.

Eiermann:1993:ZLE

[EV93] Michael Eiermann and Richard S. Varga. Zeros and local extreme points of Faber polynomials associated with hypocycloidal domains. Electronic Transactions on Numerical Analysis (ETNA), 1:49–71, 1993. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.

                                        edu/vol.1.1993/pp49-71.
                                        dir/pp49-71.pdf.

Fassbender:1996:NNA

[Faß96] Heike Faßbender. A note on Newbery’s algorithm for discrete least-squares approximation by trigonometric polynomials. Electronic Transactions on Numerical Analysis (ETNA), 4:64–74, 1996. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.4.1996/pp64-74.
                                        dir/pp64-74.pdf.

Fassbender:2007:PSA

[Faß07]  H. Faßbender.  The

parametrized SR algorithm for Hamiltonian matrices. Electronic Transactions on Numerical Analysis (ETNA), 26:121–145, 2007. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

                                        /etna.mcs.kent.edu/vol.
                                        26.2007/pp121-145.dir/
                                        pp121-145.pdf.

Fattebert:1998:BRQ

[Fat98]  Jean-Luc Fattebert.  A

block Rayleigh quotient iteration with local quadratic convergence. Electronic Transactions on Numerical Analysis (ETNA), 7:56–74, 1998. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.

edu/vol.7.1998/pp56-74.
dir/pp56-74.pdf. Large

scale eigenvalue problems (Argonne, IL, 1997).

Fairag:2021:ACM

[FCA21] Faisal Fairag, Ke Chen, and Shahbaz Ahmad. Analysis of the CCFD method for MC-based image denoising problems. Electronic Transactions on Numerical Analysis (ETNA), 54:108–127, 2021. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.54.2021/pp108-
127.dir/pp108-127.pdf;
http://etna.mcs.kent.
edu/volumes/2021-2030/
vol54/abstract.php?vol=
54&pages=108-127.

Fernandez:2005:LPB

[Fer05] Noemí Laín Fernández. Localized polynomial bases on the sphere. Electronic Transactions on Numerical Analysis (ETNA), 19:84–93, 2005. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.19.2005/pp84-
93.dir/pp84-93.pdf.

Feischl:2015:QOC

[FFK^{+}15] Michael Feischl, Thomas Führer, Michael Karkulik, J. Markus Melenk, and Dirk Praetorius. Quasi-optimal convergence rates for adaptive boundary element methods with data approximation. Part II: Hypersingular integral equation. Electronic Transactions on Numerical Analysis (ETNA), 44:153–176, 2015. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.44.2015/pp153-
176.dir/pp153-176.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol44/abstract.php?vol=
44&pages=153-176.

Friedlander:2017:EES

[FG17] Michael P. Friedlander and Gabriel Goh. Efficient evaluation of scaled proximal operators. Electronic Transactions on Numerical Analysis (ETNA), 46:1–22, 2017. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
46.2017/pp1-22.dir/pp1-
22.pdf;   http://etna.
mcs.kent.edu/volumes/
2011-2020/vol46/abstract.
php?vol=46&pages=1-22.

Frederix:2009:FRS

[FGV09] K. Frederix, L. Gemignani, and M. Van Barel. On the fast reduction of symmetric rationally generated Toeplitz matrices to tridiagonal form. Electronic Transactions on Numerical Analysis (ETNA), 35: 129–147, 2009. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
35.2009/pp129-147.dir/
pp129-147.pdf.

Frommer:2013:VSA

[FH13] Andreas Frommer and Behnam Hashemi. Verified stability analysis using the Lyapunov matrix equation. Electronic Transactions on Numerical Analysis (ETNA), 40:187–203, 2013. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu//vol.
40.2013/pp187-203.dir/
pp187-203.pdf.

Fasi:2020:SAR

[FI20] Massimiliano Fasi and Bruno Iannazzo. Substitution algorithms for rational matrix equations. Electronic Transactions on Numerical Analysis (ETNA), 53:500–521, 2020. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.53.2020/pp500-
521.dir/pp500-521.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/

                                        vol53/abstract.php?vol=
                                        53&pages=500-521.

Frommer:2020:KTM

[FJK^{+}20] Andreas Frommer, Brigit Jacob, Kartsen Kahl, Christian Wyss, and Ian Zwaan. Krylov type methods for linear systems exploiting properties of the quadratic numerical range. Electronic Transactions on Numerical Analysis (ETNA), 53:541–561, 2020. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.53.2020/pp541-
                                        561.dir/pp541-561.pdf;
                                         http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol53/abstract.php?vol=
                                        53&pages=541-561.

Fedotov:2006:AAV

[FJMS06] Igor Fedotov, Steve Joubert, Julian Marais, and Michael Shatalov. Another approach to vibration analysis of stepped structures. Electronic Transactions on Numerical Analysis (ETNA), 24:66–73, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.24.2006/pp66-
                                        73.dir/pp66-73.pdf.

Feistauer:2008:ADN

[FK08] Miloslav Feistauer and Václav Kučera. Analysis of the DGFEM for nonlinear convection-diffusion problems. Electronic Transactions on Numerical Analysis (ETNA), 32:33–48, 2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.32.2008/pp33-
48.dir/pp33-48.pdf.

Farago:2009:DMP

[FKK10] István Faragó, János Karátson, and Sergey Korotov. Discrete maximum principles for the FEM solution of some nonlinear parabolic problems. Electronic Transactions on Numerical Analysis (ETNA), 36: 149–167, 2009–2010. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
36.2009-2010/pp149-167.
dir/pp149-167.pdf.

Fontes:2007:SVD

[FKSV07] Natacha Fontes, Janice Kover, Laura Smithies, and Richard S. Varga. Singular value decomposition normally estimated Geršgorin sets. Electronic Transactions on Numerical Analysis (ETNA), 26:320–329, 2007. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.

                                        26.2007/pp320-329.dir/
                                        pp320-329.pdf.

Ford:2007:TNM

[FL07] Neville J. Ford and Patricia M. Lumb. Theory and numerics for multi-term periodic delay differential equations: small solutions and their detection. Electronic Transactions on Numerical Analysis (ETNA), 26:474–483, 2007. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        26.2007/pp474-483.dir/
                                        pp474-483.pdf.

Ferreira:2005:AAI

[FLMT05] Chelo Ferreira, José L. López, Esmeralda Mainar, and Nico M. Temme. Asymptotic approximations of integrals: an introduction, with recent developments and applications to orthogonal polynomials. Electronic Transactions on Numerical Analysis (ETNA), 19:58–83, 2005. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.19.2005/pp58-
                                        83.dir/pp58-83.pdf.

Frommer:2017:BKS

[FLS17] Andreas Frommer, Kathryn Lund, and Daniel B. Szyld. Block Krylov subspace methods for functions of matrices. Electronic Transactions on Numerical Analysis (ETNA), 47:100–126, 2017. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.47.2017/pp100-
126.dir/pp100-126.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol47/abstract.php?vol=
47&pages=100-126.

Ferreira:2020:SIH

[FLS20] Chelo Ferreira, José L.

López, and Ester Pérez Sinusía. The swallowtail integral in the highly oscillatory region II. Electronic Transactions on Numerical Analysis (ETNA), 52:88–99, 2020. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.52.2020/pp88-
99.dir/pp88-99.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol52/abstract.php?vol=
52&pages=88-99.

Freeden:2002:MDS

[FM02] W. Freeden and T. Maier.

On multiscale denoising of spherical functions: basic theory and numerical aspects. Electronic Transactions on Numerical Analysis (ETNA), 14:56–78, 2002. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.14.2002/pp56-
                                        78.dir/pp56-78.pdf.

Fassbender:2008:SPE

[FMMS08] Heike Fassbender, D. Steven Mackey, Niloufer Mackey, and Christian Schröder. Structured polynomial eigenproblems related to timedelay systems. Electronic Transactions on Numerical Analysis (ETNA), 31: 306–330, 2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        31.2008/pp306-330.dir/
                                        pp306-330.pdf.

Falgout:2021:MRT

[FMOS21] R. D. Falgout, T. A. Manteuffel, B. O’Neill, and J. B. Schroder. Multigrid reduction in time with Richardson extrapolation. Electronic Transactions on Numerical Analysis (ETNA), 54:210–233, 2021. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.54.2021/pp210-
                                        233.dir/pp210-233.pdf;
                                         http://etna.mcs.kent.
                                        edu/volumes/2021-2030/
                                        vol54/abstract.php?vol=
                                        54&pages=210-233.

Fuster:1996:NSP

[FMP96] Robert Fuster, Violeta Migallón, and José Penadés. Non-stationary parallel multisplitting AOR methods. Electronic Transactions on Numerical Analysis (ETNA), 4:1–13, 1996. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
4.1996/pp1-13.dir/pp1-
13.pdf.

Fika:2014:EBF

[FMR15a] Paraskevi Fika, Marilena Mitrouli, and Paraskevi Roupa. Estimates for the bilinear form x^{T} A^{−1}y with applications to linear algebra problems. Electronic Transactions on Numerical Analysis (ETNA), 43: 70–89, 2014–2015. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.43.2014/pp70-
89.dir/pp70-89.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol43/abstract.php?vol=
43&pages=70-89.

Ford:2015:IFD

[FMR15b] N. J. Ford, M. L. Morgado, and M. Rebelo. An implicit finite difference approximation for the solution of the diffusion equation with distributed order in time. Electronic Transactions on Numerical Analysis (ETNA), 44:289–305, 2015. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.44.2015/pp289-
305.dir/pp289-305.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol44/abstract.php?vol=
44&pages=289-305.

Fritzsche:2007:SAI

[FMSV08] David Fritzsche, Volker Mehrmann, Daniel B. Szyld, and Elena Virnik. An SVD approach to identifying metastable states of Markov chains. Electronic Transactions on Numerical Analysis (ETNA), 29: 46–69, 2007–2008. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
29.2007-2008/pp46-69.
dir/pp46-69.pdf.

Farrell:2003:FIP

[FO03] Paul A. Farrell and Hong Ong. Factors involved in the performance of computations on Beowulf clusters. Electronic Transactions on Numerical Analysis (ETNA), 15:211–224, 2003. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.

15.2003/pp211-224.dir/
pp211-224.pdf.

Forster:2002:ANS

[For02] Brigitte Forster. An algorithm for nonharmonic signal analysis using Dirichlet series on convex polygons. Electronic Transactions on Numerical Analysis (ETNA), 14:45–55, 2002. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.14.2002/pp45-
55.dir/pp45-55.pdf.

Fischer:2001:CAP

[FP01] Bernd Fischer and Franz Peherstorfer. Chebyshev approximation via polynomial mappings and the convergence behaviour of Krylov subspace methods. Electronic Transactions on Numerical Analysis (ETNA), 12:205–215, 2001. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
12.2001/pp205-215.dir/
pp205-215.pdf.

Franz:2021:CTI

[Fra21]  Sebastian Franz.  Con-

tinuous time integration for changing type systems. Electronic Transactions on Numerical Analysis (ETNA), 54:198–209, 2021. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.54.2021/pp198-
209.dir/pp198-209.pdf;
 http://etna.mcs.kent.
edu/volumes/2021-2030/
vol54/abstract.php?vol=
54&pages=198-209.

Frommer:2009:MCL

[Fro09] Andreas Frommer. Monotone convergence of the Lanczos approximations to matrix functions of Hermitian matrices. Electronic Transactions on Numerical Analysis (ETNA), 35: 118–128, 2009. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
35.2009/pp118-128.dir/
pp118-128.pdf.

Freitag:2007:CTI

[FS08] Melina A. Freitag and Alastair Spence. Convergence theory for inexact inverse iteration applied to the generalised nonsymmetric eigenproblem. Electronic Transactions on Numerical Analysis (ETNA), 28: 40–64, 2007–2008. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
28.2007-2008/pp40-64.
dir/pp40-64.pdf. Special

volume for Gene Golub.

Frommer:1997:AWA

[FSS97] Andreas Frommer, Hartmut Schwandt, and Daniel B. Szyld. Asynchronous weighted additive Schwarz methods. Electronic Transactions on Numerical Analysis (ETNA), 5:48–61, 1997. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.5.1997/pp48-61.
dir/pp48-61.pdf.

Frommer:2018:NTD

[FSS18] Andreas Frommer, Claudia Schimmel, and Marcel Schweitzer. Non-Toeplitz decay bounds for inverses of Hermitian positive definite tridiagonal matrices. Electronic Transactions on Numerical Analysis (ETNA), 48:362–372, 2018. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.48.2018/pp362-
372.dir/pp362-372.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol48/abstract.php?vol=
48&pages=362-372.

Fung:2020:ASA

[FTRH20] Samy Wu Fung, Sanna Tyrväinen, Lars Ruthotto, and Eldad Haber. ADMM-Softmax: an ADMM approach for multinomial logistic regression. Electronic Transactions on Numerical Analysis (ETNA), 52:214–229, 2020. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.52.2020/pp214-
229.dir/pp214-229.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol52/abstract.php?vol=
52&pages=214-229.

Fulton:1997:CMA

[Ful97] Scott R. Fulton. A comparison of multilevel adaptive methods for hurricane track prediction. Electronic Transactions on Numerical Analysis (ETNA), 6: 120–132, 1997. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
6.1997/pp120-132.dir/
pp120-132.pdf. Special

issue on multilevel methods (Copper Mountain, CO, 1997).

Fulton:2003:AMT

[Ful03]  Scott R. Fulton.  On

the accuracy of multigrid truncation error estimates. Electronic Transactions on Numerical Analysis (ETNA), 15:29–37, 2003. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.

edu/vol.15.2003/pp29-
37.dir/pp29-37.pdf.

Fermo:2014:PEM

[FvdMS14] Luisa Fermo, Cornelis van der Mee, and Sebastiano Seatzu. Parameter estimation of monomialexponential sums. Electronic Transactions on Numerical Analysis (ETNA), 41:249–261, 2014. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.41.2014/pp249-
261.dir/pp249-261.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol41/abstract.php?vol=
41&pages=249-261.

Feng:2019:FFS

[FXG19] Yuehua Feng, Jianwei Xiao, and Ming Gu. Flip-flop spectrum-revealing QR factorization and its applications to singular value decomposition. Electronic Transactions on Numerical Analysis (ETNA), 51: 469–494, 2019. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.51.2019/pp469-
494.dir/pp469-494.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol51/abstract.php?vol=
51&pages=469-494.

Garrido:2005:EIZ

[GAM05] A. Garrido, J. Arvesú, and F. Marcellán. An electrostatic interpretation of the zeros of the Freudtype orthogonal polynomials. Electronic Transactions on Numerical Analysis (ETNA), 19:37–47, 2005. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.19.2005/pp37-
                                        47.dir/pp37-47.pdf.

Gander:2004:FDW

[Gan04]  Martin J. Gander.  A

frequency decomposition waveform relaxation algorithm for semilinear evolution equations. Electronic Transactions on Numerical Analysis (ETNA), 17: 181–194, 2004. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        17.2004/pp181-194.dir/
                                        pp181-194.pdf.

Gander:2008:SMC

[Gan08] Martin J. Gander. Schwarz methods over the course of time. Electronic Transactions on Numerical Analysis (ETNA), 31:228–255, 2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

                                        /etna.mcs.kent.edu/vol.

31.2008/pp228-255.dir/
pp228-255.pdf.

Gatica:2004:NER

[Gat04] Gabriel N. Gatica. A note on the efficiency of residualbased a-posteriori error estimators for some mixed finite element methods. Electronic Transactions on Numerical Analysis (ETNA), 17:218–233, 2004. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
17.2004/pp218-233.dir/
pp218-233.pdf.

Gatica:2007:AMF

[Gat07] Gabriel N. Gatica. An augmented mixed finite element method for linear elasticity with non-homogeneous Dirichlet conditions. Electronic Transactions on Numerical Analysis (ETNA), 26:421–438, 2007. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
26.2007/pp421-438.dir/
pp421-438.pdf.

Gautschi:1999:OPQ

[Gau99]  Walter Gautschi.  Or-

thogonal polynomials and quadrature. Electronic Transactions on Numerical Analysis (ETNA), 9:65–76, 1999. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.9.1999/pp65-76.
                                        dir/pp65-76.pdf.

Gautschi:2002:IBC

[Gau02] Walter Gautschi. The interplay between classical analysis and (numerical) linear algebra — a tribute to Gene H. Golub. Electronic Transactions on Numerical Analysis (ETNA), 13:119–147, 2002. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        13.2002/pp119-147.dir/
                                        pp119-147.pdf.

Gautschi:2006:CTR

[Gau06] Walter Gautschi. The circle theorem and related theorems for Gauss-type quadrature rules. Electronic Transactions on Numerical Analysis (ETNA), 25: 129–137, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        25.2006/pp129-137.dir/
                                        pp129-137.pdf.

Gautschi:2009:HSB

[Gau10]  Walter Gautschi.  How

sharp is Bernstein’s inequality for Jacobi polynomials? Electronic Transactions on Numerical Analysis (ETNA), 36:1–8, 2009– 2010. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.36.2009-2010/
pp1-8.dir/pp1-8.pdf;
http://etna.mcs.kent.
edu/volumes/2001-2010/
vol36/abstract.php?vol=
36&pages=1-8.

Garcia-Caballero:1999:SOP

[GCPP99] Esther M. García-Caballero, Teresa E. Pérez, and Miguel A. Piñar. Sobolev orthogonal polynomials: interpolation and approximation. Electronic Transactions on Numerical Analysis (ETNA), 9:56–64, 1999. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.9.1999/pp56-64.
dir/pp56-64.pdf.

Gerth:2017:CRR

[Ger17] Daniel Gerth. Convergence rates for l_{1}-regularization without the help of a variational inequality. Electronic Transactions on Numerical Analysis (ETNA), 46:233–244, 2017. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.46.2017/pp233-
244.dir/pp233-244.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/

                                        vol46/abstract.php?vol=
                                        46&pages=233-244.

Garrido:2008:HCW

[GGCT09] Ricard Garrido, Pedro González-Casanova, and Elvira Torondel. Hybrid compact-WENO scheme for solving relativistic flows. Electronic Transactions on Numerical Analysis (ETNA), 34:102–118, 2008–2009. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

                                        /etna.mcs.kent.edu/vol.
                                        34.2008-2009/pp102-118.
                                        dir/pp102-118.pdf.

Gerardo-Giorda:2012:SAB

[GGM12] Luca Gerardo-Giorda and Lucia Mirabella. Spectral analysis of a blocktriangular preconditioner for the Bidomain system in electrocardiology. Electronic Transactions on Numerical Analysis (ETNA), 39:186–201, 2012. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        39.2012/pp186-201.dir/.

Gutknecht:1994:LAL

[GH94] Martin H. Gutknecht and Marlis Hochbruck. Lookahead Levinson- and Schurtype recurrences in the Padé table. Electronic Transactions on Numerical Analysis (ETNA), 2:104–129, 1994.

CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
2.1994/pp104-129.dir/
pp104-129.pdf.

Gradinaru:1999:WEP

[GH99] V. Gradinaru and R. Hiptmair. Whitney elements on pyramids. Electronic Transactions on Numerical Analysis (ETNA), 8: 154–168, 1999. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
8.1999/pp154-168.dir/
pp154-168.pdf.

Gunnel:2014:NPS

[GHS14] Andreas Günnel, Roland Herzog, and Ekkehard Sachs. A note on preconditioners and scalar products in Krylov subspace methods for selfadjoint problems in Hilbert space. Electronic Transactions on Numerical Analysis (ETNA), 41:13–20, 2014. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.41.2014/pp13-
20.dir/pp13-20.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol41/abstract.php?vol=
41&pages=13-20.

Gratton:2013:EBE

[GJV13] Serge Gratton, Pavel Jiránek, and Xavier Vasseur. Energy backward error: interpretation in numerical solution of elliptic partial differential equations and behaviour in the conjugate gradient method. Electronic Transactions on Numerical Analysis (ETNA), 40: 338–355, 2013. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu//vol.
                                        40.2013/pp338-355.dir/
                                        pp338-355.pdf.

Graf:2008:SRS

[GK08] Manuel Gräf and Stefan Kunis. Stability results for scattered data interpolation on the rotation group. Electronic Transactions on Numerical Analysis (ETNA), 31:30–39, 2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.31.2008/pp30-
                                        39.dir/pp30-39.pdf.

Guttel:2014:EHO

[GK14] Stefan Güttel and Georges Klein. Efficient high-order rational integration and deferred correction with equispaced data. Electronic Transactions on Numerical Analysis (ETNA), 41:443–464, 2014. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.41.2014/pp443-
464.dir/pp443-464.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol41/abstract.php?vol=
41&pages=443-464.

Gander:2016:DNN

[GKM16] Martin J. Gander, Felix Kwok, and Bankim C. Mandal. Dirichlet–Neumann and Neumann–Neumann waveform relaxation algorithms for parabolic problems. Electronic Transactions on Numerical Analysis (ETNA), 45:424–456, 2016. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.45.2016/pp424-
456.dir/pp424-456.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol45/abstract.php?vol=
45&pages=424-456.

Glader:2008:MDR

[Gla08] Christer Glader. Minimal degree rational unimodular interpolation on the unit circle. Electronic Transactions on Numerical Analysis (ETNA), 30:88–106, 2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

                                        /etna.mcs.kent.edu/vol.
                                        30.2008/pp88-106.dir/
                                        pp88-106.pdf.

Grubisic:2021:MDM

[GLMT21] Luka Grubišić, Matko Ljulj, Volker Mehrmann, and Josip Tambača. Modeling and discretization methods for the numerical simulation of elastic frame structures. Electronic Transactions on Numerical Analysis (ETNA), 54:1–30, 2021. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

                                        /etna.mcs.kent.edu/vol.
                                        54.2021/pp1-30.dir/pp1-
                                        30.pdf;   http://etna.
                                        mcs.kent.edu/volumes/
                                        2021-2030/vol54/abstract.
                                        php?vol=54&pages=1-30.

Gu:2004:SRR

[GM04] M. Gu and L. Miranian. Strong rank revealing Cholesky factorization. Electronic Transactions on Numerical Analysis (ETNA), 17:76–92, 2004. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.17.2004/pp76-
                                        92.dir/pp76-92.pdf.

Ganesh:2006:QFQ

[GM06] M. Ganesh and H. N. Mhaskar. Quadrature-free quasi-interpolation on the sphere. Electronic Transactions on Numerical Analysis (ETNA), 25:101–114, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
25.2006/pp101-114.dir/
pp101-114.pdf.

Gautschi:2018:BTP

[GM18] Walter Gautschi and Gradimir V.

Milovanović. Binet-type polynomials and their zeros. Electronic Transactions on Numerical Analysis (ETNA), 50:52–70, 2018. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.50.2018/pp52-
70.dir/pp52-70.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol50/abstract.php?vol=
50&pages=52-70.

Grubisic:2016:DAT

[GMT16] Luka Grubišić, Suzana Miodragović, and Ninoslav Truhar. Double angle theorems for definite matrix pairs. Electronic Transactions on Numerical Analysis (ETNA), 45:33–57, 2016. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.45.2016/pp33-
57.dir/pp33-57.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/

                                        vol45/abstract.php?vol=
                                        45&pages=33-57.

Gazzola:2013:MPA

[GN13] Silvia Gazzola and Paolo Novati. Multi-parameter Arnoldi–Tikhonov methods. Electronic Transactions on Numerical Analysis (ETNA), 40:452–475, 2013. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu//vol.
                                        40.2013/pp452-475.dir/
                                        pp452-475.pdf.

Gowrisankar:2014:RNS

[GN14] S. Gowrisankar and Srinivasan Natesan. A robust numerical scheme for singularly perturbed delay parabolic initial-boundary-value problems on equidistributed grids. Electronic Transactions on Numerical Analysis (ETNA), 41:376–395, 2014. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.41.2014/pp376-
                                        395.dir/pp376-395.pdf;
                                         http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol41/abstract.php?vol=
                                        41&pages=376-395.

Gossler:2016:AMS

[GN16] Florian Gossler and Reinhard Nabben. On AMG methods with F -smoothing based on Chebyshev polynomials and their relation to AMGr. Electronic Transactions on Numerical Analysis (ETNA), 45:146–159, 2016. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.45.2016/pp146-
159.dir/pp146-159.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol45/abstract.php?vol=
45&pages=146-159.

Gazzola:2015:KPM

[GNR15] Silvia Gazzola, Paolo Novati, and Maria Rosaria Russo. On Krylov projection methods and Tikhonov regularization. Electronic Transactions on Numerical Analysis (ETNA), 44:83–123, 2015. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.44.2015/pp83-
123.dir/pp83-123.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol44/abstract.php?vol=
44&pages=83-123.

Greif:2010:ALR

[GO10] Chen Greif and Michael L.

Overton. An analysis of low-rank modifications of preconditioners for saddle point systems. Electronic Transactions on Numerical Analysis (ETNA), 37: 307–320, 2010. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic).

Gutknecht:2007:PSV

[GOR^{+}08] Martin Gutknecht, Michael Overton, Lothar Reichel, Daniel B. Szyld, Nick Trefethen, Paul Van Dooren, and Andy Wathen. Preface: Special volume dedicated to Gene Golub on the occasion of his 75th birthday. Electronic Transactions on Numerical Analysis (ETNA), 28:vii, 2007– 2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). Special volume for Gene Golub.

Gotz:2006:EPR

[Göt06] Mario Götz. On extremal problems related to inverse balayage. Electronic Transactions on Numerical Analysis (ETNA), 23: 304–319, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        23.2006/pp304-319.dir/
                                        pp304-319.pdf.

Gonnet:2011:RRI

[GPT11] Pedro Gonnet, Ricardo Pachón, and Lloyd N. Trefethen. Robust rational interpolation and leastsquares. Electronic Transactions on Numerical Analysis (ETNA), 38:146–167, 2011. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic).

Graca:2016:MGS

[Gra16] Mário M. Graça. Maps for global separation of roots. Electronic Transactions on Numerical Analysis (ETNA), 45:241–256, 2016. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.45.2016/pp241-
256.dir/pp241-256.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol45/abstract.php?vol=
45&pages=241-256.

Greenbaum:2004:STR

[Gre04] Anne Greenbaum. Some theoretical results derived from polynomial numerical hulls of Jordan blocks. Electronic Transactions on Numerical Analysis (ETNA), 18:81–90, 2004. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.18.2004/pp81-
90.dir/pp81-90.pdf.

Gutierrez:2000:AHF

[GRS00] R. Gutiérrez, J. Rodriguez, and A. J. Sáez. Approximation of hypergeometric functions with matricial argument through their development in series of zonal polynomials. Electronic Transactions on Numerical Analysis (ETNA), 11: 121–130, 2000. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
11.2000/pp121-130.dir/
pp121-130.pdf.

Gortz:1994:RCS

[GS94] Peter Görtz and Rudolf Scherer. Reducibility and characterization of symplectic Runge–Kutta methods. Electronic Transactions on Numerical Analysis (ETNA), 2:194–204, 1994. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
2.1994/pp194-204.dir/
pp194-204.pdf.

Greif:2006:PSP

[GS06] Chen Greif and Dominik Schötzau. Preconditioners for saddle point linear systems with highly singular (1, 1) blocks. Electronic Transactions on Numerical Analysis (ETNA), 22: 114–121, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
22.2006/pp114-121.dir/
pp114-121.pdf.

Galvis:2007:NMM

[GS07] Juan Galvis and Marcus Sarkis. Non-matching mortar discretization analysis for the coupling Stokes– Darcy equations. Electronic Transactions on Numerical Analysis (ETNA), 26: 350–384, 2007. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
26.2007/pp350-384.dir/
pp350-384.pdf.

Gaul:2015:PRK

[GS15] André Gaul and Nico Schlömer. Preconditioned recycling Krylov subspace methods for self-adjoint problems. Electronic Transactions on Numerical Analysis (ETNA), 44:522–547, 2015. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.44.2015/pp522-
547.dir/pp522-547.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol44/abstract.php?vol=
44&pages=522-547.

Gander:2016:CPD

[GS16] Martin J. Gander and Kévin Santugini. Crosspoints in domain decomposition methods with a finite element discretization. Electronic Transactions on Numerical Analysis (ETNA), 45:219–240, 2016. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.45.2016/pp219-
240.dir/pp219-240.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol45/abstract.php?vol=
45&pages=219-240.

Golbabai:2012:CMB

[GSV12] Ahmad Golbabai and Ali Safdari-Vaighani. Collocation methods based on radial basis functions for the coupled Klein– Gordon–Schrödinger equations. Electronic Transactions on Numerical Analysis (ETNA), 39:22–31, 2012. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.39.2012/pp22-
31.dir/pp22-31.pdf.

Golub:2008:ASA

[GSW08] Gene H. Golub, Martin Stoll, and Andy Wathen. Approximation of the scattering amplitude and linear systems. Electronic Transactions on Numerical Analysis (ETNA), 31: 178–203, 2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.

31.2008/pp178-203.dir/
pp178-203.pdf.

Gutknecht:2009:IE

[Gut10] Martin H. Gutknecht. IDR explained. Electronic Transactions on Numerical Analysis (ETNA), 36: 126–148, 2009–2010. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
36.2009-2010/pp126-148.
dir/pp126-148.pdf;   http:
//etna.mcs.kent.edu/volumes/
2001-2010/vol36/abstract.
php?vol=36&pages=126-
148.

Gutknecht:2012:SDK

[Gut12] Martin H. Gutknecht. Spectral deflation in Krylov solvers: a theory of coordinate space based methods. Electronic Transactions on Numerical Analysis (ETNA), 39:156–185, 2012. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
39.2012/pp156-185.dir/.

Georgii:2010:SAS

[GW10] Joachim Georgii and Rüdiger Westermann. A streaming approach for sparse matrix products and its application in Galerkin multigrid methods. Electronic Transactions on Numerical Analysis (ETNA), 37:263–275, 2010. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic).

Goddard:2019:NPP

[GW19] Anthony Goddard and Andy Wathen. A note on parallel preconditioning for all-at-once evolutionary PDEs. Electronic Transactions on Numerical Analysis (ETNA), 51:135–150, 2019. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.51.2019/pp135-
                                        150.dir/pp135-150.pdf;
                                         http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol51/abstract.php?vol=
                                        51&pages=135-150.

Haase:2000:PAO

[Haa00] Gundolf Haase. A parallel AMG for overlapping and non-overlapping domain decomposition. Electronic Transactions on Numerical Analysis (ETNA), 10:41–55, 2000. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.10.2000/pp41-
                                        55.dir/pp41-55.pdf. Mul-

tilevel methods (Copper Mountain, CO, 1999).

Hanisch:2006:TLA

[Han06] M. R. Hanisch. Two-level additive Schwarz preconditioners for fourth-order mixed methods. Electronic Transactions on Numerical Analysis (ETNA), 22:1–16, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
22.2006/pp1-16.dir/pp1-
16.pdf.

Hernandez:2005:AEP

[HC05] M. Bello Hernández and J. Mínguez Ceniceros. Asymptotics for extremal polynomials with varying measures. Electronic Transactions on Numerical Analysis (ETNA), 19:29–36, 2005. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.19.2005/pp29-
36.dir/pp29-36.pdf.

Hwang:2006:PFC

[HC06] Feng-Nan Hwang and Xiao-Chuan Cai. Parallel fully coupled Schwarz preconditioners for saddle point problems. Electronic Transactions on Numerical Analysis (ETNA), 22:146–162, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.22.2006/pp146-
162.dir/pp146-162.pdf.

Hench:1994:PQS

[Hen94] J. J. Hench. On the periodic quotient singular value decomposition. Electronic Transactions on Numerical Analysis (ETNA), 2: 138–153, 1994. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        2.1994/pp138-153.dir/
                                        pp138-153.pdf.

Hanke:1993:CLS

[HH93] Martin Hanke and Marlis Hochbruck. A Chebyshevlike semiiteration for inconsistent linear systems. Electronic Transactions on Numerical Analysis (ETNA), 1:89–103, 1993. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        1.1993/pp89-103.dir/pp89-
                                        103.pdf.

Hipp:2014:EIN

[HHO14] David Hipp, Marlis Hochbruck, and Alexander Ostermann. An exponential integrator for non-autonomous parabolic problems. Electronic Transactions on Numerical Analysis (ETNA), 41:497–511, 2014. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.41.2014/pp497-

511.dir/pp497-511.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol41/abstract.php?vol=
41&pages=497-511.

Herkt:2013:CAG

[HHP13] Sabrina Herkt, Michael Hinze, and Rene Pinnau. Convergence analysis of Galerkin POD for linear second order evolution equations. Electronic Transactions on Numerical Analysis (ETNA), 40: 321–337, 2013. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu//vol.
40.2013/pp321-337.dir/
pp321-337.pdf.

Hofmann:2019:SIV

[HHP19] Christopher Hofmann, Bernd Hofmann, and Alois Pichler. Simultaneous identification of volatility and interest rate functions — a twoparameter regularization approach. Electronic Transactions on Numerical Analysis (ETNA), 51:99–117, 2019. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.51.2019/pp99-
117.dir/pp99-117.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol51/abstract.php?vol=
51&pages=99-117.

Hiptmair:1997:MMT

[Hip97] R. Hiptmair. Multigrid method for H(div) in three dimensions. Electronic Transactions on Numerical Analysis (ETNA), 6: 133–152, 1997. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        6.1997/pp133-152.dir/
                                        pp133-152.pdf. Special

issue on multilevel methods (Copper Mountain, CO, 1997).

Hansen:2008:NPR

[HJ08a] Per Christian Hansen and Toke Koldborg Jensen. Noise propagation in regularizing iterations for image deblurring. Electronic Transactions on Numerical Analysis (ETNA), 31: 204–220, 2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        31.2008/pp204-220.dir/
                                        pp204-220.pdf.

Heinrich:2008:NMP

[HJ08b] Bernd Heinrich and Beate Jung. Nitsche mortaring for parabolic initial-boundary value problems. Electronic Transactions on Numerical Analysis (ETNA), 32: 190–209, 2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
32.2008/pp190-209.dir/
pp190-209.pdf.

Heyouni:2008:EBA

[HJ09] M. Heyouni and K. Jbilou.

An extended block Arnoldi algorithm for large-scale solutions of the continuoustime algebraic Riccati equation. Electronic Transactions on Numerical Analysis (ETNA), 33:53–62, 2008– 2009. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.33.2008-2009/
pp53-62.dir/pp53-62.pdf.

Hessari:2016:LSS

[HJ16] Peyman Hessari and Bongsoo Jang. Least squares spectral method for velocityflux form of the coupled Stokes–Darcy equations. Electronic Transactions on Numerical Analysis (ETNA), 45:160–182, 2016. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.45.2016/pp160-
182.dir/pp160-182.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol45/abstract.php?vol=
45&pages=160-182.

Hassan:2018:PSC

[HJV18] Sarah Ali Hassan, Caroline Japhet, and Martin Vohralík. A posteriori stopping criteria for space-time domain decomposition for the heat equation in mixed formulations. Electronic Transactions on Numerical Analysis (ETNA), 49:151–181, 2018. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.49.2018/pp151-
                                        181.dir/pp151-181.pdf;
                                         http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol49/abstract.php?vol=
                                        49&pages=151-181.

Hyon:2010:NQM

[HK10] Yunkyong Hyon and Do Y. Kwak. New quadrilateral mixed finite elements. Electronic Transactions on Numerical Analysis (ETNA), 37:189–201, 2010. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        37.2010/pp189-201.dir/
                                        pp189-201.pdf.

Hetmaniuk:2014:EET

[HK14] Ulrich Hetmaniuk and Axel Klawonn. Error estimates for a two-dimensional special finite element method based on component mode synthesis. Electronic Transactions on Numerical Analysis (ETNA), 41:109–132, 2014. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.41.2014/pp109-
132.dir/pp109-132.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol41/abstract.php?vol=
41&pages=109-132.

Hari:2017:CCQ

[HK17] Vjeran Hari and Erna Begović Kovač. Convergence of the cyclic and quasicyclic block Jacobi methods. Electronic Transactions on Numerical Analysis (ETNA), 46:107–147, 2017. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.46.2017/pp107-
147.dir/pp107-147.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol46/abstract.php?vol=
46&pages=107-147.

Hong:2018:PRS

[HK18] Qingguo Hong and Johannes Kraus. Parameterrobust stability of classical three-field formulation of Biot’s consolidation model. Electronic Transactions on Numerical Analysis (ETNA), 48:202–226, 2018. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.48.2018/pp202-
                                        226.dir/pp202-226.pdf;
                                         http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol48/abstract.php?vol=
                                        48&pages=202-226.

Heinlein:2018:MCS

[HKKR18] Alexander Heinlein, Axel Klawonn, Jascha Knepper, and Oliver Rheinbach. Multiscale coarse spaces for overlapping Schwarz methods based on the ACMS space in 2D. Electronic Transactions on Numerical Analysis (ETNA), 48: 156–182, 2018. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.48.2018/pp156-
                                        182.dir/pp156-182.pdf;
                                         http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol48/abstract.php?vol=
                                        48&pages=156-182.

Heinlein:2020:FFDa

[HKLW20a] Alexander Heinlein, Axel Klawonn, Martin Lanser, and Janine Weber. A frugal FETI-DP and BDDC coarse space for heterogeneous problems. Electronic Transactions on Numerical Analysis (ETNA), 53:562–591, 2020. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.53.2020/pp562-
591.dir/pp562-591.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol53/abstract.php?vol=
53&pages=562-591.

Heinlein:2020:FFDb

[HKLW20b] Alexander Heinlein, Axel Klawonn, Martin Lanser, and Janine Weber. A frugal FETI-DP and BDDC coarse space for heterogeneous problems. Electronic Transactions on Numerical Analysis (ETNA), 53:562–591, 2020. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.53.2020/pp562-
591.dir/pp562-591.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol53/abstract.php?vol=
53&pages=562-591.

Hochbruck:2020:FED

[HL20] Marlis Hochbruck and Jan Leibold. Finite element discretization of semilinear acoustic wave equations with kinetic boundary conditions. Electronic Transactions on Numerical Analysis (ETNA), 53:522–540, 2020. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.53.2020/pp522-
540.dir/pp522-540.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol53/abstract.php?vol=
53&pages=522-540.

Hwang:2010:TLN

[HLC10] Feng-Nan Hwang, Hsin-Lun Lin, and Xiao-Chuan Cai. Two-level nonlinear elimination based preconditioners for inexact Newton methods with application in shocked duct flow calculation. Electronic Transactions on Numerical Analysis (ETNA), 37:239–251, 2010. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
37.2010/pp239-251.dir/
pp239-251.pdf.

Hofer:2018:TMD

[HLNT18] Christoph Hofer, Ulrich Langer, Martin Neumüller, and Ioannis Toulopoulos. Time-multipatch discontinuous Galerkin space-time isogeometric analysis of parabolic evolution problems. Electronic Transactions on Numerical Analysis (ETNA), 49:126–150, 2018. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.49.2018/pp126-
150.dir/pp126-150.pdf;

http://etna.mcs.kent.
edu/volumes/2011-2020/
vol49/abstract.php?vol=
49&pages=126-150.

Hofreither:2010:ANS

[HLP10] Clemens Hofreither, Ulrich Langer, and Clemens Pechstein. Analysis of a non-standard finite element method based on boundary integral operators. Electronic Transactions on Numerical Analysis (ETNA), 37:413–436, 2010. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic).

Hofreither:2015:BBF

[HLP15] Clemens Hofreither, Ulrich Langer, and Clemens Pechstein. BEM-based Finite Element Tearing and Interconnecting Methods. Electronic Transactions on Numerical Analysis (ETNA), 44:230–249, 2015. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.44.2015/pp230-
249.dir/pp230-249.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol44/abstract.php?vol=
44&pages=230-249.

Howell:2014:FEA

[HLX14] Jason Howell, Hyesuk Lee, and Shuhan Xu. Finite element approximation of viscoelastic flow in a moving domain. Electronic Transactions on Numerical Analysis (ETNA), 41:306–327, 2014. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.41.2014/pp306-
327.dir/pp306-327.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol41/abstract.php?vol=
41&pages=306-327.

He:2004:QCB

[HMRS04] C. He, B. Meini, N. H. Rhee, and K. Sohraby. A quadratically convergent Bernoulli-like algorithm for solving matrix polynomial equations in Markov chains. Electronic Transactions on Numerical Analysis (ETNA), 17: 151–167, 2004. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
17.2004/pp151-167.dir/
pp151-167.pdf.

Hochstenbach:2007:HJD

[HN08a] Michiel E. Hochstenbach and Yvan Notay. Homogeneous Jacobi–Davidson. Electronic Transactions on Numerical Analysis (ETNA), 29:19–30, 2007–2008. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
29.2007-2008/pp19-30.
dir/pp19-30.pdf.

Huckle:2007:PBT

[HN08b] Thomas K. Huckle and Dimitrios Noutsos. Preconditioning block Toeplitz matrices. Electronic Transactions on Numerical Analysis (ETNA), 29:31–45, 2007– 2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.29.2007-2008/
pp31-45.dir/pp31-45.pdf.

Han:2009:APB

[HNP10] Lixing Han, Michael Neumann, and Upendra Prasad. Alternating projected Barzilai– Borwein methods for nonnegative matrix factorization. Electronic Transactions on Numerical Analysis (ETNA), 36:54–82, 2009– 2010. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.36.2009-2010/
pp54-82.dir/pp54-82.pdf;
http://etna.mcs.kent.
edu/volumes/2001-2010/
vol36/abstract.php?vol=
36&pages=54-82.

Hegland:2004:MSU

[HNS04] Markus Hegland, Ole M.

Nielsen, and Zuowei Shen. Multidimensional smoothing using hyperbolic interpolatory wavelets. Electronic Transactions on Numerical Analysis (ETNA), 17:168–180, 2004. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
17.2004/pp168-180.dir/
pp168-180.pdf.

Hochstenbach:2005:GHR

[Hoc05] Michiel E. Hochstenbach. Generalizations of harmonic and refined Rayleigh–Ritz. Electronic Transactions on Numerical Analysis (ETNA), 20:235–252, 2005. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
20.2005/pp235-252.dir/
pp235-252.pdf.

Hochstenbach:2011:FVI

[Hoc11] Michiel E. Hochstenbach. Fields of values and inclusion regions for matrix pencils. Electronic Transactions on Numerical Analysis (ETNA), 38:98–112, 2011. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic).

Hohn:2006:SSE

[Hoh06] Michael H. Hohn. Solution of singular elliptic PDEs on a union of rectangles using sinc methods. Electronic Transactions on Numerical Analysis (ETNA), 23:88–104, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
23.2006/pp88-104.dir/
pp88-104.pdf.

Haasdonk:2008:RBM

[HOR08] B. Haasdonk, M. Ohlberger, and G. Rozza. A reduced basis method for evolution schemes with parameterdependent explicit operators. Electronic Transactions on Numerical Analysis (ETNA), 32:145–161, 2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
32.2008/pp145-161.dir/
pp145-161.pdf.

Hochstenbach:2007:HRR

[HP08] Michiel E. Hochstenbach and Bor Plestenjak. Harmonic Rayleigh–Ritz extraction for the multiparameter eigenvalue problem. Electronic Transactions on Numerical Analysis (ETNA), 29:81–96, 2007– 2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.29.2007-2008/
pp81-96.dir/pp81-96.pdf.

Hofmann:2020:CRL

[HP20] Bernd Hofmann and Robert Plato. Convergence results and low-order rates for nonlinear Tikhonov regularization with oversmoothing penalty term. Electronic Transactions on Numerical Analysis (ETNA), 53:313–328, 2020. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.53.2020/pp313-
328.dir/pp313-328.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol53/abstract.php?vol=
53&pages=313-328.

Hur:2006:NCP

[HR06] Youngmi Hur and Amos Ron. New constructions of piecewise-constant wavelets. Electronic Transactions on Numerical Analysis (ETNA), 25:138–157, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.25.2006/pp138-
157.dir/pp138-157.pdf.

Hu:2014:CAF

[HR14] Guanghui Hu and Andreas Rathsfeld. Convergence analysis of the FEM coupled with Fouriermode expansion for the electromagnetic scattering by biperiodic structures. Electronic Transactions on Numerical Analysis (ETNA), 41:350–375, 2014. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.41.2014/pp350-
375.dir/pp350-375.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol41/abstract.php?vol=
41&pages=350-375.

Hernandez:2008:REP

[HRT08] Vicente Hernández, José E.

Román, and Andrés Tomás. A robust and efficient parallel SVD solver based on restarted Lanczos bidiagonalization. Electronic Transactions on Numerical Analysis (ETNA), 31:68–85, 2008. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.31.2008/pp68-
85.dir/pp68-85.pdf.

Hakula:2013:CEM

[HRV13] Harri Hakula, Antti Rasila, and Matti Vuorinen. Computation of exterior moduli of quadrilaterals. Electronic Transactions on Numerical Analysis (ETNA), 40:436–451, 2013. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu//vol.
40.2013/pp436-451.dir/
pp436-451.pdf.

Heuveline:1997:AFM

[HS97] Vincent Heuveline and Miloud Sadkane. Arnoldi– Faber method for large non Hermitian eigenvalue problems. Electronic Transactions on Numerical Analysis (ETNA), 5:62–76, 1997. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.5.1997/pp62-76.
                                        dir/pp62-76.pdf.

Huckle:2002:MPT

[HS02] Thomas Huckle and Jochen Staudacher. Multigrid preconditioning and Toeplitz matrices. Electronic Transactions on Numerical Analysis (ETNA), 13:81–105, 2002. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.13.2002/pp81-
                                        105.dir/pp81-105.pdf.

Hadjidimos:2007:MSN

[HS08a] A. Hadjidimos and P. Stratis. Minimization of the spectral norm of the SOR operator in a mixed case. Electronic Transactions on Numerical Analysis (ETNA), 28:78–94, 2007–2008. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        28.2007-2008/pp78-94.

dir/pp78-94.pdf. Special

volume for Gene Golub.

Hench:2008:RMC

[HS08b] John J. Hench and Zdeněk Strakoš. The RCWA method — a case study with open questions and perspectives of algebraic computations. Electronic Transactions on Numerical Analysis (ETNA), 31:331–357, 2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
31.2008/pp331-357.dir/
pp331-357.pdf.

Henson:2012:LSE

[HS12] Van Emden Henson and Geoffrey Sanders. Locally supported eigenvectors of matrices associated with connected and unweighted power-law graphs. Electronic Transactions on Numerical Analysis (ETNA), 39:353–378, 2012. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
39.2012/pp353-378.dir/.

Imberti:2017:VYS

[IE17] David Imberti and Jocelyne Erhel. Varying the s in your s-step GM-RES. Electronic Transactions on Numerical Analysis (ETNA), 47:206–230, 2017. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.47.2017/pp206-
                                        230.dir/pp206-230.pdf;
                                         http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol47/abstract.php?vol=
                                        47&pages=206-230.

Ismail:2006:OPR

[IL06] Mourad E. H. Ismail and Xin Li. Orthogonal polynomials and Ramanujan’s qcontinued fractions. Electronic Transactions on Numerical Analysis (ETNA), 25:158–165, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        25.2006/pp158-165.dir/
                                        pp158-165.pdf.

Illan:2003:QFR

[Ill03] J. Illán. A quadrature formula of rational type for integrands with one endpoint singularity. Electronic Transactions on Numerical Analysis (ETNA), 16: 143–164, 2003. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        16.2003/pp143-164.dir/
                                        pp143-164.pdf.

Iwasaki:2011:PDM

[IN11] Masashi Iwasaki and Yoshimasa Nakamura. Positivity of dLV and mdLVs algorithms for computing singular values. Electronic Transactions on Numerical Analysis (ETNA), 38:184–201, 2011. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic).

Itai:2006:EBK

[Ita06] Uri Itai. On the eigenstructure of the Bernstein kernel. Electronic Transactions on Numerical Analysis (ETNA), 25:431–438, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
25.2006/pp431-438.dir/
pp431-438.pdf.

Jacobi:2010:RSM

[Jac10] Martin Nilsson Jacobi. A robust spectral method for finding lumpings and meta stable states of nonreversible Markov chains. Electronic Transactions on Numerical Analysis (ETNA), 37:296–306, 2010. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic).

Joubert:2006:SAM

[JC06] Wayne Joubert and Jane Cullum. Scalable algebraic multigrid on 3500 processors. Electronic Transactions on Numerical Analysis (ETNA), 23:105–128, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

                                        /etna.mcs.kent.edu/vol.
                                        23.2006/pp105-128.dir/
                                        pp105-128.pdf.

Jumaat:2017:OBM

[JC17] Abdul K. Jumaat and Ke Chen. An optimizationbased multilevel algorithm for variational image segmentation models. Electronic Transactions on Numerical Analysis (ETNA), 46:474–504, 2017. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.46.2017/pp474-
                                        504.dir/pp474-504.pdf;
                                         http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol46/abstract.php?vol=
                                        46&pages=474-504.

Jarlebring:2014:SAI

[JG14] Elias Jarlebring and Stefan Güttel. A spatially adaptive iterative method for a class of nonlinear operator eigenproblems. Electronic Transactions on Numerical Analysis (ETNA), 41:21–41, 2014. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu//vol.41.2014/pp21-
                                        41.dir/pp21-41.pdf.

Janovska:2010:ASN

[JJ10] Dáša Janovská and Vladimír Janovský. The analytic SVD: On the non-generic points on the path. Electronic Transactions on Numerical Analysis (ETNA), 37:70–86, 2010. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.37.2010/pp70-
86.dir/pp70-86.pdf.

Junghanns:2014:CSI

[JKM14] Peter Junghanns, Robert Kaiser, and Giuseppe Mastroianni. Collocation for singular integral equations with fixed singularities of particular Mellin type. Electronic Transactions on Numerical Analysis (ETNA), 41:190–248, 2014. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.41.2014/pp190-
248.dir/pp190-248.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol41/abstract.php?vol=
41&pages=190-248.

Jay:2007:EHM

[JN07] Laurent O. Jay and Dan Negrut. Extensions of the HHT-α method to differential-algebraic equations in mechanics. Electronic Transactions on Numerical Analysis (ETNA), 26:190–208, 2007. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        26.2007/pp190-208.dir/
                                        pp190-208.pdf.

Janson:2009:ZSB

[JN10] Svante Janson and Timothy S. Norfolk. Zeros of sections of the binomial expansion. Electronic Transactions on Numerical Analysis (ETNA), 36: 27–38, 2009–2010. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        36.2009-2010/pp27-38.
                                        dir/pp27-38.pdf;   http:
                                        //etna.mcs.kent.edu/volumes/
                                        2001-2010/vol36/abstract.
                                        php?vol=36&pages=27-38.

Janovska:2005:FGT

[JO05] Drahoslava Janovská and Gerhard Opfer. Fast Givens transformation for quaternion valued matrices applied to Hessenberg reductions. Electronic Transactions on Numerical Analysis (ETNA), 20:1–26, 2005. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

                                        /etna.mcs.kent.edu/vol.
                                        20.2005/pp1-26.dir/pp1-
                                        26.pdf.

Janovska:2007:CQR

[JO07] Drahoslava Janovská and Gerhard Opfer. Computing quaternionic roots by Newton’s method. Electronic Transactions on Numerical Analysis (ETNA), 26:82–102, 2007. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
26.2007/pp82-102.dir/
pp82-102.pdf.

Jung:2007:IIP

[JO08] Jin Hyuk Jung and Dianne P. O’Leary. Implementing an interior point method for linear programs on a CPU-GPU system. Electronic Transactions on Numerical Analysis (ETNA), 28: 174–189, 2007–2008. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
28.2007-2008/pp174-189.
dir/pp174-189.pdf. Spe-

cial volume for Gene Golub. Janovska:2014:ZSP

[JO14] Drahoslava Janovská and Gerhard Opfer. Zeros and singular points for onesided coquaternionic polynomials with an extension to other R^{4} algebras. Electronic Transactions on Numerical Analysis (ETNA), 41:133–158, 2014. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.41.2014/pp133-
                                        158.dir/pp133-158.pdf;
                                         http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol41/abstract.php?vol=
                                        41&pages=133-158.

Janovska:2017:NZU

[JO17] Drahoslava Janovská and Gerhard Opfer. The number of zeros of unilateral polynomials over coquaternions and related algebras. Electronic Transactions on Numerical Analysis (ETNA), 46:55–70, 2017. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.46.2017/pp55-
                                        70.dir/pp55-70.pdf;
                                        http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol46/abstract.php?vol=
                                        46&pages=55-70.

Jung:2008:ACR

[JOT08] Jin Hyuk Jung, Dianne P. O’Leary, and André L. Tits. Adaptive constraint reduction for training support vector machines. Electronic Transactions on Numerical Analysis (ETNA), 31: 156–177, 2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        31.2008/pp156-177.dir/
                                        pp156-177.pdf.

Junghanns:2002:PCM

[JR02] P. Junghanns and A. Rathsfeld. A polynomial collocation method for Cauchy singular integral equations over the interval. Electronic Transactions on Numerical Analysis (ETNA), 14:79–126, 2002. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
14.2002/pp79-126.dir/
pp79-126.pdf.

Junghanns:2004:CMC

[JR04] P. Junghanns and A. Rogozhin. Collocation methods for Cauchy singular integral equations on the interval. Electronic Transactions on Numerical Analysis (ETNA), 17:11–75, 2004. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.17.2004/pp11-
75.dir/pp11-75.pdf.

Juntunen:2008:MDG

[JS08] Mika Juntunen and Rolf Stenberg. On a mixed discontinuous Galerkin method. Electronic Transactions on Numerical Analysis (ETNA), 32:17–32, 2008. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.

                                        edu/vol.32.2008/pp17-
                                        32.dir/pp17-32.pdf.

Jansing:2018:CAE

[JS18] Georg Jansing and Achim Schädle. Convergence analysis of an explicit splitting method for laser plasma interaction simulations. Electronic Transactions on Numerical Analysis (ETNA), 48:243–263, 2018. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.48.2018/pp243-
                                        263.dir/pp243-263.pdf;
                                         http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol48/abstract.php?vol=
                                        48&pages=243-263.

Jbilou:2005:OPM

[JST05] K. Jbilou, H. Sadok, and A. Tinzefte. Oblique projection methods for linear systems with multiple right-hand sides. Electronic Transactions on Numerical Analysis (ETNA), 20: 119–138, 2005. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        20.2005/pp119-138.dir/
                                        pp119-138.pdf.

Jovanovic:2018:COS

[JST18] Aleksandar N. Jovanović, Marija P. Stanić, and Tatjana V. Tomović. Construction of the optimal set of quadrature rules in the sense of Borges. Electronic Transactions on Numerical Analysis (ETNA), 50:164–181, 2018. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.50.2018/pp164-
181.dir/pp164-181.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol50/abstract.php?vol=
50&pages=164-181.

Jeuris:2012:SCC

[JVV12] Ben Jeuris, Raf Vandebril, and Bart Vandereycken. A survey and comparison of contemporary algorithms for computing the matrix geometric mean. Electronic Transactions on Numerical Analysis (ETNA), 39: 379–402, 2012. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
39.2012/pp379-402.dir/.

Kalantzis:2020:SNS

[Kal20]  Vassilis Kalantzis.  A

spectral Newton–Schur algorithm for the solution of symmetric generalized eigenvalue problems. Electronic Transactions on Numerical Analysis (ETNA), 52:132–153, 2020. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.52.2020/pp132-
                                        153.dir/pp132-153.pdf;
                                         http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol52/abstract.php?vol=
                                        52&pages=132-153.

Kanda:2008:CMC

[Kan08] Hidesada Kanda. Calculation of minimum critical Reynolds number for laminar-turbulent transition in pipe flows. Electronic Transactions on Numerical Analysis (ETNA), 30:168–186, 2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        30.2008/pp168-186.dir/
                                        pp168-186.pdf.

Kanda:2015:LTT

[Kan15] Hidesada Kanda. Laminarturbulent transition in pipe flow: wall effects and critical Reynolds number. Electronic Transactions on Numerical Analysis (ETNA), 44:548–571, 2015. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.44.2015/pp548-
                                        571.dir/pp548-571.pdf;
                                         http://etna.mcs.kent.
                                        edu/volumes/2011-2020/

vol44/abstract.php?vol=
44&pages=548-571.

Kanda:2019:LTTa

[Kan19a] Hidesada Kanda. Laminarturbulent transition in channel flow: wall effects and critical Reynolds number. Electronic Transactions on Numerical Analysis (ETNA), 51:547–569, 2019. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.51.2019/pp547-
569.dir/pp547-569.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol51/abstract.php?vol=
51&pages=547-569.

Kanda:2019:LTTb

[Kan19b] Hidesada Kanda. Laminarturbulent transition in channel flow: wall effects and critical Reynolds number. Electronic Transactions on Numerical Analysis (ETNA), 51:547–569, 2019. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.51.2019/pp547-
569.dir/pp547-569.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol51/abstract.php?vol=
51&pages=547-569.

Kim:2018:BFD

[KCW18] Hyea Hyun Kim, Eric Chung, and Junxian Wang. BDDC and FETI-DP algorithms with a change of basis formulation on adaptive primal constraints. Electronic Transactions on Numerical Analysis (ETNA), 49:64–80, 2018. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.49.2018/pp64-
80.dir/pp64-80.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol49/abstract.php?vol=
49&pages=64-80.

Keiner:2008:GPS

[Kei08] Jens Keiner. Gegenbauer polynomials and semiseparable matrices. Electronic Transactions on Numerical Analysis (ETNA), 30:26–53, 2008. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.30.2008/pp26-
53.dir/pp26-53.pdf.

Khatamov:2006:EEB

[Kha06] Akhtam Khatamov. On the exact estimates of the best spline approximations of functions. Electronic Transactions on Numerical Analysis (ETNA), 25: 446–453, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
25.2006/pp446-453.dir/
pp446-453.pdf.

Khrushchev:2006:EDM

[Khr06] Sergey Khrushchev. On Euler’s differential methods for continued fractions. Electronic Transactions on Numerical Analysis (ETNA), 25:178–200, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
25.2006/pp178-200.dir/
pp178-200.pdf.

Kim:2007:FDP

[Kim07a] Hyea Hyun Kim. A FETI-DP preconditioner for mortar methods in three dimensions. Electronic Transactions on Numerical Analysis (ETNA), 26:103–120, 2007. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
26.2007/pp103-120.dir/
pp103-120.pdf.

Kim:2007:PBP

[Kim07b] Sang Dong Kim. Piecewise bilinear preconditioning of high-order finite element method. Electronic Transactions on Numerical Analysis (ETNA), 26: 228–242, 2007. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
26.2007/pp228-242.dir/
pp228-242.pdf.

Kindermann:2011:CAM

[Kin11] Stefan Kindermann. Convergence analysis of minimizationbased noise level-free parameter choice rules for linear ill-posed problems. Electronic Transactions on Numerical Analysis (ETNA), 38:233–257, 2011. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
38.2011/pp233-257.dir/
pp233-257.pdf.

Kindermann:2013:DIC

[Kin13] Stefan Kindermann. Discretization independent convergence rates for noise level-free parameter choice rules for the regularization of ill-conditioned problems. Electronic Transactions on Numerical Analysis (ETNA), 40:58–81, 2013. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.40.2013/pp58-
81.dir/pp58-81.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol40/abstract.php?vol=
40&pages=58-81.

Kindermann:2009:HTU

[KJ09] Stefan Kindermann and Marcin Janicki. The heat transform and its use in thermal identification problems for electronic circuits. Electronic Transactions on Numerical Analysis (ETNA), 35:164–184, 2009. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
35.2009/pp164-184.dir/
pp164-184.pdf.

Kadalbajoo:2008:PUF

[KK08] Mohan K. Kadalbajoo and Devendra Kumar. Parameter-uniform fitted operator B-spline collocation method for self-adjoint singularly perturbed twopoint boundary value problems. Electronic Transactions on Numerical Analysis (ETNA), 30:346–358, 2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
30.2008/pp346-358.dir/
pp346-358.pdf.

Klawonn:2018:AFD

[KKR18] Axel Klawonn, Martin Kühn, and Oliver Rheinbach. Adaptive FETI-DP and BDDC methods with a generalized transformation of basis for heterogeneous problems. Electronic Transactions on Numerical Analysis (ETNA), 49:1–27, 2018. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
49.2018/pp1-27.dir/pp1-
27.pdf;   http://etna.
mcs.kent.edu/volumes/
2011-2020/vol49/abstract.
php?vol=49&pages=1-27.

Klawonn:2020:CSF

[KKR20] Axel Klawonn, Martin Kühn, and Oliver Rheinbach. Coarse spaces for FETI-DP and BDDC methods for heterogeneous problems: connections of deflation and a generalized transformation-of-basis approach. Electronic Transactions on Numerical Analysis (ETNA), 52:43–76, 2020. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.52.2020/pp43-
76.dir/pp43-76.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol52/abstract.php?vol=
52&pages=43-76.

Kleefeld:2012:BEC

[KL12] Andreas Kleefeld and Tzu-Chu Lin. Boundary element collocation method for solving the exterior Neumann problem for Helmholtz’s equation in three dimensions. Electronic Transactions on Numerical Analysis (ETNA), 39:113–143, 2012. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.39.2012/pp113-
143.dir/pp113-143.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol39/abstract.php?vol=
39&pages=113-143.

Kramer:2018:CIB

[KL18] Lukas Krämer and Bruno Lang. Convergence of integration-based methods for the solution of standard and generalized Hermitian eigenvalue problems. Electronic Transactions on Numerical Analysis (ETNA), 48:183–201, 2018. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.48.2018/pp183-
201.dir/pp183-201.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol48/abstract.php?vol=
48&pages=183-201.

Kahl:2021:HEE

[KL21] Karsten Kahl and Bruno Lang. Hypergraph edge elimination — a symbolic phase for Hermitian eigensolvers based on rank-1 modifications. Electronic Transactions on Numerical Analysis (ETNA), 54:51–67, 2021. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.54.2021/pp51-
67.dir/pp51-67.pdf;
http://etna.mcs.kent.
edu/volumes/2021-2030/
vol54/abstract.php?vol=
54&pages=51-67.

Klein:2007:FOT

[Kle08] Christian Klein. Fourth order time-stepping for low dispersion Korteweg– de Vries and nonlinear Schrödinger equations. Electronic Transactions on Numerical Analysis (ETNA), 29:116–135, 2007–2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
29.2007-2008/pp116-135.
dir/pp116-135.pdf.

Klawonn:2019:PCP

[KLRW19] Axel Klawonn, Martin Lanser, Oliver Rheinbach, and Janine Weber. Preconditioning the coarse problem of BDDC methods — three-level, algebraic multigrid, and vertexbased preconditioners. Electronic Transactions on Numerical Analysis (ETNA), 51:432–450, 2019. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.51.2019/pp432-
450.dir/pp432-450.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol51/abstract.php?vol=
51&pages=432-450.

Kalantzis:2016:SSC

[KLS16] Vassilis Kalantzis, Ruipeng Li, and Yousef Saad. Spectral Schur complement techniques for symmetric eigenvalue problems. Electronic Transactions on Numerical Analysis (ETNA), 45:305–329, 2016. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.45.2016/pp305-
329.dir/pp305-329.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol45/abstract.php?vol=
45&pages=305-329.

Kressner:2017:ICB

[KLS17] Daniel Kressner, Robert Luce, and Francesco Statti. Incremental computation of block triangular matrix exponentials with application to option pricing. Electronic Transactions on Numerical Analysis (ETNA), 47:57–72, 2017. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.47.2017/pp57-

                                        72.dir/pp57-72.pdf;
                                        http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol47/abstract.php?vol=
                                        47&pages=57-72.

Kunkel:1996:LGI

[KM96] Peter Kunkel and Volker Mehrmann. Local and global invariants of linear differential-algebraic equations and their relation. Electronic Transactions on Numerical Analysis (ETNA), 4:138–157, 1996. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

                                        /etna.mcs.kent.edu/vol.
                                        4.1996/pp138-157.dir/
                                        pp138-157.pdf.

Kunkel:2007:SPD

[KM07] Peter Kunkel and Volker Mehrmann. Stability properties of differentialalgebraic equations and spin-stabilized discretizations. Electronic Transactions on Numerical Analysis (ETNA), 26:385–420, 2007. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

                                        /etna.mcs.kent.edu/vol.
                                        26.2007/pp385-420.dir/
                                        pp385-420.pdf.

Kravvaritis:2008:TCM

[KM08] Christos Kravvaritis and Marilena Mitrouli. A technique for computing minors of binary Hadamard matrices and application to the growth problem. Electronic Transactions on Numerical Analysis (ETNA), 31:49–67, 2008. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.31.2008/pp49-
67.dir/pp49-67.pdf.

Kunis:2017:FER

[KM17] Stefan Kunis and Ines Melzer. Fast evaluation of real and complex exponential sums. Electronic Transactions on Numerical Analysis (ETNA), 46:23–35, 2017. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.46.2017/pp23-
35.dir/pp23-35.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol46/abstract.php?vol=
46&pages=23-35.

Kara:2011:CTM

[KMC11] Mohamed Kara, Boubakeur Merouani, and Lahcène Chorfi. Computation of the torsional modes in an axisymmetric elastic layer. Electronic Transactions on Numerical Analysis (ETNA), 38:303–316, 2011. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.

                                        38.2011/pp303-316.dir/
                                        pp303-316.pdf.

Kuijlaars:2005:OJP

[KMFO05] A. B. J. Kuijlaars, A. Martínez-Finkelshtein, and R. Orive. Orthogonality of Jacobi polynomials with general parameters. Electronic Transactions on Numerical Analysis (ETNA), 19: 1–17, 2005. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        19.2005/pp1-17.dir/pp1-
                                        17.pdf.

Knyazev:2003:ESS

[KN03] Andrew V. Knyazev and Klaus Neymeyr. Efficient solution of symmetric eigenvalue problems using multigrid preconditioners in the locally optimal block conjugate gradient method. Electronic Transactions on Numerical Analysis (ETNA), 15:38–55, 2003. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.15.2003/pp38-
                                        55.dir/pp38-55.pdf.

Knobloch:2008:DSP

[Kno08] Petr Knobloch. On the definition of the SUPG parameter. Electronic Transactions on Numerical Analysis (ETNA), 32:76–89, 2008.

CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.32.2008/pp76-
89.dir/pp76-89.pdf.

Kehl:2019:AMK

[KNS19] René Kehl, Reinhard Nabben, and Daniel B. Szyld. Adaptive Multilevel Krylov Methods. Electronic Transactions on Numerical Analysis (ETNA), 51:512–528, 2019. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.51.2019/pp512-
528.dir/pp512-528.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol51/abstract.php?vol=
51&pages=512-528.

Knyazev:1998:PEO

[Kny98] Andrew V. Knyazev. Preconditioned eigensolvers — an oxymoron? Electronic Transactions on Numerical Analysis (ETNA), 7: 104–123, 1998. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
7.1998/pp104-123.dir/
pp104-123.pdf.  Large

scale eigenvalue problems (Argonne, IL, 1997).

Koepf:1999:SAW

[Koe99]  Wolfram Koepf.  Soft-

ware for the algorithmic work with orthogonal polynomials and special functions. Electronic Transactions on Numerical Analysis (ETNA), 9:77–101, 1999. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
9.1999/pp77-101.dir/pp77-
101.pdf.

Koepf:2007:OPR

[Koe07] Wolfram Koepf. Orthogonal polynomials and recurrence equations, operator equations and factorization. Electronic Transactions on Numerical Analysis (ETNA), 27:113–123, 2007. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
27.2007/pp113-123.dir/
pp113-123.pdf.

Kolmbauer:2012:RFB

[Kol12]  Michael Kolmbauer.  A

robust FEM–BEM Min-Res solver for distributed multiharmonic eddy current optimal control problems in unbounded domains. Electronic Transactions on Numerical Analysis (ETNA), 39:231–252, 2012. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.39.2012/pp231-

252.dir/pp231-252.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol39/abstract.php?vol=
39&pages=231-252.

Kouatchou:1997:ASP

[Kou97] Jules Kouatchou. Asymptotic stability of a 9-point multigrid algorithm for convection-diffusion equations. Electronic Transactions on Numerical Analysis (ETNA), 6:153–161, 1997. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
6.1997/pp153-161.dir/
pp153-161.pdf. Special

issue on multilevel methods (Copper Mountain, CO, 1997).

Kamgnia:2013:CED

[KP13] Emmanuel Kamgnia and Bernard Philippe. Counting eigenvalues in domains of the complex field. Electronic Transactions on Numerical Analysis (ETNA), 40:1–16, 2013. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
40.2013/pp1-16.dir/pp1-
16.pdf;   http://etna.
mcs.kent.edu/volumes/
2011-2020/vol40/abstract.
php?vol=40&pages=1-16.

Khanh:2013:CSR

[KQT13] D. C. Khanh, H. T. Quyen, and D. D. X. Thanh. On computing stabilizability radii of linear time-invariant continuous systems. Electronic Transactions on Numerical Analysis (ETNA), 40:407–413, 2013. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu//vol.
                                        40.2013/pp407-413.dir/
                                        pp407-413.pdf.

Kindermann:2020:SCM

[KR20] Stefan Kindermann and Kemal Raik. A simplified Lcurve method as error estimator. Electronic Transactions on Numerical Analysis (ETNA), 53:217–238, 2020. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.53.2020/pp217-
                                        238.dir/pp217-238.pdf;
                                         http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol53/abstract.php?vol=
                                        53&pages=217-238.

Kressner:2005:ULB

[Kre05] Daniel Kressner. On the use of larger bulges in the QR algorithm. Electronic Transactions on Numerical Analysis (ETNA), 20:50–63, 2005. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.20.2005/pp50-
63.dir/pp50-63.pdf.

Kroo:2006:NFM

[Kro06] András Kroó. On norms of factors of multivariate polynomials on convex bodies. Electronic Transactions on Numerical Analysis (ETNA), 25:201–205, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
25.2006/pp201-205.dir/
pp201-205.pdf.

Klawonn:2016:CAC

[KRR16] Axel Klawonn, Patrick Radtke, and Oliver Rheinbach. A comparison of adaptive coarse spaces for iterative substructuring in two dimensions. Electronic Transactions on Numerical Analysis (ETNA), 45:75–106, 2016. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.45.2016/pp75-
106.dir/pp75-106.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol45/abstract.php?vol=
45&pages=75-106.

Krendl:2016:DRB

[KRZ16] Wolfgang Krendl, Katharina Rafetseder, and Walter Zulehner. A decomposition result for biharmonic problems and the Hellan–Herrmann–Johnson method. Electronic Transactions on Numerical Analysis (ETNA), 45:257–282, 2016. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.45.2016/pp257-
282.dir/pp257-282.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol45/abstract.php?vol=
45&pages=257-282.

Korsawe:2000:MPM

[KS00] Johannes Korsawe and Gerhard Starke. Multilevel projection methods for nonlinear least-squares finite element computations. Electronic Transactions on Numerical Analysis (ETNA), 10:56–73, 2000. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.10.2000/pp56-
73.dir/pp56-73.pdf. Mul-

tilevel methods (Copper Mountain, CO, 1999).

Kadalbajoo:2006:PUF

[KS06] M. K. Kadalbajoo and K. K. Sharma. Parameteruniform fitted mesh method for singularly perturbed delay differential equations with layer behavior. Electronic Transactions on Numerical Analysis (ETNA), 23:180–201, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
23.2006/pp180-201.dir/
pp180-201.pdf.

Kucukoglu:2018:NES

[KS18] Irem Kucukoglu and Yilmaz Simsek. Numerical evaluation of special power series including the numbers of Lyndon words: an approach to interpolation functions for Apostol-type numbers and polynomials. Electronic Transactions on Numerical Analysis (ETNA), 50:98–108, 2018. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.50.2018/pp98-
108.dir/pp98-108.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol50/abstract.php?vol=
50&pages=98-108.

Kurschner:2019:ARM

[Kür19] Patrick Kürschner. Approximate residual-minimizing shift parameters for the lowrank ADI iteration. Electronic Transactions on Numerical Analysis (ETNA), 51:240–261, 2019. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.51.2019/pp240-
                                        261.dir/pp240-261.pdf;
                                         http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol51/abstract.php?vol=
                                        51&pages=240-261.

Krautstengl:1995:MGS

[KV95] Alan Krautstengl and Richard S. Varga. Minimal Gerschgorin sets for partitioned matrices II. The spectral conjecture. Electronic Transactions on Numerical Analysis (ETNA), 3:66–82, 1995. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.3.1995/pp66-82.
                                        dir/pp66-82.pdf.

Kostic:2013:SLI

[KV13] Aleksandra Kostić and Heinrich Voss. On Sylvester’s law of inertia for nonlinear eigenvalue problems. Electronic Transactions on Numerical Analysis (ETNA), 40:82–93, 2013. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu//vol.40.2013/pp82-
                                        93.dir/pp82-93.pdf.

Kaymakcalan:2007:LIA

[KZ07] B. Kaymakçalan and A. Zafer. Langenhop’s inequality and applications for dynamic equations. Electronic Transactions on Numerical Analysis (ETNA), 27:71–77, 2007. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.27.2007/pp71-
77.dir/pp71-77.pdf.

Lambers:2005:KSS

[Lam05] James V. Lambers. Krylov subspace spectral methods for variable-coefficient initial-boundary value problems. Electronic Transactions on Numerical Analysis (ETNA), 20:212–234, 2005. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
20.2005/pp212-234.dir/
pp212-234.pdf.

Lambers:2007:DHO

[Lam08a] James V. Lambers. Derivation of high-order spectral methods for timedependent PDE using modified moments. Electronic Transactions on Numerical Analysis (ETNA), 28: 114–135, 2007–2008. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
28.2007-2008/pp114-135.
dir/pp114-135.pdf. Spe-

cial volume for Gene Golub.

Lambers:2008:EKS

[Lam08b] James V. Lambers. Enhancement of Krylov subspace spectral methods by block Lanczos iteration. Electronic Transactions on Numerical Analysis (ETNA), 31:86–109, 2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

                                        /etna.mcs.kent.edu/vol.
                                        31.2008/pp86-109.dir/
                                        pp86-109.pdf.

Lamichhane:2010:GRO

[Lam10] Bishnu P. Lamichhane. A gradient recovery operator based on an oblique projection. Electronic Transactions on Numerical Analysis (ETNA), 37:166–172, 2010. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

                                        /etna.mcs.kent.edu/vol.
                                        37.2010/pp166-172.dir/
                                        pp166-172.pdf.

Laurie:2007:VPA

[Lau08]  Dirk Laurie.  Variable-

precision arithmetic considered perilous — a detective story. Electronic Transactions on Numerical Analysis (ETNA), 28: 168–173, 2007–2008. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        28.2007-2008/pp168-173.

dir/pp168-173.pdf. Spe-

cial volume for Gene Golub.

Levrie:1996:MCF

[LB96] P. Levrie and A. Bultheel.

Matrix continued fractions related to firstorder linear recurrence systems. Electronic Transactions on Numerical Analysis (ETNA), 4:46–63, 1996. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.4.1996/pp46-63.
dir/pp46-63.pdf.

Li:2014:SPA

[LC14] Tiexiang Li and Delin Chu.

A structure-preserving algorithm for semi-stabilizing solutions of generalized algebraic Riccati equations. Electronic Transactions on Numerical Analysis (ETNA), 41:396–419, 2014. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.41.2014/pp396-
419.dir/pp396-419.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol41/abstract.php?vol=
41&pages=396-419.

Lemarie:2013:TOGa

[LDB13a] Florian Lemarié, Laurent Debreu, and Eric Blayo. Toward an optimized globalin-time Schwarz algorithm for diffusion equations with discontinuous and spatially variable coefficients. Part 1: The constant coefficients case. Electronic Transactions on Numerical Analysis (ETNA), 40:148–169, 2013. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu//vol.
40.2013/pp148-169.dir/
pp148-169.pdf.

Lemarie:2013:TOGb

[LDB13b] Florian Lemarié, Laurent Debreu, and Eric Blayo. Toward an optimized globalin-time Schwarz algorithm for diffusion equations with discontinuous and spatially variable coefficients. Part 2: The variable coefficients case. Electronic Transactions on Numerical Analysis (ETNA), 40:170–186, 2013. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu//vol.
40.2013/pp170-186.dir/
pp170-186.pdf.

Lei:2019:TIS

[Lei19] Xiao-Qiang Lei. _{δ}^{γ}Φ-type inclusion set for eigenvalues of a tensor. Electronic Transactions on Numerical Analysis (ETNA), 51:262–273, 2019. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.

edu/vol.51.2019/pp262-
273.dir/pp262-273.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol51/abstract.php?vol=
51&pages=262-273.

Lenard:2006:WTI

[Lén06]  Margit Lénárd.  On

weighted (0, 2)-type interpolation. Electronic Transactions on Numerical Analysis (ETNA), 25: 206–223, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
25.2006/pp206-223.dir/
pp206-223.pdf.

Lenard:2010:WLI

[Lén10]  Margit Lénárd.  On

weighted lacunary interpolation. Electronic Transactions on Numerical Analysis (ETNA), 37:113–122, 2010. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
37.2010/pp113-122.dir/
pp113-122.pdf.

Leopardi:2006:PUS

[Leo06] Paul Leopardi. A partition of the unit sphere into regions of equal area and small diameter. Electronic Transactions on Numerical Analysis (ETNA), 25: 309–327, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        25.2006/pp309-327.dir/
                                        pp309-327.pdf.

Leopardi:2009:DBE

[Leo09]  Paul Leopardi.  Diam-

eter bounds for equal area partitions of the unit sphere. Electronic Transactions on Numerical Analysis (ETNA), 35:1–16, 2009. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

                                        /etna.mcs.kent.edu/vol.
                                        35.2009/pp1-16.dir/pp1-
                                        16.pdf.

Liu:2012:CMP

[LH12] Zhiyong Liu and Yinnian He. Cascadic multigrid preconditioner for elliptic equations with jump coefficients. Electronic Transactions on Numerical Analysis (ETNA), 39:333–339, 2012. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.39.2012/pp333-
                                        339.dir/pp333-339.pdf;
                                         http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol39/abstract.php?vol=
                                        39&pages=333-339.

Lu:2020:EST

[LHL^{+}20] Zuliang Lu, Fei Huang, Lin Li, Xiaoxiao Zuo, and Junhong Li. An empirical study of transboundary air pollution of the Beijing–Tianjin region. Electronic Transactions on Numerical Analysis (ETNA), 52:100–112, 2020. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.52.2020/pp100-
112.dir/pp100-112.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol52/abstract.php?vol=
52&pages=100-112.

Lisitsa:2008:ODP

[Lis08] Vadim Lisitsa. Optimal discretization of PML for elasticity problems. Electronic Transactions on Numerical Analysis (ETNA), 30:258–277, 2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
30.2008/pp258-277.dir/
pp258-277.pdf.

Lorca:2021:MSP

[LK21] José Pablo Lucero Lorca and Guido Kanschat. Multilevel Schwarz preconditioners for singularly perturbed symmetric reactiondiffusion systems. Electronic Transactions on Numerical Analysis (ETNA), 54:89–107, 2021. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.54.2021/pp89-
107.dir/pp89-107.pdf;
http://etna.mcs.kent.
edu/volumes/2021-2030/
vol54/abstract.php?vol=
54&pages=89-107.

Ludwig:2007:AFV

[LKF08] Martin Ludwig, Jochim Koch, and Bernd Fischer. An application of the finite volume method to the bio-heat-transfer-equation in premature infants. Electronic Transactions on Numerical Analysis (ETNA), 28:136–148, 2007–2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
28.2007-2008/pp136-148.
dir/pp136-148.pdf. Spe-

cial volume for Gene Golub.

Llorente:2000:BPR

[LM00] Ignacio M. Llorente and N. Duane Melson. Behavior of plane relaxation methods as multigrid smoothers. Electronic Transactions on Numerical Analysis (ETNA), 10: 92–114, 2000. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
10.2000/pp92-114.dir/
pp92-114.pdf. Multilevel

methods (Copper Mountain, CO, 1999).

Lazzaro:2020:MCM

[LM20] Damiana Lazzaro and Serena Morigi. Matrix completion for matrices with lowrank displacement. Electronic Transactions on Numerical Analysis (ETNA), 53:481–499, 2020. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.53.2020/pp481-
499.dir/pp481-499.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol53/abstract.php?vol=
53&pages=481-499.

Lorenz:2012:GDT

[LMM12] Dirk A. Lorenz, Peter Maass, and Pham Q. Muoi. Gradient descent for Tikhonov functionals with sparsity constraints: Theory and numerical comparison of step size rules. Electronic Transactions on Numerical Analysis (ETNA), 39:437–463, 2012. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.39.2012/pp437-
463.dir/pp437-463.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol39/abstract.php?vol=
39&pages=437-463.

Li:2010:AFE

[LMN10] Hengguang Li, Anna Mazzucato, and Victor Nistor. Analysis of the finite element method for transmission/mixed boundary value problems on general polygonal domains. Electronic Transactions on Numerical Analysis (ETNA), 37:41–69, 2010. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.37.2010/pp41-
69.dir/pp41-69.pdf.

Loneland:2015:EBS

[LMR15] Atle Loneland, Leszek Marcinkowski, and Talal Rahman. Edge-based Schwarz methods for the Crouzeix–Raviart finite volume element discretization of elliptic problems. Electronic Transactions on Numerical Analysis (ETNA), 44:443–461, 2015. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.44.2015/pp443-
461.dir/pp443-461.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol44/abstract.php?vol=
44&pages=443-461.

Layton:1997:TLD

[LMS97] W. J. Layton, A. J. Meir, and P. G. Schmidt. A two-level discretization method for the stationary MHD equations. Electronic Transactions on Numerical Analysis (ETNA), 6: 198–210, 1997. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
6.1997/pp198-210.dir/
pp198-210.pdf. Special

issue on multilevel methods (Copper Mountain, CO, 1997).

Lopez:2008:CBD

[LMS09] Hilda López, Brígida Molina, and José J. Salas. Comparison between different numerical discretizations for a Darcy–Forchheimer model. Electronic Transactions on Numerical Analysis (ETNA), 34:187–203, 2008– 2009. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.34.2008-2009/
pp187-203.dir/pp187-203.
pdf.

Lanza:2014:VID

[LMSY14] Alessandro Lanza, Serena Morigi, Fiorella Sgallari, and Anthony J. Yezzi. Variational image denoising while constraining the distribution of the residual. Electronic Transactions on Numerical Analysis (ETNA), 42:64–84, 2014. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.42.2014/pp64-
                                        84.dir/pp64-84.pdf;
                                        http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol42/abstract.php?vol=
                                        42&pages=64-84.

Leszczynski:2020:LFM

[LMW20] Henryk Leszczyński, Milena Matusik, and Monika Wrzosek. Leap-frog method for stochastic functional wave equations. Electronic Transactions on Numerical Analysis (ETNA), 52:77–87, 2020. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.52.2020/pp77-
                                        87.dir/pp77-87.pdf;
                                        http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol52/abstract.php?vol=
                                        52&pages=77-87.

LeBorne:2010:IAI

[LN10] Sabine Le Borne and Che Ngufor. An implicit approximate inverse preconditioner for saddle point problems. Electronic Transactions on Numerical Analysis (ETNA), 37:173–188, 2010. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

                                        /etna.mcs.kent.edu/vol.
                                        37.2010/pp173-188.dir/
                                        pp173-188.pdf.

Li:2015:IMS

[LNG15] Na Li, Carmeliza Navasca, and Christina Glenn. Iterative methods for symmetric outer product tensor decomposition. Electronic Transactions on Numerical Analysis (ETNA), 44:124–139, 2015. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.44.2015/pp124-
139.dir/pp124-139.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol44/abstract.php?vol=
44&pages=124-139.

LeBorne:2007:JDD

[LO07] Sabine Le Borne and Suely Oliveira. Joint domaindecomposition H − LU preconditioners for saddle point problems. Electronic Transactions on Numerical Analysis (ETNA), 26: 285–298, 2007. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
26.2007/pp285-298.dir/
pp285-298.pdf.

Le:2013:DPI

[LO13] Anh Ha Le and Pascal Omnes. Discrete Poincaré inequalities for arbitrary meshes in the discrete duality finite volume context. Electronic Transactions on Numerical Analysis (ETNA), 40:94–119, 2013. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu//vol.40.2013/pp94-
119.dir/pp94-119.pdf.

Lopez:2016:NGA

[Lóp16] Williams López. A new geometric acceleration of the von Neumann–Halperin projection method. Electronic Transactions on Numerical Analysis (ETNA), 45:330–341, 2016. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.45.2016/pp330-
341.dir/pp330-341.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol45/abstract.php?vol=
45&pages=330-341.

Lorente:1999:CAO

[Lor99] Miguel Lorente. Creation and annihilation operators for orthogonal polynomials of continuous and discrete variables. Electronic Transactions on Numerical Analysis (ETNA), 9: 102–111, 1999. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
9.1999/pp102-111.dir/
pp102-111.pdf.

Lopez:2008:SLM

[LP08] José L. López and Pedro J.

Pagola. A simplification of the Laplace method for double integrals. Application to the second Appell function. Electronic Transactions on Numerical Analysis (ETNA), 30:224–236, 2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
30.2008/pp224-236.dir/
pp224-236.pdf.

Lentini:2008:OGM

[LP09] Marianela Lentini and Marco Paluszny. Orthogonal grids on meander-like regions. Electronic Transactions on Numerical Analysis (ETNA), 34:1–13, 2008– 2009. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.34.2008-2009/
pp1-13.dir/pp1-13.pdf.

Lopez:2017:SSP

[LP17] José L. López and Pedro J. Pagola. A simplification of the stationary phase method: application to the Anger and Weber functions. Electronic Transactions on Numerical Analysis (ETNA), 46:148–161, 2017. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.46.2017/pp148-
161.dir/pp148-161.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol46/abstract.php?vol=
46&pages=148-161.

Lanza:2020:RWP

[LPS20] Alessandro Lanza, Monica Pragliola, and Fiorella Sgallari. Residual whiteness principle for parameter-free image restoration. Electronic Transactions on Numerical Analysis (ETNA), 53:329–351, 2020. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.53.2020/pp329-
351.dir/pp329-351.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol53/abstract.php?vol=
53&pages=329-351.

Luengo:2003:GMD

[LR03] Francisco Luengo and Marcos Raydan. Gradient method with dynamical retards for large-scale optimization problems. Electronic Transactions on Numerical Analysis (ETNA), 16:186–193, 2003. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
16.2003/pp186-193.dir/
pp186-193.pdf.

Luther:2004:MEI

[LR04] Uwe Luther and Karla Rost.

Matrix exponentials and inversion of confluent Vandermonde matrices. Electronic Transactions on Numerical Analysis (ETNA), 18:91–100, 2004. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
18.2004/pp91-100.dir/
pp91-100.pdf.

Linss:2016:RPE

[LR16] Torsten Linß and Goran Radojev. Robust a posteriori error bounds for spline collocation applied to singularly perturbed reaction– diffusion problems. Electronic Transactions on Numerical Analysis (ETNA), 45:342–353, 2016. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.45.2016/pp342-
353.dir/pp342-353.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol45/abstract.php?vol=
45&pages=342-353.

Lube:2008:LPS

[LRL08] G. Lube, G. Rapin, and J. Löwe. Local projection stabilization for incompressible flows: Equal-order vs. inf-sup stable interpolation. Electronic Transactions on Numerical Analysis (ETNA), 32:106–122, 2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

                                        /etna.mcs.kent.edu/vol.
                                        32.2008/pp106-122.dir/
                                        pp106-122.pdf.

Li:2005:CVI

[LS05] Na Li and Yousef Saad. Crout versions of ILU factorization with pivoting for sparse symmetric matrices. Electronic Transactions on Numerical Analysis (ETNA), 20:75–85, 2005. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.20.2005/pp75-
                                        85.dir/pp75-85.pdf.

Lee:1994:MCM

[LT94] Barry Lee and Manfred R. Trummer. Multigrid conformal mapping via the Szegő Kernel. Electronic Transactions on Numerical Analysis (ETNA), 2:22–43, 1994. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.2.1994/pp22-43.
                                        dir/pp22-43.pdf.

Liesen:2005:WCC

[LT05] Jörg Liesen and Petr Tichý. On the worst-case convergence of MR and CG for symmetric positive definite tridiagonal Toeplitz matrices. Electronic Transactions on Numerical Analysis (ETNA), 20:180–197, 2005. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
20.2005/pp180-197.dir/
pp180-197.pdf.

Liesen:2014:MMM

[LT14] Jörg Liesen and Petr Tichý.

Max-min and min-max approximation problems for normal matrices revisited. Electronic Transactions on Numerical Analysis (ETNA), 41:159–166, 2014. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.41.2014/pp159-
166.dir/pp159-166.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol41/abstract.php?vol=
41&pages=159-166.

Linh:2018:RKM

[LT18] Vu Hoang Linh and Nguyen Duy Truong. Runge–Kutta methods revisited for a class of structured strangenessfree differential-algebraic equations. Electronic Transactions on Numerical Analysis (ETNA), 48: 131–155, 2018. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.48.2018/pp131-
                                        155.dir/pp131-155.pdf;
                                         http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol48/abstract.php?vol=
                                        48&pages=131-155.

Lubinsky:2006:BWA

[Lub06] D. S. Lubinsky. Bernstein’s weighted approximation on R still has problems. Electronic Transactions on Numerical Analysis (ETNA), 25:166–177, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        25.2006/pp166-177.dir/
                                        pp166-177.pdf.

Lui:2011:PMT

[Lui11] S. H. Lui. Pseudospectral mapping theorem II. Electronic Transactions on Numerical Analysis (ETNA), 38:168–183, 2011. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic).

Lampe:2008:FAS

[LV08] Jörg Lampe and Heinrich Voss. A fast algorithm for solving regularized total least squares problems. Electronic Transactions on Numerical Analysis (ETNA), 31:12–24, 2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.31.2008/pp12-
24.dir/pp12-24.pdf.

Lampe:2014:LSD

[LV14] Jörg Lampe and Heinrich Voss. Large-scale dual regularized total least squares. Electronic Transactions on Numerical Analysis (ETNA), 42:13–40, 2014. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.42.2014/pp13-
40.dir/pp13-40.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol42/abstract.php?vol=
42&pages=13-40.

Livne:2006:FME

[LW06] Oren E. Livne and Grady B.

Wright. Fast multilevel evaluation of smooth radial basis function expansions. Electronic Transactions on Numerical Analysis (ETNA), 23:263–287, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
23.2006/pp263-287.dir/
pp263-287.pdf.

Li:2020:SEG

[LWH20] Junlin Li, Tongke Wang, and Yonghong Hao. The series expansions and Gauss– Legendre rule for computing arbitrary derivatives of the Beta-type functions. Electronic Transactions on Numerical Analysis (ETNA), 52:203–213, 2020. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.52.2020/pp203-
213.dir/pp203-213.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol52/abstract.php?vol=
52&pages=203-213.

Liang:2014:CSV

[LY14] Qiao Liang and Qiang Ye. Computing singular values of large matrices with an inverse-free preconditioned Krylov subspace method. Electronic Transactions on Numerical Analysis (ETNA), 42:197–221, 2014. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.42.2014/pp197-
221.dir/pp197-221.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol42/abstract.php?vol=
42&pages=197-221.

Littlejohn:2007:LDV

[LZ07] L. L. Littlejohn and A. Zettl. Left-definite variations of the classical Fourier expansion theorem. Electronic Transactions on Numerical Analysis (ETNA), 27: 124–139, 2007. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
27.2007/pp124-139.dir/
pp124-139.pdf.

Laszkiewicz:2008:AMS

[LZ08a] Beata Laszkiewicz and Krystyna Zietak. Algorithms for the matrix sector function. Electronic Transactions on Numerical Analysis (ETNA), 31: 358–383, 2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
31.2008/pp358-383.dir/
pp358-383.pdf.

Liu:2008:SBG

[LZ08b] Hualei Liu and Baojiang Zhong. Simpler Block GMRES for nonsymmetric systems with multiple right-hand sides. Electronic Transactions on Numerical Analysis (ETNA), 30:1–9, 2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
30.2008/pp1-9.dir/pp1-
9.pdf.

Miyazaki:2006:NCE

[MACI06] Yoshinori Miyazaki, Nobuyoshi Asai, Dongsheng Cai, and Yasuhiko Ikebe. Numerical computation of the eigenvalues for the spheroidal wave equation with accurate error estimation by matrix method. Electronic Transactions on Numerical Analysis (ETNA), 23: 329–338, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
23.2006/pp329-338.dir/
pp329-338.pdf.

Malek:2008:MPF

[Mál08] Josef Málek. Mathematical properties of flows of incompressible power-law-like fluids that are described by implicit constitutive relations. Electronic Transactions on Numerical Analysis (ETNA), 31:110–125, 2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
31.2008/pp110-125.dir/
pp110-125.pdf.

Mandel:2003:LAE

[Man03] Jan Mandel. Local approximation estimators for algebraic multigrid. Electronic Transactions on Numerical Analysis (ETNA), 15:56–65, 2003. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.15.2003/pp56-
65.dir/pp56-65.pdf.

Mantica:2006:FBF

[Man06] Giorgio Mantica. Fourier– Bessel functions of singular continuous measures and their many asymptotics. Electronic Transactions on Numerical Analysis (ETNA), 25:409–430, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
25.2006/pp409-430.dir/
pp409-430.pdf.

Mantica:2007:QDE

[Man08] Giorgio Mantica. Quantum dynamical entropy and an algorithm by Gene Golub. Electronic Transactions on Numerical Analysis (ETNA), 28:190–205, 2007– 2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.28.2007-2008/
pp190-205.dir/pp190-205.

pdf. Special volume for Gene Golub.

Marcinkowski:2007:ASM

[Mar07] Leszek Marcinkowski. An additive Schwarz method for mortar Morley finite element discretizations of 4th order elliptic problem in 2D. Electronic Transactions on Numerical Analysis (ETNA), 26:34–54, 2007. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.26.2007/pp34-
                                        54.dir/pp34-54.pdf.

Marcinkowski:2011:PFD

[Mar11] Leszek Marcinkowski. A preconditioner for a FETI-DP method for mortar element discretization of a 4th order problem in 2D. Electronic Transactions on Numerical Analysis (ETNA), 38:1–16, 2011. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic).

Marusic:2018:HOE

[Mar18] Miljenko Marušić. Highorder exponentially fitted difference schemes for singularly perturbed twopoint boundary value problems. Electronic Transactions on Numerical Analysis (ETNA), 48:329–347, 2018. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.48.2018/pp329-
                                        347.dir/pp329-347.pdf;
                                         http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol48/abstract.php?vol=
                                        48&pages=329-347.

Mastroianni:2002:PIF

[Mas02]  G. Mastroianni.  Poly-

nomial inequalities, functional spaces and best approximation on the real semiaxis with Laguerre weights. Electronic Transactions on Numerical Analysis (ETNA), 14:142–151, 2002. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.14.2002/pp142-
151.dir/pp142-151.pdf.

Mashele:2006:COE

[Mas06] H. P. Mashele. On convergence of orthonormal expansions for exponential weights. Electronic Transactions on Numerical Analysis (ETNA), 25: 467–479, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
25.2006/pp467-479.dir/
pp467-479.pdf.

Mavriplis:1997:DCS

[Mav97] Dimitri J. Mavriplis. Directional coarsening and smoothing for anisotropic Navier–Stokes problems. Electronic Transactions on Numerical Analysis (ETNA), 6:182–197, 1997. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
6.1997/pp182-197.dir/
pp182-197.pdf. Special

issue on multilevel methods (Copper Mountain, CO,

Maes:2006:SMA

[MB06] Jan Maes and Adhemar Bultheel. Stable multiresolution analysis on triangles for surface compression. Electronic Transactions on Numerical Analysis (ETNA), 25:224–258, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

                                        /etna.mcs.kent.edu/vol.
                                        25.2006/pp224-258.dir/
                                        pp224-258.pdf.

Mao:2005:CAR

[MC05] Shipeng Mao and Shaochun Chen. Convergence analysis of the rotated Q_{1} element on anisotropic rectangular meshes. Electronic Transactions on Numerical Analysis (ETNA), 20:154–163, 2005. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

                                        /etna.mcs.kent.edu/vol.
                                        20.2005/pp154-163.dir/
                                        pp154-163.pdf.

Meerbergen:1998:TCB

[Mee98] Karl Meerbergen. A theoretical comparison between inner products in the shiftinvert Arnoldi method and the spectral transformation Lanczos method. Electronic Transactions on Numerical Analysis (ETNA), 7:90–103, 1998. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
7.1998/pp90-103.dir/pp90-

103.pdf. Large scale eigenvalue problems (Argonne, IL, 1997).

Meurant:2001:NEA

[Meu01] Gérard Meurant. Numerical experiments with algebraic multilevel preconditioners. Electronic Transactions on Numerical Analysis (ETNA), 12:1–65, 2001. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
12.2001/pp1-65.dir/pp1-
65.pdf.

Meurant:2012:CSG

[Meu12] Gérard Meurant. The complete stagnation of GM-RES for N ≤ 4. Electronic Transactions on Numerical Analysis (ETNA), 39:75–101, 2012. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.39.2012/pp75-
101.dir/pp75-101.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol39/abstract.php?vol=
39&pages=75-101.

Meurant:2014:LRV

[Meu15] Gérard Meurant. On the location of the Ritz values in the Arnoldi process. Electronic Transactions on Numerical Analysis (ETNA), 43:188–212, 2014– 2015. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.43.2014/pp188-
                                        212.dir/pp188-212.pdf;
                                         http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol43/abstract.php?vol=
                                        43&pages=188-212.

Meurant:2017:OQK

[Meu17] Gérard Meurant. An optimal Q-OR Krylov subspace method for solving linear systems. Electronic Transactions on Numerical Analysis (ETNA), 47:127–152, 2017. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.47.2017/pp127-
                                        152.dir/pp127-152.pdf;
                                         http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol47/abstract.php?vol=
                                        47&pages=127-152.

Martinez-Finkelshtein:2006:SPV

[MF06] A. Martínez-Finkelshtein. Szegő polynomials: a view from the Riemann–Hilbert window. Electronic Transactions on Numerical Analysis (ETNA), 25:369–392, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.

edu/vol.25.2006/pp369-
392.dir/pp369-392.pdf.

Martinez-Finkelshtein:2005:APS

[MFMGO05] A. Martínez-Finkelshtein, P. Martínez-González, and R. Orive. Asymptotics of polynomial solutions of a class of generalized Lamé differential equations. Electronic Transactions on Numerical Analysis (ETNA), 19:18–28, 2005. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.19.2005/pp18-
28.dir/pp18-28.pdf.

Mokhtary:2014:CAO

[MG14] P. Mokhtary and F. Ghoreishi. Convergence analysis of the operational Tau method for Abeltype Volterra integral equations. Electronic Transactions on Numerical Analysis (ETNA), 41:289–305, 2014. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.41.2014/pp289-
305.dir/pp289-305.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol41/abstract.php?vol=
41&pages=289-305.

McDonald:2001:EAQ

[MH01] Edward J. McDonald and Desmond J. Higham. Error analysis of QR algorithms for computing Lyapunov exponents. Electronic Transactions on Numerical Analysis (ETNA), 12:234–251, 2001. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
12.2001/pp234-251.dir/
pp234-251.pdf.

Morigi:2017:SIV

[MH17] Serena Morigi and Martin Huska. Sparsity-inducing variational shape partitioning. Electronic Transactions on Numerical Analysis (ETNA), 46:36–54, 2017. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.46.2017/pp36-
54.dir/pp36-54.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol46/abstract.php?vol=
46&pages=36-54.

Mitchell:1997:PMM

[Mit97] William F. Mitchell. A parallel multigrid method using the full domain partition. Electronic Transactions on Numerical Analysis (ETNA), 6:224–233, 1997. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
6.1997/pp224-233.dir/

pp224-233.pdf. Special

issue on multilevel methods (Copper Mountain, CO, 1997).

Masjed-Jamei:2011:LCA

[MJ11] Mohammad Masjed-Jamei.

A linear constructive approximation for integrable functions and a parametric quadrature model based on a generalization of Ostrowski–Grüss type inequalities. Electronic Transactions on Numerical Analysis (ETNA), 38:218–232, 2011. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic).

Masjed-Jamei:2016:WHQ

[MJM16] Mohammad Masjed-Jamei and Gradimir V. Milovanović. Weighted Hermite quadrature rules. Electronic Transactions on Numerical Analysis (ETNA), 45:476–498, 2016. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.45.2016/pp476-
498.dir/pp476-498.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol45/abstract.php?vol=
45&pages=476-498.

Marcellan:1999:COP

[MM99] F. Marcellán and J. C. Medem. Q-classical orthogonal polynomials: a very classical approach. Electronic Transactions on Numerical Analysis (ETNA), 9:112–127, 1999. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
9.1999/pp112-127.dir/
pp112-127.pdf.

Marco:2004:NSS

[MM04] Ana Marco and José-Javier Martínez. A new source of structured singular value decomposition problems. Electronic Transactions on Numerical Analysis (ETNA), 18: 188–197, 2004. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
18.2004/pp188-197.dir/
pp188-197.pdf.

Marco:2008:USB

[MM09a] Ana Marco and José-Javier Martínez. Unique solvability in bivariate Hermite interpolation. Electronic Transactions on Numerical Analysis (ETNA), 34: 20–30, 2008–2009. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
34.2008-2009/pp20-30.
dir/pp20-30.pdf.

Mocskos:2008:IGT

[MM09b] E. Mocskos and G. Marshall. Impact of gravity on thin-layer cell electrochemical deposition. Electronic Transactions on Numerical Analysis (ETNA), 34:90–101, 2008–2009. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
34.2008-2009/pp90-101.
dir/pp90-101.pdf.

Mastroianni:2018:PAP

[MMN18] Giuseppe Mastroianni, Gradimir V.

Milovanović, and Incoronata Notarangelo. Polynomial approximation with Pollaczeck-Laguerre weights on the real semiaxis. A survey. Electronic Transactions on Numerical Analysis (ETNA), 50:36–51, 2018. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.50.2018/pp36-
51.dir/pp36-51.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol50/abstract.php?vol=
50&pages=36-51.

Marek:2009:CIT

[MMP09] Ivo Marek, Petr Mayer, and Ivana Pultarová. Convergence issues in the theory and practice of iterative aggregation/disaggregation methods. Electronic Transactions on Numerical Analysis (ETNA), 35:185–200, 2009. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.35.2009/pp185-
200.dir/pp185-200.pdf.

Mense:2008:AMM

[MN08] Christian Mense and Reinhard Nabben. On algebraic multilevel methods for non-symmetric systems — convergence results. Electronic Transactions on Numerical Analysis (ETNA), 30:323–345, 2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
30.2008/pp323-345.dir/
pp323-345.pdf.

Mokhtary:2015:HOM

[Mok15] Payam Mokhtary. Highorder modified Tau method for non-smooth solutions of Abel integral equations. Electronic Transactions on Numerical Analysis (ETNA), 44:462–471, 2015. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.44.2015/pp462-
471.dir/pp462-471.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/

vol44/abstract.php?vol=
44&pages=462-471.

Mokhtary:2016:OMG

[Mok16]  P. Mokhtary.  Opera-

tional Müntz–Galerkin approximation for Abel– Hammerstein integral equations of the second kind. Electronic Transactions on Numerical Analysis (ETNA), 45:183–200, 2016. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.45.2016/pp183-
200.dir/pp183-200.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol45/abstract.php?vol=
45&pages=183-200.

Munteanu:2008:OAS

[MP08] M. Munteanu and L. F.

Pavarino. An overlapping additive Schwarz– Richardson method for monotone nonlinear parabolic problems. Electronic Transactions on Numerical Analysis (ETNA), 30:359–376, 2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.30.2008/pp359-
376.dir/pp359-376.pdf.

Mutavdzic:2018:EBK

[MPS18] Rada M. Mutavdžić, Aleksandar V. Pejčev, and Miodrag M. Spalević. Error bounds for Kronrod extension of generalizations of Micchelli–Rivlin quadrature formula for analytic functions. Electronic Transactions on Numerical Analysis (ETNA), 50:20–35, 2018. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.50.2018/pp20-
35.dir/pp20-35.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol50/abstract.php?vol=
50&pages=20-35.

Mach:2013:CAE

[MPV13] Thomas Mach, Miroslav S. Pranić, and Raf Vandebril. Computing approximate extended Krylov subspaces without explicit inversion. Electronic Transactions on Numerical Analysis (ETNA), 40:414–435, 2013. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu//vol.
40.2013/pp414-435.dir/
pp414-435.pdf.

Mach:2014:CAB

[MPV15] Thomas Mach, Miroslav S. Pranić, and Raf Vandebril. Computing approximate (block) rational Krylov subspaces without explicit inversion with extensions to symmetric matrices. Electronic Transactions on Numerical Analysis (ETNA), 43:100–124, 2014– 2015. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.43.2014/pp100-
124.dir/pp100-124.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol43/abstract.php?vol=
43&pages=100-124.

Mehrmann:1993:NMC

[MR93] Volker Mehrmann and Werner Rath. Numerical methods for the computation of analytic singular value decompositions. Electronic Transactions on Numerical Analysis (ETNA), 1:72–88, 1993. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.1.1993/pp72-88.
dir/pp72-88.pdf.

Mitrouli:2017:VEE

[MR17] Marilena Mitrouli and Paraskevi Roupa. Vector estimates for f(A)b via extrapolation. Electronic Transactions on Numerical Analysis (ETNA), 47:179–196, 2017. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.47.2017/pp179-
196.dir/pp179-196.pdf;

                                         http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol47/abstract.php?vol=
                                        47&pages=179-196.

Marcinkowski:2018:AAS

[MR18] Leszek Marcinkowski and Talal Rahman. Additive average Schwarz with adaptive coarse spaces: scalable algorithms for multiscale problems. Electronic Transactions on Numerical Analysis (ETNA), 49:28–40, 2018. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.49.2018/pp28-
                                        40.dir/pp28-40.pdf;
                                        http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol49/abstract.php?vol=
                                        49&pages=28-40.

Martikainen:2003:MPL

[MRT03] Janne Martikainen, Tuomo Rossi, and Jari Toivanen. Multilevel preconditioners for Lagrange multipliers in domain imbedding. Electronic Transactions on Numerical Analysis (ETNA), 15:94–105, 2003. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        15.2003/pp94-105.dir/
                                        pp94-105.pdf.

Malyshev:2006:CNK

[MS06] Alexander Malyshev and Miloud Sadkane. Condition numbers of the Krylov bases and spaces associated with the truncated QZ iteration. Electronic Transactions on Numerical Analysis (ETNA), 23:5–14, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
23.2006/pp5-14.dir/pp5-
14.pdf.

Meyer-Spasche:2007:DSQ

[MS07] Rita Meyer-Spasche. On difference schemes for quasilinear evolution problems. Electronic Transactions on Numerical Analysis (ETNA), 27:78–93, 2007. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.27.2007/pp78-
93.dir/pp78-93.pdf.

Mohsen:2007:ROU

[MS08] A. Mohsen and J. Stoer.

A rank-one updating approach for solving systems of linear equations in the least squares sense. Electronic Transactions on Numerical Analysis (ETNA), 29:97–115, 2007–2008. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
29.2007-2008/pp97-115.
dir/pp97-115.pdf.

Malyshev:2014:CDQ

[MS14] Alexander Malyshev and Miloud Sadkane. On the computation of the distance to quadratic matrix polynomials that are singular at some points on the unit circle. Electronic Transactions on Numerical Analysis (ETNA), 42:165–176, 2014. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.42.2014/pp165-
                                        176.dir/pp165-176.pdf;
                                         http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol42/abstract.php?vol=
                                        42&pages=165-176.

Malyshev:2014:SIS

[MSS15] Alexander Malyshev, Miloud Sadkane, and Ahmed Salam. A subspace iteration for symplectic matrices. Electronic Transactions on Numerical Analysis (ETNA), 43:213–222, 2014–2015. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.43.2014/pp213-
                                        222.dir/pp213-222.pdf;
                                         http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol43/abstract.php?vol=
                                        43&pages=213-222.

Matthies:2008:SLP

[MST08] Gunar Matthies, Piotr Skrzypacz, and Lutz Tobiska. Stabilization of local projection type applied to convection-diffusion problems with mixed boundary conditions. Electronic Transactions on Numerical Analysis (ETNA), 32: 90–105, 2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
32.2008/pp90-105.dir/
pp90-105.pdf.

Manteuffel:1995:ASI

[MSV95] Thomas A. Manteuffel, Gerhard Starke, and Richard S. Varga. Adaptive k-step iterative methods for nonsymmetric systems of linear equations. Electronic Transactions on Numerical Analysis (ETNA), 3:50–65, 1995. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.3.1995/pp50-65.
dir/pp50-65.pdf.

Mastronardi:2015:RSC

[MV15] Nicola Mastronardi and Paul Van Dooren. Revisiting the stability of computing the roots of a quadratic polynomial. Electronic Transactions on Numerical Analysis (ETNA), 44:73–82, 2015. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.44.2015/pp73-
                                        82.dir/pp73-82.pdf;
                                        http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol44/abstract.php?vol=
                                        44&pages=73-82.

Mastronardi:2008:CNS

[MVV09] Nicola Mastronardi, Marc Van Barel, and Raf Vandebril. On the computation of the null space of Toeplitzlike matrices. Electronic Transactions on Numerical Analysis (ETNA), 33: 151–162, 2008–2009. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        33.2008-2009/pp151-162.
                                        dir/pp151-162.pdf.

Mehrmann:2002:PEP

[MW02] Volker Mehrmann and David Watkins. Polynomial eigenvalue problems with Hamiltonian structure. Electronic Transactions on Numerical Analysis (ETNA), 13:106–118, 2002. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

                                        /etna.mcs.kent.edu/vol.
                                        13.2002/pp106-118.dir/
                                        pp106-118.pdf.

Mehrmann:1996:APP

[MX96] Volker Mehrmann and Hongguo Xu. An analysis of the pole placement problem. I. The single-input case. Electronic Transactions on Numerical Analysis (ETNA), 4:89–105, 1996. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
4.1996/pp89-105.dir/pp89-
105.pdf.

Mehrmann:1997:APP

[MX97] Volker Mehrmann and Hongguo Xu. An analysis of the pole placement problem II. The multi-input case. Electronic Transactions on Numerical Analysis (ETNA), 5:77–97, 1997. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.5.1997/pp77-97.
dir/pp77-97.pdf.

Mehrmann:2015:SPD

[MX15] Volker Mehrmann and Hongguo Xu. Structure preserving deflation of infinite eigenvalues in structured pencils. Electronic Transactions on Numerical Analysis (ETNA), 44:1–24, 2015. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
44.2015/pp1-24.dir/pp1-
24.pdf;   http://etna.
mcs.kent.edu/volumes/
2011-2020/vol44/abstract.
php?vol=44&pages=1-24.

Marcellan:2002:RTA

[MY02] F. Marcellán and H. O. Yakhlef. Recent trends on analytic properties of matrix orthonormal polynomials. Electronic Transactions on Numerical Analysis (ETNA), 14:127–141, 2002. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

                                        /etna.mcs.kent.edu/vol.
                                        14.2002/pp127-141.dir/
                                        pp127-141.pdf.

Mahmoud:2018:EEM

[MYZ18] Abdrhaman Mahmoud, Bo Yu, and Xuping Zhang. Eigenfunction expansion method for multiple solutions of fourth-order ordinary differential equations with cubic polynomial nonlinearity. Electronic Transactions on Numerical Analysis (ETNA), 48:81–96, 2018. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.48.2018/pp81-
                                        96.dir/pp81-96.pdf;
                                        http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol48/abstract.php?vol=
                                        48&pages=81-96.

Meng:2014:DBF

[MZLG14] Jing Meng, Pei-Yong Zhu, Hou-Biao Li, and Xian-Ming Gu. A deflated block flexible GMRES–DR method for linear systems with multiple right-hand sides. Electronic Transactions on Numerical Analysis (ETNA), 41:478–496, 2014. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.41.2014/pp478-
496.dir/pp478-496.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol41/abstract.php?vol=
41&pages=478-496.

Nallapu:2019:BFA

[Nal19] Vijender Nallapu. Bernstein fractal approximation and fractal full Müntz theorems. Electronic Transactions on Numerical Analysis (ETNA), 51:1–14, 2019. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
51.2019/pp1-14.dir/pp1-
14.pdf;   http://etna.
mcs.kent.edu/volumes/
2011-2020/vol51/abstract.
php?vol=51&pages=1-14.

Narayan:2018:CIO

[Nar18] Akil Narayan. Computation of induced orthogonal polynomial distributions. Electronic Transactions on Numerical Analysis (ETNA), 50:71–97, 2018. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.

                                        edu/vol.50.2018/pp71-
                                        97.dir/pp71-97.pdf;
                                        http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol50/abstract.php?vol=
                                        50&pages=71-97.

Nasser:2015:FSB

[Nas15] Mohamed M. S. Nasser. Fast solution of boundary integral equations with the generalized Neumann kernel. Electronic Transactions on Numerical Analysis (ETNA), 44:189–229, 2015. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.44.2015/pp189-
                                        229.dir/pp189-229.pdf;
                                         http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol44/abstract.php?vol=
                                        44&pages=189-229.

Navascues:2005:FTA

[Nav05] M. A. Navascues. Fractal trigonometric approximation. Electronic Transactions on Numerical Analysis (ETNA), 20:64–74, 2005. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.20.2005/pp64-
                                        74.dir/pp64-74.pdf.

Neytcheva:2007:PNS

[NB08a] Maya Neytcheva and Erik Bängtsson. Preconditioning of nonsymmetric saddle point systems as arising in modelling of viscoelastic problems. Electronic Transactions on Numerical Analysis (ETNA), 29: 193–211, 2007–2008. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
29.2007-2008/pp193-211.
dir/pp193-211.pdf.

NGohisse:2008:NBS

[NB08b] Firmin K. N’Gohisse and Théodore K. Boni. Numerical blow-up solutions for some semilinear heat equations. Electronic Transactions on Numerical Analysis (ETNA), 30:247–257, 2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
30.2008/pp247-257.dir/
pp247-257.pdf.

Nicaise:2006:IAP

[NC06] Serge Nicaise and Emmanuel Creusé. Isotropic and anisotropic a posteriori error estimation of the mixed finite element method for second order operators in divergence form. Electronic Transactions on Numerical Analysis (ETNA), 23:38–62, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.

                                        edu/vol.23.2006/pp38-
                                        62.dir/pp38-62.pdf.

Neilan:2014:UAT

[Nei14] Michael Neilan. A unified analysis of three finite element methods for the Monge–Ampère equation. Electronic Transactions on Numerical Analysis (ETNA), 41:262–288, 2014. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.41.2014/pp262-
                                        288.dir/pp262-288.pdf;
                                         http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol41/abstract.php?vol=
                                        41&pages=262-288.

Neubauer:2016:NOR

[Neu16]  Andreas Neubauer.  A

note on optimal rates for Lavrentiev regularization with adjoint source conditions. Electronic Transactions on Numerical Analysis (ETNA), 45:420–423, 2016. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.45.2016/pp420-
                                        423.dir/pp420-423.pdf;
                                         http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol45/abstract.php?vol=
                                        45&pages=420-423.

Neubauer:2019:AGT

[Neu19] Andreas Neubauer. Augmented GMRES-type versus CGNE methods for the solution of linear ill-posed problems. Electronic Transactions on Numerical Analysis (ETNA), 51:412–431, 2019. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.51.2019/pp412-
431.dir/pp412-431.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol51/abstract.php?vol=
51&pages=412-431.

Ng:1994:FIM

[Ng94] Michael K. Ng. Fast iterative methods for solving Toeplitz-plus-Hankel least squares problems. Electronic Transactions on Numerical Analysis (ETNA), 2:154–170, 1994. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
2.1994/pp154-170.dir/
pp154-170.pdf.

Ng:1997:CPC

[Ng97]  Michael K. Ng.  Cir-

culant preconditioners for convolution-like integral equations with higher-order quadrature rules. Electronic Transactions on Numerical Analysis (ETNA), 5:18–28, 1997. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.5.1997/pp18-28.
                                        dir/pp18-28.pdf.

Napov:2016:EMM

[NN16] Artem Napov and Yvan Notay. An efficient multigrid method for graph Laplacian systems. Electronic Transactions on Numerical Analysis (ETNA), 45:201–218, 2016. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.45.2016/pp201-
                                        218.dir/pp201-218.pdf;
                                         http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol45/abstract.php?vol=
                                        45&pages=201-218.

Naghsh-Nilchi:2006:ISC

[NND06] Ahmad Reza Naghsh-Nilchi and Shahram Daroee. Iterative sinc-convolution method for solving radiosity equation in computer graphics. Electronic Transactions on Numerical Analysis (ETNA), 23: 251–262, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        23.2006/pp251-262.dir/
                                        pp251-262.pdf.

Natalini:1999:IMC

[NNR99] P. Natalini, S. Noschese, and P. E. Ricci. An iterative method for computing the eigenvalues of second kind Fredholm operators and applications. Electronic Transactions on Numerical Analysis (ETNA), 9:128–136, 1999. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
9.1999/pp128-136.dir/
pp128-136.pdf.

Nagy:2002:IRT

[NO02] James G. Nagy and Dianne P. O’Leary. Image restoration through subimages and confidence images. Electronic Transactions on Numerical Analysis (ETNA), 13:22–37, 2002. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.13.2002/pp22-
37.dir/pp22-37.pdf.

Notay:2010:ABA

[Not10] Yvan Notay. An aggregationbased algebraic multigrid method. Electronic Transactions on Numerical Analysis (ETNA), 37:123–146, 2010. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.37.2010/pp123-
146.dir/pp123-146.pdf.

Notaris:2016:GKQ

[Not16] Sotirios E. Notaris. Gauss– Kronrod quadrature formulae — a survey of fifty years of research. Electronic Transactions on Numerical Analysis (ETNA), 45:371–404, 2016. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.45.2016/pp371-
404.dir/pp371-404.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol45/abstract.php?vol=
45&pages=371-404.

Novario:2005:RCC

[Nov05] P. G. Novario. Recursive computation of certain integrals of elliptic type. Electronic Transactions on Numerical Analysis (ETNA), 20:198–211, 2005. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
20.2005/pp198-211.dir/
pp198-211.pdf.

Ng:1996:LNA

[NP96] Michael K. Ng and Robert J. Plemmons. LMS–Newton adaptive filtering using FFT-based conjugate gradient iterations. Electronic Transactions on Numerical Analysis (ETNA), 4:14–36, 1996. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.

edu/vol.4.1996/pp14-36.
dir/pp14-36.pdf.

Napov:2019:AAT

[NP19a] Artem Napov and Ronan Perrussel. Algebraic analysis of two-level multigrid methods for edge elements. Electronic Transactions on Numerical Analysis (ETNA), 51:387–411, 2019. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.51.2019/pp387-
411.dir/pp387-411.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol51/abstract.php?vol=
51&pages=387-411.

Napov:2019:RAB

[NP19b] Artem Napov and Ronan Perrussel. Revisiting aggregation-based multigrid for edge elements. Electronic Transactions on Numerical Analysis (ETNA), 51:118–134, 2019. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.51.2019/pp118-
134.dir/pp118-134.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol51/abstract.php?vol=
51&pages=118-134.

Nasdala:2020:TRL

[NP20] Robert Nasdala and Daniel Potts. Transformed rank-1 lattices for high-dimensional approximation. Electronic Transactions on Numerical Analysis (ETNA), 53: 239–282, 2020. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.53.2020/pp239-
282.dir/pp239-282.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol53/abstract.php?vol=
53&pages=239-282.

Noschese:2007:SDN

[NPR08] S. Noschese, L. Pasquini, and L. Reichel. The structured distance to normality of an irreducible real tridiagonal matrix. Electronic Transactions on Numerical Analysis (ETNA), 28:65–77, 2007–2008. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
28.2007-2008/pp65-77.
dir/pp65-77.pdf. Special

volume for Gene Golub. Neubauer:2014:CRQ

[NR14] Andreas Neubauer and Ronny Ramlau. On convergence rates for quasisolutions of ill-posed problems. Electronic Transactions on Numerical Analysis (ETNA), 41:81–92, 2014. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.41.2014/pp81-
92.dir/pp81-92.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol41/abstract.php?vol=
41&pages=81-92.

Niebsch:2017:SCD

[NRS17] Jenny Niebsch, Ronny Ramlau, and Kirk M. Soodhalter. Solution of coupled differential equations arising from imbalance problems. Electronic Transactions on Numerical Analysis (ETNA), 46:89–106, 2017. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.46.2017/pp89-
106.dir/pp89-106.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol46/abstract.php?vol=
46&pages=89-106.

Naumann:2005:OPE

[NU05] Uwe Naumann and Jean Utke. Optimality-preserving elimination of linearities in Jacobian accumulation. Electronic Transactions on Numerical Analysis (ETNA), 21:134–150, 2005. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
21.2005/pp134-150.dir/
pp134-150.pdf.

Nhan:2019:ATE

[NV19] Thái Anh Nhan and Relja Vulanović. Analysis of the truncation error and barrier-function technique for a Bakhvalov-type mesh. Electronic Transactions on Numerical Analysis (ETNA), 51:315–330, 2019. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.51.2019/pp315-
                                        330.dir/pp315-330.pdf;
                                         http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol51/abstract.php?vol=
                                        51&pages=315-330.

Nagele:2003:LES

[NW03] Sandra Nägele and Gabriel Wittum. Large-eddy simulation and multigrid methods. Electronic Transactions on Numerical Analysis (ETNA), 15:152–164, 2003. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

                                        /etna.mcs.kent.edu/vol.
                                        15.2003/pp152-164.dir/
                                        pp152-164.pdf.

Neymeyr:2014:BPS

[NZ14] Klaus Neymeyr and Ming Zhou. The block preconditioned steepest descent iteration for elliptic operator eigenvalue problems. Electronic Transactions on Numerical Analysis (ETNA), 41:93–108, 2014. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.41.2014/pp93-
108.dir/pp93-108.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol41/abstract.php?vol=
41&pages=93-108.

OLeary:2005:QFU

[OB05] Dianne P. O’Leary and Stephen S. Bullock. QR factorizations using a restricted set of rotations. Electronic Transactions on Numerical Analysis (ETNA), 21:20–27, 2005. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.21.2005/pp20-
27.dir/pp20-27.pdf.

Oosterlee:2003:MLC

[Oos03]  C. W. Oosterlee.  On

multigrid for linear complementarity problems with application to Americanstyle options. Electronic Transactions on Numerical Analysis (ETNA), 15: 165–185, 2003. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
15.2003/pp165-185.dir/
pp165-185.pdf.

Opfer:2009:PVM

[Opf10] Gerhard Opfer. Polynomials and Vandermonde matrices over the field of quaternions. Electronic Transactions on Numerical Analysis (ETNA), 36: 9–16, 2009–2010. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.36.2009-2010/
                                        pp9-16.dir/pp9-16.pdf;
                                         http://etna.mcs.kent.
                                        edu/volumes/2001-2010/
                                        vol36/abstract.php?vol=
                                        36&pages=9-16.

Opfer:2015:PIN

[Opf15] Gerhard Opfer. Polynomial interpolation in nondivision algebras. Electronic Transactions on Numerical Analysis (ETNA), 44:660–670, 2015. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.44.2015/pp660-
                                        670.dir/pp660-670.pdf;
                                         http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol44/abstract.php?vol=
                                        44&pages=660-670.

Olshanskii:2011:ABR

[OR11] Maxim A. Olshanskii and Leo G. Rebholz. Application of barycenter refined meshes in linear elasticity and incompressible fluid dynamics. Electronic Transactions on Numerical Analysis (ETNA), 38:258–274, 2011. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
38.2011/pp258-274.dir/
pp258-274.pdf.

Orel:2010:AGE

[Ore10] Bojan Orel. Accumulation of global error in Lie group methods for linear ordinary differential equations. Electronic Transactions on Numerical Analysis (ETNA), 37:252–262, 2010. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic).

Oleary:1998:CNR

[OS98] D. P. Oleary and G. W.

Stewart. On the convergence of a new Rayleigh quotient method with applications to large eigenproblems. Electronic Transactions on Numerical Analysis (ETNA), 7:182–189, 1998. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
7.1998/pp182-189.dir/
pp182-189.pdf.  Large

scale eigenvalue problems (Argonne, IL, 1997).

Ohashi:2014:CMM

[OS15] Asuka Ohashi and Tomohiro Sogabe. On computing maximum/minimum singular values of a generalized tensor sum. Electronic Transactions on Numerical Analysis (ETNA), 43: 244–254, 2014–2015. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
43.2014-2015/pp244-254.
dir/pp244-254.pdf;   http:
//etna.mcs.kent.edu/volumes/
2011-2020/vol43/abstract.
php?vol=43&pages=244-
254.

Osborne:2007:SLS

[Osb08] M. R. Osborne. Separable least squares, variable projection, and the Gauss– Newton algorithm. Electronic Transactions on Numerical Analysis (ETNA), 28:1–15, 2007–2008. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.28.2007-2008/
pp1-15.dir/pp1-15.pdf.

Special volume for Gene Golub.

Orive:2020:CFG

[OSLS20] Ramón Orive, Juan C. Santos-León, and Miodrag M. Spalević. Cubature formulae for the Gaussian weight. Some old and new rules. Electronic Transactions on Numerical Analysis (ETNA), 53:426–438, 2020. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.53.2020/pp426-
438.dir/pp426-438.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol53/abstract.php?vol=
53&pages=426-438.

Ostrovska:2010:AAF

[Ost10] Sofiya Ostrovska. On the approximation of analytic functions by the q-Bernstein polynomials in the case q > 1. Electronic Transactions on Numerical Analysis (ETNA), 37:105–112, 2010. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
37.2010/pp105-112.dir/
pp105-112.pdf.

Pandur:2019:PGI

[Pan19] Marija Miloloža Pandur.

Preconditioned gradient iterations for the eigenproblem of definite matrix pairs. Electronic Transactions on Numerical Analysis (ETNA), 51:331–362, 2019. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.51.2019/pp331-
362.dir/pp331-362.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol51/abstract.php?vol=
51&pages=331-362.

Park:2021:PLC

[Par21] Jongho Park. Pseudo-linear convergence of an additive Schwarz method for dual total variation minimization. Electronic Transactions on Numerical Analysis (ETNA), 54:176–197, 2021. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.54.2021/pp176-
                                        197.dir/pp176-197.pdf;
                                         http://etna.mcs.kent.
                                        edu/volumes/2021-2030/
                                        vol54/abstract.php?vol=
                                        54&pages=176-197.

Pavarino:1999:DDA

[Pav99] Luca F. Pavarino. Domain decomposition algorithms for first-order system least squares methods. Electronic Transactions on Numerical Analysis (ETNA), 8:1–14, 1999. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        8.1999/pp1-14.dir/pp1-
                                        14.pdf.

Pinar:2006:CAC

[PCP06] Ali Pinar, Edmond Chow, and Alex Pothen. Combinatorial algorithms for computing column space bases that have sparse inverses. Electronic Transactions on Numerical Analysis (ETNA), 22:122–145, 2006.

CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
22.2006/pp122-145.dir/
pp122-145.pdf.

Pechstein:2017:UFA

[PD17] Clemens Pechstein and Clark R. Dohrmann. A unified framework for adaptive BDDC. Electronic Transactions on Numerical Analysis (ETNA), 46:273–336, 2017. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.46.2017/pp273-
336.dir/pp273-336.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol46/abstract.php?vol=
46&pages=273-336.

Pearson:2015:DPR

[Pea15] John W. Pearson. On the development of parameterrobust preconditioners and commutator arguments for solving Stokes control problems. Electronic Transactions on Numerical Analysis (ETNA), 44:53–72, 2015. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.44.2015/pp53-
72.dir/pp53-72.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/

                                        vol44/abstract.php?vol=
                                        44&pages=53-72.

Pedersen:2007:PFR

[Ped07] Henrik L. Pedersen. Pick functions related to entire functions having negative zeros. Electronic Transactions on Numerical Analysis (ETNA), 27:94–112, 2007. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

                                        /etna.mcs.kent.edu/vol.
                                        27.2007/pp94-112.dir/
                                        pp94-112.pdf.

Pena:2004:LDU

[Peñ04] J. M. Peña. LDU decompositions with L and U well conditioned. Electronic Transactions on Numerical Analysis (ETNA), 18:198–208, 2004. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        18.2004/pp198-208.dir/
                                        pp198-208.pdf.

Pernice:2000:HMM

[Per00] Michael Pernice. A hybrid multigrid method for the steady-state incompressible Navier–Stokes equations. Electronic Transactions on Numerical Analysis (ETNA), 10:74–91, 2000. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.

edu/vol.10.2000/pp74-
91.dir/pp74-91.pdf. Mul-

tilevel methods (Copper Mountain, CO, 1999).

Pishbin:2011:PEE

[PGH11] S. Pishbin, F. Ghoreishi, and M. Hadizadeh. A posteriori error estimation for the Legendre collocation method applied to integralalgebraic Volterra equations. Electronic Transactions on Numerical Analysis (ETNA), 38:327–346, 2011. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
38.2011/pp327-346.dir/
pp327-346.pdf.

Pereyra:2008:FWP

[PK08] V. Pereyra and B. Kaelin.

Fast wave propagation by model order reduction. Electronic Transactions on Numerical Analysis (ETNA), 30:406–419, 2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
30.2008/pp406-419.dir/
pp406-419.pdf.

Poole:2003:AAC

[PLM03] Gene Poole, Yong-Cheng Liu, and Jan Mandel. Advancing analysis capabilities in ANSYS through solver technology. Electronic Transactions on Numerical Analysis (ETNA), 15:106–121, 2003. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
15.2003/pp106-121.dir/
pp106-121.pdf.

Paluszny:2008:BII

[PMLFT09] Marco Paluszny, Miguel Martín-Landrove, Giovanni Figueroa, and Wuilian Torres. Boosting the inverse interpolation problem by a sum of decaying exponentials using an algebraic approach. Electronic Transactions on Numerical Analysis (ETNA), 34: 163–169, 2008–2009. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
34.2008-2009/pp163-169.
dir/pp163-169.pdf.

Popa:2007:AMS

[Pop08] Constantin Popa. Algebraic multigrid smoothing property of Kaczmarz’s relaxation for general rectangular linear systems. Electronic Transactions on Numerical Analysis (ETNA), 29:150–162, 2007–2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
29.2007-2008/pp150-162.
dir/pp150-162.pdf.

Pothen:2005:DAG

[Pot05] Alex Pothen. Dedication to Alan George. Electronic Transactions on Numerical Analysis (ETNA), 21: vii–xviii, 2005. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
21.2005/dedication.pdf.

Putinar:2006:RRE

[PP06] Gabriela Putinar and Mihai Putinar. Remarks on restriction eigenfunctions in C^{n}. Electronic Transactions on Numerical Analysis (ETNA), 25:393–408, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
25.2006/pp393-408.dir/
pp393-408.pdf.

Paige:2008:MBR

[PP08] Christopher C. Paige and Ivo Panayotov. Majorization bounds for Ritz values of Hermitian matrices. Electronic Transactions on Numerical Analysis (ETNA), 31:1–11, 2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
31.2008/pp1-11.dir/pp1-
11.pdf.

Pindza:2013:IEP

[PPN13] Edson Pindza, Kailash C. Patidar, and Edgard Ngounda. Implicit-explicit predictorcorrector methods combined with improved spectral methods for pricing European style vanilla and exotic options. Electronic Transactions on Numerical Analysis (ETNA), 40: 268–293, 2013. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu//vol.
40.2013/pp268-293.dir/
pp268-293.pdf.

Pozza:2018:LAC

[PPS18] Stefano Pozza, Miroslav S. Pranić, and Zdeněk Strakoš. The Lanczos algorithm and complex Gauss quadrature. Electronic Transactions on Numerical Analysis (ETNA), 50:1–19, 2018. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
50.2018/pp1-19.dir/pp1-
19.pdf;   http://etna.
mcs.kent.edu/volumes/
2011-2020/vol50/abstract.
php?vol=50&pages=1-19.

Portilla:2006:WTW

[PQRT06] Ana Portilla, Yamilet Quintana, José M. Rodríguez, and Eva Tourís. Weierstrass’ theorem in weighted Sobolev spaces with k derivatives: announcement of results. Electronic Transactions on Numerical Analysis (ETNA), 24: 103–107, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
24.2006/pp103-107.dir/
pp103-107.pdf.

Pes:2020:MNG

[PR20] Federica Pes and Giuseppe Rodriguez. The minimalnorm Gauss–Newton method and some of its regularized variants. Electronic Transactions on Numerical Analysis (ETNA), 53:459–480, 2020. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.53.2020/pp459-
480.dir/pp459-480.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol53/abstract.php?vol=
53&pages=459-480.

Pritsker:1996:RSL

[Pri96] I. E. Pritsker. Ray sequences of Laurent-type rational functions. Electronic Transactions on Numerical Analysis (ETNA), 4: 106–124, 1996. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.

                                        4.1996/pp106-124.dir/
                                        pp106-124.pdf.

Pritsker:2006:DPW

[Pri06]  Igor E. Pritsker.  Dis-

tribution of primes and a weighted energy problem. Electronic Transactions on Numerical Analysis (ETNA), 25:259–277, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

                                        /etna.mcs.kent.edu/vol.
                                        25.2006/pp259-277.dir/
                                        pp259-277.pdf.

Pryer:2014:DGM

[Pry14]  Tristan Pryer.  Discon-

tinuous Galerkin methods for the p-biharmonic equation from a discrete variational perspective. Electronic Transactions on Numerical Analysis (ETNA), 41:328–349, 2014. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.41.2014/pp328-
                                        349.dir/pp328-349.pdf;
                                         http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol41/abstract.php?vol=
                                        41&pages=328-349.

Psarrakos:2004:ENR

[Psa04] Panayiotis J. Psarrakos. On the estimation of the qnumerical range of monic matrix polynomials. Electronic Transactions on Numerical Analysis (ETNA), 17:1–10, 2004. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
17.2004/pp1-10.dir/pp1-
10.pdf.

Pardhanani:1997:SMS

[PSC97] Anand L. Pardhanani, William F. Spotz, and Graham F. Carey. A stable multigrid strategy for convection-diffusion using high order compact discretization. Electronic Transactions on Numerical Analysis (ETNA), 6: 211–223, 1997. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
6.1997/pp211-223.dir/
pp211-223.pdf. Special

issue on multilevel methods (Copper Mountain, CO, 1997).

Persson:1999:CMM

[PSS99] Ingemar Persson, Klas Samuelsson, and Anders Szepessy. On the convergence of multigrid methods for flow problems. Electronic Transactions on Numerical Analysis (ETNA), 8:46–87, 1999. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.

                                        edu/vol.8.1999/pp46-87.
                                        dir/pp46-87.pdf.

Potts:2013:PEM

[PT13] Daniel Potts and Manfred Tasche. Parameter estimation for multivariate exponential sums. Electronic Transactions on Numerical Analysis (ETNA), 40: 204–224, 2013. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu//vol.
                                        40.2013/pp204-224.dir/
                                        pp204-224.pdf.

Pinar:2005:FNS

[PV05] Ali Pinar and Virginia Vassilevska. Finding nonoverlapping substructures of a sparse matrix. Electronic Transactions on Numerical Analysis (ETNA), 21: 107–124, 2005. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        21.2005/pp107-124.dir/
                                        pp107-124.pdf.

Perego:2009:EGR

[PV09] Mauro Perego and Alessandro Veneziani. An efficient generalization of the Rush– Larsen method for solving electro-physiology membrane equations. Electronic Transactions on Numerical Analysis (ETNA), 35: 234–256, 2009. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
35.2009/pp234-256.dir/
pp234-256.pdf.

Pearson:2013:FIS

[PW13] John W. Pearson and Andrew J. Wathen. Fast iterative solvers for convectiondiffusion control problems. Electronic Transactions on Numerical Analysis (ETNA), 40:294–310, 2013. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu//vol.
40.2013/pp294-310.dir/
pp294-310.pdf.

Qiu:2015:EPP

[QvGvW^{+}15] Yue Qiu, Martin B. van Gijzen, Jan-Willem van Wingerden, Michel Verhaegen, and Cornelis Vuik. Efficient preconditioners for PDE-constrained optimization problem with a multilevel sequentially semiseparable matrix structure. Electronic Transactions on Numerical Analysis (ETNA), 44:367–400, 2015. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.44.2015/pp367-
400.dir/pp367-400.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/

                                        vol44/abstract.php?vol=
                                        44&pages=367-400.

Rajagopal:2008:MEP

[Raj08] Kumbakonam R. Rajagopal. On the modeling of entropy producing processes. Electronic Transactions on Numerical Analysis (ETNA), 31:40–48, 2008. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.31.2008/pp40-
                                        48.dir/pp40-48.pdf.

Ramlau:2008:RPT

[Ram08] Ronny Ramlau. Regularization properties of Tikhonov regularization with sparsity constraints. Electronic Transactions on Numerical Analysis (ETNA), 30:54–74, 2008. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.30.2008/pp54-
                                        74.dir/pp54-74.pdf.

Randriambelosoa:2007:PBC

[Ran07] Germain E. Randriambelosoa. Polynomial best constrained degree reduction in strain energy. Electronic Transactions on Numerical Analysis (ETNA), 26: 312–319, 2007. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
26.2007/pp312-319.dir/
pp312-319.pdf.

Rappoport:2006:PIN

[Rap06] Juri M. Rappoport. The properties, inequalities and numerical approximation of modified Bessel functions. Electronic Transactions on Numerical Analysis (ETNA), 25:454–466, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
25.2006/pp454-466.dir/
pp454-466.pdf.

Rodriguez:2006:GWS

[RÁRP06] José M. Rodríguez, Venancio Álvarez, Elena Romera, and Domingo Pestana. Generalized weighted Sobolev spaces and applications to Sobolev orthogonal polynomials: a survey. Electronic Transactions on Numerical Analysis (ETNA), 24:88–93, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.24.2006/pp88-
93.dir/pp88-93.pdf.

Ruzicka:2008:IIF

[RFHS08] Martin Růžička, Miloslav Feistauer, Jaromír Horáček, and Petr Sváček. Interaction of incompressible flow and a moving airfoil. Electronic Transactions on Numerical Analysis (ETNA), 32:123–133, 2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
32.2008/pp123-133.dir/
pp123-133.pdf.

Rodriguez:1999:QFR

[RGVP99] F. Cala Rodriguez, P. Gonzalez-Vera, and M. Jimenez Paiz. Quadrature formulas for rational functions. Electronic Transactions on Numerical Analysis (ETNA), 9:39–52, 1999. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.9.1999/pp39-52.
dir/pp39-52.pdf.

Roosta-Khorasani:2014:DCS

[RKvdDA14] Farbod Roosta-Khorasani, Kees van den Doel, and Uri Ascher. Data completion and stochastic algorithms for PDE inversion problems with many measurements. Electronic Transactions on Numerical Analysis (ETNA), 42:177–196, 2014. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.42.2014/pp177-
196.dir/pp177-196.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/

vol42/abstract.php?vol=
42&pages=177-196.

Romanovski:2007:PPS

[Rom07] Valery G. Romanovski. Periodic points of some algebraic maps. Electronic Transactions on Numerical Analysis (ETNA), 27: 156–162, 2007. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
27.2007/pp156-162.dir/
pp156-162.pdf.

Rosca:2009:SQF

[Roş09] Daniela Roşca. Spherical quadrature formulas with equally spaced nodes on latitudinal circles. Electronic Transactions on Numerical Analysis (ETNA), 35: 148–163, 2009. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
35.2009/pp148-163.dir/
pp148-163.pdf.

Rozenblyum:2006:OPR

[Roz06] Alexander Rozenblyum. qorthogonal polynomials related to the quantum group U_{q}(so(5)). Electronic Transactions on Numerical Analysis (ETNA), 24:74–78, 2006. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.24.2006/pp74-
                                        78.dir/pp74-78.pdf.

Robnik:2007:DER

[RR07] Marko Robnik and Valery G. Romanovski. The difference equation related to the problem of the hydrogen atom in a strong magnetic field. Electronic Transactions on Numerical Analysis (ETNA), 27:163–176, 2007. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

                                        /etna.mcs.kent.edu/vol.
                                        27.2007/pp163-176.dir/
                                        pp163-176.pdf.

Ramlau:2010:CRR

[RR10] Ronny Ramlau and Elena Resmerita. Convergence rates for regularization with sparsity constraints. Electronic Transactions on Numerical Analysis (ETNA), 37:87–104, 2010. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        37.2010/pp87-104.dir/
                                        pp87-104.pdf.

Rutka:2012:EIP

[RS12] Przemysław Rutka and Ryszard Smarzewski. Extremal interpolatory problem of Fejér type for all classical weight functions. Electronic Transactions on Numerical Analysis (ETNA), 39:46–59, 2012. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.39.2012/pp46-
59.dir/pp46-59.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol39/abstract.php?vol=
39&pages=46-59.

Ramlau:2014:SVD

[RS14] Ronny Ramlau and Daniel B.

Szyld. Special volume: dedicated to Lothar Reichel on the occasion of his sixtieth birthday. Electronic Transactions on Numerical Analysis (ETNA), 42:i, 2014. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic).

Ramlau:2021:AWRa

[RS21a] Ronny Ramlau and Bernadett Stadler. An augmented wavelet reconstructor for atmospheric tomography. Electronic Transactions on Numerical Analysis (ETNA), 54:256–275, 2021. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.54.2021/pp256-
275.dir/pp256-275.pdf;
 http://etna.mcs.kent.
edu/volumes/2021-2030/
vol54/abstract.php?vol=
54&pages=256-275.

Ramlau:2021:AWRb

[RS21b] Ronny Ramlau and Bernadett Stadler. An augmented wavelet reconstructor for atmospheric tomography. Electronic Transactions on Numerical Analysis (ETNA), 54:256–275, 2021. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.54.2021/pp256-
                                        275.dir/pp256-275.pdf;
                                         http://etna.mcs.kent.
                                        edu/volumes/2021-2030/
                                        vol54/abstract.php?vol=
                                        54&pages=256-275.

Rozin:2005:LRS

[RT05] Elad Rozin and Sivan Toledo. Locality of reference in sparse Cholesky factorization methods. Electronic Transactions on Numerical Analysis (ETNA), 21:81–106, 2005. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        21.2005/pp81-106.dir/
                                        pp81-106.pdf.

Roos:2008:CFF

[RV08] Hans-Görg Roos and Reiner Vanselow. A comparison of four- and fivepoint difference approximations for stabilizing the one-dimensional stationary convection-diffusion equation. Electronic Transactions on Numerical Analysis (ETNA), 32:63–75, 2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.32.2008/pp63-
75.dir/pp63-75.pdf.

Reichel:2008:SSS

[RY09] Lothar Reichel and Qiang Ye. Simple square smoothing regularization operators. Electronic Transactions on Numerical Analysis (ETNA), 33:63–83, 2008– 2009. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.33.2008-2009/
pp63-83.dir/pp63-83.pdf.

Reichel:2014:MDT

[RY15] Lothar Reichel and Xuebo Yu. Matrix decompositions for Tikhonov regularization. Electronic Transactions on Numerical Analysis (ETNA), 43: 223–243, 2014–2015. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
43.2014-2015/pp223-243.
dir/pp223-243.pdf;   http:
//etna.mcs.kent.edu/volumes/
2011-2020/vol43/abstract.
php?vol=43&pages=223-
243.

Ramlau:2012:MTF

[RZ12] Ronny Ramlau and Clemens A.

Zarzer. On the minimization of a Tikhonov functional with a non-convex sparsity constraint. Electronic Transactions on Numerical Analysis (ETNA), 39:476–507, 2012. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.39.2012/pp476-
507.dir/pp476-507.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol39/abstract.php?vol=
39&pages=476-507.

Salam:2008:EAC

[SAA09] A. Salam and E. Al-Aidarous. Error analysis and computational aspects of SR factorization via optimal symplectic Householder transformations. Electronic Transactions on Numerical Analysis (ETNA), 33: 189–206, 2008–2009. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
33.2008-2009/pp189-206.
dir/pp189-206.pdf.

Sadkane:2014:DMM

[Sad15]  Miloud Sadkane.  The

Davison–Man method revisited. Electronic Transactions on Numerical Analysis (ETNA), 43:60–69, 2014– 2015. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.43.2014/pp60-
69.dir/pp60-69.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol43/abstract.php?vol=
43&pages=60-69.

Stanescu:2008:RCD

[SCCJC09] Dan Stanescu, Benito Chen-Charpentier, Brandi J. Jensen, and Patricia J. S. Colberg. Random coefficient differential models of growth of anaerobic photosynthetic bacteria. Electronic Transactions on Numerical Analysis (ETNA), 34:44–58, 2008–2009. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
34.2008-2009/pp44-58.
dir/pp44-58.pdf.

Schneider:1993:ALI

[Sch93] Claus Schneider. Analysis of the linearly implicit midpoint rule for differentialalgebraic equations. Electronic Transactions on Numerical Analysis (ETNA), 1:1–10, 1993. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
1.1993/pp1-10.dir/pp1-
10.pdf.

Schmeisser:2002:BVT

[Sch02] Gerhard Schmeisser. Bounds for Vandermonde type determinants of orthogonal polynomials. Electronic Transactions on Numerical Analysis (ETNA), 14: 152–164, 2002. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        14.2002/pp152-164.dir/
                                        pp152-164.pdf.

Schurz:2003:GTN

[Sch03] Henri Schurz. General theorems for numerical approximation of stochastic processes on the Hilbert space H_{2}([0, T], μ, R^{d}). Electronic Transactions on Numerical Analysis (ETNA), 16:50–69, 2003. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.16.2003/pp50-
                                        69.dir/pp50-69.pdf.

Schurz:2005:SNM

[Sch05]  Henri Schurz.  Stabil-

ity of numerical methods for ordinary stochastic differential equations along Lyapunov-type and other functions with variable step sizes. Electronic Transactions on Numerical Analysis (ETNA), 20:27–49, 2005. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.20.2005/pp27-
49.dir/pp27-49.pdf.

Schlosser:2006:NER

[Sch06] Michael Schlosser. Noncommutative extensions of Ramanujan’s _{1}ψ_{1} summation. Electronic Transactions on Numerical Analysis (ETNA), 24:94–102, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
24.2006/pp94-102.dir/
pp94-102.pdf.

Schieweck:2008:RBC

[Sch08a] Friedhelm Schieweck. On the role of boundary conditions for CIP stabilization of higher order finite elements. Electronic Transactions on Numerical Analysis (ETNA), 32:1–16, 2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
32.2008/pp1-16.dir/pp1-
16.pdf.

Schnurr:2008:ACS

[Sch08b] Marco Schnurr. The automatic computation of second-order slope tuples for some nonsmooth functions. Electronic Transactions on Numerical Analysis (ETNA), 30:203–223, 2008.

CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

                                        /etna.mcs.kent.edu/vol.
                                        30.2008/pp203-223.dir/
                                        pp203-223.pdf.

Schweitzer:2016:FCC

[Sch16] Marcel Schweitzer. Any finite convergence curve is possible in the initial iterations of restarted FOM. Electronic Transactions on Numerical Analysis (ETNA), 45:133–145, 2016. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.45.2016/pp133-
                                        145.dir/pp133-145.pdf;
                                         http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol45/abstract.php?vol=
                                        45&pages=133-145.

Siefert:2006:PMS

[SdS06] Chris Siefert and Eric de Sturler. Probing methods for saddle-point problems. Electronic Transactions on Numerical Analysis (ETNA), 22:163–183, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

                                        /etna.mcs.kent.edu/vol.
                                        22.2006/pp163-183.dir/
                                        pp163-183.pdf.

Salkuyeh:2014:PPG

[SEH14] Davod Khojasteh Salkuyeh, Vahid Edalatpour, and Davod Hezari. Polynomial preconditioning for the GeneRank problem. Electronic Transactions on Numerical Analysis (ETNA), 41:179–189, 2014. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.41.2014/pp179-
189.dir/pp179-189.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol41/abstract.php?vol=
41&pages=179-189.

Selig:2002:UPR

[Sel02] Kathi K. Selig. Uncertainty principles revisited. Electronic Transactions on Numerical Analysis (ETNA), 14:165–177, 2002. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
14.2002/pp165-177.dir/
pp165-177.pdf.

Sleijpen:1993:BLE

[SF93] Gerard L. G. Sleijpen and Diederik R. Fokkema. BiCGstab(l) for linear equations involving unsymmetric matrices with complex spectrum. Electronic Transactions on Numerical Analysis (ETNA), 1:11–32, 1993. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.

                                        edu/vol.1.1993/pp11-32.
                                        dir/pp11-32.pdf.

Shih:1994:MSP

[SG94] T. M. Shih and Junbin Gao. Modified Specht’s plate bending element and its convergence analysis. Electronic Transactions on Numerical Analysis (ETNA), 2:92–103, 1994. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        2.1994/pp92-103.dir/pp92-
                                        103.pdf.

Simoncini:1998:TFR

[SG98] V. Simoncini and E. Gallopoulos. Transfer functions and resolvent norm approximation of large matrices. Electronic Transactions on Numerical Analysis (ETNA), 7:190–201, 1998. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

                                        /etna.mcs.kent.edu/vol.
                                        7.1998/pp190-201.dir/
                                        pp190-201.pdf.  Large

scale eigenvalue problems (Argonne, IL, 1997).

Segura:1999:EAL

[SG99] Javier Segura and Amparo Gil. Evaluation of associated Legendre functions off the cut and parabolic cylinder functions. Electronic Transactions on Numerical Analysis (ETNA), 9:137–146, 1999. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
9.1999/pp137-146.dir/
pp137-146.pdf.

Schenk:2006:FFP

[SG06] Olaf Schenk and Klaus Gärtner. On fast factorization pivoting methods for sparse symmetric indefinite systems. Electronic Transactions on Numerical Analysis (ETNA), 23: 158–179, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
23.2006/pp158-179.dir/
pp158-179.pdf.

Stange:2007:EUR

[SGB07] Peter Stange, Andreas Griewank, and Matthias Bollhöfer. On the efficient update of rectangular LUfactorizations subject to low rank modifications. Electronic Transactions on Numerical Analysis (ETNA), 26:161–177, 2007. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
26.2007/pp161-177.dir/
pp161-177.pdf.

Sifuentes:2015:RMR

[SGG15] Josef Sifuentes, Zydrunas Gimbutas, and Leslie Greengard. Randomized methods for rank-deficient linear systems. Electronic Transactions on Numerical Analysis (ETNA), 44:177–188, 2015. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.44.2015/pp177-
188.dir/pp177-188.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol44/abstract.php?vol=
44&pages=177-188.

Shampine:2012:IOF

[Sha12] L. F. Shampine. Integrating Oscillatory Functions in Matlab, II. Electronic Transactions on Numerical Analysis (ETNA), 39:403–413, 2012. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.39.2012/pp403-
413.dir/pp403-413.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol39/abstract.php?vol=
39&pages=403-413.

Shekhtman:2004:HI

[She04]  Boris Shekhtman.  On

Hermite interpolation in R_{d}. Electronic Transactions on Numerical Analysis (ETNA), 18:65–72, 2004. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.18.2004/pp65-
72.dir/pp65-72.pdf.

Shekhtman:2006:OQE

[She06] Boris Shekhtman. On one question of Ed Saff. Electronic Transactions on Numerical Analysis (ETNA), 25:439–445, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
25.2006/pp439-445.dir/
pp439-445.pdf.

Shekhtman:2009:STI

[She10] Boris Shekhtman. Some tidbits on ideal projectors, commuting matrices and their applications. Electronic Transactions on Numerical Analysis (ETNA), 36:17–26, 2009–2010. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
36.2009-2010/pp17-26.
dir/pp17-26.pdf;   http:
//etna.mcs.kent.edu/volumes/
2001-2010/vol36/abstract.
php?vol=36&pages=17-26.

Solonen:2012:VEK

[SHH^{+}12] Antti Solonen, Heikki Haario, Janne Hakkarainen, Harri Auvinen, Idrissa Amour, and Tuomo Kauranne. Variational ensemble Kalman filtering using limited memory BFGS. Electronic Transactions on Numerical Analysis (ETNA), 39:271–285, 2012. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.39.2012/pp271-
285.dir/pp271-285.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol39/abstract.php?vol=
39&pages=271-285.

Sogabe:2008:WQR

[SHZF08] T. Sogabe, T. Hoshi, S.-L. Zhang, and T. Fujiwara. On a weighted quasi-residual minimization strategy for solving complex symmetric shifted linear systems. Electronic Transactions on Numerical Analysis (ETNA), 31:126–140, 2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
31.2008/pp126-140.dir/
pp126-140.pdf.

Silva:2003:CAD

[Sil03] Malik Silva. Cache aware data laying for the Gauss– Seidel smoother. Electronic Transactions on Numerical Analysis (ETNA), 15:66–77, 2003. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.

edu/vol.15.2003/pp66-
77.dir/pp66-77.pdf.

Simon:2006:FSZ

[Sim06] Barry Simon. Fine structure of the zeros of orthogonal polynomials, I. A tale of two pictures. Electronic Transactions on Numerical Analysis (ETNA), 25: 328–368, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
25.2006/pp328-368.dir/
pp328-368.pdf.

Simoncini:2010:NSC

[Sim10]  Valeria Simoncini.  On

a non-stagnation condition for GMRES and application to saddle point matrices. Electronic Transactions on Numerical Analysis (ETNA), 37:202–213, 2010. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
37.2010/pp202-213.dir/
pp202-213.pdf.

Sinap:1995:GQM

[Sin95]  Ann Sinap.  Gaussian

quadrature for matrix valued functions on the unit circle. Electronic Transactions on Numerical Analysis (ETNA), 3:96–115, 1995. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

                                        /etna.mcs.kent.edu/vol.
                                        3.1995/pp96-115.dir/pp96-
                                        115.pdf.

Schneider:2008:EFE

[SJ08] René Schneider and Peter K. Jimack. On the evaluation of finite element sensitivities to nodal coordinates. Electronic Transactions on Numerical Analysis (ETNA), 32:134–144, 2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

                                        /etna.mcs.kent.edu/vol.
                                        32.2008/pp134-144.dir/
                                        pp134-144.pdf.

Shimomukai:2006:NSN

[SK06] Kenshu Shimomukai and Hidesada Kanda. Numerical study of normal pressure distribution in entrance flow between parallel plates, I. Finite difference calculations. Electronic Transactions on Numerical Analysis (ETNA), 23:202–218, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

                                        /etna.mcs.kent.edu/vol.
                                        23.2006/pp202-218.dir/
                                        pp202-218.pdf.

Shimomukai:2008:NSN

[SK08] K. Shimomukai and H. Kanda. Numerical study of normal pressure distribution in entrance pipe flow. Electronic Transactions on Numerical Analysis (ETNA), 30:10–25, 2008. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.30.2008/pp10-
25.dir/pp10-25.pdf.

Santos-Leon:2012:FOP

[SL12] Juan-Carlos Santos-León.

Fejér orthogonal polynomials and rational modification of a measure on the unit circle. Electronic Transactions on Numerical Analysis (ETNA), 39: 340–352, 2012. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
39.2012/pp340-352.dir/.

Sete:2016:CMM

[SL16] Olivier Sète and Jörg Liesen. On conformal maps from multiply connected domains onto lemniscatic domains. Electronic Transactions on Numerical Analysis (ETNA), 45:1–15, 2016. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
45.2016/pp1-15.dir/pp1-
15.pdf;   http://etna.
mcs.kent.edu/volumes/
2011-2020/vol45/abstract.
php?vol=45&pages=1-15.

Sousedik:2008:EPD

[SM08] Bedřich Sousedík and Jan Mandel. On the equivalence of primal and dual substructuring preconditioners. Electronic Transactions on Numerical Analysis (ETNA), 31:384–402, 2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
31.2008/pp384-402.dir/
pp384-402.pdf.

Spiteri:2002:PPA

[SME02] Pierre Spiteri, Jean-Claude Miellou, and Didier El Baz. Perturbation of parallel asynchronous linear iterations by floating point errors. Electronic Transactions on Numerical Analysis (ETNA), 13:38–55, 2002. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.13.2002/pp38-
55.dir/pp38-55.pdf.

Smithies:2009:SDN

[Smi10] Laura Smithies. The structured distance to nearly normal matrices. Electronic Transactions on Numerical Analysis (ETNA), 36: 99–112, 2009–2010. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.36.2009-2010/
pp99-112.dir/pp99-112.
pdf;   http://etna.mcs.

kent.edu/volumes/2001-
2010/vol36/abstract.php?
vol=36&pages=99-112.

Sakowski:2019:CDG

[SMS^{+}19] Konrad Sakowski, Leszek Marcinkowski, Pawel Strak, Pawel Kempisty, and Stanislaw Krukowski. On the composite discontinuous Galerkin method for simulations of electric properties of semiconductor devices. Electronic Transactions on Numerical Analysis (ETNA), 51:75–98, 2019. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.51.2019/pp75-
98.dir/pp75-98.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol51/abstract.php?vol=
51&pages=75-98.

Soodhalter:2016:TRG

[Soo16] Kirk M. Soodhalter. Two recursive GMRES-type methods for shifted linear systems with general preconditioning. Electronic Transactions on Numerical Analysis (ETNA), 45:499–523, 2016. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.45.2016/pp499-
523.dir/pp499-523.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/

                                        vol45/abstract.php?vol=
                                        45&pages=499-523.

Soodhalter:2017:SBG

[Soo17] Kirk M. Soodhalter. Stagnation of block GMRES and its relationship to block FOM. Electronic Transactions on Numerical Analysis (ETNA), 46:162–189, 2017. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.46.2017/pp162-
                                        189.dir/pp162-189.pdf;
                                         http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol46/abstract.php?vol=
                                        46&pages=162-189.

Sorensen:1998:TQM

[Sor98] D. C. Sorensen. Truncated QZ methods for large scale generalized eigenvalue problems. Electronic Transactions on Numerical Analysis (ETNA), 7:141–162, 1998. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

                                        /etna.mcs.kent.edu/vol.
                                        7.1998/pp141-162.dir/
                                        pp141-162.pdf.  Large

scale eigenvalue problems (Argonne, IL, 1997).

Sprengel:2002:CMC

[Spr02] Frauke Sprengel. Comparing multilevel coarsening strategies. Electronic Transactions on Numerical Analysis (ETNA), 14:

178–194, 2002. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
14.2002/pp178-194.dir/
pp178-194.pdf.

Stathopoulos:1998:RTJ

[SS98] Andreas Stathopoulos and Yousef Saad. Restarting techniques for the (Jacobi– )Davidson symmetric eigenvalue methods. Electronic Transactions on Numerical Analysis (ETNA), 7: 163–181, 1998. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
7.1998/pp163-181.dir/
pp163-181.pdf.  Large

scale eigenvalue problems (Argonne, IL, 1997).

Schwetlick:2008:CCI

[SS08] Hubert Schwetlick and Kathrin Schreiber. A counterexample for characterizing an invariant subspace of a matrix. Electronic Transactions on Numerical Analysis (ETNA), 31: 295–305, 2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
31.2008/pp295-305.dir/
pp295-305.pdf.

Sadkane:2008:NSB

[SS09] Miloud Sadkane and Ahmed Salam. A note on symplectic block reflectors. Electronic Transactions on Numerical Analysis (ETNA), 33:45–52, 2008–2009. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        33.2008-2009/pp45-52.
                                        dir/pp45-52.pdf.

Sheehan:2010:CLP

[SSS10] Bernard N. Sheehan, Yousef Saad, and Roger B. Sidje. Computing exp(−τA)b with Laguerre polynomials. Electronic Transactions on Numerical Analysis (ETNA), 37:147–165, 2010. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        37.2010/pp147-165.dir/
                                        pp147-165.pdf.

Sekhar:2012:CFO

[SSVS12] T. V. S. Sekhar, R. Sivakumar, S. Vimala, and Y. V. S. S. Sanyasiraju. A combined fourth-order compact scheme with an accelerated multigrid method for the energy equation in spherical polar coordinates. Electronic Transactions on Numerical Analysis (ETNA), 39:32–45, 2012. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.39.2012/pp32-
45.dir/pp32-45.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol39/abstract.php?vol=
39&pages=32-45.

Strakos:2002:EEC

[ST02] Zdeněk Strakoš and Petr Tichý. On error estimation in the conjugate gradient method and why it works in finite precision computations. Electronic Transactions on Numerical Analysis (ETNA), 13:56–80, 2002. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.13.2002/pp56-
80.dir/pp56-80.pdf.

Schmelzer:2007:EMF

[ST08] Thomas Schmelzer and Lloyd N. Trefethen. Evaluating matrix functions for exponential integrators via Carathéodory–Fejér approximation and contour integrals. Electronic Transactions on Numerical Analysis (ETNA), 29:1–18, 2007– 2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.29.2007-2008/
pp1-18.dir/pp1-18.pdf.

Stahl:2002:AQH

[Sta02] Herbert Stahl. Asymptotics for quadratic Hermite– Padé polynomials associated with the exponential function. Electronic Transactions on Numerical Analysis (ETNA), 14: 195–222, 2002. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        14.2002/pp195-222.dir/
                                        pp195-222.pdf.

Stals:2003:CNL

[Sta03]  Linda Stals.  Compari-

son of non-linear solvers for the solution of radiation transport equations. Electronic Transactions on Numerical Analysis (ETNA), 15:78–93, 2003. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.15.2003/pp78-
                                        93.dir/pp78-93.pdf.

Stahl:2006:TQH

[Sta06] Herbert Stahl. From Taylor to quadratic Hermite– Padé polynomials. Electronic Transactions on Numerical Analysis (ETNA), 25:480–510, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.

25.2006/pp480-510.dir/
pp480-510.pdf.

Stewart:1995:GQD

[Ste95] G. W. Stewart. On graded QR decompositions of products of matrices. Electronic Transactions on Numerical Analysis (ETNA), 3:39–49, 1995. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.3.1995/pp39-49.
dir/pp39-49.pdf.

Sterner:1997:MSH

[Ste97a] Erik Sterner. A multigrid smoother for high Reynolds number flows. Electronic Transactions on Numerical Analysis (ETNA), 6: 234–245, 1997. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
6.1997/pp234-245.dir/
pp234-245.pdf. Special

issue on multilevel methods (Copper Mountain, CO, 1997).

Stewart:1997:NAS

[Ste97b] David E. Stewart. A new algorithm for the SVD of a long product of matrices and the stability of products. Electronic Transactions on Numerical Analysis (ETNA), 5:29–47, 1997. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.5.1997/pp29-47.
                                        dir/pp29-47.pdf.

Stempak:2002:CDF

[Ste02] Krzysztof Stempak. On convergence and divergence of Fourier–Bessel series. Electronic Transactions on Numerical Analysis (ETNA), 14:223–235, 2002. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

                                        /etna.mcs.kent.edu/vol.
                                        14.2002/pp223-235.dir/
                                        pp223-235.pdf.

Sokolov:2008:NSD

[STO08] Andriy Sokolov, Stefan Turek, and Maxim A. Olshanskii. Numerical study of a discrete projection method for rotating incompressible flows. Electronic Transactions on Numerical Analysis (ETNA), 32:49–62, 2008. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.32.2008/pp49-
                                        62.dir/pp49-62.pdf.

Stykel:2008:LRI

[Sty08]  Tatjana Stykel.  Low-

rank iterative methods for projected generalized Lyapunov equations. Electronic Transactions on Numerical Analysis (ETNA), 30:

187–202, 2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
30.2008/pp187-202.dir/
pp187-202.pdf.

Suslov:2007:AIO

[Sus07] Sergei K. Suslov. An algebra of integral operators. Electronic Transactions on Numerical Analysis (ETNA), 27:140–155, 2007. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
27.2007/pp140-155.dir/
pp140-155.pdf.

Sommariva:2010:AFP

[SV10a] A. Sommariva and M. Vianello.

Approximate Fekete points for weighted polynomial interpolation. Electronic Transactions on Numerical Analysis (ETNA), 37: 1–22, 2010. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
37.2010/pp1-22.dir/pp1-
22.pdf.

Stammberger:2009:UEP

[SV10b] Markus Stammberger and Heinrich Voss. On an unsymmetric eigenvalue problem governing free vibrations of fluid-solid structures. Electronic Transactions on Numerical Analysis (ETNA), 36: 113–125, 2009–2010. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
36.2009-2010/pp113-125.
dir/pp113-125.pdf;   http:
//etna.mcs.kent.edu/volumes/
2001-2010/vol36/abstract.
php?vol=36&pages=113-
125.

Sleijpen:1998:EES

[SvdVM98] Gerard L. G. Sleijpen, Henk A. van der Vorst, and Ellen Meijerink. Efficient expansion of subspaces in the Jacobi– Davidson method for standard and generalized eigenproblems. Electronic Transactions on Numerical Analysis (ETNA), 7:75–89, 1998. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.7.1998/pp75-89.
dir/pp75-89.pdf. Large

scale eigenvalue problems (Argonne, IL, 1997).

Shen:2008:NBS

[SY09] Yun-Qiu Shen and Tjalling J. Ypma. Numerical bifurcation of separable parameterized equations. Electronic Transactions on Numerical Analysis (ETNA), 34:31–43, 2008–2009. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
34.2008-2009/pp31-43.
dir/pp31-43.pdf.

Steinbach:2020:CST

[SZ20] Olaf Steinbach and Marco Zank. Coercive space-time finite element methods for initial boundary value problems. Electronic Transactions on Numerical Analysis (ETNA), 52:154–194, 2020. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.52.2020/pp154-
194.dir/pp154-194.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol52/abstract.php?vol=
52&pages=154-194.

Tremblay:2007:RUL

[TC08] Jean Christophe Tremblay and Tucker Carrington Jr. A refined unsymmetric Lanczos eigensolver for computing accurate eigentriplets of a real unsymmetric matrix. Electronic Transactions on Numerical Analysis (ETNA), 28: 95–113, 2007–2008. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
28.2007-2008/pp95-113.
dir/pp95-113.pdf. Special

volume for Gene Golub.

Torres:2008:NSD

[TDLC09] Carlos R. Torres, Dany De Cecchis, Germán Larrazábal, and José Castillo. Numerical study of a descending sphere in a low Reynolds number strongly stratified fluid. Electronic Transactions on Numerical Analysis (ETNA), 34: 119–124, 2008–2009. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        34.2008-2009/pp119-124.
                                        dir/pp119-124.pdf.

Tifenbach:2011:SBA

[Tif11] Ryan M. Tifenbach. On an SVD-based algorithm for identifying meta-stable states of Markov chains. Electronic Transactions on Numerical Analysis (ETNA), 38:17–33, 2011. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic).

Tifenbach:2013:CAN

[Tif13] Ryan M. Tifenbach. A combinatorial approach to nearly uncoupled Markov chains I: Reversible Markov chains. Electronic Transactions on Numerical Analysis (ETNA), 40:120–147, 2013. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu//vol.

40.2013/pp120-147.dir/
pp120-147.pdf.

Tikhonov:2005:CBN

[Tik05] Sergey Tikhonov. Characteristics of Besov–Nikol’skiĭ class of functions. Electronic Transactions on Numerical Analysis (ETNA), 19:94–104, 2005. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
19.2005/pp94-104.dir/
pp94-104.pdf.

Tichy:2007:WCG

[TLF07] Petr Tichý, Jörg Liesen, and Vance Faber. On worstcase GMRES, ideal GM-RES, and the polynomial numerical hull of a Jordan block. Electronic Transactions on Numerical Analysis (ETNA), 26:453–473, 2007. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
26.2007/pp453-473.dir/
pp453-473.pdf.

Teng:2015:PPL

[TLL15] Zhongming Teng, Linzhang Lu, and Ren-Cang Li. Perturbation of partitioned linear response eigenvalue problems. Electronic Transactions on Numerical Analysis (ETNA), 44:624–638, 2015. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.44.2015/pp624-
                                        638.dir/pp624-638.pdf;
                                         http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol44/abstract.php?vol=
                                        44&pages=624-638.

Thole:1997:FSM

[TMS97] C. A. Thole, S. Mayer, and A. Supalov. Fast solution of MSC/NASTRAN sparse matrix problems using a multilevel approach. Electronic Transactions on Numerical Analysis (ETNA), 6:246–254, 1997. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        6.1997/pp246-254.dir/
                                        pp246-254.pdf. Special

issue on multilevel methods (Copper Mountain, CO, 1997).

Toselli:2000:NNM

[Tos00] Andrea Toselli. Neumann– Neumann methods for vector field problems. Electronic Transactions on Numerical Analysis (ETNA), 11:1–24, 2000. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        11.2000/pp1-24.dir/pp1-
                                        24.pdf.

Tomeo:2005:TCN

[TT05] Venancio Tomeo and Emilio Torrano. Three cases of normality of Hessenberg’s matrix related with atomic complex distributions. Electronic Transactions on Numerical Analysis (ETNA), 19:105–112, 2005. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
19.2005/pp105-112.dir/
pp105-112.pdf.

Tu:2005:BAM

[Tu05] Xuemin Tu. A BDDC algorithm for a mixed formulation of flow in porous media. Electronic Transactions on Numerical Analysis (ETNA), 20:164–179, 2005. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
20.2005/pp164-179.dir/
pp164-179.pdf.

Tu:2007:BAF

[Tu07] Xuemin Tu. A BDDC algorithm for flow in porous media with a hybrid finite element discretization. Electronic Transactions on Numerical Analysis (ETNA), 26:146–160, 2007. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.

                                        26.2007/pp146-160.dir/
                                        pp146-160.pdf.

Tang:2007:EDM

[TV07] J. M. Tang and C. Vuik. Efficient deflation methods applied to 3D bubbly flow problems. Electronic Transactions on Numerical Analysis (ETNA), 26: 330–349, 2007. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        26.2007/pp330-349.dir/
                                        pp330-349.pdf.

Tezaur:2018:ICB

[TV18] Radek Tezaur and Petr Vaněk. Improved convergence bounds for two-level methods with an aggressive coarsening and massive polynomial smoothing. Electronic Transactions on Numerical Analysis (ETNA), 48:264–285, 2018. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.48.2018/pp264-
                                        285.dir/pp264-285.pdf;
                                         http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol48/abstract.php?vol=
                                        48&pages=264-285.

Tu:2016:BAS

[TW16] Xuemin Tu and Bin Wang. A BDDC algorithm for second-order elliptic problems with hybridizable discontinuous Galerkin discretizations. Electronic Transactions on Numerical Analysis (ETNA), 45: 354–370, 2016. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.45.2016/pp354-
370.dir/pp354-370.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol45/abstract.php?vol=
45&pages=354-370.

Teng:2017:BLM

[TZ17] Zhongming Teng and Lei-Hong Zhang. A block Lanczos method for the linear response eigenvalue problem. Electronic Transactions on Numerical Analysis (ETNA), 46:505–523, 2017. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.46.2017/pp505-
523.dir/pp505-523.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol46/abstract.php?vol=
46&pages=505-523.

Villamizar:2008:EGN

[VA09] Vianey Villamizar and Sebastian Acosta. Elliptic grids with nearly uniform cell area and line spacing. Electronic Transactions on Numerical Analysis (ETNA), 34: 59–75, 2008–2009. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        34.2008-2009/pp59-75.
                                        dir/pp59-75.pdf.

Valkonen:2019:BPM

[Val19] Tuomo Valkonen. Blockproximal methods with spatially adapted acceleration. Electronic Transactions on Numerical Analysis (ETNA), 51:15–49, 2019. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.51.2019/pp15-
                                        49.dir/pp15-49.pdf;
                                        http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol51/abstract.php?vol=
                                        51&pages=15-49.

VanDeun:2007:EGP

[Van07] Joris Van Deun. Electrostatics and ghost poles in near best fixed pole rational interpolation. Electronic Transactions on Numerical Analysis (ETNA), 26: 439–452, 2007. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        26.2007/pp439-452.dir/
                                        pp439-452.pdf.

VanZandt:2019:ECR

[Van19] James R. Van Zandt. Efficient cubature rules. Electronic Transactions on Numerical Analysis (ETNA), 51:219–239, 2019. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.51.2019/pp219-
239.dir/pp219-239.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol51/abstract.php?vol=
51&pages=219-239.

Varga:2001:GTE

[Var01] Richard S. Varga. Geršgorintype eigenvalue inclusion theorems and their sharpness. Electronic Transactions on Numerical Analysis (ETNA), 12:113–133, 2001. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
12.2001/pp113-133.dir/
pp113-133.pdf.

VanBarel:1995:OPV

[VB95] M. Van Barel and A. Bultheel.

Orthonormal polynomial vectors and least squares approximation for a discrete inner product. Electronic Transactions on Numerical Analysis (ETNA), 3:1–23, 1995. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
3.1995/pp1-23.dir/pp1-
23.pdf.

vanBarel:2014:AOP

[vBHS14] Marc van Barel, Matthias Humet, and Laurent Sorber. Approximating optimal point configurations for multivariate polynomial interpolation. Electronic Transactions on Numerical Analysis (ETNA), 42:41–63, 2014. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.42.2014/pp41-
63.dir/pp41-63.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol42/abstract.php?vol=
42&pages=41-63.

Viswanathan:2014:FRS

[VC14] Puthan Veedu Viswanathan and Arya Kumar Bedabrata Chand. α-fractal rational splines for constrained interpolation. Electronic Transactions on Numerical Analysis (ETNA), 41: 420–442, 2014. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.41.2014/pp420-
442.dir/pp420-442.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/

vol41/abstract.php?vol=
41&pages=420-442.

Viswanathan:2015:MCA

[VC15] Puthan Veedu Viswanathan and Arya Kumar Bedabrata Chand. Monotone– comonotone approximation by fractal cubic splines and polynomials. Electronic Transactions on Numerical Analysis (ETNA), 44:639–659, 2015. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.44.2015/pp639-
659.dir/pp639-659.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol44/abstract.php?vol=
44&pages=639-659.

Varga:2008:AMG

[VCK08] Richard S. Varga, Ljiljana Cvetković, and Vladimir Kostić. Approximation of the minimal Geršgorin set of a square complex matrix. Electronic Transactions on Numerical Analysis (ETNA), 30:398–405, 2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
30.2008/pp398-405.dir/
pp398-405.pdf.

Varga:2008:DMZ

[VCL08] Richard S. Varga, Amos J.

Carpenter, and Bryan W.

Lewis. The dynamical motion of the zeros of the partial sums of e^{z}, and its relationship to discrepancy theory. Electronic Transactions on Numerical Analysis (ETNA), 30:128–143, 2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
30.2008/pp128-143.dir/
pp128-143.pdf.

Veselic:1996:NAS

[Ves96] K. Veselić. A note on the accuracy of symmetric eigenreduction algorithms. Electronic Transactions on Numerical Analysis (ETNA), 4:37–45, 1996. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.4.1996/pp37-45.
dir/pp37-45.pdf.

Varga:1995:MGS

[VK95] Richard S. Varga and Alan Krautstengl. Minimal Gerschgorin sets for partitioned matrices III. Sharpness of boundaries and monotonicity as a function of the partition. Electronic Transactions on Numerical Analysis (ETNA), 3:83–95, 1995. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.

edu/vol.3.1995/pp83-95.
dir/pp83-95.pdf.

Varga:1999:GTP

[VK99] Richard S. Varga and Alan Krautstengl. On Geršgorintype problems and ovals of Cassini. Electronic Transactions on Numerical Analysis (ETNA), 8:15–20, 1999. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.8.1999/pp15-20.
dir/pp15-20.pdf.

Vandebril:2007:SLS

[VMV07] Raf Vandebril, Nicola Mastronardi, and Marc Van Barel. Solving linear systems with a Levinson-like solver. Electronic Transactions on Numerical Analysis (ETNA), 26:243–269, 2007. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
26.2007/pp243-269.dir/
pp243-269.pdf.

Voss:1999:BME

[Vos99] Heinrich Voss. Bounds for the minimum eigenvalue of a symmetric Toeplitz matrix. Electronic Transactions on Numerical Analysis (ETNA), 8:127–137, 1999. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.

                                        8.1999/pp127-137.dir/
                                        pp127-137.pdf.

Voss:2003:RSP

[Vos03] Heinrich Voss. A rational spectral problem in fluidsolid vibration. Electronic Transactions on Numerical Analysis (ETNA), 16: 94–106, 2003. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        16.2003/pp94-106.dir/
                                        pp94-106.pdf.

VanAssche:2018:MHP

[VV18] Walter Van Assche and Anton Vuerinckx. Multiple Hermite polynomials and simultaneous Gaussian quadrature. Electronic Transactions on Numerical Analysis (ETNA), 50:182–198, 2018. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.50.2018/pp182-
                                        198.dir/pp182-198.pdf;
                                         http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol50/abstract.php?vol=
                                        50&pages=182-198.

Vandebril:2008:PQF

[VVM08] Raf Vandebril, Marc Van Barel, and Nicola Mastronardi. A parallel QRfactorization/solver of quasiseparable matrices. Electronic Transactions on Numerical Analysis (ETNA), 30:144–167, 2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
30.2008/pp144-167.dir/
pp144-167.pdf.

Vandebril:2008:NIC

[VVM09] Raf Vandebril, Marc Van Barel, and Nicola Mastronardi. A new iteration for computing the eigenvalues of semiseparable (plus diagonal) matrices. Electronic Transactions on Numerical Analysis (ETNA), 33:126–150, 2008–2009. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
33.2008-2009/pp126-150.
dir/pp126-150.pdf.

VanBuggenhout:2019:BRK

[VVV19] Niel Van Buggenhout, Marc Van Barel, and Raf Vandebril. Biorthogonal rational Krylov subspace methods. Electronic Transactions on Numerical Analysis (ETNA), 51:451–468, 2019. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.51.2019/pp451-
468.dir/pp451-468.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/

                                        vol51/abstract.php?vol=
                                        51&pages=451-468.

Vassilevski:1997:CMM

[VW97] P. S. Vassilevski and J. G. Wade. A comparison of multilevel methods for total variation regularization. Electronic Transactions on Numerical Analysis (ETNA), 6:255–270, 1997. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

                                        /etna.mcs.kent.edu/vol.
                                        6.1997/pp255-270.dir/
                                        pp255-270.pdf. Special

issue on multilevel methods (Copper Mountain, CO, 1997).

Vandebril:2013:EQA

[VW13] Raf Vandebril and David S. Watkins. An extension of the QZ algorithm beyond the Hessenberg-upper triangular pencil. Electronic Transactions on Numerical Analysis (ETNA), 40:17–35, 2013. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.40.2013/pp17-
                                        35.dir/pp17-35.pdf;
                                        http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol40/abstract.php?vol=
                                        40&pages=17-35.

Winkler:2008:SLR

[WA08] Joab R. Winkler and John D. Allan. Structured low rank approximations of the Sylvester resultant matrix for approximate GCDs of Bernstein basis polynomials. Electronic Transactions on Numerical Analysis (ETNA), 31:141–155, 2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
31.2008/pp141-155.dir/
pp141-155.pdf.

Wachsmuth:2013:ARD

[Wac13] Daniel Wachsmuth. Adaptive regularization and discretization of bangbang optimal control problems. Electronic Transactions on Numerical Analysis (ETNA), 40:249–267, 2013. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu//vol.
40.2013/pp249-267.dir/
pp249-267.pdf.

Watkins:2006:CWB

[Wat06a] David S. Watkins. A case where balancing is harmful. Electronic Transactions on Numerical Analysis (ETNA), 23:1–4, 2006. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
23.2006/pp1-4.dir/pp1-
4.pdf.

Watkins:2006:RHM

[Wat06b] David S. Watkins. On the reduction of a Hamiltonian matrix to Hamiltonian Schur form. Electronic Transactions on Numerical Analysis (ETNA), 23: 141–157, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        23.2006/pp141-157.dir/
                                        pp141-157.pdf.

Wohlmuth:2008:NAS

[WD08] M. Wohlmuth and M. Dobrowolski. Numerical analysis of Stokes equations with improved LBB dependency. Electronic Transactions on Numerical Analysis (ETNA), 32:173–189, 2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

                                        /etna.mcs.kent.edu/vol.
                                        32.2008/pp173-189.dir/
                                        pp173-189.pdf.

Wakam:2013:PRG

[WE13] Desire Nuentsa Wakam and Jocelyne Erhel. Parallelism and robustness in GMRES with a Newton basis and deflated restarting. Electronic Transactions on Numerical Analysis (ETNA), 40:381–406, 2013. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu//vol.

40.2013/pp381-406.dir/
pp381-406.pdf.

Weiss:1994:MPS

[Wei94] Rüdiger Weiss. Minimization properties and short recurrences for Krylov subspace methods. Electronic Transactions on Numerical Analysis (ETNA), 2:57–75, 1994. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.2.1994/pp57-75.
dir/pp57-75.pdf.

Willems:2011:BFT

[WL11] Paul R. Willems and Bruno Lang. Block factorizations and qd-type transformations for the MR^{3} algorithm. Electronic Transactions on Numerical Analysis (ETNA), 38:363–400, 2011. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
38.2011/pp363-400.dir/
pp363-400.pdf.

Willems:2012:GAB

[WL12] Paul R. Willems and Bruno Lang. The MR^{3}–GK algorithm for the bidiagonal SVD. Electronic Transactions on Numerical Analysis (ETNA), 39:1–21, 2012. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.

39.2012/pp1-21.dir/pp1-
21.pdf;   http://etna.
mcs.kent.edu/volumes/
2011-2020/vol39/abstract.
php?vol=39&pages=1-21.

Wang:2018:FSQ

[WLHH18] Hsuan-Hao Wang, Yi-Su Lo, Feng-Tai Hwang, and Feng-Nan Hwang. A full-space quasi-Lagrange– Newton–Krylov algorithm for trajectory optimization problems. Electronic Transactions on Numerical Analysis (ETNA), 49:103–125, 2018. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.49.2018/pp103-
125.dir/pp103-125.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol49/abstract.php?vol=
49&pages=103-125.

Wollstein:2002:UAF

[WLR02] Dirk Wollstein, Torsten Linss, and Hans-Görg Roos. A uniformly accurate finite volume discretization for a convection-diffusion problem. Electronic Transactions on Numerical Analysis (ETNA), 13:1–11, 2002. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
13.2002/pp1-11.dir/pp1-
11.pdf.

Wang:2018:ARA

[WNK18] Xiaofei Wang, Carmeliza Navasca, and Stefan Kindermann. On accelerating the regularized alternating least-squares algorithm for tensors. Electronic Transactions on Numerical Analysis (ETNA), 48:1–14, 2018. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
48.2018/pp1-14.dir/pp1-
14.pdf;   http://etna.
mcs.kent.edu/volumes/
2011-2020/vol48/abstract.
php?vol=48&pages=1-14.

Washio:1997:KSA

[WO97] T. Washio and C. W. Oosterlee. Krylov subspace acceleration for nonlinear multigrid schemes. Electronic Transactions on Numerical Analysis (ETNA), 6:271–290, 1997. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
6.1997/pp271-290.dir/
pp271-290.pdf. Special

issue on multilevel methods (Copper Mountain, CO, 1997).

Wohlmuth:2000:MMS

[Woh00] Barbara I. Wohlmuth. A multigrid method for saddle point problems arising from mortar finite element discretizations. Electronic Transactions on Numerical Analysis (ETNA), 11:43–54, 2000. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.11.2000/pp43-
54.dir/pp43-54.pdf.

Wathen:2008:CSI

[WR09] Andy Wathen and Tyrone Rees. Chebyshev semiiteration in preconditioning for problems including the mass matrix. Electronic Transactions on Numerical Analysis (ETNA), 34: 125–135, 2008–2009. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
34.2008-2009/pp125-135.
dir/pp125-135.pdf.

Wu:1998:INP

[WSS98] Kesheng Wu, Yousef Saad, and Andreas Stathopoulos. Inexact Newton preconditioning techniques for large symmetric eigenvalue problems. Electronic Transactions on Numerical Analysis (ETNA), 7:202–214, 1998. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
7.1998/pp202-214.dir/
pp202-214.pdf.  Large

scale eigenvalue problems (Argonne, IL, 1997).

Wu:2003:PCM

[Wu03] Limin Wu. A parameter choice method for Tikhonov regularization. Electronic Transactions on Numerical Analysis (ETNA), 16: 107–128, 2003. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
16.2003/pp107-128.dir/
pp107-128.pdf.

Xu:2019:CRN

[XBC19] Wei-Ru Xu, Natália Bebiano, and Guo-Liang Chen. On the construction of real non-selfadjoint tridiagonal matrices with prescribed three spectra. Electronic Transactions on Numerical Analysis (ETNA), 51:363–386, 2019. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.51.2019/pp363-
386.dir/pp363-386.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol51/abstract.php?vol=
51&pages=363-386.

Xenophontos:2020:IAS

[XS20] Christos Xenophontos and Irene Sykopetritou. Isogeometric analysis for singularly perturbed problems in 1-D: error estimates. Electronic Transactions on Numerical Analysis (ETNA), 52:1–25, 2020. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
52.2020/pp1-25.dir/pp1-
25.pdf;   http://etna.
mcs.kent.edu/volumes/
2011-2020/vol52/abstract.
php?vol=52&pages=1-25.

Xu:2006:AUB

[Xu06] Yuan Xu. Analysis on the unit ball and on the simplex. Electronic Transactions on Numerical Analysis (ETNA), 25:284–301, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
25.2006/pp284-301.dir/
pp284-301.pdf.

Xu:2018:IDT

[Xu18] Yingxiang Xu. The influence of domain truncation on the performance of optimized Schwarz methods. Electronic Transactions on Numerical Analysis (ETNA), 49:182–209, 2018. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.49.2018/pp182-
209.dir/pp182-209.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol49/abstract.php?vol=
49&pages=182-209.

Yamamoto:2006:AHT

[Yam06a] Toshihiro Yamamoto. Approximation of the Hilbert transform via use of sinc convolution. Electronic Transactions on Numerical Analysis (ETNA), 23: 320–328, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
23.2006/pp320-328.dir/
pp320-328.pdf.

Yamamoto:2006:TSG

[Yam06b] Toshihiro Yamamoto. Toward the Sinc–Galerkin method for the Poisson problem in one type of curvilinear coordinate domain. Electronic Transactions on Numerical Analysis (ETNA), 23:63–75, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.23.2006/pp63-
75.dir/pp63-75.pdf.

Yang:1998:CAI

[Yan98] Chao Yang. Convergence analysis of an inexact truncated RQ-iteration. Electronic Transactions on Numerical Analysis (ETNA), 7:40–55, 1998. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.7.1998/pp40-55.

                                        dir/pp40-55.pdf. Large

scale eigenvalue problems (Argonne, IL, 1997).

Yattselev:2006:NSR

[Yat06] M. Yattselev. A note on the sharpness of the Remeztype inequality for homogeneous polynomials on the sphere. Electronic Transactions on Numerical Analysis (ETNA), 25:278–283, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

                                        /etna.mcs.kent.edu/vol.
                                        25.2006/pp278-283.dir/
                                        pp278-283.pdf.

Yang:1996:RCR

[YG96] Y. D. Yang and J. B. Gao. Remarks on the Ciarlet– Raviart mixed finite element. Electronic Transactions on Numerical Analysis (ETNA), 4:158–164, 1996. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

                                        /etna.mcs.kent.edu/vol.
                                        4.1996/pp158-164.dir/
                                        pp158-164.pdf.

Yazdi:2008:EES

[YG08] M. Hadizadeh Yazdi and Gh. Kazemi Gelian. Error estimate in the sinc collocation method for Volterra– Fredholm integral equations based on DE transformation. Electronic Transactions on Numerical Analysis (ETNA), 30:75–87, 2008.

CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.30.2008/pp75-
87.dir/pp75-87.pdf.

Yueh:2020:OSP

[YHL^{+}20] Mei-Heng Yueh, Hsiao-Han Huang, Tiexiang Li, Wen-Wei Lin, and Shing-Tung Yau. Optimized surface parameterizations with applications to Chinese virtual broadcasting. Electronic Transactions on Numerical Analysis (ETNA), 53:383–405, 2020. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.53.2020/pp383-
405.dir/pp383-405.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol53/abstract.php?vol=
53&pages=383-405.

Yamamoto:2015:REA

[YNYF15] Yusaku Yamamoto, Yuji Nakatsukasa, Yuka Yanagisawa, and Takeshi Fukaya. Roundoff error analysis of the CholeskyQR2 algorithm. Electronic Transactions on Numerical Analysis (ETNA), 44:306–326, 2015. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.44.2015/pp306-

                                        326.dir/pp306-326.pdf;
                                         http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol44/abstract.php?vol=
                                        44&pages=306-326.

Zhang:2009:RSI

[ZB10] Cheng-Yi Zhang and Michele Benzi. P -regular splitting iterative methods for non– Hermitian positive definite linear systems. Electronic Transactions on Numerical Analysis (ETNA), 36: 39–53, 2009–2010. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        36.2009-2010/pp39-53.
                                        dir/pp39-53.pdf;   http:
                                        //etna.mcs.kent.edu/volumes/
                                        2001-2010/vol36/abstract.
                                        php?vol=36&pages=39-53.

Zhong:2019:TRW

[ZC19] Hong-Xiu Zhong and Guo-Liang Chen. Thick restarting the weighted harmonic Golub–Kahan–Lanczos algorithm for the linear response eigenvalue problem. Electronic Transactions on Numerical Analysis (ETNA), 51:529–546, 2019. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.51.2019/pp529-
                                        546.dir/pp529-546.pdf;
                                         http://etna.mcs.kent.
                                        edu/volumes/2011-2020/

vol51/abstract.php?vol=
51&pages=529-546.

Zeilberger:1999:PCC

[Zei99] Doron Zeilberger. Proof of a conjecture of Chan, Robbins, and Yuen. Electronic Transactions on Numerical Analysis (ETNA), 9:147–148, 1999. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
9.1999/pp147-148.dir/
pp147-148.pdf.

Zemke:2017:VIP

[Zem17] Jens-Peter M. Zemke. Variants of IDR with partial orthonormalization. Electronic Transactions on Numerical Analysis (ETNA), 46:245–272, 2017. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.46.2017/pp245-
272.dir/pp245-272.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/
vol46/abstract.php?vol=
46&pages=245-272.

Zhang:2000:PSC

[Zha00] Jun Zhang. On preconditioning Schur complement and Schur complement preconditioning. Electronic Transactions on Numerical Analysis (ETNA), 10: 115–130, 2000. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
10.2000/pp115-130.dir/
pp115-130.pdf. Multilevel

methods (Copper Mountain, CO, 1999).

Zhang:2001:NCP

[Zha01] Hong Zhang. Numerical condition of polynomials in different forms. Electronic Transactions on Numerical Analysis (ETNA), 12:66–87, 2001. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.12.2001/pp66-
87.dir/pp66-87.pdf.

Zhao:2004:CCC

[Zha04] Jie Zhao. Convergence of V -cycle and F -cycle multigrid methods for the biharmonic problem using the Morley element. Electronic Transactions on Numerical Analysis (ETNA), 17: 112–132, 2004. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu/vol.
17.2004/pp112-132.dir/
pp112-132.pdf.

Zhou:2006:MEG

[Zho06] Ping Zhou. More examples on general order multivariate Padé approximants for pseudo-multivariate functions. Electronic Transactions on Numerical Analysis (ETNA), 25:302–308, 2006. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/vol.
25.2006/pp302-308.dir/
pp302-308.pdf.

Zitko:2008:SRR

[Zít08]  Jan Zítko.  Some re-

marks on the restarted and augmented GMRES method. Electronic Transactions on Numerical Analysis (ETNA), 31:221–227, 2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.31.2008/pp221-
227.dir/pp221-227.pdf.

Zeybek:2017:ACM

[ZK17] Halil Zeybek and S. Battal Gazi Karakoç. Application of the collocation method with Bsplines to the GEW equation. Electronic Transactions on Numerical Analysis (ETNA), 46:71–88, 2017. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.46.2017/pp71-
88.dir/pp71-88.pdf;
http://etna.mcs.kent.
edu/volumes/2011-2020/

                                        vol46/abstract.php?vol=
                                        46&pages=71-88.

Zhang:2013:SRL

[ZM13] Jianjun Zhang and Benedetta Morini. Solving regularized linear least-squares problems by the alternating direction method with applications to image restoration. Electronic Transactions on Numerical Analysis (ETNA), 40:356–372, 2013. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu//vol.
                                        40.2013/pp356-372.dir/
                                        pp356-372.pdf.

Zhou:2017:SRV

[ZN17] Ming Zhou and Klaus Neymeyr. Sharp Ritz value estimates for restarted Krylov subspace iterations. Electronic Transactions on Numerical Analysis (ETNA), 46:424–446, 2017. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.46.2017/pp424-
                                        446.dir/pp424-446.pdf;
                                         http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol46/abstract.php?vol=
                                        46&pages=424-446.

Zhong:2017:WGK

[ZX17] Hong-Xiu Zhong and Hongguo Xu. Weighted Golub– Kahan–Lanczos bidiagonalization algorithms. Electronic Transactions on Numerical Analysis (ETNA), 47:153–178, 2017. CO-DEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

http://etna.mcs.kent.
edu/vol.47.2017/pp153-
178.dir/pp153-178.pdf;
 http://etna.mcs.kent.
edu/volumes/2011-2020/
vol47/abstract.php?vol=
47&pages=153-178.

Zhang:2013:NRB

[ZY13] Xuping Zhang and Bo Yu.

A note on the relation between the Newton homotopy method and the damped Newton method. Electronic Transactions on Numerical Analysis (ETNA), 40: 373–380, 2013. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

etna.mcs.kent.edu//vol.
40.2013/pp373-380.dir/
pp373-380.pdf.

Zhai:2020:FHI

[ZZW20] Jiayin Zhai, Zhiyue Zhang, and Tongke Wang. Fractional Hermite interpolation for non-smooth functions. Electronic Transactions on Numerical Analysis (ETNA), 52:113–131, 2020. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL

                                        http://etna.mcs.kent.
                                        edu/vol.52.2020/pp113-
                                        131.dir/pp113-131.pdf;
                                         http://etna.mcs.kent.
                                        edu/volumes/2011-2020/
                                        vol52/abstract.php?vol=
                                        52&pages=113-131.